A Complete Bibliography of Publications in Future Generation Computer Systems: 1984–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

10 June 2022 Version 3.00

Title word cross-reference

10 [ABB^{+}03]. 16 [Goo01]. 2 [CEGL01, HvHAS04, JNR01, NCS04, RBS93, VAS95]. 3 [CH95, EMB98, EdBG^{+}99, IdLR01, JHL^{+}06, KCK04, Kni89, KA88, MJ98, MJ06, RICW00, SGL99, WKF03, XYZ05, YMM00]. 4 [Avg00, DMM^{+}99]. $47.50 [Teb86]. ^{2} [LG08, OSCY93, ZY04]. ^{my} [DGS09]. D [PP06, SW02]. D_{2} [DRS04]. diam log(diam) [SW02]. ℓ_{1} [Tre03]. k [Pan95b, SAKOK03]. λ [LJY04]. AVISPA [BGC^{+}03]. μ [JD94]. N [Ref87, SvAS01, Pan95b, SAKOK03]. R [CPK05].

-ary [Pan95b, SAKOK03]. -body [SvAS01]. -cube [Pan95b]. -cubes [SAKOK03]. -D [Kni89, KA88, RBS93]. -dimensional [SW02]. -Expression [Ref87]. -opt [VAS95]. -orthogonal [PP06]. -phase [CEGL01].

.NET [AW03].

06 [Igl07].

1 [LF95a, Sap88]. 10 [MGH^{+}05]. 100 [SW06]. 1000 [SSMG95]. 123 [vdR86b]. 1983 [Ano84k]. 1986 [Zad87]. 1989 [Hul89]. 1990/91 [Ano91a]. 1991/92 [Ano92a, Ano92h]. 1997 [HB98]. 1998 [Lid99]. 1999 [Wil00].

2 [BT93, BP94, FHG95a, HBJ^{+}03, LM90b, Mal94, Por95]. 2001 [LBR02]. 2002 [DBdL03, GGH^{+}03]. 2005 [SDBdL06]. 21st [Mar98b, Mar98a]. 2nd [Ano86h].

3 [YTHY84]. 3/512 [Cro95]. 3P [CGSZ95].

4 [KSY92]. 4K [SST^{+}06, SKF^{+}09].

512 [Cro95]. 57.50 [vdR87f]. 5th [BYV^{+}09].

6000 [BBSV92].

7th [BGL08].

84 [Ano84e]. ’85 [Ano85a]. ’86 [Ano87j]. 860 [FHG95a, FHG95b].

90s [Dub91]. 92 [GD93a]. ’97 [Kaa98].

AAA [GdLvOT03, KKK07]. Aachen [BLB03]. ability [PIKM02]. Abstract [CPB00, DHS00b, Ném00, AC92, BM00, DDV92, DHS00a, DK00, WG00, dB90]. abstraction [KSM^{+}07a, KSM^{+}07b, MRV01]. abstractions [Kea99]. Academic [Teb86, vdR87f, vdR87i]. Accelerating [VBLS09]. acceleration [VF01]. Access [BX04, WZC08, AS99, ABF^{+}03, BCW01, CC98, DAM08, Doḡ09, FX07, LHC03, PSVL02, PH07, Sin07, Wah84, Wal86, WC06b, WC06a, ZMN99]. accessibility [DFG^{+}00]. accident [MCA02]. accounting [PGPW09, SRG^{+}03]. accurate [BBL^{+}05, KSAOK08]. achieve [CWD04]. achievements [Ano87j, Nis93]. Achieving [GE90, LSLS05, HRJ^{+}06]. Acknowledgement [Ano07, Ano08]. ACM [BGL08]. across [BCPS03, PBC^{+}01, Sin07]. action [Cad86, GdBW06, KFP^{+}02, vVDBB98]. Activated [SHJ06]. Active [BBG^{+}05, CKK^{+}04, BGL^{+}05, HC99, MTKS00]. activities [vdR87l]. ad [GR07, LBYL08, LM07, MV09, SM01a]. ad-hoc [LM07, MV09]. adaptability [MC04]. adaptable [GL04b]. adaptation [Reu03a, SGH^{+}08]. adapted [JLU03]. Adapting [AG05]. Adaption [SLS^{+}09, FM01]. Adaptive [Aba06a, ABF93, FN00, GRH05, NP03, Aba09, AS02, BMRW01, CST92, CS05, Cho04, DZ04, HWZL08, HJK^{+}04, KSAOK03, KSAOK08, KU01, LSTV07, NWE04, RSR01, SLW01, TSK03, VR05]. adaptivity [SOR05]. Addison [Zem86, vdR87g]. Addison-Wesley [vdR87g]. address [DL00, SVC^{+}07]. addressable [De 88]. addressing [DvdHdL06]. ADIC [HNS05]. Adjoint [GKS05, HHG05]. Adjusting [YYW^{+}09]. adjustment [HPLL09, LNJ04]. administration [MSLP93]. administrator [HY09]. admission [KLM^{+}03, SMA08]. advance [CFG^{+}05, ET08]. Advanced [Gil85b, SZGbC04, ADT03, Ben99, Hab05, KKL09a, Mam09, NHG02, NHG03, Ano84a, CMZ95, EGK^{+}07]. Advances [vdR87b, vdR87a]. Advertisement [Ano05e]. aerodynamic [BBC^{+}99, LKG08]. Aerospace [LPC^{+}95, Mur95, Pet89]. AFCET [Ano85a]. AFCET-Informatique [Ano85a]. after [BORM07]. agency [NTN86]. Agent [FR08, TA96, CTT^{+}08b, HB08, LL03, MJ00, NWE04, PSA^{+}09, FM08, Gra92, LWHS07, NMC05]. agent-based [CTT^{+}08b, HB08, MJ00, NMC05]. agents [AMH02, CWD04, CSJN05, ESPP01, GWO03, HQ07, KRLR01, Sch00, UTT00, BMS05, WLB00]. aggregates [NCS04]. agreement [LM07, LKA^{+}08, YKL^{+}07]. agreements [BSCC06, LJW08]. agriculture [KH89]. ai [Pud87, Fuc93, Lau92, MJ00, Oku92, Poh87, TC92]. AI-based [Lau92]. Aided [dRSBH94]. airborne [HYS04]. airfoil [BBL^{+}05, GKS05]. Airport [RdSH^{+}00]. AIV [AFPG91, AFP^{+}92]. AKL [HJP92]. Albatross [KBM^{+}02]. algebra [BFR05, BCG05, Dal03, DHD89, Igl07, Lop03, PH94, vdV89a]. algebraic [SSC04]. Algorithm [ABF93, VSvD95, Aba06a, CCL09, DS04c, ESFD06, GKIZ05, Ger02, HZC^{+}08, HHS98, HM98, JM01, LCP04, LKG07, LYQ06, LLWN04, LLZ07, LD04, Mar02, MP02, MVS00, NP06, OVDV98, Par04, RICW00, SK04, Sch03, SMA08, SCK^{+}00, SO98, SVB07, TMT^{+}07, VAS95, SMI01]. Algorithmic [CCG07, Hab05, Yos89]. Algorithms [ABMS05, Ber06, DHD89, AT01, BPS06, CFG93, DE03, DP03a, DQ97, DRNMC09, Din99, Doḡ09, DMW04, DBT00, DDS00, ET08, Fio06, GRH05, GJS^{+}94, GODM98, GA06, HLvL^{+}97, HV03, JS89, KKP^{+}05, LR01, LKG08, LC01, ST99, TLYT05, YJA03, LOJ^{+}07]. Alignment [MFE^{+}08, CS05, KSM^{+}07a, KSM^{+}07b, Sch01, YD05]. all-by-all [BORM07]. all-optical [Pal06]. Allocation [DFC^{+}08, HPLL08, LL04c, VDPHS09, Vau93, WCC^{+}09, YBQ07, YYW^{+}09]. almost [LD04]. Alternative [BDL06, DFT92]. Alvey [Ano84a]. Alzheimer [KKP^{+}05]. ameboid [Len01]. among [ELvD^{+}96]. amorphous [Ole07]. amount [PCB99]. AMPIC [CCHW03]. amplified [SLDK03]. AMROEBA [MGYC06]. Amsterdam [vdR87c, vdR87e, Ano86i, Ano87b, Ano87c, Ano87l, Baa87]. Analysis [SAKOK03, SGY^{+}07, AS02, ABB^{+}03, AMW99, ABK94, BCT^{+}07, BPP^{+}07, BPS^{+}03, CP06, Cho04, DDL01, Fah98, Fio06, GI07, GL05, HNP05, HKM^{+}06, HJK^{+}04, JAA09, JSS^{+}99, KZLK06, KN06, KV03, KU01, LOK09, LGMV02, MOK06, MLZ^{+}00, NJW^{+}06, PH07, Qin07, RL98, SOR05, SLS^{+}09, SK06, TBNF09, VBLS09, WSS^{+}09, WBF08, Yam89, ZWJ04]. Analytical [JAA07, ZY04, KSAOK03]. analyze [BPAP92]. Analyzing [AT01]. and-or [GC94]. AND-parallelism [ZS90]. anddata [CC98]. anddata-parallel [CC98].

ANEJOS [SM01a]. aneurism [HORC04]. angiographic [RICW00]. animal [Mor01]. animation [NMA00]. animations [DK00]. anisotropic [KZC04]. Annealing [GODM98, VSvD95, SZ98]. Annealing-based [GODM98]. Announcement [Ano84b, Ano87a, Ano93b]. announcements [Ano93c]. anonymity [Opp00]. Ant [DBT00, DDS00, SH00, TRFR01, CCL09, DS04c, MCA02, Gut00, PW09]. Antipolis [Zna94]. ants [MVS00, MC00, WLB00]. AnyLogic [BKK02]. apicalis [MVS00]. appears [Hen87]. Application [BHRT98, DRNMC09, JNR01, NB04, PKC^{+}05, SBA^{+}05, TKT^{+}08, YWA^{+}89, ASTEP98, AKPN01, ABB^{+}03, BMS05, BMFC07, CGSZ95, CW93, CDRS05, DDV92, DF97, DR05, Gra92, Hab05, Hir89, LW08, Low05, MM03, PO00, PPJ95, RRS99, SM03, SAMN02, SGdMM96, Sin92, SLS^{+}09, SLZ95, WBF08, BBC^{+}99, MS01, YPF05]. application-specific [DR05, WBF08]. Applications [Ano86i, CHK98, CSP98, DDM^{+}08, KLM^{+}05, dRSBH94, AW03, AB01, ATJMZ02, AD00, Ano87b, BMRW01, BBFW03, Ben99, BPS06, BB06, BKSS02, BLAV06, BCW01, BCG05, BFK02, BFW^{+}03, BGK^{+}05, Büc05b, Car03, CSW06, CBK^{+}01, CMZ95, CCHW03, CN98, CRM05, DZ98, DHB02, DMG^{+}08, DFG^{+}00, Dub91, EMB98, EO86, FHM^{+}99, FS07, FdSC07, GHWZ94, Gil85a, GMA07, GMB^{+}05, GVD^{+}03, HY09, HRSW99, Hua05, Hum92, HMC06, Igl07, KT08, KA09, KZLK06, KST92, KS02, KMK09, KB00, KKP00, KB09b, LB03, LC04, LS08, Luk00, MTNM08, MWC^{+}03, Mam09, MVRM08, MZC08, MLD08, MVT^{+}99, MM08, MTH^{+}05, MLZ^{+}00, NHG06, OS01, Pal06, Par94, PBHK01, PA01b, Pro07, RBS93, SBHD08, SG05, Ser95, SY04, SD06, SL97, SPM86, SLJ^{+}06, SSC04, SPEW09, SSK^{+}08, SSMG95, TJLT00, WS05]. applications [WL05, YHJC05, YLC^{+}06, ZBB09, dKdOS03, vdR87c, vdR87d]. applicative [OBK88]. applied [KG01, Szu98, WAE06]. Applying [CFVP03, UB07, ACML05, NWE04, OCdAM07]. Appraisal [Tic93, vdR93b]. approach [AEGF^{+}01, AGP^{+}92, Ale97, BP01, BRS04, BAC02, BCPS03, BVDF00, CPK05, CM99, Del06, DS00, DAM08, DC00, ED04, GKIZ05, HAAH05, HBH09, Hol93, KA08, Li90, LC05, LGMV02, MD92a, MCT^{+}09, MZC08, McC96, MLD08, MC04, NSP07, Nos98, PPJ95, PA01b, SB97, TC92, Var03, Vau93, VGBLGS^{+}06, Ven08, Yos89, ZS90, ZL04b]. Approaches [HXL90, BDL06, DL03, Hab05, Meu05, Pet95, STP^{+}05, ZZN04]. approximate [MM03]. approximating [SK04]. Approximation [Tho06, Tab06]. approximations [Gue01]. architect [GLSV07]. Architectural [GS95, Niw89, SSK^{+}08]. Architecture [FP03, Ger02, Her84, HKT94, RSRV88, AKB^{+}01, AG92, AGP^{+}92, AMW99, BBG^{+}05, Bhu95, BGL^{+}05, CST91, CDF^{+}05, CBK^{+}01, CS96, ES94, GD05, HHS98, HO02, HML07, HML09, Kat04, KLH^{+}04, KKK07, Kim07a, Kim07b, Kob92, KGLA85, LG08, LHL03, MH01, Mar90, MHA08, MPPM09, Mur88, Ném00, OBK88, OFT09, PO00, Par06, PSR^{+}07, PSA^{+}09, PSS01, SA07, SZC05, SSS02, SD03, SHJ06, TJLT00, YCX05, GMEL08]. Architectures [HV84, Pfi99, BFLL99, BSCC06, BW95, BSG^{+}05, CG09, Din91, Gil85b, Her87, JBA94, Kun94, Lop96, Lop93, MSLP93, SGdMM96, TR85, UWV92, WBT05, ZMN99, vM94]. archive [Fin99, GML99, ILJ^{+}08]. archives [AMW99, HC99, SB99]. archiving [PCM99]. Ardent [LM90b]. area [GGH^{+}06, GGSZ09, LRJ^{+}06, LWSC07, OS01, PSVL02, RRS99, VBLS09, vdR87l]. areas [Hab05]. arguing [Sch00]. Argument [SWCL95]. ARIANE [BPAP92]. ARIANEXPERT [BPAP92]. Arigatoni [CCL08]. arithmetic [FGG03]. Armada [OK02]. arms [Poh87]. Arranging [Leo98]. Array [CCKW88, VV92, Mur88, Pan95a, PHL98]. arrays [Dui89, EFD00, Van92]. ARS [TSK03]. art [Van87b, VLC03]. arterial [DIK^{+}06]. arthroscopic [WWSM98]. Artificial [All92, Ano84e, Kow85, Sch85, vdR87a, Ano86l, ED04, EO86, How91, JL98, KRLR01, Kow84, Lop96, Niw89, RBC^{+}88, Wii84, ZZN04, vdR87b, Ano87b, Ano87c, vdR87c]. Arts [BHD09]. ary [Pan95b, SAKOK03]. aspects [BCPS03, LLRS94, Lop03, RGH^{+}01, vdV89b]. ASPEN [LM90a]. Assembly [KM01]. Assessment [PM04, MLZ^{+}00, ZL04b]. assignment [LHC03, MC00, RN01, THKG98, TRFR01, Var03]. ASSIST [Ueh89]. assistance [Ohy89, SOR05, Tak89b, TD95]. assistant [KFF89]. assisted [LKG08]. association [SA97]. Associative [CC98, NSI84, Ami90, WGL92]. Astronomical [SB99, GML99, OFO^{+}99]. astronomy [HJPS03, PCM99, SLJ^{+}06]. astrophysical [RBS93]. astrophysics [MGYC06, BAD^{+}05]. asymptotically [CKR04]. Asynchronous [SM01b, AT01, DOV01, ESFD06, MMR02, Pap05]. ATLAS [KMCH03, Uch86]. ATM [MK04]. Atmospheric [AKMK05]. atomic [Tor04, OB04]. Atomicity [WLF^{+}09]. ATOP [SGH^{+}08]. ATOP-Grid [SGH^{+}08]. ATOS [Lau92]. ATREX [Tak89a]. attempt [SLZ95]. Attractor [Ami90]. Attribute [KH97, SH90, GGM^{+}09, SHLB08]. attribute-based [GGM^{+}09]. attributes [VLK09]. audiences [PNH99]. audio [MSK03]. audiovisual [VWD^{+}08]. Auditing [SK97, BSCC06]. augmented [BB04]. AUGUSTUS [SLS^{+}09]. Authentication [BDFP05, LHL03, MR00]. Authenticity [CBD^{+}05]. Author [Ano85b, Ano86a, Ano87d, Ano89a, Ano90a, Ano91a, Ano92a, Ano92b, Ano93a, Ano94a, Ano94b, Ano95a, Ano98a, Ano01a, Ano02a, Ano03a, Ano04a]. authoring [Tak05]. authorities [HWW04]. Authorization [GdLvOT03, ACC^{+}05b, GXD^{+}09, GGM^{+}09, LC05]. Automata [TS99, BMPS01, Ban02a, Ban05, Bog99, CM99, DS99, FW02, GZ04, HRSW99, JM02, KCT99, Mar02, Ser98, ST99, TSZP99, Wor99, BP02, CDRS05, PIKM02]. automate [TC92]. Automated [Bea03, FM01, Sti93]. Automatic [Büc05b, Kom89b, MHA09, Reu03a, Röb05, AC01, Büc05a, Fau05, HNP05, HHG05, HNS05, SK04, SBLT05]. Automating [BJA^{+}05, MH01]. automation [DDV92]. automaton [Ban02b, DDL01, SKT02]. Automobile [Tak89a]. Autonomic [LG08, BJWZ08, FS07, FLPP05]. Autonomous [YKL^{+}07]. Autopilot [RSR01]. Autostereoscopic [PSG^{+}06]. availability [BCB^{+}07, CBD^{+}05, KFC^{+}07, LSLS05, LVH08]. available [Din03]. Avatar [GC00]. avoid [GL05]. AVS [WKF03]. AVS/Express [WKF03]. aware [HZC^{+}08, HY09, KB09a, MMVV08, Pal09, PKC04, PSR^{+}07, CG09]. axisymmetric [CASW05, de 94].

B [Zem86, vdR87e, IEG04]. B-spline [IEG04]. back [GKIZ05]. Backbones [Sch03]. Backplane [BBM^{+}03]. backplanes [Mam09]. backpropagation [JL98, RM97]. bad [WBT^{+}08]. bag [WL05, DFC^{+}08]. bag-of-tasks [WL05]. balanced [CCL09, DL03]. Balancing [OSHH96, dRSS97, ADOKM06, BCMR01, BM08, BL02, CSJN05, CY01, Cho04, KNK^{+}08, KMK09, LYMZ09, MM03, PGSM05, Qin07, SB97, SMA08, ZMS^{+}06]. balls [LGMV02]. Bandwidth [VS04, CGL08, CALN03, NF07, PBHK01]. banking [GPA00]. banks [ABCD00]. barrier [Pan95b, SO98]. Base [BMFC07, Nis93, ZY90, vdR87l]. Based [GSD95, YTHY84, ÅO06, ADM06, AMH02, AAC04, ADOKM06, ATJMZ02, AMW99, AB95, Ano86j, AN08, BvdBM^{+}93, BMPS01, BP02, BMFC07, BP01, BKG05, BGH^{+}03, BKSS02, BRMN04, BS09, Boa04, BPAP92, Bru01, BCB^{+}07, CSC^{+}05, CYLT05, CdCD07, CBK^{+}01, CO03, CEGL01, CKK^{+}04, CGH04, CTT^{+}08b, CY90, CY88, CSC^{+}92, DSS98, DVVD02, DE03, DMM^{+}99, DMG^{+}08, DCK03, DS08, DT93, DRNMC09, DV03, DR03, Dör05, DS04c, ET08, EO86, EGAQ09, FMS08a, FMR05, FLPP05, FW02, FCW01, GHWZ94, GODM98, GMC03, GdLvOT03, GBE00, Gra01, GGM^{+}09, GWO03, GZ04, Gut00, Hal88, HZC^{+}08, HB08, HAFF99, HAF00, HJS^{+}99, HHS98, HBH09, Hol93, HYS04, How91, HPLL08, HMC06, IMKB89, JLU03, JFDF09, JCSS01, JC09, KG01, KFF89, KA08, KCK04, KN06, KKV^{+}99, KB00, KLSS05, KY04, Lau92, Lee04, LKG07, LLKF09]. based [Leo01, LRW01, LB09, LY90a, LM07, LHB95, LW08, LKA^{+}08, LJPS05, LLZ07, LS01, LC03, LLSR02, LSF^{+}94, LMH^{+}09, Mae89, MPH00, MCSS00, MVRM08, MJ06, Mar98a, MLD08, MJ00, MSX00, MHA08, MHA09, Mor01, MPPM09, MK95, NKX09, NMA00, NSP07, NHG06, NHT06, NMC05, Ohy89, OCdAM07, OVK^{+}09, OP97, PO00, Pal06, Par06, PSP^{+}09, PdLS^{+}99, PSS01, PPAK99, PN09, QP08, RHB08, SM01a, SPK^{+}07, SEH99, SZC05, SH99, SGL99, SM03, SMC99, Sch01, STTK03, SBHD08, SG05, SMA08, SYT09, SNA92, Sun92, TBNF09, VVC^{+}03, Vau93, VGBLGS^{+}06, Ven08, Ven09, VLC03, VM93, WSS^{+}09, WCC^{+}09, YdOLS^{+}05, Yos89, YCX05, ZSI08, ZL04a, ZMN99, dLLA93, mM95]. Bases [Wie85, NSI84]. basic [Dek86, LM90b]. basis [HAAH05, NK05, SOR05]. batch [LH07]. batteries [BRS04]. Bayanihan [SH99]. Bayesian [Kło05]. BBS [FHG95a, Mal94, PR95, RS94, Zna94]. be [HNP05]. beam [GL04a]. bearing [FN00]. BECAUSE [Zna94, BP94, Zna94]. bed [RRS99]. Beerel [vdR87d]. began [Bur02]. behave [NMC05]. Behavior [BP01, BSG^{+}05, JM02, ZZ09]. behavioral [Ale97, MS01, NMC05]. behaviour [Sch01]. Benchmark [CGSZ95, Zna94, LM90b]. Benchmarking [EGCY^{+}06, Por95]. benchmarks [WYJ99]. benefit [BGV97, MVT^{+}99]. Best [FCW01, Gra01, IdLR01, PF01, PSS01, SM01a, LBR02, WBT^{+}08, Bru01]. between [BRR^{+}04, DMN^{+}05, HICÁFM^{+}06, LPE08, Nag86a]. Beware [Chv87]. beyond [LDSH95]. BGK [AKH^{+}04]. bi [LOK09]. bi-directional [LOK09]. biconnected [AAC04]. Bień [vdR87i]. bifurcation [AKH^{+}04]. Binary [RSRV88, CC98, LLZ07]. binary-search [CC98]. Bingham [Ned06]. Bio [ZEO98, LWHC07]. Bio-inspired [ZEO98]. Bio-STEER [LWHC07]. biochemical [KCT99]. biochemistry [Rou00]. Bioinfomatics [MFE^{+}08]. Bioinformatic [SF06, BCMA07]. bioinformatics [ABM^{+}07, CCM07, SSS02, SLS^{+}09]. Biological [MB01, YD05, Ami90, FGCM07, Kol89, MP02, WRBG94]. Biologically [Pet95, Meu05]. Biology [LGS^{+}07, AV00, CCBR98, Gue01, KG01, SJTG07]. biomedical [CSC^{+}05, Ros89]. biomedicine [ABM^{+}07]. biometric [MR00]. Biometrics [Ale97, Pol99]. biomolecular [NJW^{+}06]. BioSim [CSC^{+}05]. BioSimGrid [NJW^{+}06]. Bishop [Ano86i]. bistability [vdR87e]. bit [KKL09b]. bitmap [SK04]. BitTorrent [HWZL08, WFC07]. black [SHB89]. blast [YWA^{+}89, BPC^{+}01]. blending [ZY04]. block [BFK02, EV98, LCP04, SHJR04, VF01, VFS01, mM95, vM94]. block-preconditioned [VFS01]. block-structured [BFK02]. Blocking [BCH^{+}08]. blocks [GLSV07]. blood [HORC04]. bluff [SdSP04]. Board [Ano84g, Ano86f, Ano87i, Ano88c, Ano89e, Ano90f, Ano91d, Ano92g, Ano93h, Ano94f, Ano95g, Ano96d, Ano97c, Ano05c, Ano05d]. body [SvAS01, SdSP04]. Boltzmann [ABL04, CM99, CH04, DS04b, DY04, FPdS04, HORC04, HRJ^{+}04, IOO04, KKHS01, NCS04, SdSP04, vdS04]. bonds [BRS04]. bone [BPP^{+}07, RMM^{+}98]. Book [Ano93c, Ano93b]. Boosting [PH94]. Boston [vdR87k]. both [FB93]. bound [JS89]. boundary [BS04, DS04b, Kni89, SK04]. Boussinesq [Tab06]. brain [KZC04, MLSO01]. branch [JS89]. branch-and-bound [JS89]. branches [LTOT07]. BRAVE [RBC^{+}88]. brief [Miz89a]. Bringing [ZS05b]. broadcasting [KY04, VS90]. Brockport [Lit03]. broker [ABG02]. brokerage [GPA00]. brokering [ET08]. Brownian [vOB95]. browsing [FGRZ09]. BSP [HS98, HLSØ06, HM98, McC96, RS98]. bubble [IOO04]. Buffer [KD00, WWC^{+}97]. build [LL03]. Building [LF01, Ram95, SGP^{+}09, SW06, YPF05, EV98, FS07, GDRS04, GLSV07, Niw89, SH99, VM93, WYN^{+}90, Kos95, Zem86]. builtins [CYB90]. Bulk [GL05]. bus [AB95, Pan95a, PHL98, Shi92, UWV92, Kat04]. bus-based [AB95]. Business [EO86, KF00, CM01, MVT^{+}99, SGY^{+}07]. Butcher [MKK03]. Byte [LTOT07]. Byte-code [LTOT07]. bytecode [Ber00].

C [vdR87a, vdR87d, AC01, BRS04,

CCHW03, LC04, NK07, WYN^{+}90, ZY04]. C-oriented [WYN^{+}90]. CA [SW02]. Cache [GSD95, HMC06, BSG^{+}05, DSS98, HBCR01, Kun94, Lop93, VBLS09]. cached [Yam92]. caches [ADM06, TD95]. CAD [Dae95, TY85, TG04]. calculations [BvdHN^{+}01, Gil94, LF95b, Mal94]. Calendar [Ano84c, Ano84d, Ano85c, Ano85d, Ano85e, Ano85f, Ano86b, Ano86c, Ano86d, Ano86e, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano89b, Ano89c, Ano89d, Ano90b, Ano90c, Ano90d, Ano90e, Ano91b, Ano91c, Ano92c, Ano92d, Ano92e, Ano92f, Ano93d, Ano93e, Ano93f, Ano93g, Ano94e, Ano94c, Ano94d, Ano95b, Ano95c, Ano96a, Ano96b, Ano97a, Ano97b, Ano98b]. Calendar17 [Ano96c]. Calendar99 [Ano95d]. Calendarneous [Ano95e, Ano95f]. CALIFE [SLZ95]. CAM [Avg00, SdR99]. CAMAS [dRSBH94]. Cambridge [YHJC05]. can [BGV97]. Canada [KMCH03]. Canadian [AAB^{+}07]. cancer [LGS^{+}07]. capabilities [BBLP05, DGST09, SD06, WAD^{+}89]. Capacity [AMD08, Tho06, WWC^{+}97]. car [BBC^{+}99]. carbon [Tor04]. card [AHdJF97]. cardiac [DMM^{+}99, KE85]. cards [Ale97, BGV97, DQ97, DF97, HJCD05]. care [Pol98, PPAK99]. Carlo [BST^{+}08, TMTY05]. CARPET [ST98]. CAS [GI07]. CASA’2003 [Igl07]. case [BBSV92, EGCY^{+}06, KE85, MMVV08, RHB08, Rou00, SMI01, VM93, WX02, WM07]. case-based [VM93]. casein [vOB95]. cases [de 94]. catallactic [JRF^{+}07]. catalog [GML99]. catalogs [OFO^{+}99]. cations [DRS04]. causal [AMH04, ZCT^{+}04]. cause [DLW07]. CAVE [DDS^{+}09, MvLvW98]. CAVE^{T M} [WKF03]. cavity [CASW05, VWCV94]. Cayley [DR03]. CCGrid [LBR02]. cell [BVDF00, GL95, Pal01, SJTG07]. cell-vertex [GL95]. cells [BPP^{+}07]. Cellular [Ban02a, BP02, BMS05, Ban02b, CM99, KCT99, Mur88, PIKM02, TS99, BMPS01, Ban05, Bog99, DRS^{+}97, DS99, DDL01, FLPP05, FW02, GZ04, HRSW99, JM02, Mar02, PSL^{+}04, Pud01, Ser98, SKT02, ST99, TSZP99, Wor99, CDRS05]. cemetery [MCA02]. center [CYH04]. centered [AP96]. centers [Han03]. centre [Gal87, BT93]. centred [AMB03]. Century [Mar98b, Mar98a]. certification [JLU03]. CFD [LF95a, SLZ95, WS05, YHJC05]. chain [BBL^{+}05]. chains [JFDF09, JLMR00]. Chalk [GDRS04]. challenge [BKM03, Her87, Rum99]. Challenges [Ald89, SJTG07, Ano84h, BHD09, FS97, Fre94, Hul89, Kol89, KB09b, LGS^{+}07, Meu05, Pet89, RdLM06, Rho89, Wil89]. changes [BBJ^{+}06]. Changing [DMSS97]. channel [DS04b, SG05]. channels [AJZ^{+}02, Dua94]. chaotic [GI07]. character [CSC^{+}05]. character-based [CSC^{+}05]. Characterization [ZCT^{+}04, KB00]. Characterizing [KFC^{+}07]. characters [SK04]. Charging [SRG^{+}03]. Charlotte [BKKW99]. chat [MSK03, DS00]. Cheap [HWS07]. checking [LYT^{+}05]. checkpoint [Dal06]. checkpointing [BCH^{+}08, FC05, RG04]. Chem [GAB^{+}96]. chemical [MSKT07, SBLT05]. chemically [DY04]. Chemistry [Höf03, Bro92, GAB^{+}96, LDSH95, GMB^{+}05]. Chichester [vdR87d]. CHIMP [BCM^{+}95]. chip [SF06]. chips [MPH00]. choice [BHRT98]. Cholesky [IST04]. CHORUS [AGP^{+}92]. CHORUS/MiX [AGP^{+}92]. CIIAM86 [vdR87b]. cinema [SST^{+}06]. ciphers [SHJR04]. circuit [BJNH05, GODM98, GDRS04, IdLR01, SKT02, WVC05]. circuits [CFVP03, Dör05, WSTW87]. Circulation [HHG05, CS93]. city [vVDBB98]. class [Čie04, TKT^{+}08]. classification [CHJ^{+}04, CD99, Kło05, LLWN04]. Clause [LY90a]. client [CGL08, PA01b]. client/server [PA01b]. Cliffs [vdR87h]. climate [GP09]. clinic [PPAK99]. clinical [OCdAM07]. clock [AC92]. Cloud [BYV^{+}09]. clouds [GGSZ09]. Cluster [BJC02, DT08, OSHH96, WX02, Aba06b, ACML05, BL98, Cho04, DZ04, DVVD02, Fra08, GS05, Goo02, HO02, JAA09, LSLS05, MM03, MGLV04, NSI02, RG04, STH^{+}98, SMI01, SMA08, VSM02, YJA03, ZJWZ04]. Clustered [BHH^{+}93, PSL^{+}04]. Clustering [Mic97, DB99, FT07, KCK04, LBYL08, STP^{+}05, ZM97]. Clusters [PBM95, BBSV92, BL02, BPC^{+}01, CRE01, CG02, CKFJ06, CEGL01, ELvD^{+}96, Fer96, Gos00, GVD^{+}03, LP01, MTKS00, MLSO01, OB04, PCG^{+}06, PL96, RHB08, RT06, SVC^{+}07, STTK03, SD03, TC06, SZGbC04]. CM [Mal94, Por95]. CM-2 [Mal94, Por95]. CM2 [CH95]. CNES [BT93]. Co [PW09, Ano84i, YYW^{+}09]. Co-allocation [YYW^{+}09]. co-operative [Ano84i]. Co-scheduling [PW09]. coalition [FX07]. Coarse [Vre88, VF01, Vre89]. Coarse-Grain [Vre88, Vre89]. CoConut [Reu03a]. CoConut/J [Reu03a]. code [BP01, BST^{+}04, DR05, DMN^{+}05, GKS05, GL95, IJLC03, LTOT07, LLZ07, LN94, RBS93, SLZ95]. coded [LLWN04]. Codes [SvAS01, BBW08, BVDF00, Dal03, LPB04, LRW01, Niw89]. coding [CCM98, HJK^{+}04, SMC99]. coefficients [Tab06]. cognitive [HQ07]. Coherent [Tul04]. COHESION [SBHD08]. collaboration [JHL^{+}06, SS04, SLDK03]. Collaborative [BGJ^{+}06, BLAV06, SZP00, TWC^{+}06, BPS^{+}03, GP09, HB09, HAB^{+}06, KPB^{+}03, LPC^{+}95, LGS^{+}07, RJH^{+}09, SCY01, SLW01, VGBLGS^{+}06, VWD^{+}08, YLC^{+}06, ZL04b]. Collaboratory [BAD^{+}05, LRJ^{+}06, LJPS05, SGP^{+}09, SBdL09, KFP^{+}02]. collections [AFP07, GB99]. collective [JM02, NSI02, Szu98, Szu01]. Colonies [TRFR01]. colonoscopy [MSR98]. Colony [PW09, DS04c]. Columbus [Gra92]. COMA [SWCL95]. combination [WL05]. combinator [WGL92]. combinatorial [Sch03]. Combine [LL03]. combined [HJK^{+}04, ZS05b]. Combining [LC01, GY90]. command [KTY03]. Comments [WC06b, WC06a]. commerce [ABCD00, AS99, KF00]. commercial [Mar86]. commitment [KKL09b]. commodity [GVD^{+}03, SVB07, WTC^{+}02]. common [DLW07, FB93]. Communication [CSP98, DHS99, Kal94, Tis07, ADOKM06, BCFS02, CST92, DPP03, DT94, GS95, GD93b, HO02, HML^{+}06, HJK^{+}04, JAA07, LR06, MD92b, MTKS00, MRV01, Pag99, PR95, PPAK99, WC01, WTC^{+}02, ZDR07]. Communication-efficient [Tis07]. Communications [DVV90, CYH04, Ste94, SHJ06]. communities [RSSD02, Var00]. Community [Car86, LBB^{+}09, GP09, Mar99a, PPH^{+}09, SBG^{+}09]. community-driven [SBG^{+}09]. Commutator [CMO03]. Commutator-free [CMO03]. comparative [Bal92, KDE04, OP97, SK06, ZM97]. Comparing [KGX95]. Comparison [VSvD95, BNFZ08, CCG07, STP^{+}05]. comparisons [BORM07]. compatibility [SSB05]. compatible [DMMP98]. Competent [DD86]. Competitive [LL04c]. competitors [Ano84h]. Compilation [BM00]. compiler [CMT01, DSS98, LY90a, Por95, SO98]. compiler-based [LY90a]. compiler-controlled [DSS98]. compilers [EGCY^{+}06]. compiling [ZS90]. Complete [CC07, MCQ^{+}07]. completion [CDS03]. Complex [BKS98, CCKW88, SdR99, WH05, AB01, BJWZ08, DS99, FGCM07, Fre94, HAAH05, MB01, Meu05, SJTG07, XSM04]. compliance [Niw89]. compliant [SYT09]. complicated [LHC03]. component [AAC04, BKSS02, Fio06, GW01, Lee04, LKG07, Par06, PSS01, Reu03a, Tak05]. component-based [BKSS02, LKG07, Par06]. component-oriented [Tak05]. components [CY88, DD05, Kom89b, LRW01, PA01a, WSTW87]. composer [BGK^{+}05]. composing [Kom89b, PADD03]. composite [BRMN04, WCC^{+}09, ZMS^{+}06]. composition [GBA^{+}09, YKL^{+}07, ZZ09]. compositional [VR05]. Compositionality [dB90]. Compressible [KN06, Ano96b, LF95b, RŻDM01]. Compression [GLS99, DQ97, KCK04, KKV^{+}99, SMC99]. Comput [KSM^{+}07a, MR04b]. Computation [Ban05, CASW05, DE03, Fre94, AAC04, Amo06, BCFS02, BDHK06, DT94, EL98, GBA^{+}09, GPH^{+}94, HWS07, HHSW92, MCT^{+}09, SD02, SLO^{+}05b, dlTK92]. Computational [ABMS05, Bis96, Bro92, Bun03, CH04, Joh02, MCSS00, MGYC06, MVAS89, MR04a, MR03b, MvLvW98, Pet89, RGH^{+}01, TS99, Tan02a, VMvW97, ABG02, AAB^{+}07, Ald89, ABF^{+}03, BKM03, BFR05, BLB03, CM99, CH95, CKR04, DT08, EGK^{+}07, FGG03, FvLTT98, Han03, Hul89, Joh89, JNPY06, Kol89, LL04c, LW08, Lit03, LJPS05, Lop03, MDD89, Meu05, MHA08, MvWvL99, NSHP88, NP03, OK02, PIKM02, Rho89, SMK05, SHJR04, SL97, SGH^{+}08, TMT^{+}07, TV08, VDPHS09, WFC07, WL05, Wie03, Wil89, YD05, ZCW^{+}04, CCHW03, GMB^{+}05, KFP^{+}02]. computationally [Pet95]. Computations [VV92, BBSV92, BBJ^{+}06, BST^{+}08, KL02, LF95a, Szu98]. Compute [GGSZ09, BGR^{+}99, SBLT05]. Computer [AKW90a, Ano86i, CT09, CD99, Igl04a, Igl07, KI89, NHG03, Slo06a, Slo06b, Sti93, Zin00, vdR93b, Ais88, AKW90b, Bhu95, BLB03, Bur02, CEJK94, Dal03, DRNMC09, Fur92, Her87, HXL90, Kal94, Kas85, KB92, LC01, Lin84, LSF^{+}94, MJ00, PO00, Pri95, Pud87, SEH99, SL97, SBLT05, Tak89b, TDL05, WG91, Zad87, ZGCM00, vM94, vdR93a, dRSBH94]. computer-integrated [WG91]. Computerized [KKP^{+}05]. Computers [Her84, Ano86i, CST92, CS93, DRNMC09, DD86, Omo91, OP97, Pad92, RN01, ST99, vdR86b]. Computing [AT02, Ama88, Ama89, CDF^{+}05, FVFA98, Gen95, HB98, HY03, Kow85, Lid99, Mes02, OSHH96, SG95, WMN^{+}01, Wil00, AvLR92, Aba06b, Aba09, ABG02, ACGdT02, ACC^{+}05a, AB01, AMD08, AMH04, AFF^{+}09, AG05, Ano90g, AB03, BGL08, BL98, BBG^{+}05, BBC^{+}99, BKG05, BR92, Bis94, BCM^{+}95, BJC02, BYV^{+}09, BCB^{+}07, CGT07, CTF^{+}99, CWD^{+}08, CGH04, CC09, CTMO06, CFG93, CS09, DHB02, DLR^{+}09, DVVD02, DL03, DRS^{+}97, DKD08, DDM^{+}08, DR89, DSS07, DT08, ESFD06, EW97, ESPP01, Sou91, GJS^{+}94, GS05, Gil94, GL95, Gos00, GAB^{+}96, GNWT05, Hab05, HDC^{+}94, Her91, HMP04, HG92, IT05, JSK^{+}06, JAA07, JAA09, JNPY06, JM01, Kar01, KBM^{+}02, Kim07b, hKcF09, KV09, KB09a, Kos95, Kow84, Lau01, LSLS05, LWHC07, LYMZ09, LOJ^{+}07, LDSH95, LJW08, LLSR02, LLSH07, Mal01, Mal02]. computing [Mal05, MI01, NSS99, NCCS99, OF07, Ole07, OS01, PWY03, PP06, Ray05, RBS93, SH99, Sar02, STH^{+}98, Shi04, SSF^{+}09, SLS^{+}09, SD07, Ste94, SRCR97, SBA^{+}05, SZGbC04, SHLB08, Sun92, TKT^{+}08, TJLT00, WH05, WTK07, WZC08, Wit94, Yat88, ZMS^{+}06, ZEO01, vdR87e, vdV89b, CF09, SR03]. concatenation [HRJ^{+}06]. concept [GGM^{+}09, LL03, TG04, WAE06]. Concepts [PSS01, Ano86i, DSSU97, TBK06, ANN^{+}92]. Conceptual [FJY06]. concerns [FdSC07]. Concurrency [TG07, JK92]. Concurrent [BS91a, BP01, BS92, CLP95, EL98, LM90a, Sun92, dB90]. conditions [DS04b, KDE04]. Condor [PL96]. Condors [ELvD^{+}96]. Conference [Ano84e, CF09, Ano86j, BGL08, Kaa98, Rho89]. conferences [Ano94e, Ano96a, Ano96b]. configurable [Dör05, JNPY06]. Configuration [PA01a, BORM07, DvdHdL06, MD92a]. confirming [Niw89]. conjugate [Cro95]. connected [UWV92]. Connection [BS91a, BB84, BS92, DFSZ88, BB85, CH95, Por95]. connections [LPE08, Shi92]. connectivity [BRR^{+}04, JM02, UM02]. connectivity-preserving [UM02]. Connector [EGK^{+}07]. Conservation [IS06]. conserved [DCS^{+}07]. considerations [CGL08]. Consistency [GSD95, LLKF09, OB04]. consistent [PY00]. consisting [Shi92]. constrained [DD05, HY09, PFMC04, Ven09]. constraint [Hal88]. constraint-based [Hal88]. constraints [FX07, HZC^{+}08, LYT^{+}05, NP03, SL87]. construction [BJWZ08, CPB00, TM05, YJA03]. contact [Ned06]. container [LGMV02]. contaminant [RS99]. contamination [ST98]. Content [EGAQ09, GC00, De 88, DS08, FMS08a, FM08, FR08, LPE08, MM08, NKX09, OCdAM07, QP08, SMA08, SYT09, Zin00]. Content-based [EGAQ09, OCdAM07, QP08, SMA08]. Contention [BS09]. Contention-based [BS09]. Contents [Ano01b, Ano05a, Ano05b]. context [DMG^{+}08, FD95, LG08, PEG05, SGY^{+}07, SM01b]. contextualized [SA07]. continental [RV95]. Continuation [DP03b, PPAK99]. Continuous [DS04c]. continuum [Low01]. contracts [GPK05, LKA^{+}08]. Contributions [Ser95]. Control [ABF93, BT93, BL92, ATT96, AR98, BKM03, BX04, CSC^{+}92, Dat03, DL03, FMR05, FX07, GWO03, GY90, HV92, HXL90, KLM^{+}03, LHC03, Lop03, LC03, Mam09, RSR01, SMA08, SCP09, TG07, WC06b, WC06a, WZC08, YA02]. controllable [GZ04]. Controlled [BP02, GFR^{+}06, DSS98, KKP00]. controller [Röb05]. Controlling [HAAH05]. controls [EV96, EL98]. convection [JNR01, Tab06]. convection-diffusion [JNR01]. Convenient [BKŁ01]. convergence [CDF^{+}05, Gut00]. conversion [AC01]. conversions [Rus90b]. Convert2Java [AC01]. Convex [SSMG95]. cooperating [CWD04, Reh06]. Cooperation [Nar86, JFDF09, Smi86, CKK^{+}04]. Cooperative [FMR05, GP09, HMC06, Nag86a, PP07, Wah84, CBK^{+}01, RHB08, ZL04a]. Coordinated [FX07, BCH^{+}08]. Coordination [THN^{+}06, PA01a, Pap05, Pro07, WKT00]. copies [Dup90]. Coprocessor [SK97]. copy [DS00]. copy-hybrid [DS00]. CORBA [Lan00, LRW01, LDS06, LLSR02, RdSH^{+}00]. CORBA-based [LLSR02]. Core [HT02, CKFJ06]. correct [GPA96, Pro07]. correction [KA88]. correctness [MD92b]. Correspondence [DMN^{+}05]. cosmic [NMZC06]. cosmology [LB03]. Cost [GR07, BG05, GE90, RS98, SD02, SHJR04, SDD^{+}09, SK05, de 94]. Cost-effective [GR07]. cost-optimal [BG05]. coupled [AGP^{+}92, BDNP92, Kun94, MD92b, MR04a, Mis92, SMK05, ZCT^{+}04]. coupling [HNS05]. course [BLB03]. CPU [ZSI08]. cracked [OS06]. crash [HDC^{+}94]. Cray [MPG96, SCK^{+}00]. Creating [NWE04, RSSD02]. creation [Mar98b, Mar99a, TBD^{+}02]. credit [JFDF09]. credit-based [JFDF09]. criteria [ADKS06, WHP09]. criterion [SKJ01]. critical [MC04, SSC09]. crossed [GL04a]. CROWN [HSH^{+}07]. crust [OS06]. cryptographic [BDNN02]. cryptosystem [LL04b]. Crystal [KS02]. Crystallization [VSvD94]. CS [GD93b]. CSP [MRV92]. CT [SGL99]. CTDNet [KGW95]. cube [Pan95b, YJA03]. Cubemat [Shi92]. cubes [SAKOK03]. cultivate [AHdJF97]. cultural [Nit86]. CUMULVS [WKF03]. curation [HBH09]. Current [BHD09, Fur92, MK88, Nag86b, Miz89a, dLRW03]. curricula [MR03b]. curve [LL04b]. curves [SW05]. customers [ABCD00]. customized [RSSD02]. cuts [HSS00]. CWI [Baa87]. Cyber [YYW^{+}09]. Cyber-Transformer [YYW^{+}09]. cycle [DZJ^{+}00, Goo01]. cycles [PL96, SK05]. cyclone [VBLS09]. cylindrical [LGMV02].

D [SGL99, YMM00, Zem86, vdR87h, vdR87e, vdR87k, Avg00, CH95, DMM^{+}99, EMB98, EdBG^{+}99, GR09, GGM^{+}09, HvHAS04, IdLR01, JHL^{+}06, KCK04, Kni89, KA88, LBB^{+}09, MJ98, MJ06, NCS04, RICW00, WKF03, XYZ05, RBS93]. D-Grid [GR09, GGM^{+}09, LBB^{+}09]. DAEs [CFVP03]. DAI [SPK^{+}07]. DAROC [SHJ06]. DAS-2 [HBJ^{+}03]. Data [Ama88, Ama89, AW97, FS97, Fuk85, GGH^{+}06, KB09a, PH07, SD07, VV92, WMN^{+}01, YTHY84, ABF^{+}03, ADK^{+}09, BL92, BPS^{+}03, BPAP92, BCW01, CTT02, CGM^{+}07, CKFJ06, CHJ^{+}04, CCL07, CGL08, CY01, CCG07, CGST09, CTT07, CS97, DMR93, DMMP98, DMM^{+}99, DAM08, DFG^{+}00, DF97, DSH^{+}99, Dup90, DB99, EWG99, FSP02, FB97, FAJP99, FJT01, FM01, Fin99, Fra08, FdSC07, GGW^{+}09, GLSV07, GP09, GGH^{+}03, GB99, HC99, HAE^{+}03, HBH09, HESM99, HPS97, HJPS03, HAB^{+}06, Joh02, JRF^{+}07, KN06, KMCH03, KY04, LS07b, LGH97, LC04, Leo98, LGW07, Lop93, LWHS07, LLSH07, MTNM08, MJ98, MJ06, MFP05, MBFC99, MTH^{+}05, NMZC06, NSP07, NJW^{+}06, Nis93, OB04, OFT09, PM04, Pal09, PKC^{+}05, PSR^{+}07, PSA^{+}09, PCM99, PPH^{+}09, PCB99, RL98, Rus90b, SZC05, SBG^{+}09, Sin07]. data [SSK^{+}08, TWC^{+}06, TLTY06, TD95, TJLT00, TV08, VBLS09, WSS^{+}09, WFC07, XYZ05, YJA03, YLC^{+}06, ZLD^{+}03, ZQZZ09, vdR87l, vdR87k, BCMA07, CC07, Doḡ09, HBH09, LS07b, LVH08, NZQ07, Qin07, SYT09, SHJ06, TLYT05, TLTY06, YYW^{+}09]. data- [vdR87l]. Data-Activated [SHJ06]. Data-aware [KB09a]. Data-Driven [VV92, YTHY84, MTH^{+}05, PKC^{+}05]. data-intensive [BPS^{+}03, MFP05, Pal09]. data-parallel [BL92, FSP02, LC04]. database [ACU95, Ano84j, CYB90, CW93, FAJP99, FT07, GML99, Joh92, LY90a, LY90b, LKM91, Mur88, NSI84, PNH99, PH99, vdR86a]. databases [CYB90, CFG93, FGRZ09, KP00, KST92, Kos00, PNH99, SL87, TG07, ZMN99]. dataflow [GY90, Gur85, GBT87, HG92, KSY92, ZT90, ZT91]. dataflow/von [HG92]. DataGRID [VBP03]. DataMiningGrid [SSK^{+}08]. datasets [KPB^{+}03]. DataTAG [MFP05, MMFM^{+}05]. date [Din99]. Davidson [BvdHN^{+}01, BV04]. Davidson-type [BV04]. DDL [FB97]. debugger [CLP95]. Debugging [AW03, FSP02, RCD03, BW95, Kac00]. Decentralised [Low05, Vau93]. Decentralized [LS07b, HB08, HXL90, LHL09, YBQ07]. Decision [vdR87f, AW97, DSH^{+}99, GS05, LGW07, PP07, SB97]. Declarative [TA96]. decoder [KA88]. decomposed [SKJ01]. Decomposition [dRSS97, BH03, HND06, MvdV01, Pri95]. dedicated [DZ04, NP03]. deductive [SL87]. default [ATT96]. Defining [GMS09, MR03b]. definite [Amo06]. Definition [CGT07, HML^{+}06, LRJ^{+}06, LS05, Szu01]. deformation [Tor04]. degeneracy [DMN^{+}05]. Delaunay [LGMV02, XSM04]. delay [KV09]. Delft [DFSZ88]. Delivering [Zin00, BYV^{+}09]. Delivery [GC00, HAE^{+}03, LPE08, ZCT^{+}04]. Delphi [ACGdT02]. demand [BPS^{+}03, FMR05, KLM^{+}05, SSF^{+}09, WWC^{+}97]. demanding [MVT^{+}99]. demands [SCB04]. demonstration [CALN03, HRJ^{+}06]. dense [DS04c, HPP94]. density [DRS04]. dependable [AR07, WTK07]. dependent [CP06, Dua94, Nos98, Tab06]. Deploying [PCBD99]. deployment [BJA^{+}05, CdCD07, HSH^{+}07, PSP^{+}09, PPSS06]. deployment-based [CdCD07]. Derek [Ano87b]. Derivation [DRNMC09]. derivative [GKS05]. derivatives [BBL^{+}05, SBLT05]. derived [Del06]. deriving [CFVP03]. describe [vdHDT^{+}06]. description [HK88]. descriptions [BHK90]. Descriptive [SGdMM96]. descriptor [Var03]. Design [CCDS08, LCP04, LJY04, LL04b, LC03, MSK03, NSS99, NP06, PSR^{+}07, Qin07, ZY90, ÅO06, AMB03, BBC^{+}99, BKG05, BB06, Cur92, DMR93, DDV92, DGS09, FD02, FAJP99, Ger02, HB09, KKL09a, LMH^{+}09, MB01, MBZL09, MVT^{+}99, Ném00, PPJ95, Röb05, SCK^{+}00, TC92, VVC^{+}03, VSM02, VR00, VP94, WKT00]. design-space [SCK^{+}00]. designed [ZJWZ04]. Designing [BGL^{+}05, GBE00, ST98, TKT^{+}08]. desktop [BCB^{+}07, DSS07, KFC^{+}07, WFC07, KNK^{+}08, KLM^{+}05, LTOT07, SBHD08]. detailed [LR06]. detect [ZZN04]. detecting [SCB04, SK05]. detection [HHS98, KTV03, KKP00]. determination [Dör05]. Deterministic [Gue01, MKH06]. deterministically [LOK09]. deterministically-routed [LOK09]. Developing [AAB^{+}92, DW87, GW01, GJS^{+}94, LGS^{+}07, Reu03b]. Development [BFR05, BHH91, DDO^{+}92, GML99, IMKB89, YA02, BdCYG05, BHH92, Cas94, Dal03, Dek86, DZJ^{+}00, FJ00, FR08, HRSW99, Hen87, IMSV90, KDFL99, KV03, Li90, LWHS07, Mat89, RG04, WWSM98]. developments [Sch94]. device [DLR^{+}09, FHG95b, Kaw92, SGFS01, SLDK03, SW06]. devices [KK97, OB04]. DEVS [SZP00]. DFT [BST^{+}04]. diagnosis [BvdBM^{+}93, CD99, DFT92, FM01, Han89, KE85, MH01, Suz89, tTvH96]. diagnostic [ASTEP98]. diagonal [DL03]. diagram [KS02, SCK^{+}00]. DIALOG [ZT90, ZT91].

DICOM [EGAQ09]. difference

[ÅO06, CS93]. differences [Nit86, PBHK01]. different [SD06, vM94]. Differential [DL03, BC03, Del06, SS03]. differentiated [PEG05, SPBT07]. differentiation [Büc05a, Büc05b, Fau05, HNP05, HHG05, HNS05, Pal06, Röb05, SBLT05]. Diffusion [BKS98, vdS04, BMPS01, Ban02b, JNR01, KZC04, Mar02]. Digital [TS08, FBBW99, HESM99, KY85, NKX09, PCM99, SST^{+}06, WSTW87]. dimension [SAKOK03]. dimension-ordered [SAKOK03]. dimensional [BP02, BS04, CHJ^{+}04, DS04b, DSH^{+}99, DB99, EFD00, FHG95b, Pal01, SW02, Wes99, XSM04]. dimensionality [BFLL99]. dimensioning [TDV^{+}08]. DINCast [GNWT05]. DIPAS [TBNF09]. Direct [NEJP94, VWCV94]. Directed [WTC^{+}02, RSR01]. direction [Dör05, Fur92]. directional [LOK09, Mar02, OS06]. directions [Ano84j, LKM91]. directly [De 88, ZCT^{+}04]. directory [Ohy89]. Dirichlet [KDE04]. DIS [RdSH^{+}00]. DISC [IMSV90]. disciplinary [DFG^{+}00]. Discord [SCB04]. Discover [LJPS05]. discovering [SSC09, VGBLGS^{+}06]. discovery [CCL08, DMZ09, FMS08b, HB08, MTV05, MP02, MV09, TTP^{+}07]. Discrete [She00, NB04]. Discrete-event [She00]. DISCWorld [HJS^{+}99]. disease [KKP^{+}05]. disease-EEG [KKP^{+}05]. diseases [KE85]. disk [CFGC03, NP06]. disks [CkLC06]. display [PSG^{+}06, RJH^{+}09]. dissemination [FMS08a]. Distance [FGRZ09]. distances [MGH^{+}05]. distinguished [HWW04]. distribute [FT07]. Distributed [BDF^{+}99, BCT^{+}07, BM08, BBSV92, BM92, BKK02, BDHK06, BPS^{+}03, CTT02, CkLC06, CTT07, DVV90, FC09, FJT01, HH98, HV92, HB00, HAB^{+}06, KK00, KGX95, LWHS07, LLSH07, Mul92, SKJ01, SLW01, VVC^{+}03, WM07, AT01, AMB^{+}92, AGJN00, AB03, BMRW01, BFLL99, BKG05, BR92, BW95, BB06, BAD^{+}05, BDL06, BFW^{+}03, BGR^{+}99, CM01, CST92, CWD04, CDF^{+}05, CBK^{+}01, Cho04, CS93, CGST09, CRM05, DCL00, DBA98, DAM08, DT93, DRNMC09, DDM^{+}08, DL00, Din03, FGRZ09, FB97, FAJP99, FJ00, FBBW99, GMEL08, GA06, GW01, HC99, HAC92, Her91, HPP94, JSK^{+}06, Joh92, Kim07b, KNK^{+}08, hKcF09, KKL09b, Kob92, KB00, KKP00, LB03, Lau01, LR06, LL03, LKG08, LWSC07, Luk89, Luk00, LMH^{+}09, MCSS00, MMR02, NSI02, Ole07, Öst92, OS01, PY00, Par04, PH99, PX07, PSG^{+}06]. distributed [PCB99, PVBH05, QP08, RHB08, Reh06, RM97, SZC05, Sap88, SZP00, SBSdL06, SY04, SD02, Sin92, Slo96, SD03, SHJ06, SLO^{+}05b, SHLB08, SM96, TC06, TJLT00, TBNF09, Vau93, VR00, VLC03, VBLS09, WSH99, WB90, YA02, YHJC05, ZCT^{+}04, ZLD^{+}03, ZMS^{+}06, ZWJ04, ZCW^{+}04, dKdOS03, CT09, WAE06]. distributed-data [FB97]. Distributed-Memory [BM92]. distribution [BTM06, CYH04, DLW07, DS08, FR08, LGH97, LOK09, NKX09, NMZC06, WFC07]. distributions [BDHK06]. diverse [MC04]. diversity [BORM07, HJK^{+}04]. DLM [Pud87]. DMPP [Ste92]. DNA [CGH04, CS05, Kar01, WMN^{+}01]. doctoral [SR03]. document [FC09, LYT^{+}05, ZL04a]. documents [KFF89]. Dolev [BDNN02]. domain [BFK02, GXD^{+}09, HND06, MvdV01, Pri95, YLC^{+}06, vOHD^{+}05]. domain-specific [BFK02]. domains [BCPS03]. Donald [Ano87l]. Doores [vdR87j]. Dordrecht [vdR87f, vdR87k]. Dordrecht/Boston [vdR87k]. double [KA88]. download [CGL08]. DQP [LMH^{+}09]. drawing [Büc05a]. drawings [Niw89]. DRAxML [SLO^{+}05b]. Dreyfus [Zad87]. Driven [VV92, YTHY84, FC05, Fre94, MTH^{+}05, ONHT89, PKC^{+}05, SBG^{+}09, SLŽ^{+}09, VWCV94]. driving [LKM91]. droplets [DY04]. DSP [EFD00]. Dual [ZDR07]. Dual-Level [ZDR07]. duality [UM02]. dumps [Dal06]. Durham [Her87]. during [Niw89]. Dutch [BVP^{+}87]. DVC [CT09]. DYMOS [Suz89]. Dynamic [Ber00, DMG^{+}08, DDR^{+}07, GMM^{+}09, KMK09, LH07, OSHH96, Reh06, RN01, Sch98, SLJ^{+}06, TLYT05, vOHD^{+}05, AAC04, AFP07, AKPN01, AN08, DHB02, Doḡ09, FVFA98, GMEL08, GNWT05, Hua05, HPLL09, JM01, LCP04, LRMC94, MGYC06, MR04a, PA01a, SGdMM96, Ven08, WG00, ZDR07, Suz89]. dynamic-key [LCP04]. Dynamical [LMBCC89, GLW99, Lop03, MPQ03]. dynamically [AAB^{+}92]. Dynamics [vOB95, Ban05, CM99, FAJP99, JL03, MF93, MR00, PHM^{+}99, SCK^{+}00, Tul04, Wes99].

e-commerce [KF00]. e-Lab [BGJ^{+}06]. e-mail [LL04b]. e-Science [CF09, DGST09, HT02, ZBB09, BHD09, JHL^{+}06, SBG^{+}09]. e-Toile [BBG^{+}05]. Eager [CK00, KGW95, MSS02]. Early [ACE02]. earth [GNOY01, SSMG95, Bre89, FPX^{+}09, TWC^{+}06, WKZ^{+}03]. Easing [LP01]. easy [ABF^{+}03, MZC08]. ECA [KA08]. ECAI [Ano84e]. ECAI-84 [Ano84e]. ECG [CPK05]. economic [DS08]. economy [ABG02]. ecosystems [LMBCC89]. ECRC [Gal87]. eddy [NEJP94]. edge [HHS98]. edited [Zem86]. editor [Ano87k, KDFL99]. Editorial [AKvdR86, Ano84f, Ano86g, Ano99, BLRL03, Bru01, DHS00a, FM01, Her94, HB98, Kaa98, KG01, Mal01, ON85, OS01, RR03, SG95, VN01, Wil00, ZEO01, AB01, AD00, Ano01c, BJC02, DDS00, Flu03, HP92, KK00, LBR02, LC01, Tan02a, vdR93a, Ano84g, Ano86f, Ano87i, Ano88c, Ano89e, Ano90f, Ano91d, Ano92g, Ano93h, Ano94f, Ano95g, Ano96d, Ano97c, Ano05c, Ano05d]. Editors [Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano02b, Ano03h].

Editorship [Kaa99]. EDPEPPS [DZJ^{+}00]. Eds [vdR87a, vdR87c, vdR87e]. education [AMB03, Avg00, KZBK99, Wie03]. Educational [AD00, MJ00, Jon00, RMM^{+}98]. EEG [De 98, JSS^{+}99, KKP^{+}05]. Effect [Tor04, Nit86]. effective [BAD^{+}05, CC00, GR07, HCL07, HPLL09]. effectiveness [PM04]. effects [HSS00, Yam92]. Efficiency [ZMN99, Dua94, PR95, ZMS^{+}06]. Efficient [BK97, Ber98, BBL^{+}05, DZ98, JC08, LR01, LKG07, LSTV07, PHL98, SL87, TV08, Aba09, AMH02, BPS06, BFR05, DAM08, FJ00, GKS05, GNWT05, HHG05, HMP04, LC01, MGLV04, NF07, NP06, PY00, PWY03, PVBH05, RG04, Rum99, SLS^{+}09, SHLB08, Tis07, TM05, WFC07, ZQZZ09, dlTK92, LOJ^{+}07]. Efficiently [ABL04, BRR^{+}04, SSC09, KTY03]. efforts [Ano84i]. eigen [CHJ^{+}04]. eigendecompositions [DP03b]. eigenvalue [BV04, Del06, Prz03, SKT^{+}08]. eigenvalues [Amo06, CDS03]. EJOB [vdR87e]. Elastic [MLSO01, NF07, HvHAS04, OS06]. election [SW02]. electric [DMMP98]. electrical [DFT92]. electrocardiography [PCB99]. electromagnetic [Dae95]. electromagnetics [LW08]. electromechanical [HAAH05]. Electron [FGCM07, BRS04]. Electronic [Öst92, ABCD00, AS99, Bur02, Din03, FP03, MCSS00, TC92, Bur02]. Electrosmog [Dae95]. element [AS99, LDS06, SW05, Tab06]. elements [SP93, WAE06]. elimination [Tis07]. elliptic [LL04b]. Ellis [Ano86i, Ano87b, Ano87l, vdR87d]. EM-3 [YTHY84]. EM-4 [KSY92]. embedded [NWE04, Pud87]. Embedding [Pri95]. emergency [Gra01]. emerging [BYV^{+}09, How91]. empirical [Cur92, DS99]. enabled [ABB^{+}03, EPJ^{+}05, FS07, GLM^{+}08, JHL^{+}06, LKG08, LWSC07, MGYC06, MJ06, NJW^{+}06, PKC^{+}05]. enables [AB01]. Enabling [Mam09, RJH^{+}09, TMV^{+}07, ET08, GMEL08, KT08, MWC^{+}03, MLD08, SSK^{+}08]. encrypted [DF97]. encryption [LCP04]. end [JLU03, KT08]. endpoint [SCP09]. Energy [LBB^{+}09, ABB^{+}03]. enforcing [LHC03]. Engine [BG87]. Engineering [Kow85, AMB03, Ano87b, Ben99, Bun03, CTMO06, DFG^{+}00, Hir89, Joh02, Kim07b, Kow84, Mai91, Mat89, TY85, Van87b, VVC^{+}03, SR03]. Englewood [vdR87h]. enhance [GMB^{+}05]. Enhanced [WDD00, AV00, PH99]. enhancements [PSVL02]. Enhancing [ACML05, BBLP05, FM08]. ENS [BBD^{+}99]. enterprise [AGJN00, CM01, DDR^{+}07, GLSV07, JAA07, Kim07b, KKL09a, KFC^{+}07, NSP07]. enterprise-scale [NSP07]. EnterTheGrid [Ano05e]. entities [JLU03]. Entropy [EHMS00, Fre94]. Entropy-driven [Fre94]. enumeration [MKK03]. Environment [BKS98, BP94, CWD^{+}08, LSS94, ASTEP98, ADT03, ACC^{+}05b, BMRW01, BBBD01, BKG05, BPP^{+}07, BCB^{+}07, CLP95, CSC^{+}05, CTT^{+}08a, CCDS08, DCL00, DZJ^{+}00, DRS^{+}97, DMZ09, DT93, DL00, EMB98, EPJ^{+}05, FJ00, GR96, GGW^{+}09, HJS^{+}99, IMSV90, KDFL99, KPB^{+}03, KKL09b, Kos95, KKP00, LLKF09, LM90a, LGS^{+}07, LSF^{+}94, MCSS00, MSKT07, MVT^{+}99, MTH^{+}05, MSX00, OS92, PP07, PSA^{+}09, PPAK99, PHM^{+}99, PA01b, PBB^{+}05, RL98, RMM^{+}98, SOR05, SSKF95, SMK05, SPCL04, WDD00, WB90, YMM00, ZS05b, DGS09, FMD99]. Environments [YPF05, ACGdT02, AMD08, BJA^{+}05, BR92, BFR99, CTF^{+}99, Cas94, CRM05, CS09, FX07, HZC^{+}08, Höf03, KA09, KBM^{+}02, Kim07b, LB03, LYMZ09, LSTV07, LLSR02, LLSH07, MKT09, MR03b, MvWvL99, NMC05, Pad92, Pag99, Sch98, SLDK03, SD07, Sun92, Ven08, WH05, WSS^{+}09, XHY^{+}90, ZSI08, ZAP05]. epidemic [FMS08b]. epidemics [LMBCC89]. equational [OP95]. equations [BFLL99, BMZ01, BC03, IS03, SG04, Tab06, de 94, vdR87h]. equipment [TC92]. equivalent [PEG05]. era [Zad87]. Erratum [KSM^{+}07a, MR04b, NHG03]. Error [KTV03, DR03, KA88, SS03]. eruption [CDRS05]. ES/SDEM [Mat89]. ESPRIT [Ano87j, Cad86, Ano84h]. establishment [Mar99b]. estimate [Dal06]. estimates [HV03]. estimation [EW97, PBT02]. Ethernet [HJCD05, MGH^{+}05, WTC^{+}02]. Euclidean [DL03]. Euler [GKS05]. EURO [Kaa98]. EURO-VR [Kaa98]. Euromed [EV98]. Europe [HB98, Lid99, Wil00, Ano84h, ES94, Pol98, Wal94]. European [Ano84e, Ano86h, Zna94, Dek86, LPC^{+}95, Mur95, Nar86, SS90]. evaluate [Szu98]. Evaluating [BSCC06, HBJ^{+}03, BGL^{+}05]. Evaluation [Bal93, BY93, BP94, CG09, CW93, GSD95, HRJ^{+}04, LSS94, ADKS06, BBBD01, BS09, DSS98, Din99, FS93, HD05, KSY92, Kun94, LCP04, LNJ04, LB09, MSK03, MI01, MM03, MOK06, NP06, Nis93, OP97, Par04, Shi04, SSMG95, WGL92, YJA03, ZY90, ZGCM00]. event [ABB^{+}03, GKIZ05, HMP04, KN06, ONHT89, She00]. event-based [KN06]. event-driven [ONHT89]. events [Ano96a, RT06]. eventsensional [Ano96b]. Evolution [JCSS01, Dub91, DMN^{+}05, EL98, JL95, Moo99, RS99]. Evolution-based [JCSS01]. evolutionary [ACML05, EL98, Hen87, JC09, LKG08, Nos98, OVDV98]. Evolving [EL98, LLWN04, SW99]. EWSL [Ano86h]. exact [HHG05, Röb05]. Examination [ZMS^{+}06, ABdL^{+}03]. example [DFSZ88, KA88]. examples [GJS^{+}94, tTvH96]. exchange [AC92, GPK05]. excited [REM04]. execute [CTF^{+}99]. executed [HLvL^{+}97]. Executing [WS05]. Execution [ABF93, CM01, AKPN01, AR98, BBBD01, BKSS02, BFC02, CY90, Dup90, FX07, GR96, GMEL08, GL04b, KP00, MTNM08, OP95, Pal09, SM01b, ZT90, ZT91]. existing [GJS^{+}94, Hab05, KY85]. expectations [Nis93]. Experience [BHH^{+}93, CR92, Oku92, Luk89, RvdSB^{+}03, RRS99, YJA03]. Experiences [KB09b, MPB^{+}07, ABB^{+}03, ACE02, GMS09, Jon00]. Experiment [DGS09]. Experimental [GGH^{+}03]. Experiments [Hey90, PEG05, Vre88, Vre89, BPP^{+}07, GL04a, MCSS00, MGH^{+}05, MVT^{+}99, PKSC02, YA02, CN92]. Expert [Coo86, Hir89, Van87a, vdR86a, Ano87m, BT93, DFT92, DW87, EO86, Gil85a, Han89, Hen87, Kag89, Kom89a, Mar86, Mat89, Miz89a, OS92, Par87, Pud87, Ste85, Tak89a, Tak89b, VM93, WYN^{+}90, Yam89, YWA^{+}89, Zem86, vdR87f, vdR87d]. expertise [Zad87]. explicit [HV92]. exploitation [KLM^{+}05]. Exploiting [Joh89, NSI02, Zha93, ADK^{+}09, RG04, ZS90]. Exploration [BKS98]. explorations [SSC09]. exploratory [Tak05]. Exploring [ABdLL05, LB03, AMB03]. Exponential [EL03]. exponents [DE03]. exposing [BBW08]. Express [WKF03]. Expression [Ref87]. Expressive [Par90, Avg00]. Extended [ZYXL05, AC92, DDV92, dB90]. Extending [CS96, SY04, Ste92, TG04]. extensible [BMRW01, GB99]. extension [GPA96, SPK^{+}07]. extraction [GKIZ05, WM07].

F [Teb86, Zem86, vdR87g, vdR87i]. fabrication [SKT02]. facial [SGL99]. facilitate [SHJ06]. facility [RG04, LRJ^{+}06, SDBdL06]. factor [ADKS06]. factorization [MvdV01]. FAIL [HTV07]. failure [HWS07, vdR93a]. failures [DLW07]. fairness [KV09]. FAM [KKYK04]. FAN [CG09]. far [BBJ^{+}06]. far/near [BBJ^{+}06]. farm [BFLL99, Bro92]. farms [NP03]. Fast [HYS04, Pan95b, SO98, BM00, KKHS01, LN94, VBLS09]. Faster [BRMN04]. Fault [AFP07, LAM07, PCBD99, Xia06, AMH02, ASTEP98, BCH^{+}08, CdCD07, FD02, Han89, HTV07, KA08, KL02, LHB95, LSTV07, LS01, LS08, PWY03, SG05, Suz89, THKG98]. Fault-Tolerance [PCBD99, CdCD07]. Fault-tolerant [LAM07, Xia06, AMH02, LHB95, LS01, LS08, PWY03, THKG98]. FCI [HTV07]. FCN [FC09]. Feasibility [AKW90a, AKW90b]. feature [BM00, PdLS^{+}99, SK06, WM07]. feature-based [PdLS^{+}99]. features [DGST09, HO02, KGW95]. featuring [LLH^{+}03]. Federate [CYLT05]. Federated [SVC^{+}07, LHL09, TS08]. federation [FLPP05, RHB08]. federative [HB00, Joh92]. feedback [WWSM98]. Feedforward [JL98, RM97]. FEM [BBJ^{+}06, GNOY01, LF95b]. FEM/FVM [LF95b]. few [Amo06, DE03]. FGCS [Mes02, Nis93, Ser95]. Fiber [HICÁFM^{+}06]. field [BBJ^{+}06, PdLS^{+}99, SZ98]. fields [MJ98]. Fifth [Her84, HV84, TR85, Ais88, Fur92, Lin84, vdR86b, Ano86i, FS93, KI89, Sti93, vdR93a, vdR93b]. File [KLSS05, ACC^{+}05b, Doḡ09, GCCC^{+}07, GD05, MM08, PCG^{+}06, SPK^{+}07, WX02]. File-based [KLSS05]. file-sharing [MM08]. files [SCY01]. filesystem [MCQ^{+}07]. filling [SW05]. filter [TM05]. filtering [PCB99]. finance [BS04, Par94]. financial [Kea93, TMTY05]. finding [DvdHGdL09]. Findings [WBMP99]. fine [LC03]. fine-grained [LC03]. Fingerprint [MPH00]. Finite [Tab06, ÅO06, BVDF00, CS93, LDS06, SW05, Tho06, WAE06]. fire [ACML05]. First [Gil94, SLZ95, BT93, HRJ^{+}06, Par87, WRBG94, SO98]. First-principles [Gil94, WRBG94]. five [Bal92, Van87b]. Fixed [Fio06, CFVP03]. Fixed-point [Fio06]. fleet [OVDV98]. Flexible [dKdOS03, BBW08, BLRS98, PCG^{+}06, PKSC02, tTvH96]. Floating [FGG03]. flocculation [vOB95]. flock [ELvD^{+}96]. floor [DDV92]. Flow [Ama88, Ama89, BKS98, BDNN02, SCP09, WMN^{+}01, AKH^{+}04, CH95, ED04, GKS05, GY90, HORC04, HvHAS04, RV95, SdSP04, VWCV94, VFS01, vM94, CG09]. Flow-Aware [CG09]. flows [AJZ^{+}02, ABdL^{+}03, CDRS05, DS04b, GL95, IOO04, LF95b, Prz03, RŻDM01]. fluid [CM99, Wes99]. fluids [FPdS04, NEJP94]. flying [VLC03]. focal [Ram95]. folding [DCK03]. folding/unfolding [DCK03]. foliations [MPQ03]. forces [LKM91, PO00]. Forecasting [PSP^{+}09, LFWV05, WSH99]. forest [CY90]. Foreword [LL04a, Mes02]. FORK [HSS92]. form [Tho06]. Formal [BHH91, BHH92, AHdJF97, Szu01]. Formation [MCA02, DDL01]. forming [HW95]. forms [EL03]. formulations [SS03]. forthcoming [Ano94e]. FORTRAN [RBS93, Ben99, CMZ95, Din99, dSL98]. forums [EV99]. Foundation [FPX^{+}09, NSF87]. foundations [FLPP05]. Fourier [DDL01]. fractal [NCS04]. fragmented [CC07]. frame [KTM^{+}08]. Framework [SdR99, BvdV99, BBW08, BFL99, BORM07, Boa04, BAD^{+}05, CWD04, CSW06, CCLS09, DCS^{+}07, DPP03, EV96, FvLTT98, FB93, GR96, GPK05, HAF00, KFF89, KV09, LS07a, LS07b, Lau01, LJ07, LM07, LS08, NZQ07, OK02, PADD03, PF01, PB05, PVBH05, PSS01, QP08, SSB05, Ven09, VWD^{+}08, YD05, dNE05, AW03]. frameworks [Pap05]. France [Zna94]. Free [Ano86i, Ano87b, Ano87c, Ano87l, Zad87, BS04, BB06, CMO03, LGMV02, Ray05]. free-boundary [BS04]. frequency [CP06, MC00]. friction [Ned06]. friendly [KBB^{+}09]. FTA [VR00]. Fujitsu [Uch86]. full [AC92]. Fullerenes [She04]. fully [IST04]. Function [VV92, NK05, Omo91]. Functional [ABF93, BS91a, IEG04, BS92, DMM^{+}99, DRS04, DD07, HK88, HRJ^{+}06, MSS02]. functionality [MCQ^{+}07]. functioning [BRS04]. functions [Dua94, KDE04, vdR87h]. fundamental [Hul89]. fundamentals [Kag89]. furnace [YWA^{+}89]. Fusion [KFP^{+}02]. Future [Car03, CRM05, KSM^{+}07a, LKM91, MR04b, NHG03, Ano87b, Bis94, BHD09, DD86, Ful91, Fur92, GMS09, Hul89, Kim07a, Nag86b, SB99, Slo06a, Slo06b]. fuzzy [PB05, ZL04b, vdR87f]. FV [GL95]. FVM [LF95b]. FY [NSF87].

G [vdR87g, vdR87h, ABG02, BGH^{+}03, GRPL04, HBJ^{+}03, THN^{+}06, WBF08]. G-lambda [THN^{+}06]. G-PM [WBF08]. GA [CEGL01, LLWN04]. GA-based [CEGL01]. games [PN09]. gas [CH95, DC00]. gases [Bog99]. gate [DCK03]. Gauss [HPP94]. Gb [ABB^{+}03]. GCel [vOB95, Cro95, RV95]. GCel-3 [Cro95]. GCel-3/512 [Cro95]. GCR [VFS01]. GECEM [LW08]. gel [CGH04]. gel-based [CGH04]. GEMMA [BPP^{+}07]. GEMS [WBT^{+}08]. Gen [MR04b]. Gene [STP^{+}05, DMN^{+}05, TBD^{+}02]. Gener [KSM^{+}07a]. General [HHG05, SL97, Zna94, AKW90b, LDSH95, WHP09]. general-purpose [AKW90b]. Generalized [KY04, IS03, MvdV01, SKT^{+}08, SA97]. Generalizing [PPH^{+}09]. generated [HHG05, PO00]. Generating [GKS05, TSZP99, DK00]. Generation [BDF^{+}99, Her84, HV84, Kow85, KI89, NHG03, Slo06a, Slo06b, Sti93, vdR93b, Ais88, Ano86i, BP02, DDS^{+}09, DR05, Fur92, GODM98, GZ04, HJK^{+}04, Kim07a, Kow84, Lin84, MWC^{+}03, PO00, Par87, PFMC04, Ste85, TDL05, TR85, WYN^{+}90, ZGCM00, vdR86b, vdR93a, FS93]. generator [Tan02b]. GeneRecon [MPB^{+}07]. generic [BCW01, DVVD02, GdLvOT03, Tis07]. GENESIS [RG04]. Genetic [LOJ^{+}07, AT01, CPK05, CFG93, DMN^{+}05, GRH05, GODM98, KKP^{+}05, LYQ06, LLWN04, LS01, TMT^{+}07, SMI01]. genetic-based [LS01]. GENIUS [ABF^{+}03]. genome [SLS^{+}09]. genomic [Sin07]. genotype [MS03]. Geocomputation [XSJ04]. geodesic [DE03]. geodynamics [BN06]. geoenvironmental [CTMO06]. GeoFEM [GNOY01]. geographically [AGJN00]. geological [XSM04]. geomechanical [WAE06]. geomechanics [BN06, HND06]. Geometric [DP03a, PdLS^{+}99]. Geometries [BKS98]. geometry [DFG^{+}00]. Georgia [vdR86b]. GF11 [KB92]. GHolo [BdCYG05]. Gigabit [MGH^{+}05, GGH^{+}05, HJCD05, WTC^{+}02]. Giganet [HO02]. Gill [Ano87c]. GIS [LNJ04, LD04, TG04]. Givens [DV03]. Global [LLH^{+}03, PdLS^{+}99, SWCP03, WKZ^{+}03, DR03, Dzw97, ESFD06, GCCC^{+}07, NSS99, STH^{+}98, Shi92, SBdL09, UWV92, LRJ^{+}06, SDBdL06]. global-bus [UWV92]. globally [GS05]. Globus [BMFC07, FK99, LC05]. Globus-based [BMFC07]. GLORIAD [JHL^{+}06]. GMPLS [Pal06, THN^{+}06]. GMPLS-based [Pal06]. GMRES [BB04]. government [Nag86a]. GPS [HYS04]. GRADE [KDFL99, KKS08, WDD00]. gradient [Cro95, Ven08]. gradient-based [Ven08]. GrADS [YD05]. Graduate [HY03]. Grain [Vre88, HV92, Vre89]. grained [LC03]. GrAL [Ber06]. grammar [HK88]. grammars [BM00, Hal88, Rus90b]. Grand [Hul89, Kol89, Wil89, Rho89]. Granularity [ABF93, AR98]. GRAPE [CLP95]. Graph [KB00, Vre88, BK97, BHK90, Büc05a, DL04, Gut00, KSAOK08, RSV90, Vre89]. graph-based [Gut00]. graphical [BFL99, KDFL99, SSKF95]. Graphics [LM90b, Igl04a, MJ00, SLDK03, SW06, TDL05]. graphs [Sap88]. gravitating [KMN^{+}05]. gravitational [Fin99]. gravitational-wave [Fin99]. GRB [ACE02]. GRB-GSIFTP [ACE02]. GRED [KDFL99]. greedy [SK05]. GREMLIN [Höf03]. Grid [CF09, KSM^{+}07a, MFE^{+}08, SGH^{+}08, TJLT00, VBP03, Zhu07, Aba06b, ABG02, AAB^{+}07, AG05, BGL08, BdCYG05, BBG^{+}05, BBLP05, BGH^{+}03, BS04, BGL^{+}05, BX04, BGK^{+}05, BCB^{+}07, CTT02, CDF^{+}05, CGT07, CS05, CCLS09, DLW07, DHB02, DDM^{+}08, DSS07, FMS08b, GRH05, GHWZ94, GMB^{+}05, GLM^{+}08, GP09, HWS07, HZC^{+}08, HB08, HY09, HCL07, HML09, HMP04, KKS08, KTY03, KA08, KB09a, LL04c, LC05, LYMZ09, LWSC07, LLSH07, MCQ^{+}07, MMVV08, MJ06, NHG02, NHG03, PP07, PSR^{+}07, PEG05, PBB^{+}05, RSSD02, RvdSB^{+}03, SVC^{+}07, SAMN02, SPCL04, SPEW09, SD07, SBA^{+}05, TMT^{+}07, TMTY05, VLC03, VF01, WSS^{+}09, WL05, WHP09, YHJC05, YD05, YdOLS^{+}05, ZS05a, ZSI08, ZZ09, CCG07, FR08, MVRM08, MSKT07, AMD08, AFP07, ACC^{+}05b, ABB^{+}05, ABB^{+}03, AT02, ADF^{+}05, AN08, ADK^{+}09, BGJ^{+}06, BMFC07, BCT^{+}07, BPP^{+}07, Ber06]. Grid [BLAV06, BST^{+}08, BCMA07, CCM07, CGM^{+}07, CSJN05, CTT^{+}08b, CC09, CCHW03, DCS^{+}07, DMG^{+}08, DKD08, DS08, DMZ09, ET08, EPJ^{+}05, FGCM07, FT07, FS07, FM08, FdSC07, GR09, GRPL04, GFR^{+}06, GMA07, GMS09, GGM^{+}09, HBH09, HT02, HSH^{+}07, Hua05, HPLL08, HPLL09, HML07, ILJ^{+}08, JAA07, JF05, KT08, KA09, KBM^{+}02, Kim07a, hKcF09, KV09, KMK09, KBB^{+}09, KSM^{+}07b, LS07a, LTOT07, LWHC07, LvSW^{+}04, LGW07, LB09, LKG08, LOJ^{+}07, LW08, LSTV07, LK08, LJW08, LBB^{+}09, LS08, LWHS07, MTNM08, MLD08, MKT09, Mes02, MHA09, NSP07, NHG06, NJW^{+}06, NK07, Ole07, Pal06, PKC^{+}05, PW09, PSA^{+}09, PT05, PGPW09, PSP^{+}09, Pro07, QP08, SBHD08, Sin07, SSF^{+}09, SLS^{+}09, SSK^{+}08, SDD^{+}09, SJTG07, SVB07, SHLB08, TMV^{+}07, THN^{+}06, TLYT05, TLTY06, TG07, TKT^{+}08, TDF07, TBNF09, VGBLGS^{+}06, WH05]. Grid [WC06b, WC06a, WS05, WBT05, YLC^{+}06, ZL04b, Zhu04, Zhu07, ZDR07]. grid-aware [MMVV08, PSR^{+}07]. grid-based [BGH^{+}03, GHWZ94, AN08, DMG^{+}08, DS08, NSP07, PSP^{+}09, VGBLGS^{+}06, YCX05]. grid-enabled [GLM^{+}08, LWSC07, MJ06, ABB^{+}03, EPJ^{+}05, FS07, NJW^{+}06, PKC^{+}05]. Grid-enabling [KT08, SSK^{+}08]. Grid-like [Ole07]. Grid/distributed [hKcF09]. Grid/P2P [Kim07a]. Grid2006 [BGL08]. GridFTP [ACE02, CG09, RKSU08]. GridICE [ADF^{+}05]. Gridification [VWD^{+}08, MZC08]. GridLab [SAMN02]. gridless [BVDF00]. gridmap [ACC^{+}05b]. gridmap-file [ACC^{+}05b]. GridNetworks [SCP09]. GridR [WSS^{+}09]. GridRPC [CCDS08, SKT^{+}08]. grids [ACC^{+}05a, ABM^{+}07, ABF^{+}03, AR07, CCL07, CGL08, CD08, DT08, FLPP05, Fra08, IT05, KFC^{+}07, MPPM09, OK02, PGSM05, PX07, PPH^{+}09, PPSS06, SBG^{+}09, SMK05, SGH^{+}08, TV08, VDPHS09, WFC07, ZCW^{+}04, vOHD^{+}05, vdS04, GHWZ94, CC07, CCL09, CT09, CGST09, DFC^{+}08, DRNMC09, Doḡ09, EGK^{+}07, FMS08a, GCCC^{+}07, Joh02, KFP^{+}02, KNK^{+}08, KTM^{+}08, LS07b, LHL09, LVH08, LKA^{+}08, LK08, MFP05, MTV05, MHA08, MV09, NZQ07, OVK^{+}09, Pal09, PH07, PK08, Qin07, SYT09, TDV^{+}08, TTP^{+}07, VLK09, YYW^{+}09]. GridSim [SPBT07]. GridX1 [AAB^{+}07]. Griz [RvdSB^{+}03]. Groot [vdR87h]. Group [HWW04, BKSS02, BC03, BDL06, CMO03, CYH04, Fio06, LM07, MPQ03, NKX09, PP06]. group-key [NKX09]. Group-oriented [HWW04]. grouped [LYQ06]. grouping [GMEL08, Mor01]. groups [CWD04, LPE08, NSI02, ZDR07]. growth [SUD^{+}98]. GS1000 [LM90b]. GSC [GML99]. GSC-II [GML99]. GSiB [Hua05]. GSIFTP [ACE02]. guaranteeing [MKT09]. guarantees [KV09]. Guest [AB01, AD00, Ano01c, BJC02, DHS00a, DDS00, Flu03, HP92, HB98, Kaa98, KK00, LBR02, LC01, SG95, Tan02a, Wil00, ZEO01, vdR93a, Kaa99]. guided [EdBG^{+}99]. GuiGen [RSSD02].

H [vdR87f]. H2O [SSB05]. HA-OSCAR [LSLS05]. Hall [vdR87h]. Hallam [vdR87a]. Hamiltonian [MR04b, Amo06, MR03a]. Handling [Goo01, HSS00]. handover [PSL^{+}04]. haptic [WWSM98]. hard [KV09]. Hardware [TBK06, DSS98, HHSW92, Hum92, NWE04]. hardware-based [DSS98]. Hardware-oriented [TBK06]. harmonic [BB04, KDE04]. harness [PL96, BDF^{+}99, FD02, MS01]. Harrod [GDRS04]. hash [JFDF09]. HASTE [PPJ95]. Hawaii [Hul89]. Hayes [Zem86]. Hayes-Roth [Zem86]. HD [JHL^{+}06]. health [Pol98, PPSS06, PPAK99, Wit94]. heart [LGS^{+}07]. help [Ueh89]. HEMT [Abe92]. heterarchy [DS04c]. Heterogeneous [GBA^{+}09, Sun92, ACGdT02, BL02, CBK^{+}01, CG02, CKFJ06, CS09, CFG^{+}05, DZ04, DRNMC09, DY04, EW97, Fer96, Höf03, HCL07, HXL90, Kos95, MM03, OVDV98, Öst92, PBB^{+}05, Shi04, WC01]. Heuristic [Bal91a, WL05, JM01, MC00, THKG98, ZZ90]. heuristics [GODM98]. Hierarchical [LOJ^{+}07, ASTEP98, CFGC03, EV96, EWG99, HV92, HML09, LKG08, LLZ07, OSCY93, ZDR07]. hierarchically [CBK^{+}01]. Hierarchy [ZZ90, LHC03]. High [AB01, BBC^{+}99, Bhu95, BM92, Bis94, CST92, DRS^{+}97, Gen95, GAB^{+}96, HSS92, HML^{+}06, HJPS03, KPB^{+}03, KKK07, MFP05, SEH99, Ste94, WBF08, Aba09, AFP07, ABB^{+}03, ACU95, AMW99, AB03, BL98, BMFC07, Ber96, BCS99, BS04, CGT07, CWD^{+}08, CHJ^{+}04, DCS^{+}07, GNOY01, GGH^{+}06, GGSZ09, HHSW92, HDC^{+}94, HAE^{+}03, HG92, HAB^{+}06, JM02, KBM^{+}02, KSM^{+}07a, KSM^{+}07b, Lau01, LSLS05, LRJ^{+}06, LC01, LLSR02, MI01, MR03b, OS01, PH99, PCB99, RJH^{+}09, Reu03b, STH^{+}98, SGFS01, SB97, Sch03, SLDK03, SRG^{+}03, SRCR97, TSZP99, VBP03, Wit94, Ben99, CMZ95, Din99, HB98, Lid99, LBB^{+}09, Wil00, dSL98]. High-definition [HML^{+}06]. high-dimensional [BS04, CHJ^{+}04]. high-energy [ABB^{+}03]. high-latency [ABB^{+}03]. High-Level [HSS92, WBF08, BMFC07, KSM^{+}07a, KSM^{+}07b]. High-Performance [BM92, AB01, BBC^{+}99, Bhu95, GAB^{+}96, SEH99, Ste94, Aba09, AMW99, CWD^{+}08, GGH^{+}06, HHSW92, KBM^{+}02, LLSR02, MI01, Reu03b, STH^{+}98, SGFS01, SB97, Sch03, Lid99]. high-quality [TSZP99]. High-resolution [KPB^{+}03]. High-speed [MFP05, AB03, HG92, HAB^{+}06, SRG^{+}03]. higher [Dal06]. highlighting [SS04]. Highly [SBSdL06, BLRS98, JM01, KSY92]. Highway [HICÁFM^{+}06]. hints [dNE05]. historical [PGPW09]. HLA [BKK02, CYLT05]. HLA-based [CYLT05]. hoc [GR07, LBYL08, LM07, MV09, SM01a]. holds [PM00]. holes [SHB89]. Holland [vdR87c, vdR87e]. Home [HMC06, SLO^{+}05b]. Home-based [HMC06]. homogeneous [Höf03, HXL90]. homologous [BORM07]. homotopy [CFVP03]. hop [GNWT05]. horticulture [KH89]. Horwood [Ano86i, Ano87b, Ano87l, vdR87d]. hot [HSH^{+}07]. Householder [DV03]. HPC [BHRT98, DMSS97, PA01b]. HPCN [EV98, Mur95, Par94, Ros94, Wal94]. HPCN96 [SL97]. HPF [Ben99, BCW01, CMT01, MBFC99]. HPF-like [CMT01]. Huard [HPP94]. Hubert [Zad87]. human [DMMP98, DIK^{+}06, KZC04, Pag99, Zad87, vVDBB98]. humanities [BHD09]. humanoid [NMA00]. hundreds [KTY03]. Hybrid [SvAS01, Ban02b, Boa04, BKK02, DS00, ED04, GDRS04, HQ07, HG92, KKL06, LYMZ09, Low01, MV09, SSS02, TMT^{+}07]. Hydrodynamic [NCS04, RBS93]. HyO [DL04]. HyO-XTM [DL04]. hype [BYV^{+}09]. hyper [DL04]. hyper-graph [DL04]. hyperbolic [NB04]. hypercube [FHG95b, Shi92]. hypercubes [Xia06]. hyperspectral [PM04]. hypre [FJY06].

I/O [Aba06a, BFK02, CkLC06, Geo02, HMC06, OK02, SM96]. I/O-intensive [CkLC06]. IAN [LG08]. IBP [BBLP05]. IBPster [ASPB03]. ICOT [BY93, Kuo86, Tic93, Uch87]. ideal [FPdS04]. ideas [Pin87]. identification [DCS^{+}07, SK06]. ideology [Sap88]. IEEE [BGL08, CF09]. IEEE/ACM [BGL08]. IFC [Ano02b, Ano03h]. IGrid [GGH^{+}03, DBdL03, SDBdL06]. iGRID2002 [MWC^{+}03, CALN03, LLH^{+}03]. iGrid2005 [GdBW06]. II [vdR87c, GML99, WGL92]. III [KGW95]. illustration [Avg00]. ILU [KZC04, LN94]. ILU-relaxation [LN94]. image [BDNP92, CEGL01, CEJK94, EFD00, FP03, GML99, HYS04, KKV^{+}99, KE85, LLWN04, MGLV04, OCdAM07, PK99, Pet95, PCM99]. imagery [EdBG^{+}99]. images [BCT^{+}07, CCM98, MLSO01, RICW00, SGL99, SMC99]. imaging [DMMP98, DMM^{+}99]. Immersive [WKF03, CN98, RMM^{+}98]. IMMSIM [PKSC02]. Immune [BCS99, ED04, KRLR01, PKSC02, ZZN04]. impact [Car03, Dub91, TLTY06, WAD^{+}89]. implement [HHSW92]. Implementation [BS91a, CYB90, CCKW88, DCK03, EPJ^{+}05, GL95, GSD95, MMVV08, MD92b, Pit96, SMI01, WMN^{+}01, mM95, ABG02, AKMK05, ACGdT02, ABK94, Ber96, BS92, BG05, BL92, CMZ95, CS05, CY88, CD99, DHS00b, DvdHGdL09, GD93b, GY90, GC94, JLU03, LJY04, LL04b, LY90b, LH07, LC03, Luk00, LMH^{+}09, MVT^{+}99, Mur88, OBK88, PR95, PSR^{+}07, SM96, VVC^{+}03, VSM02, VS88, YJA03, YdOLS^{+}05]. Implementations [VSvD95, Ano86i, NSS99, Ref87]. Implementing [CS93, YJA03, HS98, Pap05, RM97]. implications [PR95, vdR87d]. implicit [ID98]. importance [MS03]. impression [vdR87l]. improve [LK08, SMC99, ZMS^{+}06]. improved [KKYK04]. improvement [Hol93]. Improving [Dua94, PKC04, SS04, TS08, AB95, HBCR01]. in-core [CKFJ06]. In-VIGO [ACC^{+}05a]. incentive [RHB08]. incentive-based [RHB08]. including [LRMC94]. incomplete [MvdV01]. incompressible [VFS01]. incorporating [SPBT07]. Incorporation [GMB^{+}05]. increase [LVH08]. Increment [SS03]. incremental [PVBH05]. increments [Tor04]. independent [Fio06, Ger02, LHL09]. independently [BSCC06]. Index [Ano00a, Ano00b, Ano01a, Ano01d, ÅO06, Ano85b, Ano86a, Ano87d, Ano89a, Ano90a, Ano91a, Ano92a, Ano92b, Ano92h, Ano92i, Ano93a, Ano93i, Ano94a, Ano94b, Ano94g, Ano95a, Ano95h, Ano98a, Ano98c, Ano02a, Ano02c, Ano03a, Ano03i, Ano04a, Ano04b, Ano05f]. indexing [KCK04, SLŽ^{+}09, ZQZZ09]. indices [TBD^{+}02]. individual [Sch01, Mor01]. Individual-based [Mor01]. Induced [vOB95]. Industrial [CHK98, Gal87, LG08, Luk00, Sch94, Sin84, Smi86, LG08]. Industry [Kaa99, Mur95, BGR^{+}99, Kaa98, LPC^{+}95, Ros94]. Inexact [BMZ01]. Inference [AP96, BB84, Her84, HV84, KKYK04, Bal93, DD86, Uch87, BB85]. inference-making [DD86]. infinite [Tab06]. INFN [ABB^{+}05]. INFN-Grid [ABB^{+}05]. informatics [KZBK99]. Information [AFP07, ABMS05, FMD99, Kim07b, SR03, Ano84a, Ano86j, BRR^{+}04, BDNP92, Car86, Dek86, Dub91, EV98, FMS08b, GB99, HYC04, JL08, JC09, KKYK04, KH89, KK97, LXLS09, LC03, MSKT07, Mar98b, Mar98a, Mar99b, MS03, Nag86a, PKC04, PGPW09, SPCL04, SNA92, Wal86, WBT05]. Informatique [Ano85a]. infrared [HYS04]. Infrastructure [AT02, BCT^{+}07, Car03, CO03, CD08, GdBW06, GNWT05, Hua05, JLU03, KBB^{+}09, NSS99, STTK03, SP93]. infrastructures [PT05]. initial [SSMG95]. Initiation [OVK^{+}09]. initiative [DR89]. initiatives [dLRW03]. initio [LDSH95]. injection [HTV07]. innovative [KTM^{+}08]. input [Aba06b]. input/output [Aba06b]. ins [MS01, HYS04]. Insights [Cur92, KF00]. inspired [ZEO98, ZEO01]. instance [Leo01]. instance-based [Leo01]. instances [BKSS02, Leo98]. instantaneous [MRV01]. Institutes [Van87b]. instruction [GD05]. instructions [Goo01]. instrumentation [BP01, MPPM09]. instrumented [GD93b]. integral [de 94]. integrate [BFLL99, PBB^{+}05]. Integrated [CT09, SDBdL06, BCMA07, DZJ^{+}00, Dui89, GDRS04, IdLR01, LY90b, Ref87, STTK03, SKT02, WG91]. Integrating [BCMR01]. Integration [AFF^{+}09, AKW90a, AMB03, CGST09, DPP03, DDR^{+}07, HYC04, MR03a, MR04b, SGY^{+}07, Sin07]. Integrative [LGS^{+}07]. integrator [JL03]. integrators [DV03, IS06, MPQ03]. integrity [OB04, SL87, SHJR04]. Intel [FHG95b]. Intelligence [Kow85, vdR87a, All92, Ano86l, Ano87b, Ano87c, Ano87l, DD07, EO86, How91, Kow84, Lop96, MCA02, Niw89, RBC^{+}88, Sch85, Szu98, Szu01, Wii84, vdR87b, vdR87c, Ano84e]. Intelligent [KH89, KK97, MTNM08, CSJN05, IMKB89, KFF89, KRLR01, MWC^{+}03, Ueh89]. intensive [BPS^{+}03, CkLC06, MFP05, Pal09, Pet95, SY04, TJLT00]. Inter [DVV90, DDR^{+}07, GS95]. inter-enterprise [DDR^{+}07]. Inter-process [DVV90]. inter-stream [GS95]. interacting [DS04c]. Interaction [ZAP05, IT05, MJ06, ZS05b]. interactions [Kni89]. Interactive [EdBG^{+}99, JHL^{+}06, PF01, PK08, RS99, BDL06, DK00, HML^{+}06, IJLC03, KLC05, KKL09b, PPAK99, SBSdL06, Sch98, dNE05]. interconnection [ADKS06, JAA09, Shi92]. interconnects [CkLC06, PBHK01]. intercontinental [PBC^{+}01]. interdisciplinary [GFGB03]. interest [SMC99]. interface [AAB^{+}92, BJWZ08, BMFC07, BCW01, DLR^{+}09, MTKS00, OFO^{+}99, DKD08]. interfaces [BFL99, Büc05a, FJY06, Kam85, PNH99, RSSD02]. Interfacing [HC99, PL96, HML07]. interference [Höf03, SdSP04]. interferometer [Fin99]. InterGroup [BAC02]. interleaving [BM08]. intermediate [CEJK94]. Intermittent [JM02]. internal [KKL09a]. International [Ano86j, CF09, SST^{+}06, BGL08, LRJ^{+}06, DBdL03]. Internet [BBM^{+}03, DRNMC09, KKYK04, NCCS99, PSVL02, PH99, PM00, She00, SZGbC04, YA02]. Internet-based [DRNMC09]. interoperability [KKS08]. interoperable [HESM99]. interpolation [DS04a]. interpretation [Hal88, NSHP88]. interstellar [DRS04]. interval [CPK05, Dal06]. Intranet [PPAK99]. Introduction [Ano86k, Fre84, Igl04b, Ano86g, MMR02, FS93]. introductory [BLB03, Ray05]. intrusion [HRJ^{+}04, MC04]. intuition [Zad87]. intuitive [GGW^{+}09]. invariant [HV03]. Invariants [JK92]. invention [Bur02]. inverse [Del06, PBT02]. Investigating [PO00]. Investigation [MS03, BP01, GDRS04]. invocation [BCFS02]. involved [GDRS04]. involving [SPM86]. ion [SBS98]. Iowa [CN98]. IP [AMH02, DvdHdL06, MK04, SWCP03, YCX05]. IP-based [AMH02]. iPSC [BP94, FHG95a, FHG95b]. iPSC/2 [BP94, FHG95a]. iPSC/860 [FHG95a, FHG95b]. IPv4 [KKK07]. IPv6 [BBLP05, LLH^{+}03]. IQ [Szu01]. iRODS [HBH09]. Irregular [VV92, BCMR01, BS04, BFK02, GCK98, SBHD08, dSL98]. ISA [Goo02]. ISDA [WBMP99]. Ising [Coo90]. Isospectral [Prz03]. Isospectral-like [Prz03]. Issue [Mes02, Kaa99]. Issues [BvdBM^{+}93, Dek86, LWHS07, CMT01, FD02, GBT87, TG07, Var00, SM96]. Item [Ano97d, Ano97e]. Iteration [BW97]. Iterative [KKHS01, RN04, BJNH05, BGC^{+}03, Čie04, HV03, dSL98]. ITIS’98 [EV99].

J [vdR87a, vdR87f, vdR87g, vdR87i, vdR87j, Reu03a]. Jacobi [BvdHN^{+}01, BV04]. Jane [PF01]. Janet [BKŁ01]. January [Hul89, Rho89]. Japan [Ano87m, Fuk85, Kag89, KA88, Miz89a, Nag86a, Sas85, Wii84, Yat88]. Japanese [Ano84h, FS93, Kas85, Lin84, Van87b]. Java [ADT03, AC01, AG05, BBBD01, BP01, Ber00, BKŁ01, BFW^{+}03, ESPP01, FW02, GW01, LTOT07, LP01, Lau01, LJ07, LRW01, MI01, MJ00, SM01a, SEH99, SH99, SM01b, YdOLS^{+}05, ZMN99]. Java-based [YdOLS^{+}05, ZMN99]. Java/CORBA [LRW01]. JavaBean [FCW01]. JavaBean-based [FCW01]. JavaBeans [PSS01]. JavaSymphony [JF05]. Javelin [NCCS99]. JCASim [FW02]. JGRIM [MZC08]. JIMM [Jon00]. Job [CCL07, Fer96, LBB^{+}09, BCB^{+}07, CCL09, GRH05, GR96, HXL90, LYQ06, RNJK09, SDD^{+}09, TLTY06]. Job- [Fer96]. jobs [KNK^{+}08, PGPW09]. John [Ano87c]. Johnson [Her87]. join [CC98, YdOLS^{+}05]. joint [Gal87]. Jorrand [vdR87c]. Josephson [Kaw92]. JSBricks [BBBD01]. JSIM [MSX00]. junction [Kaw92]. JUNET [MK88]. JXTA [HD05, SZC05]. JXTA-based [SZC05].

Kalman [TM05]. Karamjit [Ano87c]. KBS [SP93]. KDD [DSSU97, FS97]. KeLP [MBFC99]. KelpIO [BFK02]. Kerberos [MPPM09]. kernel [HKT94, RT06]. kernels [BFR05, CGSZ95, JNPY06]. Key [DCL00, ZDR07, LCP04, LBYL08, LM07, LHC03, LYT^{+}05, NKX09, PM00]. keyboard [Ale97]. Keys [EHMS00]. Keystroke [MR00]. keyword [SA07]. Kites [VLC03]. KL1 [Bal93]. Kluwer [vdR87f]. Kluzniak [Teb86, vdR87i]. knapsack [VDPHS09]. Knowledge [Chv87, DD07, How91, SPCL04, SNA92, TY85, Wal86, Wie85, Zhu07, Ano86j, BPAP92, CSC^{+}92, DL04, EO86, EdBG^{+}99, Hol93, KFF89, Kim07a, Mae89, NSI84, Nis93, Ohy89, OFT09, PP07, Qu04, SY04, Van87b, YMM00, ZY90, dLLA93, vdR87l, ZL04b, vdR87g]. Knowledge-based [How91, SNA92, Ano86j, CSC^{+}92, Hol93, KFF89, Mae89, Ohy89, dLLA93]. knowledge-guided [EdBG^{+}99]. knowledge-intensive [SY04]. KOAN [BP94]. Kogan [vdR87b]. Kowalski [Ano84k]. Kronecker [BPS06]. Kruse [vdR87k]. Krylov [Dat03]. KSR [LF95a]. KSR-1 [LF95a]. Kurganov [NB04]. Kutta [CP06].

L [Zad87]. Lab [BGJ^{+}06]. laboratories [AKB^{+}01]. laboratory [BGH^{+}03, GRPL04, GFGB03, DBdL03]. LaCoS [BHH91, BHH92]. LADY [WB90]. LAIOS [Dup90]. Lambda [LRJ^{+}06, SDBdL06, MGYC06, THN^{+}06, CT09, GdBW06, PW09, TDV^{+}08, YLC^{+}06, THN^{+}06]. LambdaGrids [NF07]. LambdaRAM [VBLS09]. LAN [HO02]. Land [LNJ04]. Land-use [LNJ04]. lands [CALN03]. Language [Ano86l, BS91a, HSS92, Ano85g, AP96, BY93, BS92, DHS00b, ES94, RBC^{+}88, ST98]. Languages [CMZ95, Bal92, BL92, JBA94, MRV92, MSS02, Omo91, OP95]. Large [BBJ^{+}06, CTT^{+}08b, CHK98, FAJP99, HKM^{+}06, LPK94, WVC05, BBSV92, BDL06, BDHK06, BCW01, BCG05, BCH^{+}08, CDF^{+}05, CR92, CTMO06, CCHW03, Dat03, DAM08, Din03, DB99, GLA88, HAB^{+}06, Joh02, KPB^{+}03, Kło05, Kos95, KTV03, Luk00, MTV05, MKH06, MLSO01, NEJP94, PBHK01, RN04, RL98, RdLM06, SKT^{+}08, SNW01, SLO^{+}05b, Var00, VSV95, WKZ^{+}03, YHJC05, YCAS03, YCX05, GPH^{+}94]. large-eddy [NEJP94]. Large-Scale [CHK98, HKM^{+}06, WVC05, BDL06, BCH^{+}08, CDF^{+}05, CR92, Dat03, DAM08, Din03, GLA88, Joh02, KTV03, MTV05, MKH06, RN04, YHJC05, YCAS03, YCX05]. laser [Fin99]. Latency [Höf03, ABB^{+}03, DHB02, HML^{+}06, PBHK01]. latency-tolerant [DHB02]. Lattice [AKH^{+}04, Bog99, SdSP04, ABL04, CM99, CH04, DS04b, DC00, FPdS04, HORC04, HRJ^{+}04, IOO04, KKHS01, DY04, NCS04, vdS04]. lattice-Boltzmann [FPdS04, HRJ^{+}04, KKHS01]. layer [CCLS09, DvdHGdL09, Kni89, KSM^{+}07a, KSM^{+}07b, PBHK01]. layered [DD07, Kim07a, MS01]. layout [IdLR01]. laziness [KGW95]. lazy [CK00, PY00]. LBGK [HvHAS04]. Leader [SW02]. learned [HS98]. Learning [BGJ^{+}06, NK05, Pud01, ZZN04, Ami90, AV00, BLAV06, GBE00, NSI84, SH90, VGBLGS^{+}06, Ven08, Ven09, Ano86h]. Lease [LLKF09]. Lease-based [LLKF09]. least [vdV89a]. lecture [Ray05]. legacy [BBW08, BKŁ01, KT08, LRW01, MLD08]. legal [Ser95]. Legion [NHG03, CKKG99, NHG02]. Legislation [Ano84i]. Lenat [Zem86]. length [CP06]. Lessons [HS98]. Letter [Ano87k]. Level [HSS92, ATT96, BSCC06, BMFC07, CLP95, CST92, KKS08, Kea99, KMK09, KSM^{+}07a, KSM^{+}07b, LKA^{+}08, LJW08, Lop96, PKC04, Rus90b, WBF08, YKL^{+}07, ZDR07]. levels [CEJK94, FB93, MKT09, SPBT07]. Leveraging [HWS07, CTT07]. Levy [NB04]. lexical [HK88]. lexical-functional [HK88]. LGF [BBW08]. Libraries [LGH97, HESM99, STH^{+}98]. library [ACE02, BFK02, FB97, FBBW99, GD93b, GLM^{+}08, LC04, PCM99, vdV89a, BMFC07, Ber06, TS08]. Lie [BC03, CMO03, MPQ03, MKK03]. Life [ABM^{+}07, DZJ^{+}00, Kom89a, LWHC07]. life-cycle [DZJ^{+}00]. Light [DS04a, SHB89]. lightpath [GXD^{+}09, JHL^{+}06, KMCH03, SLJ^{+}06]. lightpaths [GMM^{+}09, GFR^{+}06]. lightweight [BK06, EGK^{+}07, SZP00]. like [CMT01, DDL01, Ole07, Prz03]. limitations [ABdLL05, KC98, Par87]. Lindstrom [vdR87h]. line [BW95, EV98, LVH08, Ueh89]. linear [BJNH05, BFR05, BRMN04, BCG05, DHD89, FPdS04, GKS05, HPP94, Lop03, LD04, Pan95a, PHL98, PH94, SGFS01, SG04, vdV89a]. linear-time [LD04]. linearization [Röb05]. Link [CC00, DFSZ88]. Link-time [CC00]. linking [ABCD00]. links [ABdL^{+}03]. Linux [VSM02]. Lip [KLH^{+}04]. liquids [Fre94, SBS98]. Lisp [Oku92, YTHY84]. Lisp-Based [YTHY84]. list [Par04]. literature [JC08]. live [TDG^{+}06]. Load [BL02, CY01, OSHH96, dRSS97, BCMR01, BM08, CWD04, CSJN05, Cho04, KNK^{+}08, LYMZ09, MM03, PGSM05, Qin07, SB97, SMA08, ZMS^{+}06, ZSI08, ELvD^{+}96]. Load-balancing [CY01, PGSM05, ZMS^{+}06]. Local [Han03, REM04, AEGF^{+}01, BPC^{+}01]. locality [BCMR01, HBCR01, Leo98, Leo01, TSK03]. localized [WKZ^{+}03]. Locally [PFMC04]. location [MK04, PKC04]. location-aware [PKC04]. LOCCS [DT94]. LoDs [PFMC04]. Log [JD94]. Logging [RT06, AMH04, PY00]. Logic [De 88, Ano84k, Ano85g, BS96, BDNN02, CST91, CY90, DCK03, DT93, DLW86, Qu04, RBC^{+}88, SGdMM96, Yos89, ZT90, ZT91, Zha93, ZS90, dB90, vdR87h]. Logical [BB84, BB85, WMN^{+}01, Vau93]. logical-time-based [Vau93]. logically [MRV01]. Logicflow [KP00]. logistical [BBG^{+}05]. logistics [SPCL04, ZL04a]. London [vdR87b]. Long [OS06, KTV03, MGH^{+}05]. Long-range [OS06]. Looking [Büc05a]. loop [KC98]. loops [LRMC94, XHY^{+}90]. Loosely [BDNP92, Mis92, AGP^{+}92, MD92b]. Loosely-coupled [Mis92]. loss [LNJ04]. Lossless [DQ97, CCM98]. Low [SHJR04, CEJK94, CALN03, GE90, HML^{+}06, MTKS00, DT94]. low-latency [HML^{+}06]. LSI [Abe92]. Ltd [Ano87b, Ano87l]. LU [MvdV01]. Lyapunov [DE03]. Lyon [BBD^{+}99].

M [vdR87g, OSCY93]. MA [vdR87g]. Mach [CR92]. machina [PKSC02]. Machine [Ama88, Ama89, Ama86, BCL88, BVP^{+}87, BDF^{+}99, BHH^{+}93, BS91a, BB84, BB85, CCKW88, CH95, Por95, YTHY84, ASTEP98, Ano87l, Bal93, BS92, BFC02, CPB00, Gur85, KSY92, Mur86, Nag86b, NTN86, Ném00, Nit86, Ram95, Sim86, Uch86, WG00, Zad87, ZY90]. machine-room [Ram95]. Machines [SvAS01, BHK90, BM00, CC98, DHS00b, DHS00a, DK00, GBT87, KZLK06, KTY03, PM04, TDG^{+}06, Uch87, dlTK92, SM01b]. Macroeconomics [HPLL08]. macromolecules [Kol89]. macroscopic [DS99]. macrostep [Kac00]. made [SS90]. Madison [vdR86b]. Magazine [EV98]. Magnetohydrodynamic [GPH^{+}94, Ano96b]. magnetohydrodynamics [BvdHN^{+}01, MPG96]. Magnus [DR03]. mail [LL04b, Öst92]. major [Lit03]. makespan [RNJK09]. Making [CN98, HNS05, WBT^{+}08, DD86, GS05, KFF89, Niw89, PP07, vdR87f]. MAN [TDG^{+}06]. MAN/WAN [TDG^{+}06]. management [Aba06b, ADM06, Ano86j, ADK^{+}09, BJWZ08, BFL99, BPP^{+}07, BDL06, BCPS03, CO03, CKK^{+}04, CKKG99, CT09, Coo86, DCL00, DMG^{+}08, DS08, DSS07, DL04, Fer96, FHM^{+}99, Fra08, GLA88, Gra01, GR07, HY09, HKT94, HB00, HML07, HML09, KA08, KKL09a, KD00, KH89, KK97, KLSS05, LS07b, LBYL08, LSTV07, MM08, MK04, PK99, PH07, PPH^{+}09, SEH99, Slo96, TMTY05, VVC^{+}03, WZC08, WBT^{+}08, YMM00, ZDR07]. Managing [Kos00, LKA^{+}08, SCY01, ACC^{+}05b, Gos00, PSA^{+}09]. Manchester [Gur85]. Mandel [vdR87e]. MANETs [JFDF09]. Manifold [AH94]. manipulation [DMMP98]. manufacturing [HDC^{+}94, WG91, ZMS^{+}06]. Many [Len01]. Many-particle [Len01]. map [DR03, KLC05, FJ00, DL04]. MAPFS [PCG^{+}06, SPK^{+}07]. Maple [YA07]. Mapping [AEGF^{+}01, Van92, dRSS97, KTM^{+}08, LJW08, MS03, Reh06]. Market [Wii84, JRF^{+}07, SVB07, Wil86, Yam89]. Markov [BDHK06, BCB^{+}07]. marrow [BPP^{+}07]. MARS [GR96]. marshaling [PVBH05]. Massive [SG95, BORM07, KKK07, WRBG94]. Massively [DDO^{+}92, KL02, Cas94, CS96, Gil94, JBA94, KZLK06, PN09, YdOLS^{+}05]. master [Bun03, HCL07]. matacomputing [HJS^{+}99]. Matching [DDRR96, KNK^{+}08, MM03]. materials [Gil94, SYT09, SCK^{+}00, WRBG94, WAD^{+}89, ZMS^{+}06]. Mathematica [UB07]. mathematical [KSAOK08]. mathematics [AMB03]. MATLAB [BKG05, EPJ^{+}05]. MATLAB-based [BKG05]. matrices [Amo06, BH03, dSL98]. matrix [BST^{+}08, CDS03, Dat03, HBCR01, Kat04, LC04]. matter [Tof99]. MAX [SH00]. maxillo [SGL99]. maxillo-facial [SGL99]. MCC [vdR87l]. McCord [vdR87g]. mean [SZ98]. mean-field [SZ98]. means [Avg00, DS04b, PW09]. measure [Szu01]. measurement [DMM^{+}99, PSP^{+}09, VVC^{+}03]. measurement-based [VVC^{+}03]. mechanics [Ald89]. mechanism [Cho04, CC09, FD95, HY09, JFDF09, KTM^{+}08, LCP04, MV09, ZZ09]. mechanisms [CST92, CY88, GDRS04, Sar02]. media [KMN^{+}05, RS99]. MediaGRID [VWD^{+}08]. mediator [WBT05]. Medical [EV98, KE85, BCT^{+}07, CD99, CCM98, EdBG^{+}99, KKV^{+}99, LLSR02, MJ98, MJ06, Mar99a, MLZ^{+}00, SMC99, XYZ05]. medicine [AV00, CCBR98, MSR98, Wit94]. MediGRID [KBB^{+}09, SLS^{+}09]. medium [DRS04]. meets [TG04]. Meiko [CW93, GL95, LF95b, RBS93]. MeikUS [GD93a]. Mellish [vdR87a]. memories [De 88, HBCR01]. Memory [BPS06, BM92, CSP98, GSD95, KGX95, PBM95, Ami90, ABL04, BFLL99, BK97, BBSV92, Ber98, CR92, CS93, FC05, HH98, HPP94, HG92, KN06, KSM^{+}07a, KSM^{+}07b, LWSC07, Mar90, NP03, PY00, Par04, PH94, RCD03, SF06, TC06, WYJ99, WWC^{+}97, SG05]. Memory-efficient [BPS06]. menu [LC03]. merchants [ABCD00]. mercury [HRJ^{+}04]. mere [MBZL09]. Merging [Kat04]. mesh [BRR^{+}04, CH95, DS04a, KCK04, LHB95, RŻDM01, UWV92, WC01]. mesh-connected [UWV92]. meshes [JL98, PFMC04, SW05]. Message [BFR99, FLPP05, Gor02, AMH04, AC92, Ber98, CST92, Kac00, Kal94, DKD08, LDSH95]. Message-based [FLPP05]. Message-passing [BFR99]. messages [MTKS00, ZCT^{+}04]. messaging [OF07]. Meta [ATT96, CTT^{+}08a, DD05, FHM^{+}99, HML07, Lop96]. meta-applications [FHM^{+}99]. Meta-level [ATT96, Lop96]. meta-modeling [DD05]. meta-scheduling [HML07]. Metacomputer [ESPP01]. metacomputers [BvdV99]. Metacomputing [BGR^{+}99, PBC^{+}01, RRS99, BFL99, BFR99, EMB98, GR96, HAFF99, HAF00, Kea99, Lee04, MVT^{+}99, MS01, Ram95, SSB05, WSH99, BKKW99]. metadata [BJA^{+}05]. METAFOL [BS96]. metagenomics [SGP^{+}09]. metaheuristic [DOV01]. metal [HW95]. metamodel [LKG08]. metamodel-assisted [LKG08]. metaphors [MB01]. metascheduling [LGW07, MHA08]. metasystems [CWW^{+}99]. Method [DB99, BCFS02, BvdHN^{+}01, BS04, BV04, Bou95, Cro95, DS04b, DS99, FT07, HESM99, HvHAS04, HPP94, HPLL09, HRJ^{+}04, IOO04, JNR01, Kom89b, Lee04, LNJ04, SKT^{+}08, SLW01, VFS01]. Methodology [DDO^{+}92, vdR87c, BJWZ08, Mat89, OP97, SKT02]. Methods [BHH91, GLS99, PBT02, AHdJF97, BMZ01, BHH92, BC03, BST^{+}08, CP06, CMO03, CM99, Čie04, Dat03, DLP06, DR03, EL03, EWG99, HND06, IS03, KG01, Mal94, Mar90, MR03a, MR04b, NK05, PP06, Rum99, Tof99, ZM97, ZMS^{+}06]. metrics [TDF07]. metrology [SM03]. metrology-based [SM03]. metropolitan [PSVL02]. Mexico [HICÁFM^{+}06]. Meyer [vdR87k]. MGF [GLM^{+}08]. micelles [vOB95]. Michie [Ano87l]. micro [GD05, Gra01, Low01]. micro-architecture [GD05]. micro-organism [Low01]. micro-world [Gra01]. microarray [BPP^{+}07]. microbenchmarks [BBBD01]. microbial [SGP^{+}09]. Microcomputers [Mat88]. Microelectronics [Bal91b]. microkernel [DDV92, SBHD08]. Microscopic [ABdL^{+}03]. microstructures [CKR04, MVAS89]. microwave [DD05]. Middleware [PSVL02, ADK^{+}09, Car03, CD08, DVVD02, DPP03, EGK^{+}07, GMB^{+}05, GGH^{+}06, LDS06, PNH99, SZP00, SGH^{+}08, SHJ06, YJA03]. middleware-based [DVVD02]. middlewares [AFF^{+}09]. MiG [MPB^{+}07]. Migol [LS08]. Migrating [KLM^{+}05]. Migration [GWO03, PCBD99, dRSBH94, AMB^{+}92, CYLT05, SM01b, TDG^{+}06]. Military [Gil85a]. millennium [KZBK99]. Miller [vdR86b]. MIMD [CS93, DFSZ88, Hey90, Kal94, KM01, Pri95, VSV95]. MIMD-multicomputers [KM01]. MIMD-supercomputers [DFSZ88]. MIN [SH00]. Mind [Zad87]. MinEX [DHB02]. Mini [Kaa98]. Mini-conference [Kaa98]. minigrid [YBQ07]. minimising [DHS99]. minimize [RNJK09]. minimizing [GR96]. minimum [Dzw97]. Mining [DSH^{+}99, LGW07, MSKT07, SA97, AW97, CTT02, CGM^{+}07, CTT07, CS97, FS97, FJT01, FM01, FdSC07, GGH^{+}06, HPS97, JRF^{+}07, LWHS07, LLSH07, NSP07, OFT09, PSR^{+}07, Sin07, SD07, SSK^{+}08, ZLD^{+}03]. MIP [MMVV08]. MiPeG [CD08]. Miscellaneous [Ano97d, Ano97e]. miscible [FPdS04]. mission [BPAP92, Lau92, MC04, TMV^{+}07]. mission-critical [MC04]. MiX [AGP^{+}92]. MJSA [BCB^{+}07]. ML [GL05]. mobile [AMH02, AMH04, CWD04, ESPP01, FX07, GR07, GWO03, HJK^{+}04, LM07, OF07, PWY03, PKC04, SCB04, UTT00, LSTV07]. mobility [AMH02, HZC^{+}08, MK04, OVK^{+}09]. mode [FC05, HNP05, MRV01, HHG05]. Model [BHH91, LB09, LH07, BvdBM^{+}93, BMPS01, Ban02b, BFLL99, BdCYG05, BHH92, CPK05, CWW^{+}99, CY90, CYB90, CDRS05, CS09, DS08, FPdS04, GY90, GB99, HD05, HHS98, IMKB89, JL95, JAA07, JF05, KP00, KSAOK03, KSAOK08, KCK04, KRLR01, KGW95, LS07a, LR06, LHL09, LF01, LvSW^{+}04, LLZ07, Low01, MTV05, MKH06, MLZ^{+}00, OP95, PO00, Pal01, PSS01, PKSC02, RSV90, RS98, SGL99, SCB04, SD02, Shi92, SUD^{+}98, Szu98, TDL05, VDPHS09, VP94, WG00, WC06b, WHP09, ZT90, ZT91, ZYXL05, dlTK92, KLM^{+}03, Sti93, PSP^{+}09]. Model-based [LB09, BvdBM^{+}93, IMKB89]. Model-oriented [BHH91, BHH92].

Modeler [FBBW99]. Modeling

[AS02, GL04b, HBCR01, Mor01, Mun04, UTT00, CKR04, DD05, DC00, HKM^{+}06, Igl04a, JAA09, MGYC06, MVAS89, Mid01, PBT02, SZP00, SBA^{+}05, VR05, Wie03]. Modelling [Pap05, ACU95, BPS06, BN06, CGT07, CH04, DS99, DY04, FPX^{+}09, FHG95b, IJLC03, KG01, KH97, Lee04, LGS^{+}07]. Models [GSD95, KGX95, TTP^{+}07, BDHK06, CDF^{+}05, CS93, DDL01, GBA^{+}09, Joh92, KCT99, Kim07a, KM01, LMBCC89, MR04a, Mor01, Sch01, ST98]. Modern [ABMS05, SSC04]. modes [CASW05]. modified [GDRS04, LNJ04]. modifying [GL04b]. modular [CHJ^{+}04, HML07, RN04]. modulation [HJK^{+}04]. module [WG00]. Modules [HLvL^{+}97]. molecular [BRS04, FAJP99, GRPL04, MDD89, MF93, PHM^{+}99, SCK^{+}00, SBA^{+}05, Szu98, ZCW^{+}04]. molecular-dynamics [SCK^{+}00]. molecules [DRS04]. Molokai [Hul89, Rho89]. momentum [KKHS01]. MOMI [DFSZ88]. MOMI-connection [DFSZ88]. monitor [HMP04]. MOnitoring [Suz89, ACU95, BSG^{+}05, BFW^{+}03, DVV90, ADF^{+}05, BBFW03, CC09, LBB^{+}09, WDD00, ZS05a, Suz89]. monotonicity [EL03]. Monte [BST^{+}08, TMTY05]. mortals [MBZL09]. MOSIX [BL98]. motherboards [HJCD05]. motility [Len01, Low01]. motion [KLC05, SKF^{+}09, Sin92]. motivated [Pet95]. movement [Pal01, Sch01]. MP [LJY04]. MPI [ACGdT02, BCH^{+}08, CRE01, DZ98, FD02, FB97, GLM^{+}08, LR06, LRW01, LS08, MM03, NHT06, Reu03b, SG05]. MPI-based [LRW01]. MPI-DDL [FB97]. MPI-Delphi [ACGdT02]. Mpixel [SW06]. MPP [HW95, vOB95]. MPP-systems [HW95]. MR [DMM^{+}99]. mRNA [TBD^{+}02]. MSIMD [GS95]. Mt. [CDRS05]. Mu [NTN86]. Multi [DF97, GXD^{+}09, IS03, JAA09, MR03a, MR04b, SDD^{+}09, YLC^{+}06, ADKS06, CCLS09, DFG^{+}00, DvdHGdL09, DSH^{+}99, DB99, KLH^{+}04, LJ07, LHL03, MSK03, MR04a, NWE04, Par06, STTK03, SHLB08, TLYT05, TYH04, WHP09, vM94, vOHD^{+}05, Gra92]. multi-agent [NWE04, Gra92]. Multi-application [DF97]. multi-attribute [SHLB08]. multi-block [vM94]. Multi-cluster [JAA09]. Multi-cost [SDD^{+}09]. multi-criteria [ADKS06, WHP09]. multi-dimensional [DSH^{+}99, DB99]. multi-disciplinary [DFG^{+}00]. Multi-domain [GXD^{+}09, YLC^{+}06, vOHD^{+}05]. multi-layer [CCLS09, DvdHGdL09]. multi-paradigm [STTK03]. multi-physics [Par06]. multi-proxy [TYH04]. multi-resolution [KLH^{+}04]. multi-scale [MR04a]. multi-server [LHL03]. multi-signature [TYH04]. Multi-symplectic [IS03, MR03a, MR04b]. multi-tier [TLYT05]. multi-tiers [LJ07]. multi-user [MSK03]. multiagent [PCG^{+}06]. multicast [ADOKM06, BDL06, FMR05, GNWT05]. multicast-based [FMR05]. multicellular [Pal01]. multicluster [CWD^{+}08]. Multicomputer [MF93, BL98, CST91, KD00]. multicomputers [KM01]. multidestination [Pan95b]. multidimensional [BS09, ZQZZ09]. multidisciplinary [ATJMZ02, SMK05]. multifrontal [IST04]. multifunctional [DCK03]. multigrid [LLRS94, LN94]. multilayer [RM97]. multilayered [NKX09]. Multilevel [WC01, SSC04]. multimedia [BS91b, CFGC03, FGRZ09, GL04b, HML^{+}06, Mul92, Pap05]. multinets [Kło05]. multiparadigm [BdCYG05]. multiparty [HML^{+}06]. multiphysics [MTH^{+}05]. multiplayer [PN09]. Multiple [MBFC99, AFF^{+}09, BCPS03, CGL08, FD95, JCSS01, JL03, MVRM08, MI01]. multiple-context [FD95]. multiple-task [MVRM08]. Multiprocessor [SZ98, AAB^{+}92, BR92, BW95, CR92, GE90, Kat04, Kun94, LS01, Nos98, OSCY93, Par04, SKJ01, SF06, Yam92, mM95]. multiprocessors [AB95, Ber98, CK00, TC06, WYJ99]. multiprogrammed [Aba06a]. multiprogramming [Ste92]. Multiresolution [EWG99, Boa04]. Multiresolutional [PCB99]. Multiset [BCL88, MK95]. multisparc [CW93]. multispectral [LLWN04]. multistage [ADKS06]. multisymplectic [IS06]. multithreaded [BBFW03, Ném00]. multivariate [DD05]. multiway [CC98]. München [Bun03]. mutation [DMN^{+}05]. Myrinet [Geo02].

N [FJ00]. N-MAP [FJ00]. nanoscience [GFGB03]. nanotubes [Tor04]. naphtylium [DRS04]. narrow [Büc05a]. NAS [WYJ99]. National [HY03, DR89, FS93, Han03, KFP^{+}02, NSF87]. Native [MGH^{+}05]. Natural [Ano85g, BY93, LPK94, PFMC04]. nature [ZEO01]. nature-inspired [ZEO01]. NAUTA [MSLP93]. Navier [ID98, vM94]. Navigating [vVDBB98]. navigation [VS04]. navigational [SOR05]. NCSA [PCM99]. near [BBJ^{+}06]. need [WRBG94]. negative [SK05]. negotiation [DMZ09, GP09, YKL^{+}07]. neighbor [KKL06]. neighbor-selection [KKL06]. neighborhood [AEGF^{+}01, RNJK09]. NEPTUNE [OS92]. nesting [GL05]. net [LS05, LS05]. NetSEC [SM03]. NetSolve [ESFD06, PCBD99]. Network [Din91, LG08, MTKS00, OVK^{+}09, Pal09, Per86, ADOKM06, ABB^{+}03, BTM06, CTF^{+}99, CCL08, CFG^{+}05, De 88, Del06, DS08, Dui89, GPA00, GFR^{+}06, GdBW06, GMM^{+}09, JAA09, JC09, KKP^{+}05, KGLA85, LJY04, Lee04, LWSC07, LLWN04, MSLP93, OS92, RKSU08, RS98, SZP00, SM03, SRCR97, SPBT07, Sun92, VS88, VSV95, WSH99]. Network-aware [Pal09]. network-based [Sun92]. Networked [FMD99]. Networking [CG09, Gen95, HB98, Lid99, Wil00, AB01, BBG^{+}05, GdBW06, HDC^{+}94, HRJ^{+}06, LPC^{+}95, MWC^{+}03, PA01b, VBP03, dLRW03]. networks [ADKS06, Ami90, ABdLL05, AB03, BS09, BCMA07, Coo90, CBD^{+}05, CS97, DZ98, DvdHdL06, DvdHGdL09, Dör05, Sou91, FM08, FR08, FSM88, Ful91, GKIZ05, GLA88, GR07, GGH^{+}06, GGSZ09, HAE^{+}03, HAB^{+}06, IEG04, JAA07, JL98, KY85, KKK07, KKL06, KD00, LKG07, LBYL08, LM07, LAM07, LS01, MFP05, NK05, NSI02, Ole07, Pan95b, PSL^{+}04, PBC^{+}01, Pud01, RM97, SM01a, Sap88, SVC^{+}07, SNW01, SCB04, SPEW09, SDD^{+}09, SRG^{+}03, SJTG07, SK05, VBLS09, WC01, YCX05, ZLD^{+}03, vOHD^{+}05, vdHDT^{+}06, WLB00]. Neumann [HG92, KDE04]. Neural [Sou91, Ful91, GKIZ05, Ami90, Ban02b, Coo90, CS97, Del06, Fio06, JL98, KKP^{+}05, LLWN04, Pud01, RS98, RM97, VSV95]. Neurocomputing [Mat88]. neurofuzzy [HAAH05]. Neurons [Mat88]. neuroscience [FBBW99]. new-generation [ZGCM00]. Newton [BMZ01]. Next [BDF^{+}99, HJK^{+}04, KZBK99, MWC^{+}03, PO00]. next-generation [PO00]. NGSSC [HY03]. NIC [HAC92]. Nimrod [ABG02]. Nimrod-G [ABG02]. Ninf [NSS99, STH^{+}98]. Nippon [Kom89a]. NJ [vdR87h]. nodal [Bou95]. Node [LG08]. nodes [AMH04, AEGF^{+}01, Reh06]. Non [ID98, BMS05, BCH^{+}08, DCL00, FPdS04, JLU03, KKL09b, LOK09, MKH06, NP03, Tho06, TBD^{+}02]. non-blocking [BCH^{+}08]. non-dedicated [NP03]. non-deterministic [MKH06]. non-interactive [KKL09b]. non-linear [FPdS04]. Non-overlapping [ID98]. non-personal [JLU03]. non-product [Tho06]. non-redundant [TBD^{+}02]. non-trusted [DCL00]. non-uniform [BMS05, LOK09]. Nonlinear [CPK05, BMZ01, BV04, CFVP03, NB04, OS06, Röb05, WVC05]. nonlogical [CY88]. nonpolar [SBS98]. nonrepudiable [TYH04]. nonuniform [TKT^{+}08]. NOOS [AP96]. norm [DL03]. Normal [KCK04]. Normalized [LS05]. North [vdR87c, vdR87e]. North-Holland [vdR87c, vdR87e]. notation [ÅO06]. note [Ano06]. notebook [MCSS00]. notification [QP08]. Novel [VR05, WBT05, HHSW92, Mam09, PCB99]. NTTs [KY85]. nuclear [Poh87]. Number [Ano86i, vdR86b, GZ04, Tab06, Tan02b]. numbers [TSZP99]. numeric [GI07, Ref87]. Numerical [BN06, DLP06, IOO04, Kni89, PP06, DHD89, DP03a, DMW04, FB93, KM01, LR01, NB04, NHT06, Rum99, SS03, VWCV94]. numerically [Var03]. numerics [TBK06]. Nursing [KZBK99]. NW [GAB^{+}96]. NWF [LS05]. NWF-net [LS05]. Nyström [CP06].

O [Aba06a, BFK02, FC09, Geo02, HMC06, OK02, SM96]. O-FCN [FC09]. O-intensive [CkLC06]. Obituary [Kas86]. Object [CHHW91, KST92, SHJ06, ADM06, AP96, BGV97, BHK90, DC00, Gil85b, HHS98, JBA94, LL03, NSI02, ZJWZ04, Kob92]. object-based [ADM06]. object-centered [AP96]. Object-Oriented [CHHW91, KST92, BGV97, BHK90, LL03, Kob92]. objects [DL00, EGAQ09, MS01, SD02, XSM04, ZZN04, WAE06]. oblivious [KKL09b]. observable [Ven08]. observatory [Fin99, SB99]. observing [WS05]. ocean [CS93, IJLC03]. October [Zna94]. octree [Boa04]. octree-based [Boa04]. ODE [PH94]. ODEs [EL03]. Odyssée [Fau05]. Off [Geo02]. Off-Processor [Geo02]. offs [KNK^{+}08]. OGSA [LKA^{+}08, LMH^{+}09, SPK^{+}07, SSB05]. OGSA-based [LKA^{+}08]. Oh [NMC05]. old [Gol00]. OMIS [WDD00]. on-demand [BPS^{+}03, FMR05, SSF^{+}09]. On-line [BW95, EV98, LVH08, Ueh89]. One [KI89, EFD00]. one-dimensional [EFD00]. online [PN09]. only [DD86]. onto [AEGF^{+}01, Van92]. ontologies [CGM^{+}07, TDF07]. Ontology [SYT09, JC09, KSM^{+}07a, KSM^{+}07b]. Ontology-based [SYT09, JC09]. OntoZilla [JC09]. Open [BJWZ08, AJZ^{+}02, CASW05, DPP03, SP93, CWD^{+}08]. Opening [HICÁFM^{+}06]. OpenMP [NHT06]. OpenRTE [CWD^{+}08]. operating [AHdJF97, BL98, Gos00, MGLV04, RG04, WB90, BG87]. operation [CYB90, MHA09, ONHT89, Tak89b, YWA^{+}89]. operational [CPB00, FCW01, HJP92]. Operations [WMN^{+}01, DL04, Kat04, Lau92, LY90b, NSI02, dLLA93]. operative [Ano84i]. operator [Kos00]. operators [NHT06]. OPIOM [Geo02]. Opportunity [BS91a, BS92]. OPT [FCW01, VAS95]. Optical [DFSZ88, WSTW87, Yat88, CASW05, DvdHdL06, HRJ^{+}06, MWC^{+}03, Mam09, Pal06, Pal09, RvdSB^{+}03, SVC^{+}07, dLRW03, vOHD^{+}05, vdR87e]. Optimal [DLW07, GC94, KLM^{+}03, LJW08, RSRV88, BG05, GS05, TKT^{+}08]. Optimisation [AKPN01, DHS99, SO98]. optimise [RS98]. optimised [BBC^{+}99]. Optimistic [YCAS03]. Optimization [BC03, ACML05, BMRW01, BBL^{+}05, FM01, Leo01, LKG08, MVRM08, PKC^{+}05, Sch03, SKJ01, TSK03, WVC05, dNE05, PW09]. optimizations [CC00, DSS98, SBA^{+}05]. optimize [WCC^{+}09]. optimized [BFLL99, JNR01, SA07, VS90]. optimizer [FN00, LK08]. Optimizing [BCFS02, BKSS02, HWZL08, MSS02, OF07, GMEL08]. optimum [Dal06, AFPG91, AFP^{+}92]. OPTIMUM-AIV [AFPG91, AFP^{+}92]. OptIPlanet [SGP^{+}09, SBdL09].

OptIPortal [DLR^{+}09, DDS^{+}09].

OptIPuter [DLR^{+}09, Mam09, SBdL09, SW06, TWC^{+}06]. OpusJava [Lau01]. OR-forest-based [CY90]. OR-parallelism [Zha93]. Oracle [CW93]. oracles [MMR02]. orbitals [REM04]. Order [Pan95a, AAC04, BFR05, BFC02, Bou95, Dal06, JNR01]. ordered [GY90, Qu04, SAKOK03]. ordering [Kos00, ZWJ04]. organisation [EGAQ09, Gra92]. organism [Low01]. organization [HG92, SYT09]. organizational [KSM^{+}07a, KSM^{+}07b, RSV90]. organizations [RJH^{+}09, VWD^{+}08]. organizing [DD05, FS07, FSM88]. Oriented [CHHW91, BvdV99, BdCYG05, BGV97, BHH91, BHH92, BHK90, CGST09, DC00, FJ00, Gil85b, HWW04, Hua05, JBA94, KST92, Kob92, LL03, LvSW^{+}04, MHA08, OFT09, Tak05, TBK06, VSM02, WYN^{+}90, GMEL08]. Origin [LHB95]. Origin-based [LHB95]. orthogonal [Fio06, PP06]. Orthonormal [DV03]. OSCAR [LSLS05]. oscillatory [DR03]. OSI [Kob92]. out-of-order [BFR05, BFC02]. output [Aba06b]. Outputs [SK97]. overall [WCC^{+}09]. Overcoming [KC98]. overhead [DT94, MTKS00]. overlap [BCFS02]. overlapped [MTNM08]. overlapping [ID98, RŻDM01]. overlay [BTM06, CCL08, JC08, RKSU08, YCX05]. overlays [AN08]. Overview [BP94, DGST09, KY85, KE85, Lan00, Miz89a, vdR93b]. OVM [BFC02].

P [Ano86i, Ano87b, Ano87c, Ano87l,

vdR87e, KKS08, SdR99]. P-CAM [SdR99]. P-GRADE [KKS08]. P2P [DMZ09, FM08, FR08, GPK05, GNWT05, IT05, Kim07a, LLZ07, MM08, SA07, WM07, YCX05]. P2P/Grid [YCX05]. P2P/Grid-based [YCX05]. Pachycondyla [MVS00]. package [BKŁ01, De 98, SLZ95, YA07]. packing [LGMV02]. PaCMAn [ESPP01]. PadicoTM [DPP03]. Page [vdR87b]. pages [Ano86i, Teb86, vdR86b]. paging [BK06]. pairwise [Tis07]. papers [BGL08, Igl07, LBR02]. paradigm [HHSW92, HCL07, KB09a, MvdV01, PA01a, Pri95, STTK03, VR00]. Paradigms [AR07, Gol00, JK92]. Paragon [ABK94]. Parallel [AvLR92, Aba06b, ABF93, AJZ^{+}02, AMB^{+}92, Ama88, Ama89, AKW90a, Ano90g, AR98, BCL88, BMPS01, BVP^{+}87, BJNH05, BHH^{+}93, BHK90, BLB03, CN92, CHHW91, CSP98, DDO^{+}92, DFSZ88, EFD00, ES94, GL05, GS05, GA06, GCK98, IMSV90, IST04, JS89, KGX95, KU01, LOJ^{+}07, LDSH95, Luk00, Mal94, Mal01, Mal02, Mal05, MPG96, OSHH96, PPJ95, RŻDM01, SdR99, SBS98, SG95, SM96, TBD^{+}02, Uch87, VSvD95, Vre88, WGL92, ZEO01, dSL98, Aba06a, AEGF^{+}01, AT01, ADT03, AKW90b, Bal92, BBC^{+}99, BK97, BPS06, BG05, BBJ^{+}06, BGC^{+}03, BFC02, BL92, Bou95, BSG^{+}05, BCW01, BCM^{+}95, CLP95, CST91, CST92, CTF^{+}99, CBK^{+}01, Cas94, CG02, CKFJ06, CEGL01, CGL08, CY90, CC98, CS96, CTMO06, CCHW03, CFG93, DDRR96, DZJ^{+}00, DRS^{+}97, DLW86, DOV01, Dup90, EGCY^{+}06]. parallel [EL98, ESPP01, FSP02, FD95, FN00, GNOY01, GCCC^{+}07, GJS^{+}94, Ger02, GHWZ94, GODM98, Gil94, GC94, HH98, HHSW92, HHS98, HRSW99, HAC92, HHG05, Her91, HKT94, HLSØ06, Höf03, HXL90, HSS00, Hum92, HMC06, ID98, JL95, JSK^{+}06, JBA94, JL03, JM01, JLMR00, KDFL99, KK00, KP00, Kac00, KA09, KZLK06, KSY92, KMK09, Kos00, Kos95, KM01, KTV03, KB92, KL02, LC04, LP01, LR01, LLRS94, LN94, LRMC94, LSS94, MD92a, MvdV01, MP02, MRV92, McC96, MLZ^{+}00, MK95, OVDV98, Oku92, OK02, Omo91, OP95, OP97, Pad92, Par06, PCG^{+}06, PF01, PBHK01, Pri95, RBS93, RBC^{+}88, RN01, RICW00, RS98, RCD03, SKT^{+}08, SPK^{+}07, Sap88, SSKF95, STP^{+}05, SGFS01, Sch03, SBHD08, SMI01, Ser98, SD02, Shi92, Sin92, SSC04, ST98, ST99, THKG98]. parallel [Tan02b, TMT^{+}07, Tic93, TV08, Tis07, TSZP99, UM02, VAS95, VP94, Vre89, VF01, VFS01, VSV95, WKZ^{+}03, WAE06, XHY^{+}90, YCAS03, YdOLS^{+}05, ZT90, ZT91, ZGCM00, ZEO98, dKdOS03, dlTK92, mM95, vdV89a, vdV89b, BG87, Her87, NCCS99, TRFR01, vdR86b]. Parallel-Operating [BG87]. parallel/distributed [CBK^{+}01]. parallelisation [IJLC03]. parallelise [IJLC03, SLZ95]. Parallelising [Kea93]. Parallelism [Par90, Bal91a, Gos00, Hey90, Joh89, Kos95, MBFC99, WRBG94, Zha93, ZS90]. Parallelization [BST^{+}04, BPC^{+}01, BVDF00, LPB04, WYJ99, BCMR01, Fah98, KC98, vM94]. parameter [NHG02, NHG03, SBA^{+}05]. parameter-space [NHG02, NHG03]. Parameterisation [GVD^{+}03]. parameters [PBT02]. PARDIS [Kea99]. PARDISO [SGFS01, SG04]. Paris [Ano84k]. PARLIB [CGSZ95]. PARLOG [Bal91a, DT93]. parsing [BM00]. Parsytec [Cro95, vOB95]. part [PH94]. partial [TBD^{+}02, Xia06]. partially [Ven08]. partially-observable [Ven08]. Particle [KG01, Sin92, ABB^{+}03, BVDF00, CM99, Len01, Low01, MKH06, SLW01]. particle/continuum [Low01]. particles [Dzw97, DB99]. partition [DLW07]. partitioner [DHB02]. Partitioning [SW05, BW97, GODM98, HZC^{+}08, WC01]. Partridge [Ano87b]. party [CLM00, Pol98]. passage [BDHK06]. Passing [DKD08, Ber98, BFR99, Gor02, Kac00, Kal94, LDSH95]. passive [DD05]. passport [GWO03]. past [Fer84]. path [ADOKM06, DvdHGdL09, GdLvOT03, KDE04, Luk89, Mar02, THN^{+}06, Xia06]. path-based [ADOKM06]. path-planning [Mar02]. paths [Alb04, CFG^{+}05, vOHD^{+}05]. patient [CTT^{+}08a]. pattern [CD99, DDL01, MP02, Pet95, WKT00]. patterns [BP02, KL02]. payment [Pol99]. PC [MTKS00, MLSO01]. PC-clusters [MLSO01]. PCs [BL02]. PDE [BRMN04, NHT06, Par06, TBK06]. PDE-based [BRMN04]. PDEs [MR04b, MR03a]. PDG [CLP95]. peer [ADK^{+}09, BS09, FLPP05, HWZL08, JC09, KKL06, LAM07, MTV05, MCT^{+}09, PGSM05, SLŽ^{+}09, ZQZZ09, BNFZ08, CTT^{+}08a, CdCD07, TTP^{+}07, WTK07]. peer-to-peer [ADK^{+}09, BS09, FLPP05, JC09, KKL06, LAM07, PGSM05, SLŽ^{+}09, ZQZZ09, BNFZ08, CTT^{+}08a, CdCD07, TTP^{+}07, WTK07]. PEI [VP94]. perception [vVDBB98]. percolation [BMPS01]. Performance [BMRW01, Ben99, BM92, BP94, BK06, CMT01, CMZ95, Cho04, CCKW88, DZ04, Din99, FM01, Gen95, GBT87, HB98, HCL07, HJCD05, HJK^{+}04, JSK^{+}06, KZC04, Kun94, LOK09, MM03, MLZ^{+}00, Par04, RKSU08, Shi04, SD03, SvAS01, SK06, TDF07, VSV95, Wil00, Yam92, dSL98, Aba09, AB01, AFP07, ADKS06, ACU95, AMW99, AB95, BL98, BBC^{+}99, Ber96, BCS99, Bhu95, Bis94, BS09, BGC^{+}03, CRE01, CSW06, CWD^{+}08, CKFJ06, DCS^{+}07, DRS^{+}97, Doḡ09, DHS99, ET08, FD02, FJ00, FJT01, FdSC07, GNOY01, GGH^{+}06, GGSZ09, GAB^{+}96, HD05, HHSW92, HDC^{+}94, HAE^{+}03, HO02, HKM^{+}06, HWZL08, JAA09, Kal94, KZLK06, KSAOK08, KBM^{+}02, KKK07, KB92, Lau01, LSLS05, Lee04, LGW07, LB09, LLSR02, LM90b, MH01, MKH06, MI01, MOK06, OF07, OP97, OS01, PH99, PSP^{+}09, PH94]. performance [RSV90, Reu03b, RSR01, RS94, RŻDM01, SEH99, STH^{+}98, SGFS01, SB97, Sch03, Ste94, SRCR97, TLTY06, TS08, TBNF09, VSM02, VBP03, WBF08, Wit94, WSH99, WCC^{+}09, YJA03, ZGCM00, dKdOS03, Lid99]. performance-directed [RSR01]. performing [CEJK94]. perimeters [DDR^{+}07]. period [DOV01]. periodic [KY04, TKT^{+}08]. PERMIS [CO03]. Personal [EHMS00, JLU03, Wal94, PSG^{+}06]. personalitie [PNH99]. personalized [LG08]. perspective [Aig86, HDC^{+}94, HPS97, HQ07, Kob92, Mur95, SSK^{+}08, Wal94, Wil86]. Perspectives [LPC^{+}95, Nar86, SRCR97, Baa87, PT05]. pervasive [CD08, HZC^{+}08]. PESYS [DW87]. Peter [Ano86i]. petrochemical [Han89]. PETSc [HNS05]. PGGA [LYQ06]. Ph [vdR87c]. pharmaceutical [BGR^{+}99]. Phase [KI89, CEGL01, Fre94, JL95, Mur86, TC92]. Phenomena [BKS98, BMPS01, DS99, KCT99, SW99]. phenomenon [Szu01]. phenotype [MS03]. Phenylium [DRS04]. photonic [GGH^{+}03, GMM^{+}09, HAE^{+}03, ZLD^{+}03, MWC^{+}03]. phylogenetic [SLO^{+}05b]. physical [SSC04]. physically [SGL99, VLC03]. Physics [LBB^{+}09, ABB^{+}03, MDD89, Par06]. picture [Fuk85]. pictures [SKF^{+}09]. PIE64 [HKT94]. pilot [Mar99b]. Pinatubo [CDRS05]. PIOMT [Aba06b]. PIPE [CBD^{+}05]. pipelined [PHL98, WLF^{+}09]. PKI [GMC03, Var00]. PKIX [JLU03]. PKIX-based [JLU03]. placement [Low05, SO98]. Planck [TMV^{+}07]. planes [Mam09]. planned [KY85]. planning [AMD08, AFPG91, AFP^{+}92, KDE04, Mar02, OS92, RN04]. plant [Hir89, Mae89, Suz89]. plants [Han89]. plasma [BVDF00]. plasmas [GPH^{+}94]. plastic [Ned06]. Platform [GC00, CW93, DDV92, DVVD02, DBA98, Fau05, LWHS07, SBHD08, WTK07, ZJWZ04]. platforms [BYV^{+}09, MI01]. playground [GGW^{+}09]. PLFG [Tan02b]. Plug [MS01, CCDS08]. plug-in [CCDS08]. Plug-ins [MS01]. plus [BS96, CYB90]. PM [WBF08]. PMCommunication [STH^{+}98]. PODOS [VSM02]. Point [WAE06, CFVP03, FGG03, Fio06, Ram95, WTC^{+}02]. polarizabilities [Tor04]. pole [Var03]. policies [DZ04, LHC03]. Policy [Baa87, Aba09, HY09, LL04c, VVC^{+}03, YBQ07]. policy-based [VVC^{+}03]. polling [AAC04]. polygons [LD04]. polymer [JLMR00]. polyradicals [She04]. POMPC [Por95]. POOSS [CHHW91]. POP [NK07]. POP-C [NK07]. POPE [BG87]. population [Gue01]. porosimetry [HRJ^{+}04]. porous [RS99]. port [CCHW03]. portability [CN92, McC96, Reu03b]. portable [BCM^{+}95, GD93b]. portal [GL04a, KKS08, LW08, YLC^{+}06, BAD^{+}05]. portal-based [LW08]. portals [BCMA07]. portfolio [BHRT98]. Porting [FHG95b, CR92, GJS^{+}94]. positive [Amo06]. posture [KLC05]. potentials [DMMP98]. £10.95 [vdR87j]. £14.95 [vdR87i]. £19.95 [Ano87c]. £22.50 [vdR87d]. £25 [vdR87a]. £25.00 [Ano87b]. £29.95 [Ano87l]. £40 [vdR87b]. £7.90 [Ano86i]. Power [Par90, CN92, LC01, Shi04, UWV92, Zad87]. Powerful [CCL08, Pud87]. pp [Zem86, vdR87b, vdR87a, vdR87d, vdR87f, vdR87i, vdR87j, vdR87c, vdR87g, vdR87h]. pp. [vdR87e, vdR87k]. Practical [SPM86, vdV89b, ABdLL05, BvdBM^{+}93, BORM07, BAC02, BLB03, Din99, MLZ^{+}00]. practice [BGR^{+}99, DSSU97, Ger02, PT05]. practices [BHD09]. PRAMs [HSS92]. Prandtl [Tab06]. pre [HCL07, HML07]. pre-scheduling [HCL07]. pre-WS [HML07]. precomputation [LJW08]. preconditioned [VFS01]. preconditioner [VF01]. preconditioners [ID98]. preconditioning [KZC04, WKZ^{+}03]. preconditionings [MvdV01, mM95]. predicate [Qu04]. predicate-ordered [Qu04]. Predict [ZSI08, PGPW09]. Predictability [Mid01, McC96]. predictable [LYQ06]. Prediction [PN09, ACML05, BGC^{+}03, BK06, CSW06, JSK^{+}06, PKC04, TMT^{+}07, VS04, VSV95]. Prediction-based [PN09]. predictions [ET08, ZSI08]. Predictive [FHM^{+}99]. preemptive [XHY^{+}90]. Preface [Fis00, Her90, Miz89b, Par91, Rus90a, Slo05a]. prefetching [CY01, PKC04]. Preliminary [LF95a, KSY92, TC92]. Prentice [vdR87h]. Prentice-Hall [vdR87h]. preparation [dLLA93]. preprocessing [CGM^{+}07]. prescribed [CDS03]. Present [Abe92, Fer84, Bis94]. presentation [Zna94]. preservation [BKM03, HBH09]. preserving [EL03, UM02]. Press [Teb86, Zad87, vdR87i, vdR87j]. pressure [DMN^{+}05]. PREVISE [dLLA93]. Price [Ano86i, Ano87b, Ano87c, Ano87l, vdR87b, vdR87a, vdR87d, vdR87f, vdR87i, vdR87j]. pricing [SVB07, Ven08]. primitives [BDFP05]. Principles [DHS00a, DK00, Gil94, PO00, Pap05, WRBG94]. print [KLH^{+}04]. priori [GL04a]. prioritization [FD95]. Prioritized [WBT^{+}08]. Privacy [Opp00]. private [KTY03, Nag86a]. privilege [CO03]. Probabilistic [MK95, AC92, Ban05]. Problem [Jon00, Pad92, YPF05, BJA^{+}05, BHRT98, CSC^{+}05, CGH04, CD99, CRM05, Del06, DOV01, EPJ^{+}05, KKS08, Kos00, MC00, NMC05, OVDV98, Prz03, SMK05, TRFR01, Tre03, VDPHS09, WH05, WAE06, ZS05b, ZAP05, tTvH96, vdV89a, SMI01, VAS95]. Problem-Solving [YPF05, Pad92, CRM05, WH05]. Problems [Nit86, Van87a, BCMR01, BS04, BV04, CTMO06, Dat03, DHD89, DR03, ED04, Hul89, JNR01, MR04a, MKK03, NB04, Ned06, PKC^{+}05, RGH^{+}01, SKT^{+}08, Sch03, Tab06, THKG98, VFS01, ZEO98, ZEO01]. procedural [BMFC07]. procedure [FCW01]. procedures [dLLA93]. Proceedings [Rho89]. process [BBL^{+}05, CLP95, Cur92, DVV90, FJ00, HAAH05, Niw89, SGY^{+}07, ZMS^{+}06]. process-level [CLP95]. processed [De 88]. processes [BMPS01, BDNN02, CM01, EL98, Gue01, HW95, Mis92, SKT02]. Processing [FMD99, Ano86j, CN92, Chv87, CEJK94, DMR93, DF97, EFD00, GP09, Her87, Kob92, KE85, LGH97, LH07, LK08, LM90a, NMZC06, ONHT89, PK99, Ref87, Sap88, SL87, SD03, Tic93, ZEO98, vdR86b]. Processor [VV92, XHY^{+}90, AG92, Dui89, Goo01, Goo02, JCSS01, LC03, LMH^{+}09, Ser98, DFSZ88, Geo02]. processors [BFR05, DR05, FD95, SD06, SCK^{+}00]. ProCoS [BHH91, BHH92]. Procrustes [Tre03]. product [HBCR01, Tho06]. production [GML99, KH89]. productivity [LJ07]. products [Pin87]. profile [FC05, ZL04a]. profile-based [ZL04a]. profile-driven [FC05]. Program [BS96, BP94, Ano86m, Baa87, Bun03, CC00, Fah98, FHG95b, Höf03, KDFL99, Kom89b, KV03, Leo98, Li90, LDSH95, MRV92, RS94, SR03, Ser98, SLO^{+}05b, SBLT05, Yam92]. Programmable [Tof99]. Programme [HT02]. Programmer [Kea99]. Programmer-level [Kea99]. programmers [Teb86, vdR87i]. Programming [BCL88, BR92, BP94, CSP98, JBA94, KGX95, KBM^{+}02, KB92, LSF^{+}94, NK07, PBM95, ST99, ACGdT02, ADT03, Ano84k, Ano85g, Bal92, BMFC07, BSG^{+}05, BLB03, CLP95, CDF^{+}05, CPK05, DDRR96, DHS00b, DLW86, GHWZ94, HLSØ06, IMSV90, JF05, Kom89b, LP01, Luk89, MD92a, McC96, MS01, MK95, OBK88, Oku92, SSKF95, SHJ06, WDD00, WB90, vdR87h, vdR87j]. Programs [ABF93, DDO^{+}92, BP01, CLP95, CRE01, CY90, De 88, FSP02, FJ00, JCSS01, Kac00, KTV03, LTOT07, LRMC94, Mal94, Reu03b, RCD03, VP94, ZT90, ZT91, Zha93, ZS90]. Progress [HV84, DQ97, Hul89]. Progressive [CCM98, CCLS09, VS04]. ProHPC [BBD^{+}99]. Project [Zin00, Ais88, Coo94, FS93, FK99, Fur92, Jon00, Kas85, Lin84, NTN86, VBP03, vdR87e, vdR93b, BVP^{+}87, KBM^{+}02, MFP05, Sti93]. projection [BV04]. Projections [HK88, KZLK06]. projects [Mar86, NSF87]. Prolog [BG87, vdR87g, Dup90, AR98, CY88, CYB90, GC94, HSS00, LY90a, LY90b, Lop93, Teb86, vdR87i, vdR87j]. Promise [TS99, FS97, Ful91]. Promoting [FdSC07]. proof [MD92b]. propagation [GKIZ05, GCK98, TM05]. propelling [Suz89]. properties [Ban05, IS06, LS05, NCS04]. proportional [LL04c, NP06]. proportional-share [NP06]. prospect [Kaw92]. Prospects [TS99, Fer84]. Protecting [EHMS00]. Protection [HLvL^{+}97, Opp00]. protein [BORM07, CCG07, DCK03, TMT^{+}07]. proteins [BORM07, FVFA98]. protocol [AMH02, AMH04, AS02, ACE02, BDFP05, GL04b, KKL09b, Reu03a, BBM^{+}03, OVK^{+}09]. Protocols [BCH^{+}08, GSD95, BAC02, DSS98, SMC99]. prototype [GRPL04, KSY92, ZY90, ZCW^{+}04]. provably [GPA96]. provenance [WLF^{+}09]. provers [Sti93]. provide [MK04]. providers [MHA09]. providing [KV09]. proving [Sti93]. provision [KSM^{+}07a, KSM^{+}07b, YKL^{+}07]. provisioning [MKT09, PN09, SLJ^{+}06]. proxy [ADM06, TYH04]. PSE [Hua05]. PSEE [LSS94]. PSEs [VR05]. pseudo [Tan02b]. pseudo-random [Tan02b]. Pseudorandom [GZ04]. pseudospectra [BKG05]. public [MCT^{+}09, Wit94]. Publications [vdR86b]. publish [HMP04]. Published [Ano86i, Zad87, vdR86b]. Publisher [Ano87b, Ano87c, Ano87l, Ano89f, Ano06]. Publishers [vdR87f]. publishing [PCM99]. Puerto [SR03]. PUL [BCM^{+}95]. purpose [AKW90b]. PVM [Ber96, CS96, PL96]. PYR [CDRS05]. pyramid [KY04]. pyramid-based [KY04]. pyroclastic [CDRS05]. Python [BCG05, HLSØ06].

QoS [BCPS03, DMZ09, FMS08a, GdLvOT03, HB00, MKT09, MM08, Pal06, WL05]. QoS-based [FMS08a]. quadratic [EL03, TRFR01]. quality [HJCD05, NZQ07, SMC99, TSZP99, KLM^{+}03, KTM^{+}08]. Quanta [HAE^{+}03]. quantiles [BDHK06]. quantization [LFWV05]. Quantum [Höf03, LPB04, LDSH95, SBLT05]. quantum-chemical [SBLT05]. QuarkNet [BGJ^{+}06]. QuarkNet/Grid [BGJ^{+}06]. quasi [Szu98, BMZ01]. Quasi-Newton [BMZ01]. quasi-random [Szu98]. QUATRAIN [DMR93]. queries [SHLB08]. Query [SLŽ^{+}09, LK08, LMH^{+}09, SD03]. Query-driven [SLŽ^{+}09]. querying [CGST09]. questions [WBMP99]. queue [Tho06]. queueing [Lee04]. quicksort [PHL98]. quo [Kaw92].

R [Ano86i, Ano87b, Ano87c, Ano87l, vdR86b, vdR87j, vdR87k, WSS^{+}09]. R&D [Ano84i, Nag86a, Smi86]. R-based [WSS^{+}09]. race [KKP00, PM00]. radial [NK05]. radio [HJPS03, SLJ^{+}06]. radiological [BDNP92]. radiology [PK99]. radiosity [Sch98]. RAISE [BHH91, BHH92]. random [BTM06, CC98, Szu98, Tan02b, TSZP99]. range [BS09, OS06, SHLB08]. rank [TM05]. Rarefied [CH95]. rate [AS02, HJPS03, SCP09]. rationale [dLRW03]. raw [DMMP98]. ray [NMZC06]. RCT [SHLB08]. RDF [SY04, vdHDT^{+}06]. re [CCG07]. re-structuring [CCG07]. reaction [BMPS01, Ban02b]. reaction-diffusion [BMPS01]. Reactions [Meu05]. reactive [LPB04]. reactivity [DRS04]. Readers [Ano05g, Slo06a, Slo06b]. Reading [vdR87g]. ready [BFK02]. Real [KLC05, MKT09, SKF^{+}09, VLK09, XYZ05, BRR^{+}04, DT93, Doḡ09, FSP02, HMP04, KA09, LF01, LLWN04, Pal09, Pfi99, PN09, SS04, SST^{+}06, Sin92, Suz89, dlTK92]. Real-time [KLC05, MKT09, VLK09, XYZ05, BRR^{+}04, DT93, Doḡ09, HMP04, LF01, Pal09, Pfi99, PN09, SS04, SST^{+}06, Sin92]. real-world [FSP02]. realisation [DGS09]. Reality [CHK98, Kaa99, Ami90, BLRS98, BYV^{+}09, CCBR98, CN98, DDS^{+}09, DBA98, Kaa98, PSG^{+}06, SUD^{+}98, vVDBB98]. realization [Kom89b, KM01]. realizations [DL03]. reasoning [ATT96, FB93, Ser95, VM93]. receive [Gor02]. recognition [KLH^{+}04, KKP^{+}05, MPH00, Pet95]. recombination [SBS98]. recommendation [KKYK04]. recommendations [WBMP99]. recomputation [HSS00]. reconfigurable [AAB^{+}92, JL98, NWE04, Pan95a, PHL98]. reconfiguration [CCLS09, MKT09, PA01a]. reconstruction [Alb04, IEG04, RICW00, XSM04, ZS05b]. record [RCD03]. record/replay [RCD03]. recorded [HNP05]. records [CTT^{+}08a]. recovery [PY00, PWY03]. rectifying [HYS04]. recurrences [Van92]. recursive [HML09, IST04, SSC04, VS90]. recursively [VS88, YYW^{+}09]. Recursively-Adjusting [YYW^{+}09]. reduce [BORM07]. reduced [BFLL99, TM05]. reduces [PCB99]. Reducing [FC05, de 94]. Reduction [Ama88, Ama89, BVP^{+}87, BHH^{+}93, RSRV88, Vre88, BK97, BHK90, Fuk85, Höf03, KGW95, RSV90, SS03, VS04, Vre89]. Redundant [dRSS97, TBD^{+}02]. reference [AKB^{+}01]. refinement [HV03]. refinements [CPB00]. Reflection [Lop96, TA96, AP96, BS96, tTvH96]. reflective [EV96, Pit96, SGdMM96]. regime [FVFA98]. region [RV95, SMC99]. region-of-interest [SMC99]. register [GD05, MSS02]. registration [CEGL01, MLSO01]. registry [BGK^{+}05]. regular [Dör05, NHT06]. Reiblein [vdR87j]. Reidel [vdR87k]. reinforcement [Ven08, Ven09]. related [BSG^{+}05, EV98]. relational [Mur88, NSI84, ZY90]. relations [vdR87h]. relative [KCK04]. relativistic [BST^{+}04]. relaxation [KKHS01, LN94]. Relaxed [GSD95]. release [PY00]. relevant [Meu05]. Reliability [RdLM06]. reliable [Var03]. Remote [HSH^{+}07, SK97, AMW99, BCFS02, Han03, KPB^{+}03, LHL03, RvdSB^{+}03, VS04, ZMN99]. remote-sensing [AMW99]. rendering [BRR^{+}04, HH98, Igl04a, KPB^{+}03, Pfi99, XYZ05]. Reorganization [FMS08b]. replacement [WG00]. replay [RCD03]. replica [CC07, KLSS05]. Replicated [HHS98, SHJ06, Bal91a, DAM08, TV08]. replication [CCL07, CCG07, Doḡ09, LVH08, LSTV07, MC04, MHA09, TLYT05, TLTY06, ZWJ04]. replicator [WKT00]. report [CN98, DQ97, FK99, Van87b, vdR87l]. repositories [EGAQ09, HESM99]. repository [FAJP99]. representation [AP96, Qu04]. representations [LY90a, NMC05]. required [PCB99]. requirements [Ber96, DDRR96, FC05, Mur95, PB05]. requires [DHS99]. Research [DGS09, Kaa99, PSP^{+}09, Van87b, BY93, BAD^{+}05, BS84, Fur92, Gal87, LPC^{+}95, LGS^{+}07, NSF87, Sch94, Sti93, Tic93, ZL04a, vdR87l, Kaa98]. reservation [CFG^{+}05]. reservations [ET08, NF07]. reservoir [PKC^{+}05]. resilient [TDV^{+}08]. resolution [HJPS03, KPB^{+}03, KLH^{+}04, PCB99, RJH^{+}09, SMC99, SLDK03]. resonators [CASW05]. Resource [CKKG99, EGK^{+}07, EW97, HQ07, Slo96, VDPHS09, Zhu04, ABG02, AFF^{+}09, BBM^{+}03, CCL08, CCLS09, CT09, CC09, DS08, ET08, Fer96, FHM^{+}99, GPK05, HZC^{+}08, HB08, HY09, HB00, HPLL08, HML07, HML09, KA08, KFC^{+}07, LL04c, MTV05, MV09, PGPW09, PN09, TTP^{+}07, Vau93, Ven09, Wal86, WSH99, WCC^{+}09, YBQ07, ZYXL05]. resource-aware [HY09]. resource-constrained [Ven09]. resource-management [Fer96]. resource/service [MV09]. Resourceomic [CCM07]. resources [ACC^{+}05a, BJWZ08, DFC^{+}08, GGM^{+}09, Kos95, NZQ07, PSA^{+}09, SVB07]. respect [HW95]. Response [BCS99]. responses [PKSC02]. Responsive [Wes99]. restart [Dal06]. Restarted [BB04]. Results [Ano87m, KI89, Ano87j, FHG95a, LF95a]. retail [GPA00]. Retargetable [DR05]. retrieval [CC07, FC09, JC08, KY04, MS03, OCdAM07, SYT09, SLŽ^{+}09, SNA92]. retrieving [SZC05]. retry [CK00]. return [Kuo86]. reusable [CN92]. reuse [TV08]. reverse [FC05, HNP05]. reverse-mode [FC05, HNP05]. reversible [JL03]. review [Gra92, Zha93, Fuk85]. Reviewer [Ano86i, Ano87b, Ano87c, Ano87l]. Reviewers [Ano07, Ano08]. revolution [Sin84, Bur02]. RF [DD05]. rheology [Ned06]. rich [RJH^{+}09]. Rico [SR03]. Riet [Ano86i, Ano87b, Ano87c, Ano87l]. rigorous [LLZ07]. RISC [BBSV92]. risk [TMTY05]. Ritz [BB04]. RMI [AG05]. road [BBC^{+}99, BS91b, SNW01]. road-vehicle [BBC^{+}99]. Robert [Ano84k]. Robot [Shi85, Mar02]. robots [Poh87]. robust [BRMN04, CDS03, Var03]. rocks [BBJ^{+}06]. role [CO03, Wie03, Wit94]. rolling [FN00]. room [Ram95]. root [TM05]. ROST [HSH^{+}07]. Roth [Zem86]. rounding [SS03]. route [RN04]. routed [KD00, LOK09, SVC^{+}07]. routine [OCdAM07]. routines [BFR05]. Routing [Dör05, GLA88, CST92, DOV01, KLM^{+}03, KSAOK03, KSAOK08, LHB95, LS01, OVDV98, SAKOK03, SDD^{+}09, Xia06]. RP3 [CR92]. RSA [KKL09b]. Rule [HBH09, LJPS05, Yos89]. Rule-based [HBH09, LJPS05, Yos89]. rules [AW97, KA08, SA97]. Run [CWD^{+}08].

Run-Time [CWD^{+}08]. Runge [CP06]. running [CRE01, DW87, ZSI08]. runtime [Höf03, TSK03, VR05]. runtime-interference [Höf03]. runtimes [DPP03, KTV03].

S [Ano87c, Teb86, vdR87e, vdR87i, vdR87j, Ano99, LJY04]. Sabotage [DSS07, Sar02]. Sabotage-tolerance [DSS07, Sar02]. safety [HDC^{+}94]. Salesman [SMI01, VAS95]. sampling [TKT^{+}08]. Satellite [BT93]. satellites [TC92]. SATEXPERT [CSC^{+}92]. Scalability [BNFZ08, LLRS94, McC96, SMC99]. Scalable [AMH02, JL08, LXLS09, SBG^{+}09, TDV^{+}08, AG92, AR07, BW95, DLR^{+}09, DVVD02, GW01, KK97, LC05, LBYL08, MCT^{+}09, Pal06, PHL98, Pro07, RL98, SCK^{+}00, SLŽ^{+}09, VS88, ZJWZ04, FMD99]. scalar [SW99]. Scale [AKW90a, CHK98, BBSV92, BBJ^{+}06, BDL06, BCG05, BCH^{+}08, CDF^{+}05, CR92, CTT^{+}08b, CTMO06, Dat03, DAM08, Din03, FAJP99, GLA88, GPH^{+}94, HKM^{+}06, JL03, Joh02, Kos95, KTV03, LPK94, MTV05, MKH06, MR04a, NSP07, RN04, RL98, WVC05, WKZ^{+}03, YHJC05, YCAS03, YCX05]. Scaling [KZLK06, Var00]. scan [Alb04]. scanning [HYS04]. scattering [LPB04]. scenario [HND06]. scene [SOR05]. scheduler [BCB^{+}07, CFGC03, THN^{+}06]. schedulers [CCDS08, EL98]. Scheduling [LRMC94, RSRV88, Ser98, SD06, Aba06a, Aba09, AMD08, AFPG91, AFP^{+}92, CGT07, CCL07, CCL09, CK00, DZ98, DRNMC09, EV96, EL98, ED04, FD95, FN00, GRH05, GP09, HV92, HCL07, HXL90, HPLL09, HML07, JCSS01, JM01, KA09, KMK09, KB09a, LTOT07, LHL09, LKG07, LYQ06, LB09, NP06, NP03, Nos98, Pal09, Par04, PW09, RNJK09, SZ98, SKJ01, Shi04, SCP09, SDD^{+}09, TLTY06, Ven09, WL05, WHP09, XHY^{+}90]. scheme [AAC04, CG02, CHJ^{+}04, CYH04, CGL08, HWW04, LBYL08, LHC03, LHL03, MK04, NB04, PY00, PWY03, PSL^{+}04, TYH04, YYW^{+}09]. schemes [ÅO06, LLKF09, PGSM05, Shi04]. School [HY03]. schooling [Sch01]. schools [MR03b]. Schrödinger [BFLL99, IS03]. SCI [STTK03]. SCI-based [STTK03]. Science [ABMS05, Bis96, CF09, DGST09, HT02, NSF87, NTN86, Ros94, TS99, ZBB09, ABM^{+}07, BHD09, BLB03, Bun03, CH04, Coo86, FGG03, GRPL04, Hul89, JHL^{+}06, Joh89, Joh02, Kol89, Lit03, MR03b, Rho89, Ros89, SBG^{+}09, SL97, SSMG95, Tan02a, Wie03, Wil89, WAD^{+}89]. Sciences [AKMK05, SR03, Han03, Hul89, LWHC07, TWC^{+}06]. Scientific [HY03, MBZL09, AW03, AKB^{+}01, BBBD01, Bea03, Ben99, CSW06, Dal03, DRS^{+}97, Din99, DT08, Fin99, GHWZ94, GP09, GB99, JNPY06, KB09b, MCT^{+}09, PNH99, PSG^{+}06, RL98, SSC09, WLF^{+}09, WSS^{+}09, vdV89b]. scope [Par87]. SCORM [SYT09]. SCORM-compliant [SYT09]. scripting [Bea03]. SDEM [Mat89]. SDI [Poh87]. SE [BT93]. SE-TC2 [BT93]. SEAI [vdR86b]. Seamless [TDG^{+}06, DFG^{+}00]. search [Bal91a, BNFZ08, BS09, CC98, Dzw97, JC08, LAM07, MVS00, NSI84, RNJK09, ZZ90]. searches [SA07, SF06]. Searching [CFG93, Sch03, SZC05]. Second [Ste85, Bou95, WYN^{+}90]. secrecy [BDNN02]. Secret [EHMS00]. Section [CF09, GR09, GMS09, KT08, LXLS09, Aba06b, ABM^{+}07, AR07, ABMS05, BGL08, BN06, BLAV06, Büc05b, DLP06, DKD08, DDM^{+}08, DT08, FM08, GA06, GMA07, Hab05, HKM^{+}06, Igl07, JL08, Kim07b, hKcF09, MOK06, Ole07, PH07, PX07, SDBdL06, SBdL09, SPEW09, SD07, VLK09, ZBB09, Zhu07, ZAP05]. sector [Nag86a]. Secure [ABCD00, KKL09b, SSF^{+}09, BDFP05, Din03, FLPP05, GPA00, HLvL^{+}97, JFDF09, LL04b, Sin07, ZDR07, CKK^{+}04]. secured [KBB^{+}09]. Securing [PPSS06, Pol99].

Security [PX07, CDF^{+}05, CKK^{+}04, CWW^{+}99, CLM00, DMG^{+}08, DDR^{+}07, GMC03, KLH^{+}04, KKL09a, hKcF09, LS07a, Lan00, MPPM09, NZQ07, PM00, SM03]. Segmentation [DMM^{+}99]. segments [TV08]. seismic [AMB^{+}92]. Selected [BGL08, DMW04, Igl07]. selection [ATT96, CC07, Dua94, DMN^{+}05, ET08, HWZL08, KA08, KH97, KKL06]. selective [HSS00]. Self [DD05, FSM88, DRNMC09, FS07, GL04b, KMN^{+}05, XHY^{+}90, PIKM02]. self-gravitating [KMN^{+}05]. self-modifying [GL04b]. Self-organizing [DD05, FSM88, FS07]. self-scheduling [DRNMC09, XHY^{+}90]. Self-Timed [PIKM02]. Semantic [HB08, ZL04a, BX04, FC09, HK88, Hal88, JC08, VGBLGS^{+}06, LWHC07, LvSW^{+}04, WC06b, Zhu07]. Semantic-supported [HB08]. semantically [BCMA07]. Semantics [JD94, Zhu04, Ber00, BL92, CPB00, Hal88, HJP92]. semi [AC01, BDHK06, JC09, NB04]. semi-automatic [AC01]. semi-discrete [NB04]. semi-Markov [BDHK06]. semi-structured [JC09]. semiconductor [FHG95b, SGFS01]. SemreX [JC08]. send [Gor02]. sensing [AMW99]. Sensitivity [PBHK01, PBT02]. sensor [LBYL08]. sentence [Nit86]. separation [FdSC07]. September [Ano84k]. sequence [BORM07, CS05, DMR93, DCS^{+}07, YD05]. sequences [Leo01, MP02, TKT^{+}08]. Sequential [CKFJ06, JBA94, LYMZ09]. SEQUIN [BCPS03]. serial [RS94, SGFS01]. Series [Ano86i, GKIZ05, LFWV05]. server [BGR^{+}99, CWD04, CGL08, Cho04, GW01, HAC92, HJCD05, Lee04, LHL03, PA01b]. serverless [SWCP03]. servers [LPE08, SMA08, WWC^{+}97]. Service [OF07, OVK^{+}09, ADF^{+}05, AN08, BJA^{+}05, BSCC06, BPC^{+}01, BGK^{+}05, CKFJ06, CGST09, DCS^{+}07, DLW07, Din03, DDR^{+}07, GPK05, GMEL08, HJS^{+}99, HSH^{+}07, Hua05, LL03, LvSW^{+}04, LKA^{+}08, LJW08, LS08, LMH^{+}09, MHA08, MV09, Pal06, PKC04, SVC^{+}07, SGY^{+}07, SPBT07, THN^{+}06, TBNF09, WWC^{+}97, WSH99, YKL^{+}07, ZZ09, GMEL08, KLM^{+}03, KTM^{+}08, YPF05]. Service-based [OVK^{+}09, HJS^{+}99, LMH^{+}09, TBNF09]. service-oriented [CGST09, Hua05, LvSW^{+}04, MHA08, GMEL08]. Services [AFP07, BMFC07, BBW08, CdCD07, CTT07, CLM00, FLPP05, GPA00, GMC03, GGH^{+}03, HML07, JRF^{+}07, JC09, KF00, LG08, MFP05, MTV05, MS01, MC04, MHA09, MPPM09, Opp00, PK99, PH99, Pol98, Pol99, PPSS06, PEG05, PBB^{+}05, QP08, RG04, SLJ^{+}06, SZGbC04, TS08, VGBLGS^{+}06, WCC^{+}09, YCX05, ZJWZ04, AKMK05]. Session [Ano86h, OVK^{+}09]. Set [Zna94, CGH04, DL04, RNJK09, Xia06]. set-splitting [CGH04]. set-up [RNJK09, Xia06]. sets [BCW01, DB99, HAB^{+}06, LGH97, Sin07, XYZ05, ZLD^{+}03, vdR87f]. setting [WM07]. SGrid [LvSW^{+}04]. Sgurev [vdR87c]. shadows [DS04a, Sch98]. shape [BBL^{+}05, LKG08]. share [NP06]. Shared [KGX95, PBM95, BK97, BBSV92, Ber98, CR92, CYH04, CCLS09, DFC^{+}08, FP03, GB99, HH98, HHS98, LWSC07, OB04, PADD03, PY00, PH94, RCD03, SF06, TC06, TYH04, WYJ99]. shared-image [FP03]. shared-memory [BK97, CR92, SF06]. Sharing [DL00, BBM^{+}03, CTT^{+}08a, CWD04, DZ04, DGS09, ELvD^{+}96, MM08, SBG^{+}09, SCY01]. Shedding [SHB89]. shelf [RV95]. shell [KTY03, Tab06]. Shells [TA96]. ship [Suz89]. shock [Kni89]. shop [DDV92, ED04, RNJK09]. shop-floor [DDV92]. shortest [Luk89]. Should [Poh87]. shrinking [UM02]. shuttles [BRS04]. side [HSS00]. side-effects [HSS00]. Sigma [vdR87j, ANN^{+}92]. SIGMCC [CTT^{+}08a]. signaling [SJTG07]. signals [Fuk85]. signature [Ale97, CYH04, HWW04, TYH04]. significance [DFG^{+}00]. signing [HWW04]. silicon [MPH00]. SIMBEX [GL04a]. SIMD [BL92, CEJK94]. similarity [BNFZ08, NSI84]. Simple [SWCL95, Sch01, ABF^{+}03, HHSW92, TDL05, VP94]. Simpler [BB04]. simplified [SUD^{+}98]. simplify [HLSØ06]. simulate [SGL99]. Simulated [VSvD95]. Simulating [DS04b, DIK^{+}06, JLMR00, DS99, KZC04, Mor01]. Simulation [BAD^{+}05, HW95, LSS94, MVRM08, RSV90, TC06, Wor99, AJZ^{+}02, AGJN00, Ano96b, BMPS01, Ban02b, Ban05, BJNH05, BKK02, BDL06, BPS^{+}03, BVDF00, Bru01, CYLT05, CGT07, CTT^{+}08b, CTMO06, FJT01, FW02, FN00, FCW01, GL04a, GBE00, Gra01, GCK98, HDC^{+}94, HQ07, IOO04, JL03, Kni89, KMN^{+}05, LR01, LF01, Len01, LB09, Mid01, MSX00, Mun04, NJW^{+}06, NCS04, PO00, Par06, PPJ95, PSS01, RdSH^{+}00, RN01, RV95, RS99, RGH^{+}01, SGFS01, Sch85, SNW01, SBS98, She00, Sin92, SKT02, SUD^{+}98, SLW01, SdSP04, UTT00, VN01, YCAS03, ZMS^{+}06, vOB95]. Simulation-based [MVRM08]. Simulations [SdR99, ABL04, AKH^{+}04, BRMN04, CTF^{+}99, HRJ^{+}04, JNPY06, KKHS01, LPK94, NEJP94, PF01, PBB^{+}05, RN04, SMK05, SCK^{+}00, VWCV94, VSV95, WKZ^{+}03, Kol89]. simulator [BCS99, LJY04, SM01a, WWSM98, YA02]. Sinclair [Sin84]. single [AFF^{+}09, DL00, MGLV04, SF06, dlTK92]. single-chip [SF06]. singlet [Tul04]. Singular [BH03]. Sink [SO98]. Situated [BMS05]. Situation [Hal88, NSHP88, WBT^{+}08]. Sixth [Ano84e]. size [BVDF00, HV92, Tor04, Van92]. SKaMPI [Reu03b]. skeletons [BG05]. SLA [BCPS03, DMZ09, HY09]. SLA-constrained [HY09]. slave [HCL07]. small [LAM07, Van92]. small-world [LAM07]. Smart [Mar90, AHdJF97, Ale97, BGV97, CkLC06, DQ97, DF97]. SMiLE [STTK03]. Smith [vdR87e]. smooth [ZY04]. SMP [BSG^{+}05, CRE01, MTKS00]. SN1000 [CGSZ95]. SNIPE [FMD99]. snoop [Yam92]. snoop-cached [Yam92]. snooping [DSS98]. SOAs [KB09b]. social [DGS09]. society [Ano87c, Mar98b, Mar98a, Mar99b]. Soft [Hab05, DS04a]. Soft-computing [Hab05]. Software [ÅO06, BHH91, Cas94, Kow84, Kow85, Mai91, Mat89, Sch94, SK97, TD95, AAB^{+}92, AMB03, Ano87b, Bea03, BBL^{+}05, BHH92, BKŁ01, CGSZ95, Cur92, De 98, GJS^{+}94, GCK98, HZC^{+}08, HLvL^{+}97, Hum92, IMSV90, Jon00, LWSC07, LDS06, MB01, Poh87, STTK03, SLZ95, Tak05, TBK06, ZMN99]. soil [LNJ04, ST98]. solar [SZC05]. solid [GNOY01, WKZ^{+}03]. solution [CGH04, KA09, SS03]. solutions [ZEO98, ZEO01]. solvability [Ned06]. solve [Del06, ED04]. solver [BG05, ID98, MMVV08, PH94, SGFS01, SSC04, vM94, Jon00]. solvers [BJNH05, BRMN04, BGC^{+}03, NHT06, dSL98]. Solving [HPP94, KKS08, SG04, YPF05, vdV89a, BJA^{+}05, CSC^{+}05, CRM05, DHD89, EPJ^{+}05, NMC05, Pad92, PKC^{+}05, Sch03, SMK05, WH05, ZS05b, ZAP05, tTvH96]. Some [Luk89, MDD89, DS99, DR03, Mid01, Kat04]. SOME-Bus [Kat04]. Sons [Ano87c]. Sophia [Zna94]. Sophia-Antipolis [Zna94]. sorting [CG02, CKFJ06, Kat04]. source [BP01, NB04, PBT02]. source-code [BP01]. Sowa [vdR87g]. space [All92, BORM07, BW97, DZ04, DFC^{+}08, DL00, Fuc93, JL95, Lau92, NHG02, NHG03, SW05, SCK^{+}00, SGH^{+}08, SSMG95, TMV^{+}07, VLC03, ZYXL05, dLLA93]. space-filling [SW05]. space-shared [DFC^{+}08]. space-sharing [DZ04]. spacecraft [AFPG91, AFP^{+}92, CSC^{+}92]. spaces [BMS05]. spanning [Luk89]. SPARC [AG92]. sparse [BJNH05, BMZ01, HBCR01, IST04, NHT06, SKT^{+}08, SGFS01, SG04, dSL98]. sparse-matrix-vector [HBCR01]. spatial [Ban05, Coo90, FX07]. Spatio [HYC04]. Spatio-temporal [HYC04]. Special [Aba06b, ABM^{+}07, AR07, ABMS05, BGL08, BN06, BLAV06, Büc05b, CF09, DLP06, DKD08, DDM^{+}08, DT08, FM08, GR09, GA06, GMA07, GMS09, Hab05, HKM^{+}06, Igl07, JL08, KT08, Kim07b, hKcF09, LXLS09, Mes02, MOK06, Ole07, PH07, PX07, SDBdL06, SBdL09, SPEW09, SD07, VLK09, ZBB09, Zhu07, ZAP05]. specific [BFK02, DR05, PNH99, WBF08]. Specification [Pro07, ATJMZ02, GGW^{+}09, ZZ09]. Specification-correct [Pro07]. specifications [ABK94, BDFP05]. specimens [FGCM07]. spectra [CGM^{+}07]. Spectral [BvdHN^{+}01, Mal94, MLZ^{+}00]. spectroscopy [GPH^{+}94]. spectrum [GPH^{+}94]. speed [AB03, HG92, HAB^{+}06, Kas85, LC01, MFP05, SRG^{+}03]. SPEEDUP [AB03, ABMS05]. Sphere [VSvD94]. spherical [Tab06]. SPiDER [FSP02]. spills [MSS02]. spin [Coo90]. SPINEware [BvdV99]. spiral [HHS98]. spline [IEG04]. Split [PNH99]. splitting [CGH04, Čie04]. sponsored [NSF87]. SPP [SSMG95]. SPP-1000 [SSMG95]. SQL [CKFJ06]. square [TM05]. square-root [TM05]. squares [vdV89a]. St [vdR87i]. Stability [CP06]. Stable [CHHW91]. staging [MTNM08]. stalemate [Chv87]. standard [ZMN99]. standardization [Kob92]. standards [MLD08, RS94]. star [KSAOK08]. StarCAVE [DDS^{+}09]. StarPlane [GMM^{+}09]. STARS [vV85]. State [CN98, BDHK06, Van87b]. statement [Leo01]. statements [Leo98]. states [REM04, Tul04, Wii84]. Static [HPLL09, GL05, OS06]. statistical [Ald89, HPS97, MCA02, ZZ90]. statistics [Pan95a, vdR87k]. Status [MK88, Abe92, FK99, Fur92, Kaw92, Miz89a, Nag86b]. STEER [LWHC07]. steered [PHM^{+}99]. steering [LJPS05, LBB^{+}09, MvLvW98, MvWvL99, VMvW97]. Stellar [LM90b]. stem [BPP^{+}07]. stented [HORC04]. step [CP06, Kos95]. stigmergy [DBT00]. stimulating [JFDF09]. stochastic [CKR04, Gue01, HWS07, BP02]. stock [Yam89]. Stokes [ID98, vM94]. storage [ADM06, DLR^{+}09, GGSZ09, KK97, NJW^{+}06, NZQ07, OB04, WBT^{+}08, YCAS03, ZJWZ04, CHHW91]. Strategic [vdR87d, BM08]. Strategies [PR95, DFC^{+}08, GJS^{+}94, LPB04, LB09, MM03]. strategy [Car86, DMZ09, FN00, HCL07, HWZL08, HPLL09, KKL06, LVH08, LYMZ09, LS01, Qin07, RM97, SK05]. stream [GS95, LM90a]. streaming [FMR05, SST^{+}06, SKF^{+}09, SLDK03]. StReD [NZQ07]. stress [BBJ^{+}06, OS06]. string [MM03]. stripes [vV85]. Structural [HJP92, Lop03]. Structure [BKM03, KKL09a, ANN^{+}92, BM00, CCG07, Coo90, DLW86, Lin84, LYT^{+}05, Nit86, TMT^{+}07, ZQZZ09, ABK94]. Structured [HB09, ADT03, BFK02, DLP06, JC09, LLZ07, PGSM05, SS03]. structures [BNFZ08, BL92, KN06, Yam92]. Structuring [Leo01, LPE08, CCG07]. Stuart [Zad87]. studies [BBSV92, Cur92, Fuk85, GGH^{+}03, MMVV08, MDD89, NHG02, NHG03, PKC^{+}05, RSV90]. Study [CY88, KI89, Bal92, BBM^{+}03, DRS04, Doḡ09, EGCY^{+}06, HORC04, KDE04, KE85, Rou00, SMI01, TC06, WX02, WM07, ZM97, vVDBB98]. studying [SH99]. Style [BCL88, Bal91a]. styles [MS01]. Subgraphs [WLB00]. Subject [Ano92h, Ano92i, Ano93i, Ano94g, Ano95h, Ano98c, Ano01d, Ano02c, Ano03i, Ano04b, Ano05f]. subroutines [DT94]. subset [OP95]. subset-equational [OP95]. subspace [CHJ^{+}04, Dat03]. subspaces [HV03]. Substation [ONHT89]. substitutable [SVB07]. subsystem [WTC^{+}02]. Success [vdR93a]. suggest [MVS00]. Suitability [KKP00]. suite [BBBD01, Reu03a, SCK^{+}00]. summary [DR89]. SUNY [Lit03]. super [Kas85, MTV05, MCT^{+}09]. super-peer [MTV05, MCT^{+}09]. super-speed [Kas85]. Supercomputer [KMN^{+}05, KGLA85, LM90b, SS90]. Supercomputers [Fer84, DFSZ88, DHD89, LM90b, Per86]. Supercomputing [Ros89, AS99, Han03, LPC^{+}95, Mun04, WTC^{+}02, WAD^{+}89]. SuperJANET [Coo94]. supernode [PR95, CGSZ95, AAB^{+}92]. superscalar [BFR05, Goo01]. supervisor [DMR93]. Support [BCM^{+}95, DDO^{+}92, Gra92, BK97, CY88, DMR93, DSH^{+}99, GS95, Hum92, LGH97, LGW07, NHG02, NHG03, ONHT89, OVK^{+}09, PM04, Pal06, Pal09, Per86, SB97, SGP^{+}09, VR05, WLF^{+}09, YCX05, dLLA93]. supported [HB08, KSM^{+}07a, KSM^{+}07b, LNJ04]. Supporting [BMS05, EMB98, Fra08, ADT03, KT08, LG08, Pag99, SZP00, SHLB08]. SUPRENUM [SS90]. Surface [Alb04, CFG93, GL95, IEG04, RBS93]. surfaces [DY04, ZY04]. surgery [SGL99]. Surgical [EKB00, WWSM98]. survey [Ano84a, Ano87m, FS93, Igl04a, MvWvL99]. SVD [PP06]. swarm [MCA02]. Swedish [HY03]. sweeps [SBA^{+}05]. SWIG [Bea03]. switched [CkLC06]. switching [CFVP03, MGYC06, SKF^{+}09]. Symbiosis [CSP98]. Symbolic [Fah98, GI07, FB93, Ref87]. Symbolic/numeric [GI07]. symmetric [AKH^{+}04, TC06]. SYMPATIX [CEJK94]. symplectic [IS03, MR03a, MR04b]. Symposium [Bis96]. synchronization [AC92, BB06, EV96, GI07, GE90, Pan95b]. Synchronous [Dui89, GL05, AT01, MD92b]. syndrome [KA88]. synergistic [AMB03].

SyNRAC [YA07]. syntactical [KL02]. synthesis [GPA96, Yos89]. synthetic [JLMR00, PFMC04]. Sys [MR04b]. Syst [KSM^{+}07a]. System [CCKW88, DLW86, HRSW99, LSS94, OSHH96, ACC^{+}05a, ANN^{+}92, AGJN00, AFPG91, AFP^{+}92, BBFW03, BL98, BT93, BG05, BLRS98, BPAP92, BDNP92, CM01, CST91, CTT^{+}08a, CGT07, CGST09, CSC^{+}92, DLW07, DGST09, DBA98, Din03, DIK^{+}06, ESFD06, ED04, FPX^{+}09, FW02, GCCC^{+}07, GS95, GHWZ94, GPA96, Gos00, GE90, HWS07, Han89, HPP94, HKM^{+}06, IMKB89, Kea93, KFF89, KKL09a, KRLR01, Kom89a, Lau92, LL04b, LL03, LC05, LWSC07, LDSH95, LF95b, Mae89, MD92b, MS01, MGLV04, Mur86, NWE04, Nis93, OFO^{+}99, Ohy89, ONHT89, Öst92, PHL98, PK99, PH99, PCG^{+}06, Pit96, PKSC02, Reh06, RSR01, RG04, SKT^{+}08, SPK^{+}07, SZP00, SLW01, Suz89, Tak89a, Tak89b, Tak05, TSK03, TDL05, Uch86, Ueh89, Yam89, Yam92, YCAS03, YWA^{+}89, dLLA93, Ama86, BBSV92, Bre89, Gut00, SH00, Suz89]. System [dRSBH94, vdR93b]. System/6000 [BBSV92]. Systematic [Kac00, Hol93]. Systems [DVV90, DSSU97, MOK06, NHG03, PW09, SdR99, Sti93, SJTG07, Van87a, Aba06a, ACML05, AMH02, AMH04, Ais88, AHdJF97, AKPN01, ACU95, ADF^{+}05, Ano84j, Ano86j, Ano87m, BJWZ08, BdCYG05, BJNH05, BPS06, BMZ01, Bis94, BKK02, BDL06, CDF^{+}05, CFGC03, Cas94, CCLS09, CY88, Cho04, Coo86, Dal03, DZ04, DLP06, DAM08, DFT92, DRNMC09, DW87, DD86, EO86, Fer96, FMR05, GS05, GHWZ94, Gil85a, GA06, GLW99, HH98, Hen87, HW95, Hir89, HPP94, Hol93, How91, HCL07, HXL90, HM98, HJK^{+}04, Igl07, JSK^{+}06, JL08, Joh92, JC09, JM01, KK00, Kag89, KLH^{+}04, Kim07a, hKcF09, KH89, KK97, Kun94, LR06, LXLS09, LC01, Lop03, LC03, LSS94, Mar86, MB01, MKH06, Mat89, MRV92, MPQ03, Meu05, Miz89a, MMR02, Mul92, Nit86, Pal01, PY00, PWY03, Par04, Par87]. systems [PX07, PB05, PVBH05, PH94, Pud87, QP08, RdLM06, RG04, SH99, Sar02, SSKF95, SG04, SB97, Ser98, SH90, Slo96, SS03, Ste85, SM96, Tho06, TTP^{+}07, VR05, Var03, Vau93, VS04, VGBLGS^{+}06, VR00, Ven09, VM93, VSV95, Wah84, WX02, WRBG94, WBT05, WYN^{+}90, WB90, ZS05a, Zem86, ZCT^{+}04, ZBB09, ZWJ04, dB90, vdR86a, vdR87c, vdR87d, vdR87f, vdR87l, vdR87g, vdR93a, Gra92, KI89, Slo06a, Slo06b]. Systolic [CCKW88]. Szpakowicz [Teb86, vdR87i].

T3D [MPG96]. T3E [SCK^{+}00]. table [FCW01]. tabling [DS00]. tactics [BS96]. TAF [GKS05]. tag [DCS^{+}07]. tailor [GVD^{+}03]. tailorable [BvdV99]. tangent [GKS05]. Target [KI89]. targeting [PNH99]. Task [BTM06, PCBD99, BFLL99, DLW07, HPLL09, JM01, LSTV07, MVRM08, SBHD08, ZSI08]. task-farm [BFLL99]. task-parallel [SBHD08]. Tasks [DFC^{+}08, AEGF^{+}01, LHL09, LYMZ09, Nos98, Reh06, Ser98, WL05]. taxonomy [ZS05a]. TC2 [BT93]. TCP [ABdL^{+}03, ABdLL05]. teaching [FGG03, SYT09]. Technical [vdR86b, CdCD07, GHWZ94, Yam89]. technique [DD05, HRJ^{+}06, HMP04, PCB99, RŻDM01, VS90, VM93, de 94]. techniques [Aba06b, ACML05, ATT96, AB95, ADK^{+}09, CY01, CTMO06, Fah98, Gra92, Hab05, KZC04, Mic97, NWE04, PB05, PdLS^{+}99, WKZ^{+}03, ZAP05, tTvH96]. Technische [Bun03]. technological [Aig86]. Technologies [DMSS97, Sim86, BGV97, Dub91, FM08, FR08, GMB^{+}05, Mal01, Mal02, Mal05, Nag86a, PKC^{+}05, SJTG07, TMV^{+}07]. Technology [BBD^{+}99, Abe92, Ano84a, BLAV06, Car86, Dek86, DvdHdL06, DD07, GMA07, How91, Kaw92, KM01, KA88, LKM91, NTN86, Poh87, Ser95, WAD^{+}89, vVDBB98, vdR86b, ABMS05]. tele [RMM^{+}98]. tele-immersive [RMM^{+}98]. Telecom [BT93]. telecommunications [Car86, Mun04]. TeleEEG [De 98]. telemedical [De 98, JSS^{+}99, Mar98b, Mar98a, Mar99b, Pol99]. telemedicine [EV98, FvLTT98, FP03, Moo99, PK99]. Telescience [LLH^{+}03]. telescope [BFK02]. telescope-ready [BFK02]. temperature [Tab06]. temperature-dependent [Tab06]. temporal [FX07, HYC04, RMM^{+}98, dB90]. Terabit [HRJ^{+}06]. Teraflows [GGH^{+}05]. TeraGrid [DIK^{+}06]. terascale [ZLD^{+}03]. TeraScope [ZLD^{+}03]. TeraStream [MWC^{+}03]. TeraVision [SLDK03]. terminals [Kam85]. terms [NB04]. terrain [PFMC04]. Test [BP94, Cho04, GODM98, KKL09a, Mal94, RRS99, RS94]. test-bed [RRS99]. testbed [MMFM^{+}05, SAMN02, ABB^{+}05]. tests [CDF^{+}05, PR95]. text [Kło05, SLŽ^{+}09]. texture [ZZN04]. their [Ano86i, Car03, CMZ95, DRS04, Igl07, PR95, PIKM02]. Thematic [Kaa99]. theorem [Sti93, UM02, WC06a]. theorem-proving [Sti93]. theoretic [Li90]. theoretical [BORM07, BRS04, GDRS04]. theories [WRBG94]. theory [DL03, Ger02, JK92, MH01, MKK03, NSHP88, PBT02, SH90]. Thermal [Tab06, DS04b]. thermo [Ned06]. thermo-visco-plastic [Ned06]. third [CLM00, DDS^{+}09, Pol98, Sin84, Bis96, CF09]. third-generation [DDS^{+}09]. thoughts [Mid01]. Thread [FD95]. Three [Wes99, FHG95b, Pal01, PPH^{+}09, XSM04]. Three-dimensional [Wes99, FHG95b, Pal01, XSM04]. threshold [CYH04, TYH04]. throughput [CGT07, CGL08]. Tibet [NMZC06]. tier [SA07, TLYT05]. tiers [LJ07]. tightly [Kun94]. Tim [Her87]. Time [CWD^{+}08, SGH^{+}08, BH03, BRR^{+}04, BDHK06, CCLS09, CC00, DT93, Doḡ09, Dua94, GKIZ05, GR96, HMP04, JL03, KA09, KSAOK03, KLC05, KKP00, LF01, LFWV05, LD04, MKT09, Pal09, Pfi99, PN09, SS04, SD02, SST^{+}06, SKF^{+}09, Sin92, SW02, Suz89, VLK09, Vau93, XYZ05, ZSI08]. time-dependent [Dua94]. time-out [KSAOK03]. time-scale [JL03]. time-series [LFWV05]. time-shared [CCLS09]. time-varying [BH03]. timed [Pap05, PIKM02]. times [RNJK09]. TIN [TG04]. Titan [LM90b]. Titan-2 [LM90b]. TODS [ZJWZ04]. Toile [BBG^{+}05]. tokens [GXD^{+}09]. Tolerance [PCBD99, CdCD07, DSS07, LSTV07, MC04, Sar02, SG05]. tolerant [AMH02, AFP07, BCH^{+}08, DHB02, FD02, KL02, LHB95, LAM07, LS01, LS08, PWY03, THKG98, Xia06]. tomography [FGCM07]. tomorrow [vdR87j]. Tony [Her87]. Tool [LLSR02, DMR93, FS07, KZLK06, LWHC07, OS92, Reu03a, WSS^{+}09, WBF08, WYN^{+}90]. toolkit [CN92, HAE^{+}03, HBJ^{+}03, IJLC03, LWHS07, SAMN02, MFE^{+}08]. toolkits [YPF05]. Tools [KV03, TA96, ACU95, BR92, BSG^{+}05, GD93b, MCSS00, TC06, Wal86]. toolset [RSSD02]. TOP [CCHW03]. TOP-C [CCHW03]. Topic [DL04]. Topological [AAC04]. Topological-order [AAC04]. topologies [VS90]. tort [Kag89]. torus [KD00, LOK09]. Toshiba [Ama86]. totalistic [JM02]. trace [KN06, SK06]. tracing [BK06]. Trade [KNK^{+}08]. Trade-offs [KNK^{+}08]. trading [HQ07]. traditional [KC98]. Traffic [Mid01, RN01, RGH^{+}01, SNW01, VN01, ADOKM06, AS02, CG09, LOK09, VVC^{+}03]. training [BLB03, EKB00, Gra01, Han03, IMKB89, RS98]. trajectory [FAJP99, SOR05]. Transaction [Joh92]. transatlantic [ABdL^{+}03, ABdLL05, MMFM^{+}05, RRS99]. transcoding [GFR^{+}06]. transcript [STP^{+}05]. transcripts [TBD^{+}02]. Transfer [NMZC06, BBD^{+}99, KKL09b, KMCH03, OS06]. Transformation [BCL88, MK95, RM97]. transformations [DV03, UM02]. Transformer [YYW^{+}09]. transient [BDHK06]. TRANSIMS [RN01]. transitions [DCK03, Fre94, KLC05]. Translation [Ama86, Mur86, Nag86b, NTN86, Nit86, Sim86, Uch86]. transmission [HJPS03, KY85, SKF^{+}09, VS04]. transmit [HJK^{+}04]. transonic [GL95]. transparent [CWD^{+}08]. Transport [HDC^{+}94, AS02, CM99, JHL^{+}06, MKH06]. transportation [RN04]. transputer [CFG93, Dui89, LSF^{+}94, MSLP93, mM95]. transputer-based [mM95]. trapezoidation [LD04]. Trapper [SSKF95]. travel [Van87b, vdR87l]. Traveling [SMI01, VAS95]. treatment [Mae89]. Tree [LM07, LLZ07, SHLB08, ZQZZ09]. Tree-based [LM07]. Treecode [Pri95]. Trees [RSRV88, AW97, Luk89, SLO^{+}05b, WLB00]. Trends [Her91, Sas85, WG91, AMB03, Ano87m, Bal91b, CRM05, GMS09, Nag86b, Wii84]. trial [KMCH03]. triangles [BRR^{+}04]. triangular [vdS04]. triangulation [XSM04]. tridiagonal [BG05, HM98]. triplet [Tul04]. tropical [VBLS09]. troubleshooting [Tak89a]. truly [HNS05]. trust [AHdJF97, DSS07]. Trusted [Pol98, SK97, CLM00, DCL00, KF00]. trustworthiness [HSH^{+}07]. TTN [BBD^{+}99]. TTPs [Pol99]. tube [HvHAS04]. tubulin [Tul04]. tumour [SUD^{+}98]. tuning [FMS08b, WVC05]. turbulence [NEJP94]. turbulent [GL95, Kni89, VWCV94]. Turing [DDL01]. Turing-like [DDL01]. Two [AB95, GJS^{+}94, KI89, Rus90b, BP02, DS04b, HM98, LM90b, PBHK01, SA07, Ser98, UM02, BBSV92, Mur86]. two-dimensional [BP02, DS04b]. two-layer [PBHK01]. Two-level [Rus90b].

Two-phase [Mur86]. two-processor [Ser98]. two-tier [SA07]. two-way [HM98]. type [BV04, Čie04, Li90]. type-theoretic [Li90]. typed [BM00]. Typing [Goo02].

U.K. [Ano84a]. U.S. [Ano84h, Smi86, WAD^{+}89]. U.S.A. [Ano84i, BS84]. ubiquitous [WZC08]. UCLP [JHL^{+}06]. UCLP-enabled [JHL^{+}06]. UDT [GGH^{+}05]. UK [HT02]. ultra [LRJ^{+}06, VBLS09]. ultra-fast [VBLS09]. ultra-high-definition [LRJ^{+}06]. uncertainty [Ano86j, FB93, SPM86, TM05, ZMS^{+}06]. uncompressed [SKF^{+}09]. undergraduates [BLB03]. Understanding [CRE01, Ano85g, SJTG07]. underwriting [Kom89a]. unexpected [SCB04]. unfolding [DCK03]. UNICORE [AS99]. unified [OFO^{+}99]. uniform [AS99, BMS05, KLC05, LOK09, Van92]. unifying [VP94]. United [Wii84]. universal [ADKS06, Bur02]. Universität [Bun03]. University [Ano86i, Ano87b, Ano87c, Ano87l, BLB03, CN98, HICÁFM^{+}06, SR03]. UNIX [AGP^{+}92, Ber98]. Unsteady [HvHAS04]. Unstructured [LF95a, CH95, vdS04]. unsymmetric [SG04]. up-to-date [Din99]. UPC [EGCY^{+}06]. Update [GSD95, ZWJ04]. Update-Based [GSD95]. Upgrading [WWC^{+}97]. USA [vdR86b]. usability [GMB^{+}05]. usage [FD02, PGPW09]. Use [CTMO06, Niw89, BKŁ01, HHSW92, JSK^{+}06, JNPY06, Jon00, LNJ04, MJ00, Ser98, VM93]. useful [CN98]. User [Gra92, dNE05, AMB03, BvdV99, BFL99, DDRR96, Kam85, KV09, KMK09, KBB^{+}09, LHL03, MSK03, RSSD02, Tak89b, GFR^{+}06]. user-centred [AMB03]. user-level [KMK09]. user-oriented [BvdV99]. USERNET [KGLA85]. users [KT08, OS92]. uses [Ano86i]. Using [AC92, AHdJF97, BB06, BCG05, CGM^{+}07, CCHW03, CS97, Dal03, DvdHdL06, FR08, HLSØ06, IJLC03, LDS06, MRV92, PGPW09, PBB^{+}05, Reu03b, RV95, RS98, SK97, SSC04, SBLT05, tTvH96, vVDBB98, vdHDT^{+}06, AAC04, ABB^{+}03, ACE02, AKW90b, AKW90a, ABL04, Bal91a, BvdHN^{+}01, BRR^{+}04, BKSS02, BRMN04, BG05, Büc05a, BK06, CWD04, CSJN05, CPK05, CGSZ95, CFG93, FC05, FC09, Gil94, GXD^{+}09, GD93b, GGH^{+}03, GGSZ09, HWS07, Han03, HBH09, HQ07, HWZL08, HAB^{+}06, IdLR01, JFDF09, KZLK06, KLM^{+}03, KKYK04, KKP^{+}05, KKL09b, LCP04, LL04b, LOJ^{+}07, LLWN04, LYT^{+}05, Luk89, MCSS00, MJ98, MC04, NKX09, NSI84, NCS04, PA01a, Pap05, Pri95, RdSH^{+}00, RCD03, SKT^{+}08, SZ98, SH99, SW05, SVC^{+}07, SF06, VDPHS09, VBLS09, WDD00, WBF08, Xia06, XSM04, YJA03, YD05, ZMS^{+}06, vdV89a, vdS04]. utilisation [EPJ^{+}05]. utility [BYV^{+}09, MSLP93, VDPHS09, Ven09, KLM^{+}03]. utility-based [Ven09]. utilization [DFC^{+}08, NF07]. utilizing [KTY03]. Utrecht [Bis96].

V [vdR87c]. Vadera [vdR87j]. vague

[vdR87k]. Validating [LYT^{+}05]. validation [Hol93]. value [BH03]. variable [RNJK09]. variance [DHS99]. variant [JCSS01]. variation [CPK05]. Varrier [PSG^{+}06]. varying [BH03]. vascular [ZS05b]. VDM [BHH91, BHH92]. VE [DBA98]. Vector [LFWV05, VSvD95, HBCR01, KCK04, PM04, PdLS^{+}99]. Vectorization [vM94]. vectorized [LN94]. vectors [BB04]. vehicle [BBC^{+}99, DOV01, HDC^{+}94, OVDV98]. VENUS [Mur86]. verifiable [KKL09b]. verification [Hol93, TYH04, ZZ09, dLLA93]. Versatile [HTV07]. version [JL95, LFWV05]. versus [BVDF00, BM00, GJS^{+}94, NHT06]. vertex [GL95]. Very [Kło05, GLA88]. VHSIC [Ano86m]. VI [HO02]. via [HHG05, MHA09]. Video [ASPB03, GFR^{+}06, JHL^{+}06, WWC^{+}97]. video-on-demand [WWC^{+}97]. videoconferencing [SWCP03]. videos [KY04]. videostreaming [PK08]. view [VS04, YCAS03]. viewers [WKF03]. VIGO [ACC^{+}05a]. Virtual [BDF^{+}99, BKS98, CCBR98, CT09, CHK98, DBdL03, DGS09, DFG^{+}00, Dzw97, Kaa98, Kaa99, KTY03, KGX95, MSR98, PBM95, Rou00, SOR05, SM01b, Zin00, ACC^{+}05a, AKB^{+}01, ASTEP98, BGH^{+}03, BBSV92, BLRS98, BFC02, CM01, CN98, DDS^{+}09, DBA98, Dua94, EGAQ09, GRPL04, GFGB03, HRJ^{+}06, JSS^{+}99, Lee04, Mar99a, MSK03, PADD03, PSG^{+}06, PHM^{+}99, Ram95, RL98, RMM^{+}98, RJH^{+}09, Sch98, SUD^{+}98, TDG^{+}06, vVDBB98, EV98, SB99]. virtualized [ACC^{+}05a, BJWZ08]. visa [GWO03]. visco [Ned06]. visco-plastic [Ned06]. visibility [DS04a]. Vision [BYV^{+}09, CCKW88, LPK94, Shi85]. visionary [Pin87]. visit [Kuo86]. Visual [Kam85, ACGdT02, ATJMZ02, DB99, ZLD^{+}03]. Visualisation [PB05, CTMO06, DMM^{+}99, ZAP05]. Visualization [GLW99, GLS99, LRJ^{+}06, MJ98, Pag99, DLR^{+}09, DMMP98, EWG99, EdBG^{+}99, HAB^{+}06, IdLR01, LGH97, LJPS05, PSG^{+}06, PK08, PdLS^{+}99, RvdSB^{+}03, RJH^{+}09, Rum99, SBSdL06, Sch00, SUD^{+}98, TWC^{+}06, Tak05, Wes99, ZCW^{+}04, dKdOS03]. Visualizing [CTF^{+}99, DK00, SW99, BGC^{+}03, DIK^{+}06]. vitro [Kar01, SUD^{+}98]. ViVa [Hol93]. vivo [Kar01]. VizieR [OFO^{+}99]. VLADYMIR [LC04]. VLAM [BGH^{+}03, HBJ^{+}03]. VLAM-G [BGH^{+}03, HBJ^{+}03]. vlbiGRID [HJPS03]. VLSI [BS84, KA88, OBK88, Sas85, TR85, VS88]. VMSLab [GRPL04]. VMSLab-G [GRPL04]. VoIP [YCX05]. volume [Ano85b, Ano86a, Ano87d, Ano89a, Ano90a, Ano91a, Ano92a, Ano92b, Ano92h, Ano92i, Ano93a, Ano93i, Ano94a, Ano94g, Ano95a, Ano95h, Ano98a, Ano98c, Ano03a, Ano03i, Ano05f, EWG99, HH98, LGMV02, Pfi99]. volumes [Ano94b]. volunteer [SH99, Sar02]. VOMS [ACC^{+}05b]. Voronoi [KS02, LGMV02]. voting [Din03]. VR [EMB98, Kaa98, MSR98]. VRML [Zin00, AD00, AV00, Avg00, IdLR01, Jon00, MJ98, MJ00]. VRML-enhanced [AV00]. vs [BCH^{+}08, Ger02]. VTK [WKF03]. VTK/CAVE^{T M} [WKF03].

W [vdR87g]. Wafer [AKW90a]. Wait [Ray05, BB06]. Wait-free [Ray05, BB06]. wakes [SdSP04]. Walker [vdR87g]. WAM [CY88, CYB90]. WAM-based [CY88]. WAM-Plus [CYB90]. WAMM [BFL99]. WAN [TDG^{+}06]. WANs [GGH^{+}05]. wastewater [Mae89]. water [Igl04a]. Waterman [Zem86]. wave [Fin99, GCK98, Kni89, Sap88]. WAVE-1 [Sap88]. wave-turbulent [Kni89]. Wavelet [KKV^{+}99, SMC99, KU01]. Wavelet-based [KKV^{+}99, SMC99]. waves [GPH^{+}94]. way [ABF^{+}03, HM98]. weather [WSH99, PSP^{+}09]. Web [Opp00, YPF05, Qu04, ADM06, AKMK05, AFP07, AMW99, BJA^{+}05, BKKW99, Bru01, BCMA07, CWD04, CBK^{+}01, DCS^{+}07, EKB00, GBE00, Gra01, HAFF99, HAF00, Lan00, LWHC07, LJ07, LLKF09, LC03, MCSS00, MJ06, Mar98a, MSX00, MPPM09, NKX09, NMA00, OF07, OVK^{+}09, PO00, PPSS06, PPAK99, PBB^{+}05, QP08, SH99, SMA08, She00, SGY^{+}07, YMM00]. Web-based [AMW99, Bru01, CBK^{+}01, GBE00, Gra01, LC03, MCSS00, MJ06, Mar98a, MSX00, PO00, PPAK99, SH99]. WebFlow [HAF00]. Webservices [OFT09]. Websim99 [Bru01, FCW01, Gra01, IdLR01, PF01, PSS01, SM01a]. weighted [LFWV05]. Welcome [vdR85]. Wesley [Zem86, vdR87g]. Wherrett [vdR87e]. whiteboard [FP03]. whole [CC00]. whole-program [CC00]. Wide [Opp00, OS01, GGH^{+}06, GGSZ09, LRJ^{+}06, LWSC07, RRS99, VBLS09]. Wide-area [OS01, GGH^{+}06, LRJ^{+}06, VBLS09]. wildland [ACML05]. Wiley [Ano87c, vdR87a, vdR87d]. Wilson [vdR87g]. Windows [FPX^{+}09]. wireless [LBYL08, MK04, PSVL02]. within [BBD^{+}99, PO00, Pag99, PA01b]. without [CYH04, Gor02]. WK [VS90]. WK-recursive [VS90]. WonderSpace [NMA00]. work [Hen87]. workbench [dRSBH94, Wes99]. worker [Bal91a]. Workflow [BGK^{+}05, FPX^{+}09, NHG06, WCC^{+}09, CS09, DGST09, GGW^{+}09, GMEL08, HB09, KKS08, Kim07a, KTM^{+}08, LWHC07, LS05, MBZL09, SD06, WHP09, ZBB09, AKMK05]. Workflow-based [NHG06, WCC^{+}09]. Workflows [DGST09, DGS09, GP09, LH07, SSC09, TDF07, TBNF09, WLF^{+}09]. Working [Ano86h]. workload [KMK09, OP97]. workload-based [OP97]. Workloads [ILJ^{+}08, BK06, CkLC06]. Workshop [ABMS05, AB03, WBMP99, Zna94]. Workspace [FBBW99]. Workstation [BM92, Hum92, PBM95, BPC^{+}01, Bro92, ELvD^{+}96, MRV92, PL96, Sin92, SD03]. workstation-farm [Bro92]. workstations [BDNP92, DZ98, MM03, SMI01]. world [FSP02, Gra01, JSS^{+}99, LAM07, Zin00, Opp00]. Worlds [EV98, Mar99a, PADD03]. worldwide [ELvD^{+}96]. wormhole [KSAOK03, KD00, Pan95b]. wormhole-routed [KD00]. worms [Pan95b]. Worst [HND06]. Wrapping [LRW01]. Writes [OB04]. writing [BMFC07]. WS [HML07]. WSI [AKW90b]. WSRF [CTT07, MHA08, MHA09]. WSRF-based [MHA08, MHA09]. WWW [Opp00, CLM00, DMMP98, DMM^{+}99, SCY01].

X.509 [CO03]. XML [ATJMZ02, AMW99, DL04, HYC04, LJ07, LYT^{+}05]. XML-based [ATJMZ02]. XPath [LYT^{+}05]. XTM [DL04]. XtremWeb [CDF^{+}05]. xvi [Zem86].

Yao [BDNN02]. Ygdrasil [PADD03]. York [Zad87, vdR87a, vdR87i].

zero [DvdHdL06]. Zimmermann [vdR87f]. Zone [CKK^{+}04].

References

Adamo:1992:DSI

[AAB^{+}92] J. M. Adamo, N. Alhafez, J. Bonneville, C. Bonello, P. Moukeli, and L. Trejo. Developing a software interface for the dynamically reconfigurable SuperNode multiprocessor. Future Generation Computer Systems, 8(1– 3):183–189, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Agarwal:2007:GCC

[AAB^{+}07] A. Agarwal, M. Ahmed, A. Berman, B. L. Caron, A. Charbonneau, D. Deatrich, R. Desmarais, A. Dimopoulos, I. Gable, L. S. Groer, R. Haria, R. Impey, L. Klektau, C. Lindsay, G. Mateescu, Q. Matthews, A. Norton, W. Podaima, D. Quesnel, R. J. Sobie R. Simmonds, B. St. Arnaud, C. Usheri, D. C. Vandersteri, M. Vetterlid, R. Walker, and M. Yuen.

GridX1: a Canadian computational grid. Future Generation Computer Systems, 23 (5):680–687, June 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ahn:2004:TOB

[AAC04] Byungoh Ahn, Seongjin Ahn, and Jinwook Chung. Topologicalorder based dynamic polling scheme using biconnected component computation. Future Generation Computer Systems, 20(2):275–282, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anderson:1995:TTI

[AB95] Craig Anderson and Jean-Loup Baer. Two techniques for improving performance on bus-based multiprocessors. Future Generation Computer Systems, 11(6):537–551, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Afsarmanesh:2001:GEH

[AB01] Hamideh Afsarmanesh and Marian Bubak. Guest editorial: High-performance computing and networking enables complex applications. Future Generation Computer Systems, 17(8):v–vi, June 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

                                        elsevier.com/gej-ng/10/
                                        19/19/45/35/25/abstract.
                                        html.

Arbenz:2003:SWD

[AB03] Peter Arbenz and Torsten Braun. SPEEDUP workshop on distributed computing and high-speed networks. Future Generation Computer Systems, 19(1):99–100, January 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abawajy:2006:APS

[Aba06a] J. H. Abawajy. Adaptive parallel I/O scheduling algorithm for multiprogrammed systems. Future Generation Computer Systems, 22(5): 611–619, April 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abawajy:2006:SSP

[Aba06b] J. H. Abawajy. Special section: Parallel input/output management techniques (PI-OMT) in cluster and grid computing. Future Generation Computer Systems, 22(5):600, April 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abawajy:2009:EAS

[Aba09] J. H. Abawajy. An efficient adaptive scheduling policy for high-performance computing. Future Generation Computer Systems, 25(3):

364–370, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Allcock:2003:GEP

[ABB^{+}03] W. Allcock, J. Bresnahan, J. Bunn, S. Hegde, J. Insley, R. Kettimuthu, H. Newman, S. Ravot, T. Rimovsky, C. Steenberg, and L. Winkler. Grid-enabled particle physics event analysis: experiences using a 10 Gb, high-latency network for a high-energy physics application. Future Generation Computer Systems, 19(6): 983–997, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alfieri:2005:IGT

[ABB^{+}05] R. Alfieri, R. Barbera, P. Belluomo, A. Cavalli, R. Cecchini, A. Chierici, V. Ciaschini, L. Dell’Agnello, F. Donno, E. Ferro, et al. The INFN-Grid Testbed. Future Generation Computer Systems, 21 (2):249–258, February 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alexandris:2000:SLC

[ABCD00] N. Alexandris, M. Burmester, V. Chrissikopoulos, and Y. Desmedt. Secure linking of customers, merchants and banks in electronic commerce. Future Generation Computer Systems, 16 (4):393–401, February 2000.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/27/34/abstract.
                                        html.

Antony:2003:MET

[ABdL^{+}03] Antony Antony, Johan Blom, Cees de Laat, Jason Lee, and Wim Sjouw. Microscopic examination of TCP flows over transatlantic links. Future Generation Computer Systems, 19(6):1017–1029, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Antony:2005:EPL

[ABdLL05] Antony Antony, Johan Blom, Cees de Laat, and Jason Lee. Exploring practical limitations of TCP over transatlantic networks. Future Generation Computer Systems, 21 (4):489–499, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abe:1992:PSH

[Abe92]  Masayuki Abe.  Present

status of HEMT LSI technology. Future Generation Computer Systems, 7(2–3): 283–291, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aharoni:1993:AGC

[ABF93] Gad Aharoni, Amnon Barak, and Yaron Farber. An adaptive granularity control algorithm for the parallel execution of functional programs. Future Generation Computer Systems, 9(3):163– 174, September 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Andronico:2003:GSE

[ABF^{+}03] A. Andronico, R. Barbera, A. Falzone, P. Kunszt, G. Lo Re, A. Pulvirenti, and A. Rodolico. GENIUS: a simple and easy way to access computational and data grids. Future Generation Computer Systems, 19(6):805–813, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abramson:2002:CEG

[ABG02] David Abramson, Rajkumar Buyya, and Jonathan Giddy. A computational economy for grid computing and its implementation in the Nimrod-G resource broker. Future Generation Computer Systems, 18 (8):1061–1074, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anderson:1994:PSS

[ABK94] Paul Anderson, David Bolton, and Paul Kelly. Paragon specifications: Structure, analysis and implementation. Future Generation Computer Systems, 10(1):137–148, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Argentini:2004:EUM

[ABL04] R. Argentini, A. F. Bakker, and C. P. Lowe. Efficiently using memory in lattice Boltzmann simulations. Future Generation Computer Systems, 20(6):973–980, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aloisio:2007:SSL

[ABM^{+}07] Giovanni Aloisio, Vincent Breton, Maria Mirto, Almerico Murli, and Tony Solomonides. Special section: Life science grids for biomedicine and bioinformatics. Future Generation Computer Systems, 23 (3):367–370, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Arbenz:2005:SSS

[ABMS05] P. Arbenz, H. Burkhart, E. Maehle, and O. Schenk. Special section: SPEEDUP Workshop on Modern Algorithms in Computational Science and Information Technology. Future Generation Computer Systems, 21(8):1249– 1250, October 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alari:1992:UFM

[AC92] G. Alari and A. Ciuffoletti.

Using full message exchange for probabilistic clock synchronization (extended abstract). Future Generation Computer Systems, 8(1–3): 249–252, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Allan:2001:CSA

[AC01] V. H. Allan and X. Chen. Convert2Java: semi-automatic conversion of C to Java. Future Generation Computer Systems, 18(2):201– 211, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/28/abstract.
html.

Adabala:2005:VRV

[ACC^{+}05a] Sumalatha Adabala, Vineet Chadha, Puneet Chawla, Renato Figueiredo, José Fortes, Ivan Krsul, Andrea Matsunaga, Mauricio Tsugawa, Jian Zhang, Ming Zhao, et al. From virtualized resources to virtual computing grids: the In-VIGO system. Future Generation Computer Systems, 21 (6):896–909, June 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alfieri:2005:GFV

[ACC^{+}05b] R. Alfieri, R. Cecchini, V. Ciaschini, L. dell’Agnello, Á. Frohner, K. Lőrentey, and F. Spataro. From gridmapfile to VOMS: managing authorization in a Grid environment. Future Generation Computer Systems, 21 (4):549–558, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aloisio:2002:EEG

[ACE02] Giovanni Aloisio, Massimo Cafaro, and Italo Epicoco. Early experiences with the GridFTP protocol using the GRB-GSIFTP library. Future Generation Computer Systems, 18(8):1053–1059, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Acacio:2002:MDM

[ACGdT02] M. Acacio, O. Cánovas, J. M. García, and P. E. López de Teruel. MPI-Delphi: an MPI implementation for visual programming environments and heterogeneous computing. Future Generation Computer Systems, 18(3): 317–333, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/28/abstract.
html.

Abdalhaq:2005:EWF

[ACML05] Baker Abdalhaq, Ana Cortés, Tomàs Margalef, and Emilio Luque. Enhancing wildland fire prediction on cluster systems applying evolutionary optimization techniques. Future Generation Computer Systems, 21(1):61–67, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Allen:1995:MMT

[ACU95] Paul Allen, Iain Cramb, and Colin Upstill. Monitoring and modelling tools for high performance database systems. Future Generation Computer Systems, 11(2): 193–198, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amon:2000:GEE

[AD00] Tomaz Amon and Stephan Diehl. Guest editorial: Educational applications of VRML. Future Generation Computer Systems, 17(1):vii– ix, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/24/24/abstract.
html.

Andreozzi:2005:GMS

[ADF^{+}05] Sergio Andreozzi, Natascia De Bortoli, Sergio Fantinel, Antonia Ghiselli, Gian Luca Rubini, Gennaro Tortone, and Maria Cristina Vistoli. GridICE: a monitoring service for Grid systems. Future Generation Computer Systems, 21 (4):559–571, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Asiki:2009:GMD

[ADK^{+}09] Athanasia Asiki, Katerina Doka, Ioannis Konstantinou, Antonis Zissimos, Dimitrios Tsoumakos, Nectarios Koziris, and Panayiotis Tsanakas. A Grid middleware for data management exploiting peerto-peer techniques. Future Generation Computer Systems, 25(4):426–435, April 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aljundi:2006:UPF

[ADKS06] A. Chadi Aljundi, Jean-Luc Dekeyser, M-Tahar Kechadi, and Isaac D. Scherson. A universal performance factor for multi-criteria evaluation of multistage interconnection networks. Future Generation Computer Systems, 22(7): 794–804, August 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abhari:2006:WOB

[ADM06] Abdolreza Abhari, Sivarama P. Dandamudi, and Shikharesh Majumdar. Web object-based storage management in proxy caches. Future Generation Computer Systems, 22(1–2): 16–31, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Al-Dubai:2006:BNT

[ADOKM06] Ahmed Al-Dubai, Mohamed Ould-Khaoua, and Lewis Mackenzie. On balancing network traffic in path-based multicast communication. Future Generation Computer Systems, 22(7):805–811, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aldinucci:2003:AES

[ADT03] M. Aldinucci, M. Danelutto, and P. Teti. An advanced environment supporting structured parallel programming in Java. Future Generation Computer Systems, 19(5):611–626, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alaoui:2001:MTN

[AEGF^{+}01] S. Mounir Alaoui, T. El-Ghazawi, O. Frieder, A. Bellachia, and A. Bensaid. Mapping tasks onto nodes: a parallel local neighborhood approach. Future Generation Computer Systems, 17 (4):397–403, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

                                        elsevier.com/gej-ng/10/
                                        19/19/45/29/30/abstract.
                                        html.

Alef:2009:IMM

[AFF^{+}09] M. Alef, T. Fieseler, S. Freitag, A. Garcia, C. Grimm, W. Gürich, H. Mehammed, L. Schley, O. Schneider, and G. L. Volpato. Integration of multiple middlewares on a single computing resource. Future Generation Computer Systems, 25(3): 268–274, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Arentoft:1992:OAP

[AFP^{+}92] M. M. Arentoft, J. J. Fuchs, Y. Parrod, A. Gasquet, J. Stader, I. Stokes, and H. Vadon. OPTIMUM-AIV: a planning and scheduling system for spacecraft AIV. Future Generation Computer Systems, 7(4):403–412, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aktas:2007:FTH

[AFP07] Mehmet S. Aktas, Geoffrey C. Fox, and Marlon Pierce. Fault tolerant high performance information services for dynamic collections of Grid and Web services. Future Generation Computer Systems, 23 (3):317–337, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Arentoft:1991:OAP

[AFPG91] M. M. Arentoft, J. J. Fuchs, Y. Parrod, and A. Gasquet. OPTIMUM-AIV: a planning and scheduling system for spacecraft AIV. Future Generation Computer Systems, 7 (4):403–412, May 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Agrawal:1992:SSP

[AG92] Anant Agrawal and Robert B.

Garner. SPARC: a scalable processor architecture. Future Generation Computer Systems, 7(2–3):303–309, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alt:2005:AJR

[AG05] Martin Alt and Sergei Gorlatch. Adapting Java RMI for grid computing. Future Generation Computer Systems, 21 (5):699–707, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ammerlahn:2000:GDE

[AGJN00] Heidi R. Ammerlahn, Michael E.

Goldsby, Michael M. Johnson, and David M. Nicol. A geographically distributed enterprise simulation system. Future Generation Computer Systems, 17(2):135– 146, October 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/27/28/abstract.
html.

Albinson:1992:ULC

[AGP^{+}92] Lawrence Albinson, Dominique Grabas, Pascal Piovesan, Michel Tombroff, Christian Tricot, and Hossein Yassaie. UNIX on a loosely coupled architecture: The CHORUS/MiX approach. Future Generation Computer Systems, 8(1–3):67–81, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Arbab:1994:M

[AH94] F. Arbab and I. Herman. Manifold. Future Generation Computer Systems, 10(2–3): 273–277, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alberda:1997:UFM

[AHdJF97] Marjan I. Alberda, Pieter H. Hartel, and Eduard K. de Jong Frz. Using formal methods to cultivate trust in smart card operating systems. Future Generation Computer Systems, 13(1):39– 54, June 20, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
19/19/28/17/19/abstract.
html.

Aigrain:1986:TP

[Aig86] M. Aigrain. The technological perspective. Future Generation Computer Systems, 2 (1):23–25, March 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aiso:1988:FGC

[Ais88] Hideo Aiso. The fifth generation computer systems project. Future Generation Computer Systems, 4(3):159– 175, October 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aliabadi:2002:PSF

[AJZ^{+}02] Shahrouz Aliabadi, Andrew Johnson, Bruce Zellars, Ade Abatan, and Charlie Berger. Parallel simulation of flows in open channels. Future Generation Computer Systems, 18 (5):627–637, April ??, 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/30/abstract.
html.

Afsarmanesh:2001:RAS

[AKB^{+}01] H. Afsarmanesh, E. C.

Kaletas, A. Benabdelkader, C. Garita, and L. O. Hertzberger. A reference architecture for scientific virtual laboratories. Future Generation Computer Systems, 17 (8):999–1008, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/36/abstract.
html.

Artoli:2004:LBS

[AKH^{+}04] A. M. Artoli, D. Kandhai, H. C. J. Hoefsloot, A. G. Hoekstra, and P. M. A. Sloot. Lattice BGK simulations of flow in a symmetric bifurcation. Future Generation Computer Systems, 20(6): 909–916, August 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Abramson:2005:ASW

[AKMK05] David Abramson, Jagan Kommineni, John L. McGregor, and Jack Katzfey. An Atmospheric Sciences Workflow and its implementation with Web Services. Future Generation Computer Systems, 21(1): 69–78, January 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alkindi:2001:OAE

[AKPN01] A. M. Alkindi, D. J. Kerbyson, E. Papaefstathiou, and G. R. Nudd. Optimisation of application execution on dynamic systems. Future Generation Computer Systems, 17 (8):941–949, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/30/abstract.
html.

Aiso:1986:E

[AKvdR86] H. Aiso, F. Kuo, and R. P.

van de Riet. Editorial. Future Generation Computer Systems, 2(1):3–4, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anderson:1990:FPC

[AKW90a] P. Anderson, P. Kelly, and P. Winterbottom. The feasibility of a parallel computer using wafer scale integration. Future Generation Computer Systems, ??(??):??, May 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anderson:1990:FGP

[AKW90b] Paul Anderson, Paul Kelly, and Phil Winterbottom. The feasibility of a general-purpose parallel computer using WSI. Future Generation Computer Systems, 6(3):241–253, December 1990. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alberts:2004:SRS

[Alb04] Claus-Peter Alberts. Surface reconstruction from scan paths. Future Generation Computer Systems, 20(8): 1285–1298, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alder:1989:CCS

[Ald89] Berni J. Alder. Challenges in computational statistical mechanics. Future Generation Computer Systems, 5(2– 3):191–195, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Alexandre:1997:BSC

[Ale97] Thomas J. Alexandre. Biometrics on smart cards: an approach to keyboard behavioral signature. Future Generation Computer Systems, 13 (1):19–26, June 20, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/17/18/abstract.
                                        html.

Allard:1992:AIS

[All92] F. Allard. Artificial intelligence and space. Future Generation Computer Systems, 7 (4):341–342, May 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amano:1986:TMT

[Ama86] S. Amano. The Toshiba Machine Translation System.

Future Generation Computer Systems, 2(2):121–123, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amamiya:1988:DFC

[Ama88] Makoto Amamiya. Data flow computing and parallel reduction machine. Future Generation Computer Systems, 4 (1):53–67, August 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amamiya:1989:DFC

[Ama89] M. Amamiya. Data flow computing and parallel reduction machine. Future Generation Computer Systems, 4 (??):53–67, ???? 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Almasi:1992:PDS

[AMB^{+}92] G. S. Almasi, T. McLuckie, J. Bell, A. Gordon, and D. Hale. Parallel distributed seismic migration. Future Generation Computer Systems, 8(1–3):9–26, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Angelov:2003:SIM

[AMB03] C. Angelov, R. V. N. Melnik, and J. Buur. The synergistic integration of mathematics, software engineering, and user-centred design: exploring new trends in education. Future Generation Computer Systems, 19(8):1299– 1307, November 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Afzal:2008:CPS

[AMD08] Ali Afzal, A. Stephen Mc-Gough, and John Darlington. Capacity planning and scheduling in Grid computing environments. Future Generation Computer Systems, 24 (5):404–414, May 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ahn:2002:SEF

[AMH02] JinHo Ahn, Sung-Gi Min, and Chong-Sun Hwang. Scalable and efficient faulttolerant protocol for mobility agents in mobile IP-based systems. Future Generation Computer Systems, 18(5): 613–625, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/29/abstract.
html.

Ahn:2004:CML

[AMH04] JinHo Ahn, Sung-Gi Min, and Chong-Sun Hwang. A causal message logging protocol for mobile nodes in mobile computing systems. Future Generation Computer Systems, 20 (4):663–686, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amit:1990:ANN

[Ami90] Daniel J. Amit. Attractor neural networks and biological reality: associative memory and learning. Future Generation Computer Systems, 6(2):111– 119, November 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amodio:2006:CFE

[Amo06] Pierluigi Amodio. On the computation of few eigenvalues of positive definite Hamiltonian matrices. Future Generation Computer Systems, 22 (4):403–411, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aloisio:1999:XAH

[AMW99] G. Aloisio, G. Milillo, and R. D. Williams. An XML architecture for high-performance Web-based analysis of remotesensing archives. Future Generation Computer Systems, 16 (1):91–100, November 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/34/abstract.
html.

Ardaiz:2008:GBD

[AN08] O. Ardaiz and L. Navarro. Grid-based dynamic service overlays. Future Generation Computer Systems, 24(8): 813–823, October 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Akima:1992:SSC

[ANN^{+}92] Noboru Akima, Shuji Nakata, Akiyoshi Nishijima, Tetsuya Kubo, and Yasuo Konishi. Sigma system: Concepts and structure. Future Generation Computer Systems, 7(2– 3):293–302, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:ASA

[Ano84a] Anonymous. An Alvey survey: Advanced information technology in the U.K. Future Generation Computer Systems, 1(1):69–78, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:A

[Ano84b] Anonymous. Announcement. Future Generation Computer Systems, 1(1):91–92, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:Ca

[Ano84c] Anonymous. Calendar. Future Generation Computer Systems, 1(1):89–90, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:Cb

[Ano84d] Anonymous. Calendar. Future Generation Computer Systems, 1(2):145–149, November 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:ESE

[Ano84e] Anonymous. ECAI-84 — Sixth European Conference on Artificial Intelligence. Future Generation Computer Systems, 1(2):135–144, November 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:E

[Ano84f] Anonymous. Editorial. Future Generation Computer Systems, 1(1):1, July 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:EB

[Ano84g] Anonymous. Editorial Board.

Future Generation Computer Systems, 1(1):??, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:EEC

[Ano84h] Anonymous. Esprit: Europe challenges U.S. and Japanese competitors. Future Generation Computer Systems, 1 (1):61–68, July 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:LCO

[Ano84i] Anonymous. Legislation of co-operative R&D efforts in the U.S.A. Future Generation Computer Systems, 1 (1):57–60, July 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:NDD

[Ano84j] Anonymous. New directions for database systems. Future Generation Computer Systems, 1(1):85–88, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1984:RKL

[Ano84k] Anonymous. Robert Kowalski on logic programming: Paris, 21 September 1983. Future Generation Computer Systems, 1(1):79–83, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1985:AI

[Ano85a] Anonymous. AFCET-Informatique ’85. Future Generation Computer Systems, 1(5):377– 381, September 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1985:AIV

[Ano85b] Anonymous. Author index to volume 1. Future Generation Computer Systems, 1 (6):423, December 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1985:Ca

[Ano85c] Anonymous. Calendar. Future Generation Computer Systems, 1(3):195–200, February 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1985:Cb

[Ano85d] Anonymous. Calendar. Future Generation Computer Systems, 1(4):249–254, June 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1985:Cc

[Ano85e] Anonymous. Calendar. Future Generation Computer Systems, 1(5):383–386, September 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1985:Cd

[Ano85f] Anonymous. Calendar. Future Generation Computer Systems, 1(6):417–421, December 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1985:NLU

[Ano85g] Anonymous. Natural language understanding and logic programming. Future Generation Computer Systems, 1(3): 189–193, February 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:AIV

[Ano86a] Anonymous. Author index to volume 2. Future Generation Computer Systems, 2 (4):269, December 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:Ca

[Ano86b] Anonymous. Calendar. Future Generation Computer Systems, 2(1):69–76, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:Cb

[Ano86c] Anonymous. Calendar. Future Generation Computer Systems, 2(2):145–150, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:Cc

[Ano86d] Anonymous. Calendar. Future Generation Computer Systems, 2(3):205–209, September 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:Cd

[Ano86e] Anonymous. Calendar. Future Generation Computer Systems, 2(4):265–268, December 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:EB

[Ano86f] Anonymous. Editorial Board.

Future Generation Computer Systems, 2(1):i–ii, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:EI

[Ano86g] Anonymous. Editorial introduction. Future Generation Computer Systems, 2(3): 151, September 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:EEW

[Ano86h] Anonymous. EWSL 87 2nd European Working Session on Learning. Future Generation Computer Systems, 2(3):211– 212, September 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:FGC

[Ano86i] Anonymous. Fifth generation computers, concepts, implementations and uses: Peter Bishop. Published by: Ellis Horwood (Series Computer and their Applications). Price:

£7.90. Number of pages: 166. Reviewer: R. P. van de Riet, Free University, Amsterdam. Future Generation Computer Systems, 2(3):201– 203, September 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:ICI

[Ano86j]  Anonymous.  International

conference on information processing and management of uncertainty in knowledgebased systems. Future Generation Computer Systems, 2 (4):245–259, December 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:I

[Ano86k] Anonymous. Introduction. Future Generation Computer Systems, 2(1):5–11, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:LAI

[Ano86l] Anonymous. Language and artificial intelligence. Future Generation Computer Systems, 2(2):141–143, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1986:VP

[Ano86m] Anonymous. The VHSIC program. Future Generation Computer Systems, 2(3):187– 190, September 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:A

[Ano87a]  Anonymous.  Announce-

ment. Future Generation Computer Systems, 3(3):231– 232, September 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:AIA

[Ano87b] Anonymous. Artificial intelligence, applications in the future of software engineering: Derek Partridge. Publisher: Ellis Horwood Ltd. (1986). Price: £25.00 Reviewer: R. P. van de Riet, Free University Amsterdam. Future Generation Computer Systems, 3(1):51–52, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:AIS

[Ano87c] Anonymous. Artificial intelligence for society: Karamjit S. Gill Publisher: John Wiley & Sons (1986). Price: £19.95 Reviewer: R. P. van de Riet, Free University Amsterdam. Future Generation Computer Systems, 3(1):53–57, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:AIV

[Ano87d] Anonymous. Author index to volume 3. Future Generation Computer Systems, 3 (4):311, December 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:Ca

[Ano87e] Anonymous. Calendar. Future Generation Computer Systems, 3(1):75–82, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:Cb

[Ano87f] Anonymous. Calendar. Future Generation Computer Systems, 3(2):155–159, May 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:Cc

[Ano87g] Anonymous. Calendar. Future Generation Computer Systems, 3(3):227–230, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:Cd

[Ano87h] Anonymous. Calendar. Future Generation Computer Systems, 3(4):307–310, December 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:EB

[Ano87i] Anonymous. Editorial Board.

Future Generation Computer Systems, 3(1):ii, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:ERA

[Ano87j] Anonymous. ESPRIT ’86 — results and achievements. Future Generation Computer Systems, 3(1):59–73, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:LE

[Ano87k] Anonymous. Letter to the editor. Future Generation Computer Systems, 3(2):153, May 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:MID

[Ano87l] Anonymous. On machine intelligence: Donald Michie Publisher: Ellis Horwood Ltd. (1986). Price: £29.95. Reviewer: R. P. van de Riet, Free University Amsterdam. Future Generation Computer Systems, 3(1):52–53, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1987:RST

[Ano87m] Anonymous. Results of survey on trends in expert systems in Japan. Future Generation Computer Systems, 3(1):

17–36, February 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1988:Ca

[Ano88a] Anonymous. Calendar. Future Generation Computer Systems, 4(1):77–79, August 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1988:Cb

[Ano88b] Anonymous. Calendar. Future Generation Computer Systems, 4(2):157–158, September 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1988:EB

[Ano88c] Anonymous. Editorial Board. Future Generation Computer Systems, 4(1):i, August 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1989:AIV

[Ano89a] Anonymous. Author index to volume 4. Future Generation Computer Systems, 4(4): 333–334, March 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1989:Ca

[Ano89b] Anonymous. Calendar. Future Generation Computer Systems, 4(4):327–332, March 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1989:Cb

[Ano89c] Anonymous. Calendar. Future Generation Computer Systems, 5(1):163–166, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1989:Cc

[Ano89d]  Anonymous.  Calendar.

Future Generation Computer Systems, 5(2–3):347– 350, September 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1989:EB

[Ano89e] Anonymous. Editorial Board.

Future Generation Computer Systems, 5(1):CO2, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1989:P

[Ano89f]  Anonymous.  From the

publisher. Future Generation Computer Systems, 5 (1):1, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1990:AIV

[Ano90a] Anonymous. Author index to volume 5. Future Generation Computer Systems, 5(4): 399–400, January 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1990:Ca

[Ano90b] Anonymous. Calendar. Future Generation Computer Systems, 5(4):395–398, January 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1990:Cb

[Ano90c] Anonymous. Calendar. Future Generation Computer Systems, 6(1):105–109, June 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1990:Cc

[Ano90d] Anonymous. Calendar. Future Generation Computer Systems, 6(2):181–182, November 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1990:Cd

[Ano90e] Anonymous. Calendar. Future Generation Computer Systems, 6(3):301–302, December 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1990:EB

[Ano90f] Anonymous. Editorial Board. Future Generation Computer Systems, 6(1):CO2, June 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1990:PC

[Ano90g] Anonymous. Parallel computing. Future Generation Computer Systems, 6(3): 183, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1991:AIV

[Ano91a] Anonymous. Author index to volume 6 (1990/91). Future Generation Computer Systems, 6(4):389–390, September 1991. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1991:Ca

[Ano91b] Anonymous. Calendar. Future Generation Computer Systems, 6(4):391–392, September 1991. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1991:Cb

[Ano91c] Anonymous. Calendar. Future Generation Computer Systems, 7(1):109–110, October 1991. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1991:EB

[Ano91d] Anonymous. Editorial Board.

Future Generation Computer Systems, 7(1):CO2, October 1991. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:AIVa

[Ano92a]  Anonymous.  Author in-

dex to volume 7 (1991/92). Future Generation Computer Systems, 7(4):465–467, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:AIVb

[Ano92b] Anonymous. Author index to volume 8 (1992). Future Generation Computer Systems, 8 (4):437–438, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:Ca

[Ano92c] Anonymous. Calendar. Future Generation Computer Systems, 7(2–3):339–340, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:Cb

[Ano92d] Anonymous. Calendar. Future Generation Computer Systems, 7(4):473–474, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:Cc

[Ano92e] Anonymous. Calendar. Future Generation Computer Systems, 8(1–3):263–266, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:Cd

[Ano92f] Anonymous. Calendar. Future Generation Computer Systems, 8(4):441–442, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:EB

[Ano92g] Anonymous. Editorial Board.

Future Generation Computer Systems, 8(1–3):CO2, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:SIVa

[Ano92h]  Anonymous.  Subject in-

dex of volume 7 (1991/92). Future Generation Computer Systems, 7(4):469–471, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1992:SIVb

[Ano92i] Anonymous. Subject index to volume 8 (1992). Future Generation Computer Systems, 8 (4):439–440, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:AIV

[Ano93a] Anonymous. Author index to volume 9 (1993). Future Generation Computer Systems, 9 (4):381–382, December 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:BAb

[Ano93b]  Anonymous.  Book an-

nouncement. Future Generation Computer Systems, 9(4): 379, December 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:BAa

[Ano93c] Anonymous. Book announcements. Future Generation Computer Systems, 9 (1):73–74, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:Ca

[Ano93d] Anonymous. Calendar. Future Generation Computer Systems, 9(1):75–78, May 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:Cb

[Ano93e] Anonymous. Calendar. Future Generation Computer Systems, 9(2):159–161, July 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:Cc

[Ano93f] Anonymous. Calendar. Future Generation Computer Systems, 9(3):281–283, September 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:Cd

[Ano93g] Anonymous. Calendar. Future Generation Computer Systems, 9(4):385–387, December 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:EB

[Ano93h] Anonymous. Editorial Board.

Future Generation Computer Systems, 9(1):CO2, May 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1993:SIV

[Ano93i] Anonymous. Subject index to volume 9 (1993). Future Generation Computer Systems, 9 (4):383–384, December 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1994:AIVa

[Ano94a] Anonymous. Author index to volume 10 (1994). Future Generation Computer Systems, 10 (4):443–445, November 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1994:AIVb

[Ano94b]  Anonymous.  Author in-

dex to volumes 1–10 (1984– 1994). Future Generation Computer Systems, 10(4): 451–464, November 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1994:Ca

[Ano94c] Anonymous. Calendar. Future Generation Computer Systems, 10(2–3):363–364, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1994:Cb

[Ano94d] Anonymous. Calendar. Future Generation Computer Systems, 10(4):441–442, November 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1994:CFC

[Ano94e]  Anonymous.  Calendar of

forthcoming conferences. Future Generation Computer Systems, 10(1):149–150, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1994:EB

[Ano94f] Anonymous. Editorial Board. Future Generation Computer Systems, 10(1):ii, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1994:SIV

[Ano94g] Anonymous. Subject index to volume 10 (1994). Future Generation Computer Systems, 10 (4):447–449, November 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:AIV

[Ano95a] Anonymous. Author index to volume 11 (1995). Future Generation Computer Systems, 11 (6):631–633, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:Ca

[Ano95b] Anonymous. Calendar. Future Generation Computer Systems, 11(1):111–112, February 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:Cc

[Ano95c] Anonymous. Calendar. Future Generation Computer Systems, 11(3):345–346, June 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:Cd

[Ano95d]  Anonymous.  Calendar99.

Future Generation Computer Systems, 11(4–5):499– 500, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:Cb

[Ano95e]  Anonymous.  Calendar-

neous. Future Generation Computer Systems, 11(2): 245–246, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:Ce

[Ano95f] Anonymous. Calendarneous. Future Generation Computer Systems, 11(6):639–640, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:EB

[Ano95g] Anonymous. Editorial Board. Future Generation Computer Systems, 11(1):ii, February 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1995:SIV

[Ano95h] Anonymous. Subject index to volume 11 (1995). Future Generation Computer Systems, 11 (6):635–637, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1996:CCEa

[Ano96a] Anonymous. Calendar of conferences and events. Future Generation Computer Systems, 12(2–3):257–261, September 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1996:CCEb

[Ano96b] Anonymous. Calendar of conferences and eventsensional compressible magnetohydrodynamic simulation. Future Generation Computer Systems, 12(4):325–330, December 1996. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1996:C

[Ano96c] Anonymous. Calendar17. Future Generation Computer Systems, 12(1):117–118, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1996:EB

[Ano96d] Anonymous. Editorial Board.

Future Generation Computer Systems, 12(1):CO2, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1997:Ca

[Ano97a] Anonymous. Calendar. Future Generation Computer Systems, 12(5):475–479, April 1, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1997:Cb

[Ano97b] Anonymous. Calendar. Future Generation Computer Systems, 13(1):91–94, June 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/17/21/abstract.
html.

Anonymous:1997:EB

[Ano97c] Anonymous. Editorial Board.

Future Generation Computer Systems, 13(1):CO2, June 1997. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/17/21/abstract.
html.

Anonymous:1997:MIa

[Ano97d] Anonymous. Miscellaneous item. Future Generation Computer Systems, 12(6):579– 580, June 15, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1997:MIb

[Ano97e] Anonymous. Miscellaneous item. Future Generation Computer Systems, 12(6): 581, June 15, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1998:AIV

[Ano98a] Anonymous. Author index to volume 14 (1998). Future Generation Computer Systems, 14(5–6):453–455, December 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1998:C

[Ano98b] Anonymous. Calendar. Future Generation Computer Systems, 14(5–6):459–462, December 1998. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1998:SIV

[Ano98c] Anonymous. Subject index to volume 14 (1998). Future Generation Computer Systems, 14(5–6):457–458, December 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:1999:E

[Ano99]  Anonymous.  Editorial(s).

Future Generation Computer Systems, 16(1):vii– viii, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/25/abstract.
html.

Anonymous:2000:Ia

[Ano00a] Anonymous. Index. Future Generation Computer Systems, 16(8):947–951, June 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.elsevier.com/gej-ng/
10/19/19/41/31/32/abstract.
html.

Anonymous:2000:Ib

[Ano00b] Anonymous. Index. Future Generation Computer Systems, 16(8):953–955, June 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.elsevier.com/gej-ng/
10/19/19/41/31/33/abstract.
html.

Anonymous:2001:AI

[Ano01a]  Anonymous.  Author in-

dex. Future Generation Computer Systems, 17(8):1059– 1064, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/41/abstract.
html.

Anonymous:2001:C

[Ano01b] Anonymous. Contents. Future Generation Computer Systems, 17(5):iii, March 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/24/abstract.
html.

Anonymous:2001:GE

[Ano01c]  Anonymous.  Guest edi-

torial. Future Generation Computer Systems, 18(2): v–vi, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/27/abstract.
html.

Anonymous:2001:SI

[Ano01d]  Anonymous.  Subject in-

dex. Future Generation Computer Systems, 17(8):1065– 1068, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
19/19/45/35/42/abstract.
html.

Anonymous:2002:AI

[Ano02a]  Anonymous.  Author in-

dex. Future Generation Computer Systems, 18(8):1155– 1161, October 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2002:IE

[Ano02b] Anonymous. IFC editors.

Future Generation Computer Systems, 18(8):CO2, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2002:SI

[Ano02c]  Anonymous.  Subject in-

dex. Future Generation Computer Systems, 18(8):1163– 1166, October 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:AIV

[Ano03a] Anonymous. Author index to volume 19(2003). Future Generation Computer Systems, 19(8):1375–1383, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:Ea

[Ano03b] Anonymous. Editors. Future Generation Computer Systems, 19(1):CO2, January 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:Eb

[Ano03c] Anonymous. Editors. Future Generation Computer Systems, 19(2):CO2, February 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:Ec

[Ano03d] Anonymous. Editors. Future Generation Computer Systems, 19(3):CO2, April 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:Ed

[Ano03e] Anonymous. Editors. Future Generation Computer Systems, 19(6):CO2, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:Ee

[Ano03f] Anonymous. Editors. Future Generation Computer Systems, 19(7):CO2, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:Ef

[Ano03g] Anonymous. Editors. Future Generation Computer Systems, 19(8):CO2, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:IE

[Ano03h] Anonymous. IFC editors.

Future Generation Computer Systems, 19(4):CO2, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2003:SIV

[Ano03i] Anonymous. Subject index to volume 19(2003). Future Generation Computer Systems, 19(8):1385–1389, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2004:AI

[Ano04a] Anonymous. Author index.

Future Generation Computer Systems, 20(8):??, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2004:SI

[Ano04b] Anonymous. Subject index.

Future Generation Computer Systems, 20(8):??, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2005:Ca

[Ano05a] Anonymous. Contents. Future Generation Computer Systems, 21(6):iii, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2005:Cb

[Ano05b] Anonymous. Contents. Future Generation Computer Systems, 21(7):iii–v, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2005:EBa

[Ano05c] Anonymous. Editorial Board. Future Generation Computer Systems, 21(6):CO2, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2005:EBb

[Ano05d] Anonymous. Editorial Board. Future Generation Computer Systems, 21(7):CO2, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2005:EA

[Ano05e] Anonymous. EnterTheGrid: Advertisement. Future Generation Computer Systems, 21(4):EX2–EX3, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2005:SIV

[Ano05f] Anonymous. Subject index to volume 21. Future Generation Computer Systems, 21 (8):1439–1451, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2005:R

[Ano05g] Anonymous. To the readers.

Future Generation Computer Systems, 21(7):987, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2006:PN

[Ano06] Anonymous. Publisher’s note.

Future Generation Computer Systems, 22(1–2):34, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2007:AR

[Ano07] Anonymous. Acknowledgement of reviewers over 2006. Future Generation Computer Systems, 23(6):735–736, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anonymous:2008:AR

[Ano08] Anonymous. Acknowledgement of reviewers over 2007. Future Generation Computer Systems, 24(5):329–330, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Aahlander:2006:SDF

[ÅO06] Krister Åhlander and Kurt Otto. Software design for finite difference schemes based on index notation. Future Generation Computer Systems, 22(1–2):102–109, January 2006. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Arcos:1996:IRO

[AP96] Josep Lluís Arcos and Enric Plaza. Inference and reflection in the objectcentered representation language NOOS. Future Generation Computer Systems, 12(2– 3):173–188, September 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Araujo:1998:PEP

[AR98] Lourdes Araujo and Jose J. Ruz. Parallel execution of Prolog with granularity control. Future Generation Computer Systems, 13(6):421– 441, May 20, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/20/17/abstract.
html.

Andrzejak:2007:SSP

[AR07] Artur Andrzejak and Alexander Reinefeld. Special section: Paradigms for scalable and dependable grids. Future Generation Computer Systems, 23(7):861–863, August 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Almond:1999:UUA

[AS99] Jim Almond and Dave Snelling. UNICORE: uniform access to supercomputing as an element of electronic commerce. Future Generation Computer Systems, 15(5–6): 539–548, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/17/abstract.
html.

Alekseev:2002:MTA

[AS02] Igor V. Alekseev and Valery A.

Sokolov. Modeling and traffic analysis of the adaptive rate transport protocol. Future Generation Computer Systems, 18(6):813–827, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Atchley:2003:VI

[ASPB03] Scott Atchley, Stephen Soltesz, James S. Plank, and Micah Beck. Video IBPster. Future Generation Computer Systems, 19(6):861–870, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ala-Siuru:1998:HVE

[ASTEP98] Pekka Ala-Siuru, Juha Takalo, Jari Ensomaa, and Johan Plomp. A hierarchical virtual environment for a machine fault diagnostic application. Future Generation Computer Systems, 14(3–4): 179–183, July 31, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/18/20/abstract.
html.

Alba:2001:ASA

[AT01] Enrique Alba and José M. Troya. Analyzing synchronous and asynchronous parallel distributed genetic algorithms. Future Generation Computer Systems, 17 (4):451–465, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/35/abstract.
html.

Aloisio:2002:GCT

[AT02] Giovanni Aloisio and Domenico Talia. Grid computing: Towards a new computing infrastructure. Future Generation Computer Systems, 18 (8):v–vi, October 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Al-Theneyan:2002:XBV

[ATJMZ02] Ahmed Al-Theneyan, Amol Jakatdar, Piyush Mehrotra, and Mohammad Zubair. XML-based visual specification of multidisciplinary applications. Future Generation Computer Systems, 18(4): 539–548, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
19/19/60/33/35/abstract.
html.

Allis:1996:MLS

[ATT96] Victor Allis, Yao-Hua Tan, and Jan Treur. Meta-level selection techniques for the control of default reasoning. Future Generation Computer Systems, 12(2–3):189– 201, September 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Amon:2000:VEL

[AV00] Tomaz Amon and Vojko Valencic. VRML-enhanced learning in biology and medicine. Future Generation Computer Systems, 17 (1):1–6, September 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/45/24/25/41/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/45/24/25/abstract.
html.

Avgoustinov:2000:VME

[Avg00] Nikolay Avgoustinov. VRML as means of expressive 4D illustration in CAM education. Future Generation Computer Systems, 17(1):39– 48, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        19/19/45/24/29/abstract.
                                        html.

Aarts:1992:PC

[AvLR92] Emile Aarts, Jan van Leeuwen, and Martin Rem. Parallel computing. Future Generation Computer Systems, 8(4):267– 268, September 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Apte:1997:DMD

[AW97] Chidanand Apté and Sholom M. Weiss. Data mining with decision trees and decision rules. Future Generation Computer Systems, 13(2–3): 197–210, November 14, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/18/23/abstract.
                                        html.

Abramson:2003:DSA

[AW03] David Abramson and Greg Watson. Debugging scientific applications in the .NET Framework. Future Generation Computer Systems, 19 (5):665–678, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baayen:1987:PPP

[Baa87] P. C. Baayen. Policy, program and perspectives of CWI, Amsterdam. Future Generation Computer Systems, 3(4):271– 277, December 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Berket:2002:PAI

[BAC02] K. Berket, D. A. Agarwal, and O. Chevassut. A practical approach to the Inter-Group protocols. Future Generation Computer Systems, 18 (5):709–719, April ??, 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/38/abstract.
html.

Bondarescu:2005:ASC

[BAD^{+}05] Ruxandra Bondarescu, Gabrielle Allen, Gregory Daues, Ian Kelley, Michael Russell, Edward Seidel, John Shalf, and Malcolm Tobias. The Astrophysics Simulation Collaboratory Portal: a framework for effective distributed research. Future Generation Computer Systems, 21(2): 259–270, February 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bal:1991:HSP

[Bal91a]  Henri E. Bal.  Heuristic

search in PARLOG using replicated worker style parallelism. Future Generation Computer Systems, 6(4):303– 315, September 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baldi:1991:MT

[Bal91b]  L. Baldi.  Microelectron-

ics trends. Future Generation Computer Systems, 7(1): 3–13, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bal:1992:CSF

[Bal92] Henri E. Bal. A comparative study of five parallel programming languages. Future Generation Computer Systems, 8 (1–3):121–135, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bal:1993:EKI

[Bal93] Henri E. Bal. Evaluation of KL1 and the inference machine. Future Generation Computer Systems, 9(2): 119–125, July 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bandini:2002:CA

[Ban02a]  S. Bandini.  Cellular au-

tomata. Future Generation Computer Systems, 18 (7):v–vi, August 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bandman:2002:CNA

[Ban02b] Olga Bandman. Cellular– neural automaton: a hybrid model for reaction–diffusion simulation. Future Generation Computer Systems, 18 (6):737–745, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bandman:2005:CPS

[Ban05] Olga L. Bandman. Computation properties of spatial dynamics simulation by probabilistic cellular automata. Future Generation Computer Systems, 21(5):633–643, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bibel:1984:TCM

[BB84] Wolfgang Bibel and Bruno Buchberger. Towards a connection machine for logical inference. Future Generation Computer Systems, ?? (??):177–188, ???? 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). Also: Technical Report 84-19, RISC-Linz, Johannes Kepler University, Linz, Austria, 1984.

Bibel:1985:TCM

[BB85] W. Bibel and B. Buchberger.

Toward a Connection Machine for Logical Inference. Future Generation Computer Systems, 1(3):177–188, February 1985. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Boojhawon:2004:RSG

[BB04] Ravindra Boojhawon and Muddun Bhuruth. Restarted Simpler GMRES augmented with harmonic Ritz vectors. Future Generation Computer Systems, 20(3):389–397, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Berrios:2006:UWF

[BB06] Joseph S. Berrios and Manuel E. Bermudez. Using wait-free synchronization in the design of distributed applications. Future Generation Computer Systems, 22(1–2):46–56, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Barisone:2001:JSM

[BBBD01] A. Barisone, F. Bellotti, R. Berta, and A. De Gloria. JSBricks: a suite of microbenchmarks for the evaluation of Java as a scientific execution environment. Future Generation Computer Systems, 18(2): 293–306, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/31/35/abstract.
                                        html.

Beccaria:1999:HPR

[BBC^{+}99] Matteo Beccaria, Guido Buresti, Alberto Ciampa, Giovanni Lombardi, Wolfgang Gentzsch, Hans-Georg Paap, and Andrea Viceré. Highperformance road-vehicle optimised aerodynamic design: Application of parallel computing to car design. Future Generation Computer Systems, 15(3):323– 332, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/18/abstract.
html.

Barberet:1999:TTW

[BBD^{+}99] Christophe Barberet, Lionel Brunie, Frédéric Desprez, Gilles Lebourgeois, Raymond Namyst, Yves Robert, Stéphane Ubéda, and Karine Van Heumen. Technology transfer within the ProHPC TTN at ENS Lyon. Future Generation Computer Systems, 15(3): 309–321, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/17/abstract.
html.

Balis:2003:MSM

[BBFW03] Bartosz Baliś, Marian Bubak, Włodzimierz Funika, and Roland Wismüller. A monitoring system for multithreaded applications. Future Generation Computer Systems, 19(5):641–650, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bassi:2005:ALN

[BBG^{+}05] Alessandro Bassi, Micah Beck, Jean-Patrick Gelas, Laurent Lefèvre, Terry Moore, James Plank, and Pascale Vicat-Blanc Primet. Active and logistical networking for grid computing: the e-Toile architecture. Future Generation Computer Systems, 21 (1):199–208, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Blaheta:2006:LSP

[BBJ^{+}06] R. Blaheta, P. Byczanski, O. Jakl, R. Kohut, A. Kolcun, K. Krečmer, and J. Starý. Large scale parallel FEM computations of far/near stress field changes in rocks. Future Generation Computer Systems, 22(4):449–459, March 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bischof:2005:EAD

[BBL^{+}05] C. H. Bischof, H. M. Bücker, B. Lang, A. Rasch, and E. Slusanschi. Efficient and accurate derivatives for a software process chain in airfoil shape optimization. Future Generation Computer Systems, 21 (8):1333–1344, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bassi:2005:EGC

[BBLP05] Alessandro Bassi, Micah Beck, Julien Laganier, and Gabriella Paolini. Enhancing grid capabilities: IBP over IPv6. Future Generation Computer Systems, 21(2):303–313, February 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bassi:2003:IBP

[BBM^{+}03] Alessandro Bassi, Micah Beck, Terry Moore, James S. Plank, Martin Swany, Rich Wolski, and Graham Fagg. The Internet Backplane Protocol: a study in resource sharing. Future Generation Computer Systems, 19(4):551–561, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bernaschi:1992:DSV

[BBSV92] M. Bernaschi, M. Blount, P. Sguazzero, and M. Vitaletti. Distributed shared virtual memory on RISC System/6000 clusters and large scale computations: Two case studies. Future Generation Computer Systems, 8(1–3): 235–242, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Balis:2008:LFF

[BBW08] Bartosz Baliś, Marian Bubak, and Michal Wegiel. LGF: a flexible framework for exposing legacy codes as services. Future Generation Computer Systems, 24(7):711–719, July 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Blanes:2003:OLG

[BC03] S. Blanes and F. Casas. Optimization of Lie group methods for differential equations. Future Generation Computer Systems, 19(3):331–339, April 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Byun:2007:MMJ

[BCB^{+}07] EunJoung Byun, SungJin Choi, MaengSoon Baik, Joon-Min Gil, ChanYeol Park, and ChongSun Hwang. MJSA: Markov job scheduler based on availability in desktop grid computing environment. Future Generation Computer Systems, 23(4):616–622, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baude:2002:ORM

[BCFS02] Françoise Baude, Denis Caromel, Nathalie Furmento, and David Sagnol. Optimizing remote method invocation with communication–computation overlap. Future Generation Computer Systems, 18 (6):769–778, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Broker:2005:UPL

[BCG05] Oliver Bröker, Oscar Chinellato, and Roman Geus. Using Python for large scale linear algebra applications. Future Generation Computer Systems, 21(6):969–979, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Buntinas:2008:BVN

[BCH^{+}08] Darius Buntinas, Camille Coti, Thomas Herault, Pierre Lemarinier, Laurence Pilard, Ala Rezmerita, Eric Rodriguez, and Franck Cappello. Blocking vs. non-blocking coordinated checkpointing for large-scale fault tolerant MPI protocols. Future Generation Computer Systems, 24(1): 73–84, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Banatre:1988:PMM

[BCL88] Jean-Pierre Banâtre, Anne Coutant, and Daniel Le Metayer. A parallel machine for multiset transformation and its programming style. Future Generation Computer Systems, 4(2):133– 144, September 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bruce:1995:CPS

[BCM^{+}95] R. A. A. Bruce, S. Chapple, N. B. MacDonald, A. S. Trew, and S. Trewin. CHIMP and PUL: Support for portable parallel computing. Future Generation Computer Systems, 11(2):211–219, March 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Budura:2007:BWP

[BCMA07] Adriana Budura, Philippe Cudré-Mauroux, and Karl Aberer. From bioinformatic Web portals to semantically integrated Data Grid networks. Future Generation Computer Systems, 23(3): 485–496, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baiardi:2001:ILB

[BCMR01] Fabrizio Baiardi, Sarah Chiti, Paolo Mori, and Laura Ricci. Integrating load balancing and locality in the parallelization of irregular problems. Future Generation Computer Systems, 17(8): 969–975, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/33/abstract.
html.

Bouras:2003:QSA

[BCPS03] Christos Bouras, Mauro Campanella, Michal Przybylski, and Afrodite Sevasti. QoS and SLA aspects across multiple management domains: the SEQUIN approach. Future Generation Computer Systems, 19(2):313–326, February 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bernaschi:1999:HPS

[BCS99] M. Bernaschi, F. Castiglione, and S. Succi. A high performance simulator of the immune response. Future Generation Computer Systems, 15 (3):333–342, April 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/19/abstract.
html.

Bellotti:2007:DMI

[BCT^{+}07] R. Bellotti, P. Cerello, S. Tangaro, V. Bevilacqua, M. Castellano, G. Mastronardi, F. De Carlo, S. Bagnasco, U. Bottigli, R. Cataldo, E. Catanzariti, S. C. Cheran, P. Delogu, I. De Mitri, G. De Nunzio, M. E. Fantacci, F. Fauci, G. Gargano, B. Golosio, and P. L. Indovina. Distributed medical images analysis on a Grid infrastructure. Future Generation Computer Systems, 23(3):475–484, March 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brezany:2001:GIP

[BCW01] Peter Brezany, Przemysław Czerwiński, and Marianne Winslett. A generic interface for parallel access to large data sets from HPF applications. Future Generation Computer Systems, 17 (8):977–985, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/35/34/abstract.
                                        html.

Barbosa:2005:GMM

[BdCYG05] Jorge Luis Victória Barbosa, Cristiano André da Costa, Adenauer Corrêa Yamin, and Cláudio Fernando Resin Geyer. GHolo: a multiparadigm model oriented to development of grid systems. Future Generation Computer Systems, 21(1): 227–237, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Beck:1999:HNG

[BDF^{+}99] Micah Beck, Jack J. Dongarra, Graham E. Fagg, G. Al Geist, Paul Gray, James Kohl, Mauro Migliardi, Keith Moore, Terry Moore, Philip Papadopoulous, Stephen L. Scott, and Vaidy Sunderam. HARNESS: a next generation distributed virtual machine. Future Generation Computer Systems, 15(5–6): 571–582, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/20/abstract.
html;   http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/harness2.ps.

Bodei:2005:APS

[BDFP05] C. Bodei, P. Degano, R. Focardi, and C. Priami. Authentication primitives for secure protocol specifications. Future Generation Computer Systems, 21(5):645–653, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bradley:2006:DCT

[BDHK06] Jeremy T. Bradley, Nicholas J.

Dingle, Peter G. Harrison, and William J. Knottenbelt. Distributed computation of transient state distributions and passage time quantiles in large semi-Markov models. Future Generation Computer Systems, 22(7):828–837, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Boukerche:2006:AAM

[BDL06] Azzedine Boukerche, Caron Dzermajko, and Kaiyuan Lu. Alternative approaches to multicast group management in large-scale distributed interactive simulation systems. Future Generation Computer Systems, 22(7):755–763, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bodei:2002:FLD

[BDNN02] C. Bodei, P. Degano, F. Nielson, and H. Riis Nielson. Flow logic for Dolev–Yao secrecy in cryptographic processes. Future Generation Computer Systems, 18(6):747–756, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bucci:1992:LCW

[BDNP92] G. Bucci, R. Detti, S. Nativi, and V. Pasqui. Loosely coupled workstations in a radiological image information system. Future Generation Computer Systems, 8(1–3):31–42, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Beazley:2003:ASS

[Bea03] D. M. Beazley. Automated scientific software scripting with SWIG. Future Generation Computer Systems, 19 (5):599–609, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Benkner:1999:HHP

[Ben99] Siegfried Benkner. HPF+: High Performance Fortran for advanced scientific and engineering applications. Future Generation Computer Systems, 15(3):381–391, April 1, 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/23/abstract.
html.

Bernaschi:1996:RHP

[Ber96] Massimo Bernaschi. The requirements of a high performance implementation of PVM. Future Generation Computer Systems, 12(1): 3–11, May 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bernaschi:1998:EMP

[Ber98]  Massimo Bernaschi.  Ef-

ficient message passing on UNIX shared memory multiprocessors. Future Generation Computer Systems, 13(6): 443–449, May 20, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/20/18/abstract.
html.

Bertelsen:2000:DSJ

[Ber00] Peter Bertelsen. Dynamic semantics of Java bytecode. Future Generation Computer Systems, 16(7): 841–850, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/30/32/abstract.
                                        html.

Berti:2006:GGA

[Ber06] Guntram Berti. GrAL — the Grid Algorithms Library. Future Generation Computer Systems, 22(1–2): 110–122, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bosilca:2002:OOE

[BFC02] George Bosilca, Gilles Fedak, and Franck Cappello. OVM: Out-of-order execution parallel virtual machine. Future Generation Computer Systems, 18(4):525–537, March 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.elsevier.com/gej-ng/
                                        10/19/19/60/33/34/abstract.
                                        html.

Broom:2002:KTR

[BFK02] Bradley Broom, Rob Fowler, and Ken Kennedy. KelpIO: a telescope-ready domainspecific I/O library for irregular block-structured applications. Future Generation Computer Systems, 18(4): 449–460, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

19/19/60/33/29/abstract.
html.

Baraglia:1999:WFG

[BFL99] R. Baraglia, R. Ferrini, and D. Laforenza. WAMM in the framework of graphical user interfaces for metacomputing management. Future Generation Computer Systems, 15(5– 6):687–698, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/29/abstract.
html.

Baraglia:1999:OTF

[BFLL99] R. Baraglia, R. Ferrini, D. Laforenza, and A. Laganà. An optimized task-farm model to integrate reduced dimensionality Schrödinger equations on distributed memory architectures. Future Generation Computer Systems, 15(4): 497–512, July 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/20/20/abstract.
html.

Brune:1999:MPE

[BFR99] Matthias A. Brune, Graham E. Fagg, and Michael M. Resch. Message-passing environments for metacomputing. Future Generation Computer Systems, 15(5–6):699– 712, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/21/30/abstract.
                                        html.

Bessonov:2005:DEC

[BFR05] O. Bessonov, D. Fougère, and B. Roux. Development of efficient computational kernels and linear algebra routines for out-of-order superscalar processors. Future Generation Computer Systems, 21 (5):743–748, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bubak:2003:MDJ

[BFW^{+}03] Marian Bubak, Włodzimierz Funika, Roland Wismüller, Piotr Mȩtel, and Rafał Orłowski. Monitoring of distributed Java applications. Future Generation Computer Systems, 19 (5):651–663, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Beer:1987:PPO

[BG87] J. Beer and W. K. Giloi. POPE — a Parallel-Operating Prolog Engine. Future Generation Computer Systems, 3 (2):83–92, May 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bischof:2005:COP

[BG05] Holger Bischof and Sergei Gorlatch. A cost-optimal parallel implementation of a tridiagonal system solver using skeletons. Future Generation Computer Systems, 21 (5):737–742, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Blanco:2003:VPP

[BGC^{+}03] V. Blanco, P. González, J. C.

Cabaleiro, D. B. Heras, T. F. Pena, J. J. Pombo, and F. F. Rivera. AVISPA: visualizing the performance prediction of parallel iterative solvers. Future Generation Computer Systems, 19(5):721–733, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Belloum:2003:VGG

[BGH^{+}03] Adam S. Z. Belloum, David L.

Groep, Zeger W. Hendrikse, Bob L. O. Hertzberger, Vladimir Korkhov, Cees T. A. M. de Laat, and Dmitry Vasunin. VLAM-G: a grid-based virtual laboratory. Future Generation Computer Systems, 19(2): 209–217, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bardeen:2006:QGC

[BGJ^{+}06] Marjorie Bardeen, Eric Gilbert, Thomas Jordan, Paul Nepywoda, Elizabeth Quigg, Mike Wilde, and Yong Zhao. The QuarkNet/Grid Collaborative Learning e-Lab. Future Generation Computer Systems, 22 (6):700–708, May 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bubak:2005:WCS

[BGK^{+}05] Marian Bubak, Tomasz Gubała, Michał Kapałka, Maciej Malawski, and Katarzyna Rycerz. Workflow composer and service registry for grid applications. Future Generation Computer Systems, 21(1):79–86, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bouhafs:2005:DEA

[BGL^{+}05] F. Bouhafs, J. P. Gelas, L. Lefèvre, M. Maimour, C. Pham, P. Vicat-Blanc Primet, and B. Tourancheau. Designing and evaluating an active grid architecture. Future Generation Computer Systems, 21(2):315–330, February 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Badia:2008:SSS

[BGL08] Rosa M. Badia, Dennis Gannon, and Craig Lee. Special section: Selected papers from the 7th IEEE/ACM international conference on grid computing (Grid2006). Future Generation Computer Systems, 24(5):402–403, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bywater:1999:MPD

[BGR^{+}99] Robert Bywater, Jörn Gehring, Alexander Reinefeld, Friedrich Rippmann, and Anke Weber. Metacomputing in practice: a distributed compute server for pharmaceutical industry. Future Generation Computer Systems, 15(5–6): 769–785, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/36/abstract.
html.

Biget:1997:HSC

[BGV97] Patrick Biget, Patrick George, and Jean-Jacques Vandewalle. How smart cards can benefit from object-oriented technologies. Future Generation Computer Systems, 13 (1):75–90, June 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/17/21/abstract.
html.

Baumann:2003:SVD

[BH03] Markus Baumann and Uwe Helmke. Singular value decomposition of time-varying matrices. Future Generation Computer Systems, 19(3): 353–361, April 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Blanke:2009:AHS

[BHD09] Tobias Blanke, Mark Hedges, and Stuart Dunn. Arts and humanities e-science — current practices and future challenges. Future Generation Computer Systems, 25(4): 474–480, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bjorner:1991:FMO

[BHH91] D. Bjørner, A. Haxthausen, and K. Havelund. Formal, model-oriented software development methods — from VDM to ProCoS & from RAISE to LaCoS. Future Generation Computer Systems, ?? (7):??, ???? 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bjorner:1992:FMO

[BHH92] Dines Bjørner, Anne Elisabeth Haxthausen, and Klaus Havelund. Formal, modeloriented software development methods: From VDM to ProCoS & from RAISE to LaCoS. Future Generation Computer Systems, 7(2–3): 111–138, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Beemster:1993:ECP

[BHH^{+}93] M. Beemster, P. H. Hartel, L. O. Hertzberger, R. F. H. Hofman, K. G. Langendoen, L. L. Li, R. Milikowski, W. G. Vree, H. P. Barendregt, and J. C. Mulder. Experience with a clustered parallel reduction machine. Future Generation Computer Systems, 9(3):175– 200, September 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bolton:1990:POO

[BHK90] David Bolton, Chris Hankin, and Paul Kelly. Parallel object-oriented descriptions of graph reduction machines. Future Generation Computer Systems, 6(3):225– 239, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Breitler:1998:AHP

[BHRT98] Marc Breitler, Stéphane Hegi, Jean-Daniel Reymond, and Nils S. Tuchschmid. Application of HPC to a portfolio choice problem. Future Generation Computer Systems, 13(4–5):269–278, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/20/abstract.
html.

Bhuyan:1995:HPC

[Bhu95] Laxmi N. Bhuyan. Highperformance computer architecture. Future Generation Computer Systems, 11(6): 501–502, October 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bisiani:1994:HPC

[Bis94] R. Bisiani. High performance computing systems: Present and future. Future Generation Computer Systems, 10(2– 3):241–248, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bisseling:1996:TUC

[Bis96] Rob Bisseling. Third Utrecht Computational Science Symposium. Future Generation Computer Systems, 12(4): 263–264, December 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Balsoy:2005:AMW

[BJA^{+}05] Ozgur Balsoy, Ying Jin, Galip Aydin, Marlon Pierce, and Geoffrey Fox. Automating metadata Web service deployment for problem solving environments. Future Generation Computer Systems, 21 (6):910–919, June 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Buyya:2002:GEC

[BJC02] Rajkumar Buyya, Hai Jin, and Toni Cortes. Guest editorial: Cluster computing. Future Generation Computer Systems, 18(3):v– viii, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/27/abstract.
html.

Basermann:2005:PIS

[BJNH05] A. Basermann, U. Jaekel, M. Nordhausen, and K. Hachiya. Parallel iterative solvers for sparse linear systems in circuit simulation. Future Generation Computer Systems, 21 (8):1275–1284, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Balos:2008:OIA

[BJWZ08] Kazimierz Bałos, Marcin Jarza̧b, Damian Wieczorek, and Krzysztof Zieliński. Open interface for autonomic management of virtualized resources in complex systems — construction methodology. Future Generation Computer Systems, 24(5):390–401, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bennett:1997:ESM

[BK97] Andrew J. Bennett and Paul H. J. Kelly. Efficient sharedmemory support, for parallel graph reduction. Future Generation Computer Systems, 12 (6):481–503, June 15, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/18/17/abstract.
html.

Burton:2006:PPP

[BK06] Ariel N. Burton and Paul H. J. Kelly. Performance prediction of paging workloads using lightweight tracing. Future Generation Computer Systems, 22(7):784–793, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bekas:2005:DDM

[BKG05] C. Bekas, E. Kokiopoulou, and E. Gallopoulos. The design of a distributed MATLABbased environment for computing pseudospectra. Future Generation Computer Systems, 21(6):930–941, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Borshchev:2002:DSH

[BKK02] Andrei Borshchev, Yuri Karpov, and Vladimir Kharitonov. Distributed simulation of hybrid systems with AnyLogic and HLA. Future Generation Computer Systems, 18 (6):829–839, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Baratloo:1999:CMW

[BKKW99] A. Baratloo, M. Karaul, Z. M.

Kedem, and P. Wijckoff. Charlotte: Metacomputing on the Web. Future Generation Computer Systems, 15(5– 6):559–570, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/19/abstract.
html.

Bubak:2001:CUL

[BKŁ01] Marian Bubak, Dawid Kurzyniec, and Piotr Łuszczek. Convenient use of legacy software in Java with Janet package. Future Generation Computer Systems, 17 (8):987–997, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/35/abstract.
html.

Benner:2003:SPC

[BKM03] Peter Benner, Daniel Kressner, and Volker Mehrmann. Structure preservation: a challenge in computational control. Future Generation Computer Systems, 19 (7):1243–1252, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Belleman:1998:VEE

[BKS98] Robert G. Belleman, Jaap A. Kaandorp, and Peter M. A. Sloot. A virtual environment for the exploration of diffusion and flow phenomena in complex geometries. Future Generation Computer Systems, 14(3–4):209–214, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/29/18/24/28/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/29/18/24/abstract.
html.

Beynon:2002:OEC

[BKSS02] Michael D. Beynon, Tahsin Kurc, Alan Sussman, and Joel Saltz. Optimizing execution of component-based applications using group instances. Future Generation Computer Systems, 18(4): 435–448, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/33/28/abstract.
html.

Bouge:1992:CSD

[BL92] Luc Bougé and Jean-Luc Levaire. Control structures for data-parallel SIMD languages: semantics and implementation. Future Generation Computer Systems, 8(4):363– 378, September 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Barak:1998:MMO

[BL98] Amnon Barak and Oren La’adan. The MOSIX multicomputer operating system for high performance cluster computing. Future Generation Computer Systems, 13(4– 5):361–372, March 11, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/28/abstract.
html.

Bohn:2002:LBH

[BL02] Christopher A. Bohn and Gary B. Lamont. Load balancing for heterogeneous clusters of PCs. Future Generation Computer Systems, 18(3): 389–400, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/32/abstract.
html.

Bote-Lorenzo:2006:SSC

[BLAV06] Miguel L. Bote-Lorenzo and Oscar Ardaiz-Villanueva. Special section: Collaborative and learning applications of Grid technology. Future Generation Computer Systems, 22(6):699, May 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bucker:2003:PPC

[BLB03] H. M. Bücker, B. Lang, and C. H. Bischof. Parallel programming in computational science: an introductory practical training course for computer science undergraduates at Aachen University. Future Generation Computer Systems, 19(8):1309– 1319, November 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bal:2003:E

[BLRL03] Henri Bal, Klaus-Peter Löhr, Alexander Reinefeld, and Craig Lee. Editorial. Future Generation Computer Systems, 19(4):503–504, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Blach:1998:HFV

[BLRS98] Roland Blach, Jürgen Landauer, Angela Rösch, and Andreas Simon. A highly flexible virtual reality system. Future Generation Computer Systems, 14(3–4):167– 178, July 31, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/29/18/19/abstract.
                                        html.

Bisiani:1992:DMH

[BM92] R. Bisiani and O. Martin. A distributed-memory, highperformance workstation. Future Generation Computer Systems, 8(1–3):83–91, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brown:2000:CVA

[BM00] John C. Brown and Suresh Manandhar. Compilation versus abstract machines for fast parsing of typed feature structure grammars. Future Generation Computer Systems, 16 (7):771–791, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/30/28/abstract.
html.

Bergstra:2008:DSI

[BM08] J. A. Bergstra and C. A. Middelburg. Distributed strategic interleaving with load balancing. Future Generation Computer Systems, 24(6):530–548, June 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bazinet:2007:GSB

[BMFC07] Adam L. Bazinet, Daniel S.

Myers, John Fuetsch, and Michael P. Cummings. Grid Services Base Library: a high-level, procedural application programming interface for writing Globus-based Grid services. Future Generation Computer Systems, 23(3): 517–522, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bandini:2001:PSR

[BMPS01] Stefania Bandini, Giancarlo Mauri, Giulio Pavesi, and Carla Simone. Parallel simulation of reaction-diffusion phenomena in percolation processes — a model based on cellular automata. Future Generation Computer Systems, 17 (6):679–688, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/27/abstract.
html.

Bakic:2001:POD

[BMRW01] Aleksandar Bakić, Matt W. Mutka, Diane T. Rover, and Abdul Waheed. Performance optimization of distributed applications in an extensible, adaptive environment. Future Generation Computer Systems, 18 (1):131–145, September 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/40/abstract.
html.

Bandini:2005:SAS

[BMS05] Stefania Bandini, Sara Manzoni, and Carla Simone. Supporting the application of Situated Cellular Agents in nonuniform spaces. Future Generation Computer Systems, 21 (5):627–631, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bergamaschi:2001:IQN

[BMZ01] Luca Bergamaschi, Igor Moret, and Giovanni Zilli. Inexact Quasi-Newton methods for sparse systems of nonlinear equations. Future Generation Computer Systems, 18(1): 41–53, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/31/abstract.
html.

Blaheta:2006:SSN

[BN06] Radim Blaheta and Jiří Nedoma. Special section: Numerical modelling in geomechanics and geodynamics. Future Generation Computer Systems, 22(4):447– 448, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Batko:2008:SCP

[BNFZ08] Michal Batko, David Novak, Fabrizio Falchi, and Pavel Zezula. Scalability comparison of Peer-to-Peer similarity search structures. Future Generation Computer Systems, 24 (8):834–848, October 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Boada:2004:OBM

[Boa04] Imma Boada. An octreebased multiresolution hybrid framework. Future Generation Computer Systems, 20(8): 1275–1284, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Boghosian:1999:LGC

[Bog99] Bruce M. Boghosian. Lattice gases and cellular automata. Future Generation Computer Systems, 16(2–3): 171–185, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/25/27/abstract.
                                        html.

Bastien:2007:CSH

[BORM07] Olivier Bastien, Philippe Ortet, Sylvaine Roy, and Eric Maréchal. The configuration space of homologous proteins: a theoretical and practical framework to reduce the diversity of the protein sequence space after massive all-by-all sequence comparisons. Future Generation Computer Systems, 23(3): 410–427, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bousri:1995:PNM

[Bou95]  Sebti Bousri.  A paral-

lel nodal method of second order. Future Generation Computer Systems, 11(2): 153–160, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bodin:1994:OKP

[BP94] François Bodin and Thierry Priol. Overview of the KOAN programming environment for the iPSC/2 and performance evaluation of the BECAUSE test program 2.51. Future Generation Computer Systems, 10(4):391–401, November 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bechini:2001:BIC

[BP01] Alessio Bechini and Cosimo Antonio Prete. Behavior investigation of concurrent Java programs: an approach based on source-code instrumentation. Future Generation Computer Systems, 18(2):307– 316, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/36/abstract.
html.

Bandini:2002:CGT

[BP02] Stefania Bandini and Giulio Pavesi. Controlled generation of two-dimensional patterns based on Stochastic Cellular Automata. Future Generation Computer Systems, 18(7):

973–981, August 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brenot:1992:AKB

[BPAP92] Jean-Marc Brenot, Yann Parrod, Didier Aubin, and Christian Parquet. ARIANEX-PERT: a knowledge based system to analyze ARIANE’s mission data. Future Generation Computer Systems, 7 (4):365–377, May 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Braun:2001:PLB

[BPC^{+}01] R. C. Braun, K. T. Pedretti, T. L. Casavant, T. E. Scheetz, C. L. Birkett, and C. A. Roberts. Parallelization of local BLAST service on workstation clusters. Future Generation Computer Systems, 17 (6):745–754, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/33/33/abstract.
                                        html.

Beltrame:2007:GGE

[BPP^{+}07] Francesco Beltrame, Adam Papadimitropoulos, Ivan Porro, Silvia Scaglione, Andrea Schenone, Livia Torterolo, and Federica Viti. GEMMA — a Grid environment for microarray management and analysis in bone marrow stem cells experiments. Future Generation Computer Systems, 23 (3):382–390, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Breckenridge:2003:DDD

[BPS^{+}03] Arthurine Breckenridge, Lyndon Pierson, Sergiu Sanielevici, Joel Welling, Rainer Keller, Uwe Woessner, and Juergen Schulze. Distributed, on-demand, data-intensive and collaborative simulation analysis. Future Generation Computer Systems, 19(6): 849–859, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Benoit:2006:MEK

[BPS06] Anne Benoit, Brigitte Plateau, and William J. Stewart. Memory-efficient Kronecker algorithms with applications to the modelling of parallel systems. Future Generation Computer Systems, 22(7): 838–847, August 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bemmerl:1992:PTD

[BR92] Thomas Bemmerl and Bernhard Ries. Programming tools for distributed multiprocessor computing environments. Future Generation Computer Systems, 8(1–3):221–234, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bretherton:1989:ES

[Bre89] Francis P. Bretherton. The Earth System. Future Generation Computer Systems, 5(2– 3):259–264, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bhowmick:2004:FPB

[BRMN04] S. Bhowmick, P. Raghavan, L. McInnes, and B. Norris. Faster PDE-based simulations using robust composite linear solvers. Future Generation Computer Systems, 20(3): 373–387, April 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brode:1992:CCW

[Bro92] Stefan Brode. Computational chemistry with a workstationfarm. Future Generation Computer Systems, 8(1–3): 27–29, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Belmonte:2004:EUC

[BRR^{+}04] Ó. Belmonte, I. Remolar, J. Ribelles, M. Chover, and M. Fernández. Efficiently using connectivity information between triangles in a mesh for real-time rendering. Future Generation Computer Systems, 20(8):1263– 1273, November 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Belanzoni:2004:TAM

[BRS04] P. Belanzoni, M. Rosi, and A. Sgamellotti. A theoretical approach to molecular batteries: C — C bonds functioning as electron shuttles. Future Generation Computer Systems, 20(5):793–805, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bruzzone:2001:EWB

[Bru01] Agostino G. Bruzzone. Editorial: Web-based simulation — Best of Websim99. Future Generation Computer Systems, 17(5): 501–502, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/25/abstract.
html.

Burger:1984:VRU

[BS84] R. M. Burger and L. W.

Sumney. VLSI research in the U.S.A. Future Generation Computer Systems, 1 (1):31–38, July 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bettini:1991:CMO

[BS91a] C. Bettini and L. Spampinato.

The connection machine opportunity for the implementation of a concurrent functional language. Future Generation Computer Systems, 7 (??):231–245, ???? 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brackenbury:1991:RM

[BS91b] Ian Brackenbury and Howard Sachar. The road to multimedia. Future Generation Computer Systems, 7(1):91– 96, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bettini:1992:CMO

[BS92] Claudio Bettini and Luca Spampinato. The connection machine opportunity for the implementation of a concurrent functional language. Future Generation Computer Systems, 7(2–3): 231–245, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Benerecetti:1996:MPT

[BS96] Massimo Benerecetti and Luca Spalazzi. METAFOL: Program tactics and logic tactics plus reflection. Future Generation Computer Systems, 12(2–3):139–156, September 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Berridge:2004:IGM

[BS04] Steffan Berridge and J. M.

Schumacher. An irregular grid method for high-dimensional free-boundary problems in finance. Future Generation Computer Systems, 20(3): 353–362, April 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Blanas:2009:CBP

[BS09] Spyros Blanas and Vasilis Samoladas. Contention-based performance evaluation of multidimensional range search in peer-to-peer networks. Future Generation Computer Systems, 25(1):100–108, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Barbosa:2006:EAI

[BSCC06] Ana Carolina Barbosa, Jacques Sauvé, Walfredo Cirne, and Mirna Carelli. Evaluating architectures for independently auditing service level agreements. Future Generation Computer Systems, 22(7): 721–731, August 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brandes:2005:MCB

[BSG^{+}05] Thomas Brandes, Helmut Schwamborn, Michael Gerndt, Jürgen Jeitner, Edmond Kereku, Martin Schulz, Holger Brunst, Wolfgang Nagel, Reinhard Neumann, Ralph Müller-Pfefferkorn, Bernd Trenkler, Wolfgang Karl, Jie Tao, and Hans-Christian Hoppe. Monitoring cache behavior on parallel SMP architectures and related programming tools. Future Generation Computer Systems, 21 (8):1298–1311, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Belpassi:2004:PRD

[BST^{+}04] Leonardo Belpassi, Loriano Storchi, Francesco Tarantelli, Antonio Sgamellotti, and Harry M. Quiney. Parallelization of a relativistic DFT code. Future Generation Computer Systems, 20(5): 739–747, June 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Branford:2008:MCM

[BST^{+}08] S. Branford, C. Sahin, A. Thandavan, C. Weihrauch, V. N. Alexandrov, and I. T. Dimov. Monte Carlo methods for matrix computations on the Grid. Future Generation Computer Systems, 24 (6):605–612, June 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bastien-Thiry:1993:STT

[BT93]  Christophe  Bastien-Thiry.

SE-TC2: Telecom 2 expert system (the first expert system in a CNES Satellite Control Centre). Future Generation Computer Systems, 9(4): 303–309, December 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Boloni:2006:TDR

[BTM06] Ladislau Bölöni, Damla Turgut, and Dan C. Marinescu. Task distribution with a random overlay network. Future Generation Computer Systems, 22 (6):676–687, May 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bucker:2005:LNI

[Büc05a] H. Martin Bücker. Looking for narrow interfaces in automatic differentiation using graph drawing. Future Generation Computer Systems, 21 (8):1418–1425, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bucker:2005:SSA

[Büc05b]  H. Martin Bücker.  Spe-

cial section: Automatic differentiation and its applications. Future Generation Computer Systems, 21 (8):1322–1323, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bungartz:2003:CSE

[Bun03] Hans-Joachim Bungartz. Computational science and engineering: a new master’s program at the Technische Universität München. Future Generation Computer Systems, 19(8):1267–1274, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Burks:2002:IUE

[Bur02] Arthur W. Burks. The invention of the universal electronic computer — how the Electronic Computer Revolution began. Future Generation Computer Systems, 18(7): 871–892, August 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Betcke:2004:JDT

[BV04] Timo Betcke and Heinrich Voss. A Jacobi–Davidson-type projection method for nonlinear eigenvalue problems. Future Generation Computer Systems, 20(3):363–372, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bakker:1993:IPM

[BvdBM^{+}93] R. R. Bakker, P. C. A. van den Bempt, N. J. I. Mars, D.-J. Out, and D. C. van Soest. Issues in practical model-based diagnosis. Future Generation Computer Systems, 9(4):329– 337, December 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Briguglio:2000:PPS

[BVDF00] S. Briguglio, G. Vlad, B. Di Martino, and G. Fogaccia. Parallelization of plasma simulation codes: gridless finite size particle versus particle in cell approach. Future Generation Computer Systems, 16 (5):541–552, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/37/abstract.
html.

Belien:2001:SCM

[BvdHN^{+}01] A. J. C. Beliën, B. van der Holst, M. Nool, A. van der Ploeg, and J. P. Goedbloed. Spectral calculations in magnetohydrodynamics using the Jacobi–Davidson method. Future Generation Computer Systems, 17 (8):919–924, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/27/abstract.
html.

Baalbergen:1999:SFU

[BvdV99] E. H. Baalbergen and H. van der Ven. SPINEware — a framework for user-oriented and tailorable metacomputers. Future Generation Computer Systems, 15(5–6):549– 558, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/18/abstract.
html.

Barendregt:1987:DPR

[BVP^{+}87] H. P. Barendregt, M. C. J. D. Van Eekelen, M. J. Plasmeijer, P. H. Hartel, L. O. Hertzberger, and W. G. Vree. The Dutch Parallel Reduction Machine Project. Future Generation Computer Systems, 3 (4):261–270, December 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL ftp://ftp.cs.
kun.nl/pub/CSI/SoftwEng.
FunctLang/papers/barh87-
PRMprojekt.abs;   ftp:/
/ftp.cs.kun.nl/pub/CSI/
SoftwEng.FunctLang/papers/
barh87-PRMprojekt.ps.gz.

Bemmerl:1995:LDD

[BW95] Thomas Bemmerl and Roland Wismüller. On-line distributed debugging on scalable multiprocessor architectures. Future Generation Computer Systems, 11(4–5): 375–385, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bik:1997:ISP

[BW97] Aart J. C. Bik and Harry A. G. Wijshoff. Iteration space partitioning. Future Generation Computer Systems, 12(5): 421–429, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/24/abstract.
html.

Bu:2004:ACS

[BX04] Guanying Bu and Zhiwei Xu.

Access control in semantic grid. Future Generation Computer Systems, 20(1):113–122, January 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Barnett:1993:EIN

[BY93] James Barnett and Kenji Yamada. Evaluation of ICOT’s natural language research. Future Generation Computer Systems, 9(2):137–142, July 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Buyya:2009:CCE

[BYV^{+}09] Rajkumar Buyya, Chee Shin Yeo, Srikumar Venugopal, James Broberg, and Ivona Brandic. Cloud computing and emerging IT platforms: Vision, hype, and reality for delivering computing as the 5th utility. Future Generation Computer Systems, 25 (6):599–616, June 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cadiou:1986:EA

[Cad86] Jean-Marie Cadiou. ESPRIT in action. Future Generation Computer Systems, 2(1): 51–60, March 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cottrell:2003:IDB

[CALN03] R. Les Cottrell, Antony Antony, Connie Logg, and Jiri Navratil. iGrid2002 demonstration: bandwidth from the low lands. Future Generation Computer Systems, 19(6): 825–837, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carpentier:1986:CSI

[Car86]  M. Carpentier.  Commu-

nity strategy in information technology and telecommunications. Future Generation Computer Systems, 2(1): 19–22, March 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carpenter:2003:FAM

[Car03] Brian E. Carpenter. Future applications and middleware, and their impact on the infrastructure. Future Generation Computer Systems, 19(2): 191–197, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Castelli:1994:SDE

[Cas94]  Gianluigi Castelli.  Soft-

ware development environments for massively parallel systems. Future Generation Computer Systems, 10(2–3): 249–256, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chinellato:2005:COM

[CASW05] O. Chinellato, P. Arbenz, M. Streiff, and A. Witzig. Computation of optical modes in axisymmetric open cavity resonators. Future Generation Computer Systems, 21 (8):1263–1274, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cooper:2005:AAP

[CBD^{+}05] Brian F. Cooper, Mayank Bawa, Neil Daswani, Sergio Marti, and Hector Garcia-Molina. Authenticity and availability in PIPE networks. Future Generation Computer Systems, 21(3): 391–400, March 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Casavant:2001:PDA

[CBK^{+}01] Thomas L. Casavant, Terry A.

Braun, Sureshkumar Kaliannan, Todd E. Scheetz, Kyle J.

Munn, and Clayton L. Birkett. A parallel/distributed architecture for hierarchically heterogeneous Webbased cooperative applications. Future Generation Computer Systems, 17(6): 783–793, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/37/abstract.
html.

Cho:1998:ARA

[CC98] Ok-Hyeong Cho and Robert M. Colomb. Associative random access machines anddataparallel multiway binarysearch join. Future Generation Computer Systems, 13(6): 451–467, May 20, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/20/19/abstract.
html.

Cilio:2000:LTE

[CC00] Andrea G. M. Cilio and Henk Corporaal. Link-time effective whole-program optimizations. Future Generation Computer Systems, 16(5): 503–511, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/33/abstract.
html.

Chang:2007:CFR

[CC07] Ruay-Shiung Chang and Po-Hung Chen. Complete and fragmented replica selection and retrieval in Data Grids. Future Generation Computer Systems, 23(4):536–546, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chung:2009:NMR

[CC09] Wu-Chun Chung and Ruay-Shiung Chang. A new mechanism for resource monitoring in Grid computing. Future Generation Computer Systems, 25(1):1–7, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Camp:1998:VRM

[CCBR98] J. J. Camp, B. M. Cameron, D. Blezek, and R. A. Robb. Virtual reality in medicine and biology. Future Generation Computer Systems, 14(1–2): 91–108, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/25/abstract.
html.

Caron:2008:DPS

[CCDS08] E. Caron, A. Chis, F. Desprez, and A. Su. Design of plugin schedulers for a GridRPC environment. Future Generation Computer Systems, 24 (1):46–57, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ciriello:2007:ARS

[CCG07] G. Ciriello, M. Comin, and C. Guerra. Algorithmic restructuring and data replication for protein structure comparison on a GRID. Future Generation Computer Systems, 23(3):391–397, March 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cooperman:2003:UTC

[CCHW03] Gene Cooperman, Henri Casanova, Jim Hayes, and Thomas Witzel. Using TOP-C and AMPIC to port large parallel applications to the Computational Grid. Future Generation Computer Systems, 19 (4):587–596, May 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Clune:1988:IPC

[CCKW88] Ed Clune, Jill D. Crisman, Gudrun J. Klinker, and Jon A. Webb. Implementation and performance of a complex vision system on a systolic array machine. Future Generation Computer Systems, 4 (1):15–29, August 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2007:JSD

[CCL07] Ruay-Shiung Chang, Jih-Sheng Chang, and Shin-Yi Lin. Job scheduling and data replication on data grids. Future Generation Computer Systems, 23(7):846–860, August 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chand:2008:PRD

[CCL08] Raphael Chand, Michel Cosnard, and Luigi Liquori. Powerful resource discovery for Arigatoni overlay network. Future Generation Computer Systems, 24(1):31–38, January 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2009:AAB

[CCL09] Ruay-Shiung Chang, Jih-Sheng Chang, and Po-Sheng Lin. An ant algorithm for balanced job scheduling in Grids. Future Generation Computer Systems, 25(1):20–27, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2009:PML

[CCLS09] Po-Cheng Chen, Jyh-Biau Chang, Tyng-Yeu Liang, and Ce-Kuen Shieh. A progressive multi-layer resource reconfiguration framework for time-shared grid systems. Future Generation Computer Systems, 25(6):662–673, June 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cristea:1998:PLC

[CCM98] Paul Cristea, Jan Cornelis, and Adrian Munteanu. Progressive lossless coding of medical images. Future Generation Computer Systems, 14 (1–2):23–32, June 15, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/19/abstract.
html.

Cannata:2007:RGB

[CCM07] Nicola Cannata, Flavio Corradini, and Emanuela Merelli. A Resourceomic Grid for bioinformatics. Future Generation Computer Systems, 23(3): 510–516, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Conforti:1999:CIM

[CD99] Domenico Conforti and Luigi De Luca. Computer implementation of a medical diagnosis problem by pattern classification. Future Generation Computer Systems, 15 (2):287–292, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X98000739.

Coronato:2008:MMI

[CD08] Antonio Coronato and Giuseppe De Pietro. MiPeG: a middleware infrastructure for pervasive grids. Future Generation Computer Systems, 24 (1):17–29, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Caromel:2007:PPF

[CdCD07] Denis Caromel, Alexandre di Costanzo, and Christian Delbé. Peer-to-Peer and fault-tolerance: Towards deployment-based technical services. Future Generation Computer Systems, 23(7): 879–887, August 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cappello:2005:CLS

[CDF^{+}05] Franck Cappello, Samir Djilali, Gilles Fedak, Thomas Herault, Frédéric Magniette, Vincent Néri, and Oleg Lodygensky. Computing on largescale distributed systems: XtremWeb architecture, programming models, security, tests and convergence with grid. Future Generation Computer Systems, 21(3):417– 437, March 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Crisci:2005:PCA

[CDRS05] Gino Mirocle Crisci, Salvatore Di Gregorio, Rocco Rongo, and William Spataro. PYR: a Cellular Automata model for pyroclastic flows and application to the 1991 Mt. Pinatubo eruption. Future Generation Computer Systems, 21(7): 1019–1032, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chu:2003:RMC

[CDS03] M. T. Chu, F. Diele, and I. Sgura. On robust matrix completion with prescribed eigenvalues. Future Generation Computer Systems, 19 (7):1139–1153, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chalermwat:2001:PGB

[CEGL01] Prachya Chalermwat, Tarek El-Ghazawi, and Jacqueline LeMoigne. 2-phase GA-based image registration on parallel clusters. Future Generation Computer Systems, 17 (4):467–476, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/29/36/abstract.
                                        html.

Collette:1994:SSC

[CEJK94] T. Collette, H. Essafi, D. Juvin, and J. Kaiser. SYM-PATIX: a SIMD computer performing the low and intermediate levels of image processing. Future Generation Computer Systems, 10 (1):3–13, April 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chiu:2009:SST

[CF09] Kenneth Chiu and Geoffrey C.

Fox. Special section: Third IEEE International Conference on e-Science and Grid Computing. Future Generation Computer Systems, 25 (4):444–445, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cui:1993:SDU

[CFG93] Jun Cui, Terence C. Fogarty, and John G. Gammack. Searching databases using parallel genetic algorithms on a transputer computing surface. Future Generation Computer Systems, 9 (1):33–40, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Curti:2005:ARH

[CFG^{+}05] Chiara Curti, Tiziana Ferrari, Leon Gommans, S. van Oudenaarde, Elisabetta Ronchieri, Francesco Giacomini, and Cristina Vistoli. On advance reservation of heterogeneous network paths. Future Generation Computer Systems, 21 (4):525–538, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carretero:2003:HDS

[CFGC03] Jesús Carretero, Javier Fernández, Félix García, and Alok Choudhary. A hierarchical disk scheduler for multimedia systems. Future Generation Computer Systems, 19 (1):23–35, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chiarantoni:2003:AFP

[CFVP03] E. Chiarantoni, G. Fornarelli, S. Vergura, and T. Politi. Applying fixed point homotopy to nonlinear DAEs deriving from switching circuits. Future Generation Computer Systems, 19(3):431–441, April 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cerin:2002:SPS

[CG02] Christophe Cérin and Jean-Luc Gaudiot. On a scheme for parallel sorting on heterogeneous clusters. Future Generation Computer Systems, 18(3): 353–372, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/32/30/abstract.
                                        html.

Cardenas:2009:EFA

[CG09] César Cárdenas and Maurice Gagnaire. Evaluation of Flow-Aware Networking (FAN) architectures under GridFTP traffic. Future Generation Computer Systems, 25 (8):895–903, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2004:TSSa

[CGH04] Weng-Long Chang, Minyi Guo, and Michael Ho. Towards solution of the setsplitting problem on gel-based DNA computing. Future Generation Computer Systems, 20 (5):875–885, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2008:MPD

[CGL08] Ruay-Shiung Chang, Ming-Huang Guo, and Hau-Chin Lin. A multiple parallel download scheme with server throughput and client bandwidth considerations for data grids. Future Generation Computer Systems, 24(8): 798–805, October 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cannataro:2007:UOP

[CGM^{+}07] M. Cannataro, P. H. Guzzi, T. Mazza, G. Tradigo, and P. Veltri. Using ontologies for preprocessing and mining spectra data on the Grid. Future Generation Computer Systems, 23(1):55–60, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Comito:2009:SOS

[CGST09] Carmela Comito, Anastasios Gounaris, Rizos Sakellariou, and Domenico Talia. A service-oriented system for distributed data querying and integration on Grids. Future Generation Computer Systems, 25(5):511–524, May 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cornubert:1995:BAS

[CGSZ95] R. Cornubert, G. Gruez, P. Steinfeld, and E. Znaty. Benchmark of application software kernels on the SU-PERNODE SN1000 using the 3P PARLIB. Future Generation Computer Systems, 11(1): 87–109, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Caron:2007:DMS

[CGT07] Eddy Caron, Vincent Garonne, and Andreï Tsaregorodtsev. Definition, modelling and simulation of a grid computing scheduling system for high throughput computing. Future Generation Computer Systems, 23(8):968– 976, November 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Coron:1995:RGF

[CH95] François Coron and Philippe Homsi. Rarefied gas flow computational with a 3D unstructured mesh on a Connection Machine (CM2). Future Generation Computer Systems, 11 (1):1–6, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chopard:2004:CSL

[CH04] Bastien Chopard and Alfons Hoekstra. Computational science of lattice Boltzmann modelling. Future Generation Computer Systems, 20(6): 907–908, August 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chengzheng:1991:PPO

[CHHW91] Sun Chengzheng, L. O.

Hertzberger, B. J. A. Hulshof, and Rogier Wester. POOSS: A Parallel Object-Oriented Stable Storage. Future Generation Computer Systems, 6 (4):333–349, September 1991. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2004:MES

[CHJ^{+}04] Yang-Lang Chang, Chin-Chuan Han, Fan-Di Jou, Kuo-Chin Fan, K. S. Chen, and Jeng-Horng Chang. A modular eigen subspace scheme for high-dimensional data classification. Future Generation Computer Systems, 20(7): 1131–1143, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cook:1998:VRL

[CHK98] Jon Cook, Roger Hubbold, and Martin Keates. Virtual reality for large-scale industrial applications. Future Generation Computer Systems, 14(3–4):157–166, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/29/18/18/28/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/29/18/18/abstract.
html.

Choi:2004:PTA

[Cho04] Eunmi Choi. Performance test and analysis for an adaptive load balancing mechanism on distributed server cluster systems. Future Generation Computer Systems, 20(2): 237–247, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chvalovsky:1987:KPB

[Chv87] Václav Chvalovský. Knowledge processing: Beware of stalemate. Future Generation Computer Systems, 3(2): 93–101, May 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ciegis:2004:NCS

[Čie04] R. Čiegis. On a new class of splitting type iterative methods. Future Generation Computer Systems, 20(3): 399–407, April 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2000:ESL

[CK00] Huey-Ling Chen and Chung-Ta King. Eager scheduling with lazy retry in multiprocessors. Future Generation Computer Systems, 17(3): 215–226, November 1, 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/25/abstract.
html.

Cerin:2006:SCS

[CKFJ06] Christophe Cérin, Michel Koskas, Hazem Fkaier, and Mohamed Jemni. Sequential in-core sorting performance for a SQL data service and for parallel sorting on heterogeneous clusters. Future Generation Computer Systems, 22(7): 776–783, August 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2004:ASM

[CKK^{+}04] Beom-Hwan Chang, Dong-Soo Kim, Hyun-Ku Kim, Jung-Chan Na, and Tai-Myoung Chung. Active security management based on Secure Zone Cooperation. Future Generation Computer Systems, 20(2):283–293, February 16, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chapin:1999:RML

[CKKG99] Steve J. Chapin, Dimitrios Katramatos, John Karpovich, and Andrew Grimshaw. Resource management in Legion. Future Generation Computer Systems, 15(5–6):583– 594, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/21/abstract.
html.

Chiu:2006:DSD

[CkLC06] Steve C. Chiu, Wei keng Liao, and Alok N. Choudhary. Distributed smart disks for I/O-intensive workloads on switched interconnects. Future Generation Computer Systems, 22(5):643–656, April 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cox:2004:ASC

[CKR04] Dennis D. Cox, Petr Klouček, and Daniel R. Reynolds. On the asymptotically stochastic computational modeling of microstructures. Future Generation Computer Systems, 20 (3):409–424, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Crispo:2000:WST

[CLM00] Bruno Crispo, Peter Landrock, and Václav Matyáš, Jr. WWW security and trusted third party services. Future Generation Computer Systems, 16(4): 331–341, February 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/28/abstract.
html.

Caerts:1995:PPL

[CLP95] Chris Caerts, Rudy Lauwereins, and J. A. Peperstraete. PDG: a processlevel debugger for concurrent programs in the GRAPE parallel programming environment. Future Generation Computer Systems, 11(2): 199–210, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chopard:1999:CAL

[CM99] Bastien Chopard and Alexandre Masselot. Cellular automata and lattice Boltzmann methods: a new approach to computational fluid dynamics and particle transport. Future Generation Computer Systems, 16(2–3):249– 257, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/32/abstract.
html.

Camarinha-Matos:2001:ESD

[CM01] Luis M. Camarinha-Matos. Execution system for distributed business processes in a virtual enterprise. Future Generation Computer Systems, 17(8):1009–1021, June 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/37/abstract.
html.

Celledoni:2003:CFL

[CMO03] Elena Celledoni, Arne Marthinsen, and Brynjulf Owren. Commutator-free Lie group methods. Future Generation Computer Systems, 19 (3):341–352, April 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Calzarossa:2001:PIH

[CMT01] Maria Calzarossa, Luisa Massari, and Daniele Tessera. Performance issues of an HPFlike compiler. Future Generation Computer Systems, 18 (1):147–156, September 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/41/abstract.
html.

Chapman:1995:HPF

[CMZ95] Barbara Chapman, Piyush Mehrotra, and Hans Zima. High Performance Fortran languages: Advanced applications and their implementation. Future Generation Computer Systems, 11(4–5): 401–407, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carmichael:1992:PPP

[CN92] Neil Carmichael and Michael Norman. Parallel processing: the power and the portability. Experiments with ‘reusable toolkits’. Future Generation Computer Systems, 8 (1–3):3–8, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cruz-Neira:1998:MVR

[CN98] Carolina Cruz-Neira. Making virtual reality useful: a report on immersive applications at Iowa State University. Future Generation Computer Systems, 14(3–4):147–155, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/19/19/29/18/17/28/
                                        index.htt;   http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/29/18/17/abstract.
                                        html.

Chadwick:2003:PXR

[CO03] David W. Chadwick and Alexander Otenko. The PER-MIS X.509 role based privilege management infrastructure. Future Generation Computer Systems, 19(2):277– 289, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cooper:1986:ESM

[Coo86] Philip Cooper. Expert systems in management science. Future Generation Computer Systems, 2(4):217– 223, December 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Coolen:1990:ISN

[Coo90] A. C. C. Coolen. Ising spin neural networks with spatial structure. Future Generation Computer Systems, 6(2):121– 130, November 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cooper:1994:SP

[Coo94] R. Cooper. The SuperJANET project. Future Generation Computer Systems, 10(2–3): 233–240, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carpentieri:2006:SAF

[CP06] M. Carpentieri and B. Paternoster. Stability analysis of frequency and step length dependent Runge– Kutta–Nyström methods. Future Generation Computer Systems, 22(4):395–402, March 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cabestre:2000:AMC

[CPB00] Frédéric Cabestre, Christian Percebois, and Jean-Paul Bodeveix. Abstract machine construction through operational semantics refinements. Future Generation Computer Systems, 16(7): 753–769, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/30/27/abstract.
html.

Chang:2005:NME

[CPK05] Yun Seok Chang, Kwang Suk Park, and Bo Yeon Kim. Nonlinear model for ECG R–R interval variation using genetic programming approach. Future Generation Computer Systems, 21(7): 1117–1123, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:1992:EPM

[CR92] Henry H. Y. Chang and Bryan Rosenburg. Experience porting Mach to the RP3 largescale shared-memory multiprocessor. Future Generation Computer Systems, 7(2– 3):259–267, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cappello:2001:UPS

[CRE01] Franck Cappello, Olivier Richard, and Daniel Etiemble. Understanding performance of SMP clusters running MPI programs. Future Generation Computer Systems, 17 (6):711–720, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/30/abstract.
html.

Cunha:2005:FTD

[CRM05] José C. Cunha, Omer F. Rana, and Pedro D. Medeiros. Future trends in distributed applications and problemsolving environments. Future Generation Computer Systems, 21(6):843–855, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Crone:1995:CGM

[Cro95] Lianne G. C. Crone. The conjugate gradient method on the Parsytec GCel-3/512. Future Generation Computer Systems, 11(2):161–166, March 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Clare:1993:IFD

[CS93] A. R. Clare and D. P. Stevens.

Implementing finite difference ocean circulation models on MIMD, distributed memory computers. Future Generation Computer Systems, 9 (1):11–18, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ciampolini:1996:EPM

[CS96] A. Ciampolini and C. Stefanelli. Extending PVM to a massively parallel architecture. Future Generation Computer Systems, 12(1):13–23, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Craven:1997:UNN

[CS97] Mark W. Craven and Jude W.

Shavlik. Using neural networks for data mining. Future Generation Computer Systems, 13(2–3):211–229, November 14, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/24/abstract.
html.

Chen:2005:AGI

[CS05] Chunxi Chen and Bertil Schmidt. An adaptive grid implementation of DNA sequence alignment. Future Generation Computer Systems, 21(7):988–1003, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Curran:2009:WMH

[CS09] Oisín Curran and Andy Shearer. A workflow model for heterogeneous computing environments. Future Generation Computer Systems, 25 (4):414–425, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Courtin:1992:SKB

[CSC^{+}92] Jean-Pierre Courtin, Pierre-Yves Schmerber, Jean Christophe Chautard, Sylvain Martin, Pascal Marrot, François Arlabosse, and Nathalie Porté. SATEXPERT: a knowledgebased system for spacecraft control. Future Generation Computer Systems, 7(4): 453–464, May 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cai:2005:BBC

[CSC^{+}05] Yang Cai, Ingo Snel, Betty Cheng, B. Suman Bharathi, Clementine Klein, and Judith Klein-Seetharaman. BioSim — a biomedical characterbased problem solving environment. Future Generation Computer Systems, 21(7): 1145–1156, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cao:2005:GLB

[CSJN05] Junwei Cao, Daniel P. Spooner, Stephen A. Jarvis, and Graham R. Nudd. Grid load balancing using intelligent agents. Future Generation Computer Systems, 21(1): 135–149, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cordsen:1998:SMC

[CSP98] J. Cordsen and W. Schröder-Preikschat. On the symbiosis of memory and communication in the programming of parallel applications. Future Generation Computer Systems, 13(4–5): 373–383, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/29/abstract.
html.

Cannataro:1991:PLS

[CST91] M. Cannataro, G. Spezzano, and D. Talia. A parallel logic system on a multicomputer architecture. Future Generation Computer Systems, 6(4):317– 331, September 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cannataro:1992:HLC

[CST92] M. Cannataro, G. Spezzano, and D. Talia. High level communication mechanisms for distributed parallel computers from an adaptive message routing. Future Generation Computer Systems, 8(1– 3):253–255, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carrington:2006:PPF

[CSW06] Laura Carrington, Allan Snavely, and Nicole Wolter. A performance prediction framework for scientific applications. Future Generation Computer Systems, 22(3): 336–346, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chien:2009:IRM

[CT09] Andrew A. Chien and Nut Taesombut. Integrated resource management for Lambda–Grids: The Distributed Virtual Computer (DVC). Future Generation Computer Systems, 25(2): 147–152, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Carothers:1999:VPS

[CTF^{+}99] C. D. Carothers, Brad Topol, R. M. Fujimoto, J. T. Stasko, and Vaidy Sunderam. Visualizing parallel simulations that execute in network computing environments. Future Generation Computer Systems, 15(4): 513–529, July 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/20/21/abstract.
html.

Cleall:2006:UPC

[CTMO06] P. J. Cleall, H. R. Thomas, T. A. Melhuish, and D. H. Owen. Use of parallel computing and visualisation techniques in the simulation of large scale geoenvironmental engineering problems. Future Generation Computer Systems, 22(4):460–467, March 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cannataro:2002:DDM

[CTT02] Mario Cannataro, Domenico Talia, and Paolo Trunfio. Distributed data mining on the grid. Future Generation Computer Systems, 18(8):1101– 1112, October 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Congiusta:2007:DDM

[CTT07] Antonio Congiusta, Domenico Talia, and Paolo Trunfio. Distributed data mining services leveraging WSRF. Future Generation Computer Systems, 23(1):34–41, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cannataro:2008:SSS

[CTT^{+}08a] Mario Cannataro, Domenico Talia, Giuseppe Tradigo, Paolo Trunfio, and Pierangelo Veltri. SIGMCC: a system for sharing meta patient records in a Peer-to-Peer environment. Future Generation Computer Systems, 24(3): 222–234, March 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:2008:LSA

[CTT^{+}08b] Dan Chen, Georgios K. Theodoropoulos, Stephen J. Turner, Wentong Cai, Robert Minson, and Yi Zhang. Large scale agent-based simulation on the Grid. Future Generation Computer Systems, 24 (7):658–671, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Curtis:1992:IES

[Cur92] Bill Curtis. Insights from empirical studies of the software design process. Future Generation Computer Systems, 7(2– 3):139–149, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cramb:1993:EDA

[CW93] Iain Cramb and Norman Winterbottom. Evaluation of a database application on a Meiko multisparc Oracle platform. Future Generation Computer Systems, 9 (1):25–31, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cao:2004:FUC

[CWD04] Jiannong Cao, Xianbing Wang, and Sajal K. Das. A framework of using cooperating mobile agents to achieve load sharing in distributed Web server groups. Future Generation Computer Systems, 20(4):591–603, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Castain:2008:ORT

[CWD^{+}08] R. H. Castain, T. S. Woodall, D. J. Daniel, J. M. Squyres, B. Barrett, and G. E. Fagg. The Open Run-Time Environment (OpenRTE): a transparent multicluster environment for high-performance computing. Future Generation Computer Systems, 24(2): 153–157, February 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chapin:1999:NMS

[CWW^{+}99] Steve J. Chapin, Chenxi Wang, William A. Wulf, Frederick Knabe, and Andrew Grimshaw. A new model of security for metasystems. Future Generation Computer Systems, 15(5–6): 713–722, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/31/abstract.
html.

Chenxi:1988:SMS

[CY88] Zhang Chenxi and Tzu Yungui. Study of mechanisms that support the implementation of nonlogical components of Prolog in WAM-based systems. Future Generation Computer Systems, 4(3):217–224, October 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chengzheng:1990:FBP

[CY90] Sun Chengzheng and Ci Yungui. The OR-forest-based parallel execution model of logic programs. Future Generation Computer Systems, 6 (1):25–34, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chi:2001:LBD

[CY01] Chi-Hung Chi and Jun-Li Yuan. Load-balancing data prefetching techniques. Future Generation Computer Systems, 17(6):733– 744, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/32/abstract.
html.

Chenxi:1990:IPD

[CYB90] Zhang Chenxi, Ci Yungui, and Liu Bo. Implementation of Prolog databases and database operation builtins in the WAM-Plus model. Future Generation Computer Systems, 6(1):3–9, June 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chang:2004:TSSb

[CYH04] Ting-Yi Chang, Chou-Chen Yang, and Min-Shiang Hwang. A threshold signature scheme for group communications without a shared distribution center. Future Generation Computer Systems, 20(6): 1013–1021, August 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Cai:2005:FMH

[CYLT05] Wentong Cai, Zijing Yuan, Malcolm Yoke Hean Low, and Stephen J. Turner. Federate migration in HLA-based simulation. Future Generation Computer Systems, 21(1):87– 95, January 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Daehn:1995:EEC

[Dae95]  Wilfried Daehn.  Elec-

trosmog & electromagnetic CAD. Future Generation Computer Systems, 11(4–5): 439–444, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DallOsso:2003:UCA

[Dal03] Aldo Dall’Osso. Using computer algebra systems in the development of scientific computer codes. Future Generation Computer Systems, 19(2): 143–160, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Daly:2006:HOE

[Dal06] J. T. Daly. A higher order estimate of the optimum checkpoint interval for restart dumps. Future Generation Computer Systems, 22 (3):303–312, February 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Deris:2008:ERD

[DAM08] Mustafa Mat Deris, Jemal H.

Abawajy, and Ali Mamat. An efficient replicated data access approach for large-scale distributed systems. Future Generation Computer Systems, 24 (1):1–9, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Datta:2003:KSM

[Dat03] Biswa Nath Datta. Krylov subspace methods for largescale matrix problems in control. Future Generation Computer Systems, 19(7):1253– 1263, October 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

deBoer:1990:CTL

[dB90] F. S. de Boer. Compositionality in the temporal logic of concurrent systems (extended abstract). Future Generation Computer Systems, 6(3):287– 299, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dzwinel:1999:MPV

[DB99] Witold Dzwinel and Jan Blasiak. Method of particles in visual clustering of multi-dimensional and large data sets. Future Generation Computer Systems, 15(3): 365–379, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/22/abstract.
html.

Demuynck:1998:VPS

[DBA98] Kris Demuynck, Jan Broeckhove, and Frans Arickx. The VE platform system: a system for distributed virtual reality. Future Generation Computer Systems, 14(3–4):193– 198, July 31, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/18/22/abstract.
html.

DeFanti:2003:IIV

[DBdL03] Thomas A. DeFanti, Maxine D. Brown, and Cees de Laat. iGrid 2002: The International Virtual Laboratory. Future Generation Computer Systems, 19(6): 803–804, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dorigo:2000:AAS

[DBT00] Marco Dorigo, Eric Bonabeau, and Guy Theraulaz. Ant algorithms and stigmergy. Future Generation Computer Systems, 16(8):851– 871, June 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
19/19/41/31/26/abstract.
html.

Dupuis:2000:OOA

[DC00] Alexandre Dupuis and Bastien Chopard. An object oriented approach to lattice gas modeling. Future Generation Computer Systems, 16(5): 523–532, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/35/abstract.
html.

Deonarine:2003:IML

[DCK03] Andrew S. Deonarine, Sonya M.

Clark, and Lars Konermann. Implementation of a multifunctional logic gate based on folding/unfolding transitions of a protein. Future Generation Computer Systems, 19 (1):87–97, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dawson:2000:KMN

[DCL00] Ed Dawson, Andrew Clark, and Mark Looi. Key management in a non-trusted distributed environment. Future Generation Computer Systems, 16(4):319–329, February 2000. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

19/19/41/27/27/abstract.
html.

DOnorioDeMeo:2007:HPG

[DCS^{+}07] Paolo D’Onorio De Meo, Danilo Carrabino, Nico Sanna, Tiziana Castrignano, Giorgio Grillo, Flavio Licciulli, Sabino Liuni, Matteo Re, Flavio Mignone, and Graziano Pesole. A high performance Grid–Web service framework for the identification of ‘conserved sequence tags’. Future Generation Computer Systems, 23(3):371–381, March 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dreyfus:1986:CSO

[DD86] Hubert L. Dreyfus and Stuart E. Dreyfus. Competent systems: The only future for inference-making computers. Future Generation Computer Systems, 2(4):233– 243, December 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dhaene:2005:SOM

[DD05] Tom Dhaene and Jan De Geest. Self-organizing multivariate constrained metamodeling technique for passive microwave and RF components. Future Generation Computer Systems, 21(7): 1040–1046, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dimkovski:2007:KTT

[DD07] Martin Dimkovski and Kevin Deeb. Knowledge technology through functional layered intelligence. Future Generation Computer Systems, 23(3): 295–303, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dormann:2001:FAT

[DDL01] Sabine Dormann, Andreas Deutsch, and Anna T. Lawniczak. Fourier analysis of Turinglike pattern formation in cellular automaton models. Future Generation Computer Systems, 17(7): 901–909, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/34/33/abstract.
html.

Dimov:2008:SSA

[DDM^{+}08] Ivan Dimov, Jack Dongarra, Kaj Madsen, Jerzy Wasniewski, and Zahari Zlatev. Special section: Applications of distributed and grid computing. Future Generation Computer Systems, 24(6): 582–584, June 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Danelutto:1992:MDS

[DDO^{+}92] Marco Danelutto, Roberto Di Meglio, Salvatore Orlando, Susanna Pelagatti, and Marco Vanneschi. A methodology for the development and the support of massively parallel programs. Future Generation Computer Systems, 8(1– 3):205–220, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). Also appears as “The P^{3}L language: an introduction”, Hewlett–Packard Report HPL-PSC-91-29, December 1991.

Djordjevic:2007:DSP

[DDR^{+}07] I. Djordjevic, T. Dimitrakos, N. Romano, D. Mac Randal, and P. Ritrovato. Dynamic security perimeters for interenterprise service integration. Future Generation Computer Systems, 23(4):633–657, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Decker:1996:MUR

[DDRR96] K. M. Decker, J. J. Dvorak, R. M. Rehmann, and R. Rühl. Matching user requirements in parallel programming. Future Generation Computer Systems, 12(4): 291–305, December 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dorigo:2000:GEA

[DDS00] Marco Dorigo, Gianni Di Caro, and Thomas Stützle. Guest editorial: Ant algorithms. Future Generation Computer Systems, 16 (8):v–vii, June 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/31/25/abstract.
html.

DeFanti:2009:STG

[DDS^{+}09] Thomas A. DeFanti, Gregory Dawe, Daniel J. Sandin, Jurgen P. Schulze, Peter Otto, Javier Girado, Falko Kuester, Larry Smarr, and Ramesh Rao. The StarCAVE, a thirdgeneration CAVE and virtual reality OptIPortal. Future Generation Computer Systems, 25(2):169–178, February 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Damiani:1992:DMA

[DDV92] E. Damiani, O. D’Antona, and Huu Le Van. The design of a microkernel application platform for shop-floor automation extended abstract. Future Generation Computer Systems, 8(1–3):257–259, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeCallatay:1988:LPD

[De 88] Armand De Callatay. Logic programs directly processed in a network of content addressable memories. Future Generation Computer Systems, 4 (2):117–131, September 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

deLaBourdonnaye:1994:NTR

[de 94] Armel de La Bourdonnaye. A new technique for reducing the cost of integral equations in axisymmetric cases. Future Generation Computer Systems, 10(4):411–417, November 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeLuca:1998:TTS

[De 98] Luigi De Luca. TeleEEG: a telemedical software package for EEG. Future Generation Computer Systems, 14(1–2): 61–66, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/29/17/22/abstract.
                                        html.

DelBuono:2003:CFL

[DE03] N. Del Buono and C. Elia. Computation of few Lyapunov exponents by geodesic based algorithms. Future Generation Computer Systems, 19(3): 425–430, April 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dekker:1986:IBD

[Dek86] W. Dekker. Issues basic to the development of a European information technology.

Future Generation Computer Systems, 2(1):33–37, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DelBuono:2006:DAS

[Del06]  N. Del Buono.  A dif-

ferential approach to solve the inverse eigenvalue problem derived from a neural network. Future Generation Computer Systems, 22(4): 441–446, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Domingo-Ferrer:1997:MAS

[DF97] Josep Domingo-Ferrer. Multiapplication smart cards and encrypted data, processing. Future Generation Computer Systems, 13(1):65– 74, June 20, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/17/21/abstract.
html.

DeRose:2008:ASU

[DFC^{+}08] César A. F. De Rose, Tiago Ferreto, Rodrigo N. Calheiros, Walfredo Cirne, Lauro B. Costa, and Daniel Fireman. Allocation strategies for utilization of space-shared resources in Bag of Tasks Grids. Future Generation Computer Systems, 24(5):331–341, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Deshpande:2000:VEM

[DFG^{+}00] Vaibhav Deshpande, Luciano Fornasier, Edgar A. Gerteisen, Nils Hilbrink, Andrey Mezentsev, Silvio Merazzi, and Thomas Wöhler. Virtual engineering of multi-disciplinary applications and the significance of seamless accessibility of geometry data. Future Generation Computer Systems, 16 (5):435–444, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/28/26/abstract.
                                        html.

Dekker:1988:OLD

[DFSZ88] L. Dekker, E. E. E. Frietman, W. Smit, and J. C. Zuidervaart. Optical link in the Delft Parallel Processor — an example of MOMI-connection in MIMD-supercomputers. Future Generation Computer Systems, 4(3):189–203, October 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Deves:1992:AES

[DFT92] Philippe Deves, Cécile Fischer, and Patrick Taillibert. An alternative to expert systems for electrical diagnosis. Future Generation Computer Systems, 7(4):343–352, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeRoure:2009:DRV

[DGS09] David De Roure, Carole Goble, and Robert Stevens. The design and realisation of the ^{my}Experiment Virtual Research Environment for social sharing of workflows. Future Generation Computer Systems, 25(5):561–567, May 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Deelman:2009:WSO

[DGST09] Ewa Deelman, Dennis Gannon, Matthew Shields, and Ian Taylor. Workflows and escience: An overview of workflow system features and capabilities. Future Generation Computer Systems, 25 (5):528–540, May 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Das:2002:MLT

[DHB02] Sajal K. Das, Daniel J.

Harvey, and Rupak Biswas. MinEX: a latency-tolerant dynamic partitioner for grid computing applications. Future Generation Computer Systems, 18(4):477–489, March 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.elsevier.com/gej-ng/
10/19/19/60/33/31/abstract.
html.

Dekker:1989:ASN

[DHD89] T. J. Dekker, W. Hoffmann, and P. P. M. De Rijk. Algorithms for solving numerical linear algebra problems on supercomputers. Future Generation Computer Systems, 4 (4):255–263, March 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Donaldson:1999:CPO

[DHS99] Stephen R. Donaldson, Jonathan M. D. Hill, and David B. Skillicorn. Communication performance optimisation requires minimising variance. Future Generation Computer Systems, 15(3):453– 459, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/19/30/abstract.
                                        html.

Diehl:2000:GEP

[DHS00a] S. Diehl, P. H. Hartel, and P. Sestoft. Guest editorial: Principles of abstract machines. Future Generation Computer Systems, 16 (7):v–vi, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/30/25/abstract.
                                        html.

Diehl:2000:AMP

[DHS00b] Stephan Diehl, Pieter Hartel, and Peter Sestoft. Abstract machines for programming language implementation. Future Generation Computer Systems, 16(7): 739–751, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/30/26/abstract.
html.

Dong:2006:SVH

[DIK^{+}06] Suchuan Dong, Joseph Insley, Nicholas T. Karonis, Michael E. Papka, Justin Binns, and George Karniadakis. Simulating and visualizing the human arterial system on the TeraGrid. Future Generation Computer Systems, 22(8):1011–1017, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dinn:1991:NA

[Din91] Neil F. Dinn. Network architectures. Future Generation Computer Systems, 7(1): 79–89, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ding:1999:HPF

[Din99] Chris H. Q. Ding. High Performance Fortran for practical scientific algorithms: an up-to-date evaluation. Future Generation Computer Systems, 15(3):343–352, April 1, 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/20/abstract.
html.

Dini:2003:SAE

[Din03] Gianluca Dini. A secure and available electronic voting service for a large-scale distributed system. Future Generation Computer Systems, 19 (1):69–85, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Diehl:2000:VPA

[DK00] Stephan Diehl and Thomas Kunze. Visualizing principles of abstract machines by generating interactive animations. Future Generation Computer Systems, 16 (7):831–839, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/30/31/abstract.
html.

DiMartino:2008:SSG

[DKD08] Beniamino Di Martino, Dieter Kranzlmüller, and Jack Dongarra. Special section: Grid computing and the Message Passing Interface. Future Generation Computer Systems, 24 (2):119–120, February 2008.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

deKergommeaux:2003:FPV

[dKdOS03] J. Chassin de Kergommeaux and B. de Oliveira Stein. Flexible performance visualization of parallel and distributed applications. Future Generation Computer Systems, 19 (5):735–747, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dini:2000:SOD

[DL00] Gianluca Dini and Lanfranco Lopriore. Sharing objects in a distributed, single address space environment. Future Generation Computer Systems, 17(3): 247–264, November 1, 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/27/abstract.
html.

DelBuono:2003:DAC

[DL03] N. Del Buono and L. Lopez.

Differential approaches for computing Euclidean diagonal norm balanced realizations in control theory. Future Generation Computer Systems, 19 (7):1155–1163, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dong:2004:HXS

[DL04] Ying Dong and Mingshu Li. HyO-XTM: a set of hypergraph operations on XML Topic Map toward knowledge management. Future Generation Computer Systems, 20 (1):81–100, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

deSaintVincent:1993:PKB

[dLLA93] A. de Saint Vincent, F. Lecouat, G. Leonis, and F. Allard. PREVISE: a knowledge-based system to support the preparation and verification of space operations procedures. Future Generation Computer Systems, 9(4):287–301, December 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DelBuono:2006:SSN

[DLP06] N. Del Buono, L. Lopez, and T. Politi. Special section: Numerical methods for structured systems. Future Generation Computer Systems, 22 (4):393–394, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeFanti:2009:OSV

[DLR^{+}09] Thomas A. DeFanti, Jason Leigh, Luc Renambot, Byungil Jeong, Alan Verlo, Lance Long, Maxine Brown, Daniel J. Sandin, Venkatram Vishwanath, Qian Liu, Mason J. Katz, Philip Papadopoulos, Joseph P. Keefe, Gregory R. Hidley, Gregory L. Dawe, Ian Kaufman, Bryan Glogowski, Kai-Uwe Doerr, Rajvikram Singh, Javier Girado, Jurgen P. Schulze, Falko Kuester, and Larry Smarr. The OptIPortal, a scalable visualization, storage, and computing interface device for the OptIPuter. Future Generation Computer Systems, 25(2): 114–123, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

deLaat:2003:RCO

[dLRW03] Cees de Laat, Erik Radius, and Steven Wallace. The rationale of the current optical networking initiatives. Future Generation Computer Systems, 19(6):999–1008, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

delaTorre:1992:TSM

[dlTK92] Pilar de la Torre and Clyde P.

Kruskal. Towards a single model of efficient computation in real parallel machines. Future Generation Computer Systems, 8(4):395– 408, September 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Diel:1986:SSP

[DLW86] H. Diel, N. Lenz, and H. M.

Welsch. System structure for parallel logic programming. Future Generation Computer Systems, 2(4):225– 231, December 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dai:2007:OTP

[DLW07] Yuan-Shun Dai, Gregory Levitin, and Xiaolong Wang. Optimal task partition and distribution in grid service system with common cause failures. Future Generation Computer Systems, 23(2):209– 218, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Demchenko:2008:DSC

[DMG^{+}08] Yuri Demchenko, Olle Mulmo, Leon Gommans, Cees de Laat, and Alfred Wan. Dynamic security context management in Grid-based applications. Future Generation Computer Systems, 24(5):434–441, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Delibasis:1999:MFC

[DMM^{+}99] K. K. Delibasis, N. Mouravliansky, G. K. Matsopoulos, K. S. Nikita, and A. Marsh. MR functional cardiac imaging: Segmentation, measurement and WWW based visualisation of 4D data. Future Generation Computer Systems, 15 (2):185–193, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/17/21/abstract.
html.

Delibasis:1998:RDW

[DMMP98] K. Delibasis, Ch. Michael, N. Mouravliansky, and K. Papaodysseas. From raw data to WWW compatible visualization and manipulation imaging human electric potentials. Future Generation Computer Systems, 14(1–2): 67–77, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/23/abstract.
html.

Dudkiewicz:2005:CBM

[DMN^{+}05] M. Dudkiewicz, P. Mackiewicz, A. Nowicka, M. Kowalczuk, D. Mackiewicz, N. Polak, K. Smolarczyk, J. Banaszak, M. R. Dudek, and S. Cebrat. Correspondence between mutation and selection pressure and the genetic code degeneracy in the gene evolution. Future Generation Computer Systems, 21(7): 1033–1039, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DHeygere:1993:QDS

[DMR93] François D’Heygère, Pierre Mariot, and Jean-Baptiste Renard. QUATRAIN — a design support tool and a data processing sequence supervisor. Future Generation Computer Systems, 9(4):321– 328, December 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dongarra:1997:CTH

[DMSS97] J. J. Dongarra, H. W. Meuer, H. D. Simon, and E. Strohmaier. Changing technologies of HPC. Future Generation Computer Systems, 12 (5):461–474, April 1, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/27/17/28/abstract.
                                        html;   http://www.netlib.
                                        org/utk/people/JackDongarra/
                                        PAPERS/ct-hpc.ps;   http:
                                        //www.netlib.org/utk/people/
                                        JackDongarra/pdf/ct-hpc.
                                        pdf.

Dongarra:2004:SNA

[DMW04] Jack Dongarra, Kaj Madsen, and Jerzy Waśniewski. Selected numerical algorithms. Future Generation Computer Systems, 20(3):349–351, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DiStefano:2009:PSQ

[DMZ09] Antonella Di Stefano, Giovanni Morana, and Daniele Zito. A P2P strategy for QoS discovery and SLA negotiation in Grid environment. Future Generation Computer Systems, 25(8): 862–875, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

doNascimento:2005:UHF

[dNE05] Hugo A. D. do Nascimento and Peter Eades. User hints: a framework for interactive optimization. Future Generation Computer Systems, 21(7): 1177–1191, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dogan:2009:SPD

[Doḡ09] Atakan Doḡan. A study on performance of dynamic file replication algorithms for real-time file access in Data Grids. Future Generation Computer Systems, 25(8): 829–839, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Doring:2005:RDD

[Dör05] Andreas C. Döring. Routing direction determination in regular networks based on configurable circuits. Future Generation Computer Systems, 21(8):1312–1321, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Drummond:2001:APM

[DOV01] Lúcia M. A. Drummond, Luiz S. Ochi, and Dalessandro S. Vianna. An asynchronous parallel metaheuristic for the period vehicle routing problem. Future Generation Computer Systems, 17 (4):379–386, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/29/28/abstract.
                                        html.

DelBuono:2003:GNA

[DP03a] N. Del Buono and T. Politi. Geometric numerical algorithms. Future Generation Computer Systems, 19(3): 327–329, April 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dieci:2003:CE

[DP03b] L. Dieci and A. Papini. Continuation of eigendecompositions. Future Generation Computer Systems, 19 (7):1125–1137, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Denis:2003:POI

[DPP03] Alexandre Denis, Christian Pérez, and Thierry Priol. PadicoTM: an open integration framework for communication middleware and runtimes. Future Generation Computer Systems, 19 (4):575–585, May 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dhem:1997:LCA

[DQ97] J. F. Dhem and J. J.

Quisquater. Lossless compression algorithms for smart cards: A progress report. Future Generation Computer Systems, 13(1):27–38, June 1997. URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X97891091.

Dixon:1989:NCI

[DR89] David A. Dixon and Harold J.

Raveché. A national computing initiative: a summary. Future Generation Computer Systems, 5(2–3): 339–345, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Diele:2003:GEM

[DR03] F. Diele and S. Ragni. The global error of Magnus methods based on the Cayley map for some oscillatory problems. Future Generation Computer Systems, 19(3):385–393, April 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Doroshenko:2005:RCG

[DR05] A. Doroshenko and D. Ragozin.

Retargetable code generation for application-specific processors. Future Generation Computer Systems, 21(5):679–685, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Diaz:2009:DSS

[DRNMC09] Javier Díaz, Sebastián Reyes, Alfonso Niño, and Camelia Muñoz-Caro. Derivation of self-scheduling algorithms for heterogeneous distributed computer systems: Application to Internet-based Grids of computers. Future Generation Computer Systems, 25 (6):617–626, June 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DiGregorio:1997:HPS

[DRS^{+}97] S. Di Gregorio, R. Rongo, W. Spataro, G. Spezzano, and D. Talia. High performance scientific computing by a parallel cellular environment. Future Generation Computer Systems, 12(5): 357–369, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/20/abstract.
html.

DiStefano:2004:PNC

[DRS04] Marco Di Stefano, Marzio Rosi, and Antonio Sgamellotti. Phenylium and naphtylium cations in the interstellar medium: a density functional study on their reactivity towards D_{2} molecules. Future Generation Computer Systems, 20(5):807–819, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

deRonde:1994:CWC

[dRSBH94] J. F. de Ronde, P. M. A.

Sloot, M. Beemster, and L. O. Hertzberger. The CA-MAS workbench: Computer Aided Migration of Applications System. Future Generation Computer Systems, 10(2– 3):305–308, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

deRonde:1997:LBR

[dRSS97] J. F. de Ronde, A. Schoneveld, and P. M. A. Sloot. Load balancing by redundant decomposition and mapping. Future Generation Computer Systems, 12(5):391– 406, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/22/abstract.
html.

DiGregorio:1999:EMM

[DS99] Salvatore Di Gregorio and Roberto Serra. An empirical method for modelling and simulating some complex macroscopic phenomena by cellular automata. Future Generation Computer Systems, 16(2–3): 259–271, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/33/abstract.
html.

Demoen:2000:CCH

[DS00] Bart Demoen and Konstantinos Sagonas. CHAT: the copy-hybrid approach to tabling. Future Generation Computer Systems, 16(7): 809–830, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/30/30/abstract.
html.

Debelov:2004:LMS

[DS04a] Victor A. Debelov and Igor Sevastianov. Light mesh: soft shadows as interpolation of visibility. Future Generation Computer Systems, 20(8): 1299–1315, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DOrazio:2004:STD

[DS04b] Annunziata D’Orazio and Sauro Succi. Simulating twodimensional thermal channel flows by means of a lattice Boltzmann method with new boundary conditions. Future Generation Computer Systems, 20(6):935–944, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dreo:2004:CIA

[DS04c] J. Dréo and P. Siarry. Continuous interacting ant colony algorithm based on dense heterarchy. Future Generation Computer Systems, 20(5): 841–856, June 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DiStefano:2008:EMR

[DS08] Antonella Di Stefano and Corrado Santoro. An economic model for resource management in a Grid-based content distribution network. Future Generation Computer Systems, 24(3):202–212, March 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Donato:1999:MMD

[DSH^{+}99] June M. Donato, Jack C.

Schryver, Gregory C. Hinkel, Richard L. Schmoyer, Jr., Michael R. Leuze, and Nancy W. Grandy. Mining multi-dimensional data for decision support. Future Generation Computer Systems, 15 (3):433–441, April 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/28/abstract.
html.

deSturler:1998:PIS

[dSL98] E. de Sturler and D. Loher. Parallel iterative solvers for irregular sparse matrices in High Performance Fortran. Future Generation Computer Systems, 13(4–5):315– 325, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/24/abstract.
html.

Dahlgren:1998:EHB

[DSS98] Fredrik Dahlgren, Jonas Skeppstedt, and Per Stenström. An evaluation of hardware-based and compilercontrolled optimizations of snooping cache protocols. Future Generation Computer Systems, 13(6):469–487, May 20, 1998. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/20/20/abstract.
html.

Domingues:2007:STT

[DSS07] Patricio Domingues, Bruno Sousa, and Luis Moura Silva. Sabotage-tolerance and trust management in desktop grid computing. Future Generation Computer Systems, 23(7): 904–912, August 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dunkel:1997:SKC

[DSSU97] Brian Dunkel, Nandit Soparkar, John Szaro, and Ramasamy Uthurusamy. Systems for KDD: From concepts to practice. Future Generation Computer Systems, 13(2–3):231– 242, November 14, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/25/abstract.
html.

Diaz:1993:PBR

[DT93] M. Díaz and J. M. Troya.

A Parlog based real-time distributed logic environment. Future Generation Computer Systems, 9(3):201– 218, September 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Desprez:1994:LLO

[DT94] F. Desprez and B. Tourancheau.

LOCCS: Low overhead communication and computation subroutines. Future Generation Computer Systems, 10(2– 3):279–284, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dongarra:2008:SSC

[DT08] Jack Dongarra and Bernard Tourancheau. Special section: Cluster and computational grids for scientific computing. Future Generation Computer Systems, 24 (1):30, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Duato:1994:IEV

[Dua94] José Duato. Improving the efficiency of virtual channels with time-dependent selection functions. Future Generation Computer Systems, 10 (1):45–58, April 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Duby:1991:EIT

[Dub91] Jean-Jacques Duby. The evolution of information technologies in the 90s and its impact on applications. Future Generation Computer Systems, 7 (1):15–21, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Duinker:1989:SPA

[Dui89] W. Duinker. Synchronous processor arrays integrated in a transputer network. Future Generation Computer Systems, 4(4):265–269, March 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Duprat:1990:LPE

[Dup90] Jean Duprat. LAIOS: a parallel execution of PROLOG by data copies. Future Generation Computer Systems, 6(2): 131–144, November 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dieci:2003:OIB

[DV03] Luca Dieci and Erik S. Van Vleck. Orthonormal integrators based on Householder and Givens transformations. Future Generation Computer Systems, 19(3):363–373, April 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dijkstra:2006:UZC

[DvdHdL06] Freek Dijkstra, Jeroen J.

van der Ham, and Cees T. A. M. de Laat. Using zero configuration technology for IP addressing in optical networks. Future Generation Computer Systems, 22(8): 908–914, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dijkstra:2009:PFI

[DvdHGdL09] Freek Dijkstra, Jeroen van der Ham, Paola Grosso, and Cees de Laat. A path finding implementation for multi-layer networks. Future Generation Computer Systems, 25(2): 142–146, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeCarlini:1990:MIP

[DVV90] U. De Carlini, R. Vaccaro, and U. Villano. The monitoring of inter-process communications in distributed systems. Future Generation Computer Systems, 5(4):359–363, January 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

DeTurck:2002:GMB

[DVVD02] F. De Turck, S. Vanhastel, B. Volckaert, and P. Demeester. A generic middleware-based platform for scalable cluster computing. Future Generation Computer Systems, 18(4):549– 560, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/33/36/abstract.
                                        html.

Doukidis:1987:DRE

[DW87] Georgios I. Doukidis and Edgar A. Whitley. Developing and running expert systems with PESYS. Future Generation Computer Systems, 3 (3):189–199, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dupuis:2004:LBM

[DY04] A. Dupuis and J. M. Yeomans. Lattice Boltzmann modelling of droplets on chemically heterogeneous surfaces.

Future Generation Computer Systems, 20(6):993–1001, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dantas:1998:ESM

[DZ98] M. A. R. Dantas and E. J. Zaluska. Efficient scheduling of MPI applications on networks of workstations. Future Generation Computer Systems, 13 (6):489–499, May 20, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/20/21/abstract.
html.

Dandamudi:2004:PAS

[DZ04] Sivarama P. Dandamudi and Zhengao Zhou. Performance of adaptive space-sharing policies in dedicated heterogeneous cluster systems. Future Generation Computer Systems, 20(5):895–906, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Delaitre:2000:EIE

[DZJ^{+}00] T. Delaitre, M. J. Zemerly, G. R. Justo, O. Audo, and S. C. Winter. EDPEPPS: an integrated environment for the parallel development life-cycle. Future Generation Computer Systems, 16 (6):585–595, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/27/abstract.
html.

Dzwinel:1997:VPS

[Dzw97] W. Dzwinel. Virtual particles and search for global minimum. Future Generation Computer Systems, 12(5): 371–389, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/21/abstract.
html.

Engin:2004:NAS

[ED04] Orhan Engin and Alper Döyen. A new approach to solve hybrid flow shop scheduling problems by artificial immune system. Future Generation Computer Systems, 20(6):1083–1095, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ezquerra:1999:IKG

[EdBG^{+}99] N. Ezquerra, L. de Braal, E. Garcia, C. Cooke, and E. Krawczynska. Interactive, knowledge-guided visualization of 3D medical imagery. Future Generation Computer Systems, 15(1): 59–73, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/18/21/29/

index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/18/21/abstract.
html.

Ercan:2000:PIP

[EFD00] M. Fikret Ercan, Yu-Fai Fung, and M. Suleyman Demokan. Parallel image processing with one-dimensional DSP arrays. Future Generation Computer Systems, 17(3):197–214, November 1, 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/24/abstract.
html.

Espert:2009:CBO

[EGAQ09] Ignacio Blanquer Espert, Vicente Hernández García, Fco. Javier Meseguer Anastásio, and J. Damià Segrelles Quilis. Content-based organisation of virtual repositories of DI-COM objects. Future Generation Computer Systems, 25 (6):627–637, June 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

El-Ghazawi:2006:BPC

[EGCY^{+}06] Tarek A. El-Ghazawi, François Cantonnet, Yiyi Yao, Smita Annareddy, and Ahmed S. Mohamed. Benchmarking parallel compilers: a UPC case study. Future Generation Computer Systems, 22(7): 764–775, August 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ellert:2007:ARC

[EGK^{+}07] M. Ellert, M. Grønager, A. Konstantinov, B. Kónya, J. Lindemann, I. Livenson, J. L. Nielsen, M. Niinimäki, O. Smirnova, and A. Wäänänen. Advanced Resource Connector middleware for lightweight computational Grids. Future Generation Computer Systems, 23(2): 219–240, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ellison:2000:PSK

[EHMS00] Carl Ellison, Chris Hall, Randy Milbert, and Bruce Schneier. Protecting secret keys with personal entropy. Future Generation Computer Systems, 16 (4):311–318, February 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.counterpane.
                                        com/personal-entropy.pdf;
                                         http://www.elsevier.com/
                                        gej-ng/10/19/19/41/27/26/
                                        abstract.html.

El-Khalili:2000:STW

[EKB00] Nuha H. El-Khalili and Ken W. Brodlie. Surgical training on the Web. Future Generation Computer Systems, 17(2):147–158, October 2000. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/27/29/abstract.
html.

Elrad:1998:EPE

[EL98] Tzilla Elrad and Jinlong Lin.

Evolving processes and evolution schedulers for concurrent scheduling controls and parallel evolutionary computation. Future Generation Computer Systems, 14(5–6): 365–381, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Elia:2003:EMQ

[EL03] C. Elia and L. Lopez. Exponential monotonicity of quadratic forms in ODEs and preserving methods. Future Generation Computer Systems, 19(7):1187–1195, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Epema:1996:WFC

[ELvD^{+}96] D. H. J. Epema, M. Livny, R. van Dantzig, X. Evers, and J. Pruyne. A worldwide flock of Condors: Load sharing among workstation clusters. Future Generation Computer Systems, 12(1):53–65, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Emmen:1998:SVA

[EMB98] Ad Emmen, Martijn Mulder, and Ion Barosan. Supporting 3D and VR applications in a metacomputing environment. Future Generation Computer Systems, 14(3–4): 185–192, July 31, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/18/21/abstract.
html.

Ernst:1986:BAA

[EO86] Martin L. Ernst and Helen Ojha. Business applications of artificial intelligence knowledge based expert systems. Future Generation Computer Systems, 2(3):173– 185, September 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Eres:2005:IUG

[EPJ^{+}05] M. Hakki Eres, Graeme E. Pound, Zhouan Jiao, Jasmin L. Wason, Fenglian Xu, Andy J. Keane, and Simon J. Cox. Implementation and utilisation of a Grid-enabled problem solving environment in Matlab. Future Generation Computer Systems, 21 (6):920–929, June 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Etiemble:1994:PAL

[ES94] Daniel Etiemble and Jean Claude Syre. Parallel architecture and language in Europe. Future Generation Computer Systems, 10(1):1–2, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Emad:2006:AAN

[ESFD06] Nahid Emad, S.-A. Shahzadeh-Fazeli, and Jack Dongarra. An asynchronous algorithm on the NetSolve global computing system. Future Generation Computer Systems, 22 (3):279–290, February 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Evripidou:2001:PMP

[ESPP01] Paraskevas Evripidou, George Samaras, Christoforos Panayiotou, and Evaggelia Pitoura. The PaCMAn Metacomputer: parallel computing with Java mobile agents. Future Generation Computer Systems, 18(2): 265–280, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/33/abstract.
html.

Elmroth:2008:GRB

[ET08] Erik Elmroth and Johan Tordsson. Grid resource brokering algorithms enabling advance reservations and resource selection based on performance predictions. Future Generation Computer Systems, 24(6):585–593, June 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Elrad:1996:HRF

[EV96] Tzilla Elrad and Ufuk Verün. A hierarchical and reflective framework for synchronization and scheduling controls. Future Generation Computer Systems, 12(2–3):243– 256, September 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Emmen:1998:VMW

[EV98] Ad Emmen and Leslie Versweyveld. Virtual Medical Worlds Magazine — a Euromed on-line information building block for HPCN related telemedicine. Future Generation Computer Systems, 14(1–2):131–136, June 15, 1998. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/28/abstract.
html.

Emmen:1999:IF

[EV99] Ad Emmen and Leslie Versweyveld. The ITIS’98 forums. Future Generation Computer Systems, 15(2): 157–169, March 11, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/17/19/abstract.
html.

Eshaghian:1997:REH

[EW97] Mary M. Eshaghian and Ying-Chieh Wu. Resource estimation for heterogeneous computing. Future Generation Computer Systems, 12(6): 505–520, June 15, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/18/18/abstract.
html.

Ertl:1999:MHM

[EWG99] T. Ertl, R. Westermann, and R. Grosso. Multiresolution and hierarchical methods for the visualization of volume data. Future Generation Computer Systems, 15 (1):31–42, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/18/19/29/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/18/19/abstract.
html.

Fahringer:1998:SAT

[Fah98]  Thomas Fahringer.  Sym-

bolic analysis techniques for program parallelization. Future Generation Computer Systems, 13(4–5):385–396, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/19/30/abstract.
                                        html.

Feig:1999:LSD

[FAJP99] Michael Feig, Matin Abdullah, Lennart Johnsson, and B. Montgomery Pettitt. Large scale distributed data repository: design of a molecular dynamics trajectory database. Future Generation Computer Systems, 16 (1):101–110, November 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/32/35/abstract.
                                        html.

Faure:2005:ADP

[Fau05]  Christèle Faure.  An au-

tomatic differentiation platform: Odyssée. Future Generation Computer Systems, 21 (8):1391–1400, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fualdes:1993:CFR

[FB93] Thierry D. Fualdes and Claude J. Barrouil. A common framework for reasoning on uncertainty both at symbolic and numerical levels. Future Generation Computer Systems, 9(4):339–347, December 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fang:1997:MDD

[FB97] Niandong Fang and Helmar Burkhart. MPI-DDL: a distributed-data library for MPI. Future Generation Computer Systems, 12(5):407– 419, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/23/abstract.
html.

Forss:1999:MWD

[FBBW99] Jenny Forss, David Beeman, James M. Bower, and Rogene M. Eichler West. The Modeler’s Workspace: a distributed digital library for neuroscience. Future Generation Computer Systems, 16 (1):111–121, November 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/36/abstract.
html.

Fagan:2005:RRM

[FC05] Mike Fagan and Alan Carle.

Reducing reverse-mode memory requirements by using profile-driven checkpointing. Future Generation Computer Systems, 21(8):1380– 1390, October 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Farina:2009:DSD

[FC09] Fabio Farina and Silvia Calegari. Distributed semantic document retrieval using O-FCN. Future Generation Computer Systems, 25 (8):853–861, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fukunari:2001:BWJ

[FCW01] Miki Fukunari, Yu-Liang Chi, and Philip M. Wolfe. Best of Websim99: JavaBeanbased simulation with operational procedure table (OPT). Future Generation Computer Systems, 17(5): 513–523, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/30/27/abstract.
                                        html.

Fiske:1995:TPT

[FD95] Stuart Fiske and William J. Dally. Thread prioritization: a thread scheduling mechanism for multiple-context parallel processors. Future Generation Computer Systems, 11(6): 503–518, October 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fagg:2002:HFT

[FD02] Graham E. Fagg and Jack J.

Dongarra. HARNESS fault tolerant MPI design, usage and performance issues. Future Generation Computer Systems, 18(8):1127– 1142, October 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Furtado:2007:PPS

[FdSC07] Vasco Furtado, Francisco Flávio de Souza, and Walfredo Cirne. Promoting performance and separation of concerns for data mining applications on the Grid. Future Generation Computer Systems, 23 (1):100–106, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fernbach:1984:SPP

[Fer84]  Sidney Fernbach.  Super-

computers — past, present, prospects. Future Generation Computer Systems, 1 (1):23–30, July 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ferstl:1996:JRM

[Fer96] F. Ferstl. Job- and resourcemanagement systems in heterogeneous clusters. Future Generation Computer Systems, 12(1):39–51, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fernandez:2007:ETC

[FGCM07] José-Jesús Fernández, Inmaculada García, Jose-María Carazo, and Roberto Marabini. Electron tomography of complex biological specimens on the Grid. Future Generation Computer Systems, 23 (3):435–446, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fernandez:2003:FPA

[FGG03] José-Jesús Fernández, Inmaculada García, and Ester M. Garzón. Floating point arithmetic teaching for computational science. Future Generation Computer Systems, 19(8): 1321–1334, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Falchi:2009:DBD

[FGRZ09] Fabrizio Falchi, Claudio Gennaro, Fausto Rabitti, and Pavel Zezula. Distance browsing in distributed multimedia databases. Future Generation Computer Systems, 25(1): 64–76, January 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fowler:1995:BRI

[FHG95a] R. F. Fowler, B. W. Henderson, and C. Greenough. BBS results for the iPSC/2 and iPSC/860. Future Generation Computer Systems, 11(1): 49–59, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fowler:1995:PTD

[FHG95b] R. F. Fowler, B. W. Henderson, and C. Greenough. Porting a three-dimensional semiconductor device modelling program to the Intel iPSC/860 hypercube. Future Generation Computer Systems, 11(1): 61–70, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Floros:1999:PRM

[FHM^{+}99] N. Floros, A. J. G. Hey, K. E. Meacham, J. Papay, and M. Surridge. Predictive resource management for metaapplications. Future Generation Computer Systems, 15(5– 6):723–734, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/32/abstract.
html.

Finn:1999:LIG

[Fin99] Lee Samuel Finn. The laser interferometer gravitationalwave observatory scientific data archive. Future Generation Computer Systems, 16 (1):123–134, November 1999.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/32/37/abstract.
                                        html.

Fiori:2006:FPN

[Fio06] Simone Fiori. Fixed-point neural independent component analysis algorithms on the orthogonal group. Future Generation Computer Systems, 22(4):430–440, March 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fishwick:2000:P

[Fis00]  Paul A. Fishwick.  Pref-

ace. Future Generation Computer Systems, 17(2): v, October 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/27/24/abstract.
                                        html.

Ferscha:2000:MEP

[FJ00] Alois Ferscha and James Johnson. N-MAP — an environment for the performance oriented development process of efficient distributed programs. Future Generation Computer Systems, 16 (6):571–584, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

19/19/41/29/26/abstract.
html.

Ferscha:2001:DSP

[FJT01] Alois Ferscha, James Johnson, and Stephen J. Turner. Distributed simulation performance data mining. Future Generation Computer Systems, 18(1):157–174, September 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/42/abstract.
html.

Falgout:2006:CIH

[FJY06] Robert D. Falgout, Jim E.

Jones, and Ulrike Meier Yang. Conceptual interfaces in hypre. Future Generation Computer Systems, 22(1–2): 239–251, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Foster:1999:GPS

[FK99] Ian Foster and Carl Kesselman. The Globus project: a status report. Future Generation Computer Systems, 15(5– 6):607–621, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/23/abstract.
html.

Fox:2005:MBC

[FLPP05] Geoffrey Fox, Sang Lim, Shrideep Pallickara, and Marlon Pierce. Message-based cellular peer-to-peer grids: foundations for secure federation and autonomic services. Future Generation Computer Systems, 21(3): 401–415, March 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fluckiger:2003:GE

[Flu03] François Fluckiger. Guest editorial. Future Generation Computer Systems, 19(2): 187–189, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ferscha:2001:EPD

[FM01] Alois Ferscha and Allen D. Malony. Editorial: Performance data mining — automated diagnosis, adaption, and optimization. Future Generation Computer Systems, 18 (1):127–130, September 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/27/39/abstract.
                                        html.

Fortino:2008:SSE

[FM08] Giancarlo Fortino and Carlo Mastroianni. Special section: Enhancing content networks with P2P, Grid and Agent technologies. Future Generation Computer Systems, 24(3): 177–179, March 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fagg:1999:SNI

[FMD99] Graham E. Fagg, Keith Moore, and Jack J. Dongarra. Scalable Networked Information Processing Environment (SNIPE). Future Generation Computer Systems, 15(5–6): 595–605, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/22/abstract.
html;   http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/snipe-fgcs.ps.

Fortino:2005:CCM

[FMR05] Giancarlo Fortino, Carlo Mastroianni, and Wilma Russo. Cooperative control of multicast-based streaming ondemand systems. Future Generation Computer Systems, 21 (5):823–839, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Forestiero:2008:QBD

[FMS08a] Agostino Forestiero, Carlo Mastroianni, and Giandomenico Spezzano. QoS-based dissemination of content in Grids. Future Generation Computer Systems, 24(3):

235–244, March 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Forestiero:2008:RDG

[FMS08b] Agostino Forestiero, Carlo Mastroianni, and Giandomenico Spezzano. Reorganization and discovery of grid information with epidemic tuning. Future Generation Computer Systems, 24(8):788–797, October 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fritzson:2000:ASS

[FN00] Dag Fritzson and Patrik Nordling. Adaptive scheduling strategy optimizer for parallel rolling bearing simulation. Future Generation Computer Systems, 16(5): 563–570, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/28/39/abstract.
                                        html.

Fromme:2003:ASI

[FP03] Michael Fromme and Helmut Pralle. Architecture of a shared-image electronic whiteboard in telemedicine. Future Generation Computer Systems, 19(2):241–249, February 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Facin:2004:NLL

[FPdS04] Paulo C. Facin, Paulo C.

Philippi, and Luís O. E. dos Santos. A non-linear lattice-Boltzmann model for ideal miscible fluids. Future Generation Computer Systems, 20(6):945–949, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fairman:2009:ESM

[FPX^{+}09] Matthew J. Fairman, Andrew R. Price, Gang Xue, Marc Molinari, Denis A. Nicole, Timothy M. Lenton, Robert Marsh, Kenji Takeda, and Simon J. Cox. Earth system modelling with Windows Workflow Foundation. Future Generation Computer Systems, 25(5):586–597, May 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fortino:2008:UPG

[FR08] Giancarlo Fortino and Wilma Russo. Using P2P, GRID and agent technologies for the development of content distribution networks. Future Generation Computer Systems, 24 (3):180–190, March 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Frattolillo:2008:SDM

[Fra08] Franco Frattolillo. Supporting data management on cluster grids. Future Generation Computer Systems, 24(2): 166–176, February 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fredriksson:1984:I

[Fre84] E. H. Fredriksson. Introduction. Future Generation Computer Systems, 1(1):3–7, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Frenkel:1994:CCC

[Fre94]  Daan Frenkel.  Computa-

tion challenges in complex liquids: Entropy-driven phase transitions. Future Generation Computer Systems, 10(2– 3):207–212, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Feigenbaum:1993:JNF

[FS93] Edward Feigenbaum and Howard Shrobe. The Japanese national Fifth Generation project: Introduction, survey, and evaluation. Future Generation Computer Systems, 9 (2):105–117, July 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fayyad:1997:DMK

[FS97] Usama Fayyad and Paul Stolorz. Data mining and KDD Promise and challenges. Future Generation Computer Systems, 13(2–3): 99–115, November 14, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/17/abstract.
html.

Folino:2007:ATB

[FS07] Gianluigi Folino and Giandomenico Spezzano. An autonomic tool for building selforganizing Grid-enabled applications. Future Generation Computer Systems, 23(5): 671–679, June 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Frankel:1988:SON

[FSM88] Michael Frankel, Nachum Shacham, and James E. Mathis. Self-organizing networks. Future Generation Computer Systems, 4(2):95– 115, September 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fahringer:2002:DRW

[FSP02] T. Fahringer and K. Sowa-Piekło. Debugging real-world data-parallel programs with SPiDER. Future Generation Computer Systems, 18 (6):779–788, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fiolet:2007:CMD

[FT07] Valerie Fiolet and Bernard Toursel. A clustering method to distribute a database on a Grid. Future Generation Computer Systems, 23 (8):997–1002, November 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fuchs:1993:AS

[Fuc93] J. Fuchs. AI and space. Future Generation Computer Systems, 9(4):285–286, December 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fukinuki:1985:DRP

[Fuk85]  Takahiko Fukinuki.  Data

reduction of picture signals: Review on the studies in Japan. Future Generation Computer Systems, 1(5):279– 308, September 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fulcher:1991:NNP

[Ful91]  J. Fulcher.  Neural net-

works: Promise for the future? Future Generation Computer Systems, 6(4):351– 354, September 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Furukawa:1992:CRS

[Fur92] Koichi Furukawa. Current research status and future direction of the fifth generation computer project. Future Generation Computer Systems, 7(2–3):269–273, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fernandez-Villacanas:1998:CPD

[FVFA98] J. L. Fernández-Villacañas, J. M. Fatah, and S. Amin. Computing with proteins in a dynamic regime. Future Generation Computer Systems, 14(5–6):275–283, December 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Foster:1998:CFT

[FvLTT98] Ian Foster, Gregor von Laszewski, George K. Thiruvathukal, and Brian Toonen. A computational framework for telemedicine. Future Generation Computer Systems, 14 (1–2):109–123, June 15, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/26/abstract.
html.

Freiwald:2002:JBC

[FW02] Uwe Freiwald and Jörg R.

Weimar. The Java based cellular automata simulation system — JCASim. Future Generation Computer Systems, 18 (7):995–1004, August 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fu:2007:CAC

[FX07] Song Fu and Cheng-Zhong Xu. Coordinated access control with temporal and spatial constraints on mobile execution in coalition environments. Future Generation Computer Systems, 23(6):804–815, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gravvanis:2006:SSP

[GA06] George A. Gravvanis and Hamid R. Arabnia. Special section: Parallel and distributed algorithms and systems. Future Generation Computer Systems, 22(1–2): 32–33, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guest:1996:HPC

[GAB^{+}96] M. F. Guest, E. Apra, D. E. Bernholdt, H. A. Früchtl, R. J. Harrison, R. A. Kendall, R. A. Kutteh, X. Long, J. B. Nicholas, J. A. Nichols, H. L. Taylor, A. T. Wong, G. I. Fann, R. J. Littlefield, and J. Nieplocha. Highperformance computing in chemistry: NW Chem. Future Generation Computer Systems, 12(4):273–289, December 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gallaire:1987:EJI

[Gal87]  Hervé Gallaire.  ECRC:

a joint industrial research centre. Future Generation Computer Systems, 3(4):279– 283, December 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gupta:1999:EIM

[GB99] Amarnath Gupta and Chaitanya Baru. An extensible information model for shared scientific data collections. Future Generation Computer Systems, 16(1):9– 20, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/27/abstract.
html.

Goderis:2009:HCM

[GBA^{+}09] Antoon Goderis, Christopher Brooks, Ilkay Altintas, Edward A. Lee, and Carole Goble. Heterogeneous composition of models of computation. Future Generation Computer Systems, 25(5):552–560, May 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Granlund:2000:DWB

[GBE00] Rego Granlund, Erik Berglund, and Henrik Eriksson. Designing Web-based simulation for learning. Future Generation Computer Systems, 17(2):

171–185, October 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/27/31/abstract.
                                        html.

Gurd:1987:PID

[GBT87] John Gurd, Wim Bohm, and Yong Meng Teo. Performance issues in dataflow machines. Future Generation Computer Systems, 3(4):285– 297, December 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gupta:1994:OIP

[GC94] Gopal Gupta and Vítor Santos Costa. Optimal implementation of and-or parallel Prolog. Future Generation Computer Systems, 10(1):71–92, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guynup:2000:ACD

[GC00] Steve Guynup and Kyle Carlson. Avatar as content delivery platform. Future Generation Computer Systems, 17(1): 65–71, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/24/32/abstract.
                                        html.

Garcia-Carballeira:2007:GPF

[GCCC^{+}07] Félix García-Carballeira, Jesús Carretero, Alejandro Calderón, J. Daniel García, and Luis. M. Sanchez. A global and parallel file system for Grids. Future Generation Computer Systems, 23(1):116–122, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guidec:1998:PIS

[GCK98] Frédéric Guidec, Patrice Calégari, and Pierre Kuonen. Parallel irregular software for wave propagation simulation. Future Generation Computer Systems, 13(4– 5):279–289, March 11, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/21/abstract.
html.

Grant:1993:M

[GD93a] Andrew Grant and Peter Dew.

MeikUS 92. Future Generation Computer Systems, 9 (1):3–4, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grant:1993:IPI

[GD93b] Andrew Grant and Robert Dickens. An implementation of a portable instrumented communication library using CS tools. Future Generation Computer Systems, 9 (1):53–61, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Goossens:2005:IRF

[GD05] Bernard Goossens and David Defour. The instruction register file micro-architecture. Future Generation Computer Systems, 21(5):767–773, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grosso:2006:NII

[GdBW06] Paola Grosso, Pieter de Boer, and Linda Winkler. The network infrastructure at iGrid2005: Lambda networking in action. Future Generation Computer Systems, 22(8): 915–919, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gommans:2003:AQP

[GdLvOT03] Leon Gommans, Cees de Laat, Bas van Oudenaarde, and Arie Taal. Authorization of a QoS path based on generic AAA. Future Generation Computer Systems, 19(6): 1009–1016, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Giorgi:2004:TIC

[GDRS04] G. Giorgi, F. De Angelis, N. Re, and A. Sgamellotti. A theoretical investigation of the Chalk–Harrod and modified Chalk–Harrod mechanisms involved in hybrid integrated circuit building. Future Generation Computer Systems, 20 (5):781–791, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gupta:1990:ALC

[GE90] Rajiv Gupta and Michael Epstein. Achieving low cost synchronization in a multiprocessor system. Future Generation Computer Systems, 6(3): 255–269, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gentzsch:1995:HPC

[Gen95] W. Gentzsch. High performance computing and networking. Future Generation Computer Systems, 11(4–5): 347–349, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Geoffray:2002:OPM

[Geo02] Patrick Geoffray. OPIOM:

Off-Processor I/O with Myrinet. Future Generation Computer Systems, 18(4):491– 499, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/33/32/abstract.
html.

Gerbessiotis:2002:AIP

[Ger02] Alexandros V. Gerbessiotis. Architecture independent parallel algorithm design: theory vs practice. Future Generation Computer Systems, 18(5): 573–593, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/36/27/abstract.
                                        html.

Guggisberg:2003:IVL

[GFGB03] M. Guggisberg, P. Fornaro, T. Gyalog, and H. Burkhart. An interdisciplinary virtual laboratory on nanoscience. Future Generation Computer Systems, 19(1):133–141, January 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grasa:2006:VTG

[GFR^{+}06] Eduard Grasa, Sergi Figuerola, Joaquim Recio, Albert López, Marc de Palol, Lluis Ribes, Vicente Díaz, Ramon Sangüesa, Gabriel Junyent, and Michel Savoie. Video transcoding in a Grid network with User Controlled LightPaths. Future Generation Computer Systems, 22(8):920–928, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grossman:2003:ESU

[GGH^{+}03] Robert L. Grossman, Yunhong Gu, Don Hamelburg, Dave Hanley, Xinwei Hong, Jorge Levera, Dave Lillethun, Marco Mazzucco, Joe Mambretti, and Jeremy Weinberger. Experimental studies using photonic data services at IGrid 2002. Future Generation Computer Systems, 19(6): 945–955, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grossman:2005:TGW

[GGH^{+}05] Robert L. Grossman, Yunhong Gu, Xinwei Hong, Antony Antony, Johan Blom, Freek Dijkstra, and Cees de Laat. Teraflows over Gigabit WANs with UDT. Future Generation Computer Systems, 21(4):501–513, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grossman:2006:DMM

[GGH^{+}06] Robert L. Grossman, Yunhong Gu, David Hanley, Michal Sabala, Joe Mambretti, Alex Szalay, Ani Thakar, Kazumi Kumazoe, Oie Yuji, Minsun Lee, Yoonjoo Kwon, and Woojin Seok. Data mining middleware for wide-area high-performance networks. Future Generation Computer Systems, 22(8): 940–948, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Groeper:2009:CAB

[GGM^{+}09] Ralf Groeper, Christian Grimm, Siegfried Makedanz, Hans Pfeiffenberger, Wolfgang Ziegler, Peter Gietz, and Michael Schiffers. A concept for attribute-based authorization on D-Grid resources. Future Generation Computer Systems, 25(3):275–280, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grossman:2009:CSC

[GGSZ09] Robert L. Grossman, Yunhong Gu, Michael Sabala, and Wanzhi Zhang. Compute and storage clouds using wide area high performance networks. Future Generation Computer Systems, 25 (2):179–183, February 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gibson:2009:DPI

[GGW^{+}09] Andrew Gibson, Matthew Gamble, Katy Wolstencroft, Tom Oinn, Carole Goble, Khalid Belhajjame, and Paolo Missier. The data playground: An intuitive workflow specification environment. Future Generation Computer Systems, 25(4):453–459, April 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Geuder:1994:GPS

[GHWZ94] U. Geuder, M. Härdtner, B. Wörner, and R. Zink. GRIDS — a programming system for grid-based technical and scientific applications on parallel systems. Future Generation Computer Systems, 10(2–3):285–289, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Galvez:2007:SNA

[GI07] A. Gálvez and A. Iglesias.

Symbolic/numeric analysis of chaotic synchronization with a CAS. Future Generation Computer Systems, 23(5): 727–733, June 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gilmore:1985:MAE

[Gil85a]  John F. Gilmore.  Mili-

tary applications of expert systems. Future Generation Computer Systems, 1(6):403– 410, December 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Giloi:1985:AOO

[Gil85b] Wolfgang K. Giloi. Advanced object oriented architectures. Future Generation Computer Systems, 1(3):169–175, February 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gillan:1994:FPC

[Gil94] M. J. Gillan. First-principles calculations on materials using massively parallel computing. Future Generation Computer Systems, 10(2–3): 213–222, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gartel:1994:TSP

[GJS^{+}94] U. Gärtel, W. Joppich, A. Schüller, H. Schwichtenberg, U. Trottenberg, and G. Winter. Two strategies in parallel computing: Porting existing software versus developing new parallel algorithms — two examples. Future Generation Computer Systems, 10(2–3):257–262, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gao:2005:NNE

[GKIZ05] D. Gao, Y. Kinouchi, K. Ito, and X. Zhao. Neural networks for event extraction from time series: a back propagation algorithm approach. Future Generation Computer Systems, 21(7):1096–1105, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Giering:2005:GED

[GKS05] R. Giering, T. Kaminski, and T. Slawig. Generating efficient derivative code with TAF: Adjoint and tangent linear Euler flow around an airfoil. Future Generation Computer Systems, 21(8):1345– 1355, October 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Golby:1995:ICV

[GL95] D. Golby and M. A. Leschziner.

Implementation of a cellvertex FV code for turbulent transonic flows on a Meiko computing surface. Future Generation Computer Systems, 11(2):145–152, March 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gervasi:2004:SPP

[GL04a] O. Gervasi and A. Laganà.

SIMBEX: a portal for the a priori simulation of crossed beam experiments. Future Generation Computer Systems, 20(5):703–715, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guan:2004:MAM

[GL04b] Sheng-Uei Guan and Sok-Seng Lim. Modeling adaptable multimedia and self-modifying protocol execution. Future Generation Computer Systems, 20(1):123–143, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gava:2005:SAB

[GL05] F. Gava and F. Loulergue. A static analysis for Bulk Synchronous Parallel ML to avoid parallel nesting. Future Generation Computer Systems, 21 (5):665–671, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Garcia-Luna-Aceves:1988:RMV

[GLA88] J. J. Garcia-Luna-Aceves. Routing management in very large-scale networks. Future Generation Computer Systems, 4(2):81–93, September 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gregoretti:2008:MGE

[GLM^{+}08] F. Gregoretti, G. Laccetti, A. Murli, G. Oliva, and U. Scafuri. MGF: a gridenabled MPI library. Future Generation Computer Systems, 24(2):158–165, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gross:1999:CMV

[GLS99] Markus H. Gross, Lars Lippert, and Oliver G. Staadt. Compression methods for visualization. Future Generation Computer Systems, 15 (1):11–29, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//visinfo.zib.de/EVlib/
Show?EVL-1999-101;   http:
//www.elsevier.com/gej-
ng/10/19/19/30/18/18/abstract.
html.

Goethals:2007:DBB

[GLSV07] Frank G. Goethals, Wilfried Lemahieu, Monique Snoeck, and Jacques A. Vandenbulcke. The data building blocks of the enterprise architect. Future Generation Computer Systems, 23(2): 269–274, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Groller:1999:VDS

[GLW99] Eduard Gröller, Helwig Löffelmann, and Rainer Wegenkittl. Visualization of dynamical systems. Future Generation Computer Systems, 15(1): 75–86, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/18/22/29/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/18/22/abstract.
html.

Gravvanis:2007:SSG

[GMA07] George A. Gravvanis, John P.

Morrison, and Hamid R. Arabnia. Special section: Grid technology and applications. Future Generation Computer Systems, 23(4):523–524, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Greenberg:2005:IMG

[GMB^{+}05] Jerry P. Greenberg, Steve Mock, Karan Bhatia, Mason Katz, Greg Bruno, Federico Sacerdoti, Phil Papadopoulos, and Kim K. Baldridge. Incorporation of middleware and grid technologies to enhance usability in Computational Chemistry applications. Future Generation Computer Systems, 21(1):3–10, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gomez:2003:NSS

[GMC03] Antonio F. Gómez, Gregorio Martínez, and Óscar Cánovas. New security services based on PKI. Future Generation Computer Systems, 19(2): 251–262, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Glatard:2008:SOA

[GMEL08] Tristan Glatard, Johan Montagnat, David Emsellem, and Diane Lingrand. A Service-Oriented Architecture enabling dynamic service grouping for optimizing distributed workflow execution. Future Generation Computer Systems, 24(7):720–730, July 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Greene:1999:DAI

[GML99] Gretchen Greene, Brian McLean, and Barry Lasker. Development of the astronomical image archive and catalog database for production of GSC-II. Future Generation Computer Systems, 16(1):29– 38, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/29/abstract.
html.

Grosso:2009:DPL

[GMM^{+}09] Paola Grosso, Damien Marchal, Jason Maassen, Eric Bernier, Li Xu, and Cees de Laat. Dynamic photonic lightpaths in the Star-Plane network. Future Generation Computer Systems, 25 (2):132–136, February 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gravvanis:2009:SSD

[GMS09] George A. Gravvanis, John P.

Morrison, and Heinz Stockinger. Special section: Defining the Grid, experiences and future trends. Future Generation Computer Systems, 25 (4):399–400, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Garatani:2001:GHP

[GNOY01] K. Garatani, H. Nakamura, H. Okuda, and G. Yagawa. GeoFEM: high performance parallel FEM for solid earth. Future Generation Computer Systems, 18 (1):107–114, September 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/37/abstract.
html.

Guo:2005:DHE

[GNWT05] H. Q. Guo, L. H. Ngoh, W. C. Wong, and J. G. Tan. DIN-Cast: a hop efficient dynamic multicast infrastructure for P2P computing. Future Generation Computer Systems, 21 (3):361–375, March 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gil:1998:ABH

[GODM98] C. Gil, J. Ortega, A. F. Díaz, and M. D. G. Montoya. Annealing-based heuristics and genetic algorithms for circuit partitioning in parallel test generation. Future Generation Computer Systems, 14(5–6):439–451, December 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gollmann:2000:NPO

[Gol00]  Dieter Gollmann.  New

paradigms — old paradigms? Future Generation Computer Systems, 16(4):343– 349, February 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/29/abstract.
html.

Goossens:2001:HIP

[Goo01]  Bernard Goossens.  Han-

dling 16 instructions per cycle in a superscalar processor. Future Generation Computer Systems, 17(6): 699–709, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/29/abstract.
html.

Goossens:2002:TIC

[Goo02] Bernard Goossens. Typing the ISA to cluster the processor. Future Generation Computer Systems, 18(6):789–796, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gorlatch:2002:MPS

[Gor02] Sergei Gorlatch. Message passing without send–receive. Future Generation Computer Systems, 18(6):797–805, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Goscinski:2000:TOS

[Gos00] Andrzej M. Goscinski. Towards an operating system managing parallelism of computing on clusters. Future Generation Computer Systems, 17(3):293–314, November 1, 2000. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/30/abstract.
html.

Grimme:2009:CNS

[GP09] Christian Grimme and Alexander Papaspyrou. Cooperative negotiation and scheduling of scientific workflows in the collaborative climate community data and processing grid. Future Generation Computer Systems, 25(3): 301–307, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Giunchiglia:1996:TPC

[GPA96] Fausto Giunchiglia, Paolo Pecchiari, and Alessandro Armando. Towards provably correct system synthesis and extension. Future Generation Computer Systems, 12(2– 3):123–137, September 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gamvroulas:2000:SBN

[GPA00] S. Gamvroulas, D. Polemi, and M. Anagnostou. A secure brokerage network for retail banking services. Future Generation Computer Systems, 16 (4):423–430, February 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/37/abstract.
html.

Goedbloed:1994:MSL

[GPH^{+}94] J. P. Goedbloed, S. Poedts, G. T. A. Huysmans, G. Halberstadt, H. Holties, and A. J. C. Beliën. Magnetohydrodynamic spectroscopy: Large scale computation of the spectrum of waves in plasmas. Future Generation Computer Systems, 10(2–3): 339–343, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ghosal:2005:PCF

[GPK05] Dipak Ghosal, Benjamin K.

Poon, and Keith Kong. P2P contracts: a framework for resource and service exchange. Future Generation Computer Systems, 21(3): 333–347, March 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gehring:1996:MFM

[GR96] Jörn Gehring and Alexander Reinefeld. MARS — a framework for minimizing the job execution time in a metacomputing environment. Future Generation Computer Systems, 12(1):87–99, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grigoras:2007:CEM

[GR07] Dan Grigoras and Mark Riordan. Cost-effective mobile ad hoc networks management. Future Generation Computer Systems, 23(8): 990–996, November 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gentzsch:2009:SSD

[GR09] Wolfgang Gentzsch and Alexander Reinefeld. Special section: D-Grid. Future Generation Computer Systems, 25 (3):266–267, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grant:1992:RMA

[Gra92]  T. J. Grant.  A review

of Multi-Agent Systems techniques, with application to Columbus User Support Organisation. Future Generation Computer Systems, 7(4): 413–437, May 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Granlund:2001:BWW

[Gra01]  Rego Granlund.  Best of

Websim99: Web-based microworld simulation for emergency management training. Future Generation Computer Systems, 17(5):561– 572, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/30/abstract.
html.

Gao:2005:AGJ

[GRH05] Yang Gao, Hongqiang Rong, and Joshua Zhexue Huang. Adaptive grid job scheduling with genetic algorithms. Future Generation Computer Systems, 21(1):151–161, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gervasi:2004:VGV

[GRPL04] O. Gervasi, A. Riganelli, L. Pacifici, and A. Laganà. VMSLab-G: a virtual laboratory prototype for molecular science on the Grid. Future Generation Computer Systems, 20(5):717–726, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Garg:1995:ASI

[GS95] Vivek Garg and David E.

Schimmel. Architectural support for inter-stream communication in an MSIMD system. Future Generation Computer Systems, 11(6):617–629, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gergel:2005:PCG

[GS05] V. P. Gergel and R. G. Strongin. Parallel computing for globally optimal decision making on cluster systems. Future Generation Computer Systems, 21(5):673–678, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Grahn:1995:IEU

[GSD95] Håkan Grahn, Per Stenström, and Michel Dubois. Implementation and evaluation of update-based cache protocols under relaxed memory consistency models. Future Generation Computer Systems, 11 (3):247–271, June 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). Accepted for publication.

Gueron:2001:DAS

[Gue01] Shay Gueron. Deterministic approximations for stochastic processes in population biology. Future Generation Computer Systems, 17(7): 893–899, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/34/32/abstract.
                                        html.

Gurd:1985:MDM

[Gur85] J. R. Gurd. The Manchester dataflow machine. Future Generation Computer Systems, 1(4):201–212, June 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gutjahr:2000:GBA

[Gut00] Walter J. Gutjahr. A graphbased Ant System and its convergence. Future Generation Computer Systems, 16 (8):873–888, June 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/31/27/abstract.
html.

Gruber:2003:PTC

[GVD^{+}03] Ralf Gruber, Pieter Volgers, Alessandro De Vita, Massimiliano Stengel, and Trach-Minh Tran. Parameterisation to tailor commodity clusters to applications. Future Generation Computer Systems, 19(1): 111–120, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guo:2001:DDS

[GW01] Yike Guo and Patrick Wendel. Developing a distributed scalable Java component server. Future Generation Computer Systems, 17(8): 1051–1057, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/40/abstract.
html.

Guan:2003:MCM

[GWO03] Sheng-Uei Guan, Tianhan Wang, and Sim-Heng Ong. Migration control for mobile agents based on passport and visa. Future Generation Computer Systems, 19(2): 173–186, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gommans:2009:MDL

[GXD^{+}09] Leon Gommans, Li Xu, Yuri Demchenko, Alfred Wan, Mihai Cristea, Robert Meijer, and Cees de Laat. Multidomain lightpath authorization, using tokens. Future Generation Computer Systems, 25(2):153–160, February 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guizhong:1990:OMC

[GY90] Liu Guizhong and Ci Yungui. An ordered model combining dataflow with control flow and its implementation. Future Generation Computer Systems, 6(1):85–89, June 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Guan:2004:PNG

[GZ04] Sheng-Uei Guan and Shu Zhang. Pseudorandom number generation based on controllable cellular automata. Future Generation Computer Systems, 20(4):627–641, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Haber:2005:CCE

[HAAH05] Rodolfo E. Haber, J. R.

Alique, A. Alique, and R. H. Haber. Controlling a complex electromechanical process on the basis of a neurofuzzy approach. Future Generation Computer Systems, 21 (7):1083–1095, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Haber:2005:SSS

[Hab05] Rodolfo E. Haber. Special section: “Soft-computing and advanced techniques in new algorithmic approaches to existing application areas”. Future Generation Computer Systems, 21(7):1015–1018, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hutanu:2006:DCV

[HAB^{+}06] Andrei Hutanu, Gabrielle Allen, Stephen D. Beck, Petr Holub, Hartmut Kaiser, Archit Kulshrestha, Miloš Liška, Jon MacLaren, Luděk Matyska, Ravi Paruchuri, Steffen Prohaska, Ed Seidel, Brygg Ullmer, and Shalini Venkataraman. Distributed and collaborative visualization of large data sets using high-speed networks. Future Generation Computer Systems, 22(8):1004–1010, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Heilper:1992:DPN

[HAC92] Andrei Heilper, Neta Amit, and Doron Cohen. The distributed and parallel NIC server. Future Generation Computer Systems, 8(1–3): 165–181, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

He:2003:QTH

[HAE^{+}03] Eric He, Javid Alimohideen, Josh Eliason, Naveen K. Krishnaprasad, Jason Leigh, Oliver Yu, and Thomas A. De-Fanti. Quanta: a toolkit for high performance data delivery over photonic networks. Future Generation Computer Systems, 19(6):919–933, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Haupt:2000:WFW

[HAF00] Tomasz Haupt, Erol Akarsu, and Geoffrey Fox. WebFlow: a framework for Web based metacomputing. Future Generation Computer Systems, 16 (5):445–451, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

19/19/41/28/27/abstract.
html.

Haupt:1999:WBM

[HAFF99] Tomasz Haupt, Erol Akarsu, Geoffrey Fox, and Wojtek Furmanski. Web based metacomputing. Future Generation Computer Systems, 15(5– 6):735–743, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/33/abstract.
html.

Halvorsen:1988:SSS

[Hal88] Per-Kristian Halvorsen. Situation semantics and semantic interpretation in constraintbased grammars. Future Generation Computer Systems, 3 (??):??, ???? 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hanakuma:1989:ESF

[Han89] Yoshitomo Hanakuma. An expert system for fault diagnosis at petrochemical plants. Future Generation Computer Systems, 5(1):103–107, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hanson:2003:LST

[Han03]  Floyd B. Hanson.  Local

supercomputing training in the computational sciences using remote national centers. Future Generation Computer Systems, 19(8):1335– 1347, November 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hoekstra:1998:GEH

[HB98] Alfons G. Hoekstra and Arndt Bode. Guest editorial: High Performance Computing and Networking Europe 1997. Future Generation Computer Systems, 13(4– 5):247–249, March 11, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/17/abstract.
html.

Holzl:2000:DFQ

[HB00] Günther Hölzl and László Böszörményi. Distributed federative QoS resource management. Future Generation Computer Systems, 16 (6):717–725, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/37/abstract.
html.

Han:2008:SSA

[HB08] Liangxiu Han and Dave Berry. Semantic-supported and agent-based decentralized grid resource discovery. Future Generation Computer Systems, 24(8):806–812, October 2008. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Held:2009:SCW

[HB09] Markus Held and Wolfgang Blochinger. Structured collaborative workflow design. Future Generation Computer Systems, 25(6):638–653, June 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Heras:2001:MIL

[HBCR01] D. B. Heras, V. Blanco, J. C.

Cabaleiro, and F. F. Rivera. Modeling and improving locality for the sparse-matrix-vector product on cache memories. Future Generation Computer Systems, 18(1):55– 67, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/32/abstract.
html.

Hedges:2009:RBC

[HBH09] Mark Hedges, Tobias Blanke, and Adil Hasan. Rule-based curation and preservation of data: a Data Grid approach using iRODS. Future Generation Computer Systems, 25 (4):446–452, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hendrikse:2003:EVG

[HBJ^{+}03] Zeger W. Hendrikse, Adam S. Z. Belloum, Philip M. R. Jonkergouw, Gert B. Eijkel, Ron M. A. Heeren, Bob L. O. Hertzberger, Vladimir Korkhov, Cees T. A. M. de Laat, and Dmitry Vasunin. Evaluating the VLAM-G toolkit on the DAS-2. Future Generation Computer Systems, 19(6):815–824, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hawick:1999:IDA

[HC99] K. A. Hawick and P. D. Coddington. Interfacing to distributed active data archives. Future Generation Computer Systems, 16(1):73– 89, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/33/abstract.
html.

Hsu:2007:PEP

[HCL07] Ching-Hsien Hsu, Tai-Lung Chen, and Kuan-Ching Li. Performance effective prescheduling strategy for heterogeneous grid systems in the master slave paradigm. Future Generation Computer Systems, 23(4):569–579, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Halepovic:2005:JPM

[HD05] Emir Halepovic and Ralph Deters. The JXTA performance model and evaluation. Future Generation Computer Systems, 21(3):377– 390, March 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Haug:1994:TVC

[HDC^{+}94] E. Haug, J. Dubois, J. Clinckemaillie, S. Vlachoutsis, and G. Lonsdale. Transport vehicle crash, safety and manufacturing simulation in the perspective of high performance computing and networking. Future Generation Computer Systems, 10(2–3): 173–181, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Heng:1987:WED

[Hen87] Michael S. H. Heng. Why evolutionary development of expert systems appears to work. Future Generation Computer Systems, 3(2):103–109, May 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hertzberger:1984:AFG

[Her84] L. O. Hertzberger. The architecture of fifth generation inference computers. Future Generation Computer Systems, 1(1):9–21, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hertzberger:1987:PPC

[Her87] L. O. Hertzberger. Parallel processing, the challenge of new computer architectures: Tim Johnson and Tony Durham. Future Generation Computer Systems, 3(2): 147–148, May 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hertzberger:1990:P

[Her90] L. O. Hertzberger. Preface. Future Generation Computer Systems, 6(1):1–2, June 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hertzberger:1991:TPD

[Her91] L. O. Hertzberger. Trends in parallel and distributed computing. Future Generation Computer Systems, 7(1): 31–40, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hertzberger:1994:E

[Her94] L. O. Hertzberger. Editorial. Future Generation Computer Systems, 10(2–3):151, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Helly:1999:MID

[HESM99] John J. Helly, T. Todd Elvins, Don Sutton, and David Martinez. A method for interoperable digital libraries and data repositories. Future Generation Computer Systems, 16(1): 21–28, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/28/abstract.
html.

Hey:1990:EMP

[Hey90] Anthony J. G. Hey. Experiments in MIMD parallelism. Future Generation Computer Systems, 6(3):185– 196, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hum:1992:HSM

[HG92] Herbert H. J. Hum and Guang R. Gao. A high-speed memory organization for hybrid dataflow/von Neumann computing. Future Generation Computer Systems, 8(4):287– 301, September 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hancock:1998:DPV

[HH98] D. J. Hancock and R. J.

Hubbold. Distributed parallel volume rendering on shared memory systems. Future Generation Computer Systems, 13(4–5):251–259, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

                                        elsevier.com/gej-ng/10/
                                        19/19/28/19/18/abstract.
                                        html.

Heimbach:2005:EEA

[HHG05] Patrick Heimbach, Chris Hill, and Ralf Giering. An efficient exact adjoint of the parallel MIT General Circulation Mode, generated via automatic differentiation. Future Generation Computer Systems, 21(8):1356–1371, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

He:1998:RSO

[HHS98] Xiangjian He, Tom Hintz, and Ury Szewcow. Replicated shared object model for parallel edge detection algorithm based on spiral architecture. Future Generation Computer Systems, 14(5–6): 341–350, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hartenstein:1992:NPP

[HHSW92] R. W. Hartenstein, A. G. Hirschbiel, K. Schmidt, and M. Weber. A novel paradigm of parallel computation and its use to implement simple high-performance hardware. Future Generation Computer Systems, 7(2–3): 181–198, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hazas-Izquierdo:2006:OUF

[HICÁFM^{+}06] Raúl G. Hazas-Izquierdo, Salvador Castañeda-Ávila, Luis M. Farfán-Molina, Julián Delgado-Jiménez, Daniel García-Gradilla, and José L. Rodríguez-Navarro. Opening a University Fiber Highway between Mexico and the US. Future Generation Computer Systems, 22(8):892–895, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hiraide:1989:ESA

[Hir89] Masataka Hiraide. Expert systems application in plant engineering. Future Generation Computer Systems, 5(1): 77–85, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hughes-Jones:2005:PGE

[HJCD05] Richard Hughes-Jones, Peter Clarke, and Steven Dallison. Performance of 1 and 10 Gigabit Ethernet cards with server quality motherboards. Future Generation Computer Systems, 21(4):469–488, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hwang:2004:PAA

[HJK^{+}04] Intae Hwang, Taewon Jang, Mingoo Kang, Sangmin No, Jungyoung Son, Daesik Hong, and Changeon Kang. Performance analysis of adaptive modulation and coding combined with transmit diversity in next generation mobile communication systems. Future Generation Computer Systems, 20(2): 189–196, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Haridi:1992:SOS

[HJP92] Seif Haridi, Sverker Janson, and Catuscia Palamidessi. Structural operational semantics for AKL. Future Generation Computer Systems, 8 (4):409–421, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hughes-Jones:2003:HDR

[HJPS03] Richard Hughes-Jones, Steve Parsley, and Ralph Spencer. High data rate transmission in high resolution radio astronomy — vlbiGRID. Future Generation Computer Systems, 19(6):883–896, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hawick:1999:DES

[HJS^{+}99] K. A. Hawick, H. A. James, A. J. Silis, D. A. Grove, C. J. Patten, J. A. Mathew, P. D. Coddington, K. E. Kerry, J. F. Hercus, and F. A. Vaughan. DISCWorld: an environment for service-based matacomputing. Future Generation Computer Systems, 15(5–6): 623–635, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/24/abstract.
html.

Halvorsen:1988:PSD

[HK88] Per-Kristian Halvorsen and Ronald M. Kaplan. Projections and semantic description in lexical-functional grammar. Future Generation Computer Systems, 3(??):00, ???? 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2009:SSG

[hKcF09] Tai hoon Kim and Wai chi Fang. Special section: Grid/ distributed computing systems security. Future Generation Computer Systems, 25 (3):351, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hoisie:2006:SSL

[HKM^{+}06] A. Hoisie, D. J. Kerbyson, C. Mendes, D. Reed, and A. Snavely. Special section: Large-scale system performance modeling and analysis. Future Generation Computer Systems, 22(3):291– 292, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hidaka:1994:APM

[HKT94] Yasuo Hidaka, Hanpei Koike, and Hidehiko Tanaka. Architecture of parallel management kernel for PIE64. Future Generation Computer Systems, 10(1):29–43, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hinsen:2006:UBP

[HLSØ06] Konrad Hinsen, Hans Petter Langtangen, Ola Skavhaug, and Åsmund Ødegård. Using BSP and Python to simplify parallel programming. Future Generation Computer Systems, 22(1–2):123–157, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hollmann:1997:PSA

[HLvL^{+}97] H. D. L. Hollmann, J. P. M. G. Linnartz, J. H. van Lint, C. P. M. J. Baggen, and L. M. G. Tolhuizen. Protection of software algorithms executed on secure, Modules. Future Generation Computer Systems, 13 (1):55–63, June 20, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/17/20/abstract.

html. Special Issue on Smart Cards.

Huang:1998:TWB

[HM98] Yuguang Huang and William F. McColl. A two-way BSP algorithm for tridiagonal systems. Future Generation Computer Systems, 13(4–5): 337–347, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/26/abstract.
html.

Hwang:2006:HBC

[HMC06] In-Chul Hwang, Seung-Ryoul Maeng, and Jung-Wan Cho. Home-based cooperative cache for parallel I/O applications. Future Generation Computer Systems, 22(5):633–642, April 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Holub:2006:HDM

[HML^{+}06] Petr Holub, Luděk Matyska, Miloš Liška, Lukáš Hejtmánek, Jiří Denemark, Tomáš Rebok, Andrei Hutanu, Ravi Paruchuri, Jan Radil, and Eva Hladká. High-definition multimedia for multiparty lowlatency interactive communication. Future Generation Computer Systems, 22(8): 856–861, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huedo:2007:MMS

[HML07] Eduardo Huedo, Rubén S.

Montero, and Ignacio M. Llorente. A modular metascheduling architecture for interfacing with pre-WS and WS Grid resource management services. Future Generation Computer Systems, 23(2): 252–261, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huedo:2009:RAH

[HML09] Eduardo Huedo, Rubén S. Montero, and Ignacio M. Llorente. A recursive architecture for hierarchical grid resource management. Future Generation Computer Systems, 25(4):401–405, April 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huh:2004:EEP

[HMP04] Eui-Nam Huh, Y. Mun, and Hyoung-Woo Park. An efficient event publish technique for real-time monitor on real-time grid computing. Future Generation Computer Systems, 20(2):257–263, February 16, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hlavacek:2006:WSD

[HND06] I. Hlaváček, J. Nedoma, and J. Daněk. Worst scenario and domain decomposition methods in geomechanics. Future Generation Computer Systems, 22(4):468–483, March 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hascoet:2005:RAR

[HNP05] Laurent Hascoët, Uwe Naumann, and Valérie Pascual. “to be recorded” analysis in reverse-mode automatic differentiation. Future Generation Computer Systems, 21 (8):1401–1417, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hovland:2005:MAD

[HNS05] Paul Hovland, Boyana Norris, and Barry Smith. Making automatic differentiation truly automatic: coupling PETSc with ADIC. Future Generation Computer Systems, 21 (8):1426–1438, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hellwagner:2002:VAC

[HO02] Hermann Hellwagner and Matthias Ohlenroth. VI architecture communication features and performance on the Giganet cluster LAN. Future Generation Computer Systems, 18(3):421– 433, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/34/abstract.
html.

Hofinger:2003:LRR

[Höf03]  Siegfried Höfinger.  La-

tency reduction from runtimeinterference to the parallel quantum chemistry program GREMLIN in heterogeneous and homogeneous environments. Future Generation Computer Systems, 19 (5):777–788, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hollnagel:1993:VSA

[Hol93] Erik Hollnagel. ViVa: a systematic approach to verification, improvement and validation of knowledge-based systems. Future Generation Computer Systems, 9(4):371– 378, December 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hirabayashi:2004:LBS

[HORC04] Miki Hirabayashi, Makoto Ohta, Daniel A. Rüfenacht, and Bastien Chopard. A lattice Boltzmann study of blood flow in stented aneurism. Future Generation Computer Systems, 20(6):925–934, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Howe:1991:KBS

[How91] Jim Howe. Knowledge-based systems and artificial intelligence: emerging technology. Future Generation Computer Systems, 7(1):55–68, October 1991. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hey:1992:GE

[HP92] A. J. G. Hey and E. Paris.

Guest editorial. Future Generation Computer Systems, 8 (1–3):1–2, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huang:2008:MBG

[HPLL08] Peijie Huang, Hong Peng, Piyuan Lin, and Xuezhen Li. Macroeconomics based Grid resource allocation. Future Generation Computer Systems, 24(7):694–700, July 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huang:2009:SSD

[HPLL09] Peijie Huang, Hong Peng, Piyuan Lin, and Xuezhen Li. Static strategy and dynamic adjustment: An effective method for Grid task scheduling. Future Generation Computer Systems, 25 (8):884–892, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hoffmann:1994:SDL

[HPP94] Walter Hoffmann, Kitty Potma, and Gera G. Pronk. Solving dense linear systems by Gauss–Huard’s method on a distributed memory system. Future Generation Computer Systems, 10 (2–3):321–325, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.fwi.uva.nl/research/
pwrs/papers/archive/Hoffmann94_
2.ps.gz.

Hosking:1997:SPD

[HPS97] Jonathan R. M. Hosking, Edwin P. D. Pednault, and Madhu Sudan. A statistical perspective on data mining. Future Generation Computer Systems, 13(2–3):117– 134, November 14, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/18/abstract.
html.

Huang:2007:RTU

[HQ07] Zhixing Huang and Yuhui Qiu. Resource trading using cognitive agents: a hybrid perspective and its simulation. Future Generation Computer Systems, 23(7):837–845, August 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hyvaluoma:2004:ELB

[HRJ^{+}04] J. Hyväluoma, P. Raiskinmäki, A. Jäsberg, A. Koponen, M. Kataja, and J. Timonen. Evaluation of a lattice-Boltzmann method for mercury intrusion porosimetry simulations. Future Generation Computer Systems, 20(6): 1003–1011, August 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hirano:2006:FFD

[HRJ^{+}06] Akira Hirano, Luc Renambot, Byungil Jeong, Jason Leigh, Alan Verlo, Venkatram Vishwanath, Rajvikram Singh, Julieta Aguilera, Andrew Johnson, Thomas A. De-Fanti, Lance Long, Nicholas Schwarz, Maxine Brown, Naohide Nagatsu, Yukio Tsukishima, Masahito Tomizawa, Yutaka Miyamoto, Masahiko Jinno, Yoshihiro Takigawa, and Osamu Ishida. The first functional demonstration of optical virtual concatenation as a technique for achieving Terabit networking. Future Generation Computer Systems, 22(8):876–883, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hecker:1999:SDP

[HRSW99] C. Hecker, D. Roytenberg, J.- R. Sack, and Z. Wang. System development for parallel cellular automata and its applications. Future Generation Computer Systems, 16(2–3): 235–247, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        19/19/41/25/31/abstract.
                                        html.

Hill:1998:LLI

[HS98] Jonathan M. D. Hill and David B. Skillicorn. Lessons learned from implementing BSP. Future Generation Computer Systems, 13(4–5):327– 335, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/19/25/abstract.
                                        html.

Huai:2007:RRH

[HSH^{+}07] Jinpeng Huai, Hailong Sun, Chunming Hu, Yanmin Zhu, Yunhao Liu, and Jianxin Li. ROST: Remote and hot service deployment with trustworthiness in CROWN Grid. Future Generation Computer Systems, 23(6):825–835, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hagerup:1992:FHL

[HSS92] T. Hagerup, A. Schmitt, and H. Seidl. FORK: a high-level language for PRAMs. Future Generation Computer Systems, 8(4):379– 393, September 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www-
                                        wjp.cs.uni-sb.de/sbpram/
                                        papers/forkpram.ps.gz.

Huang:2000:HSE

[HSS00] Zhiyi Huang, Chengzheng Sun, and Abdul Sattar. Handling side-effects and cuts with selective recomputation in parallel Prolog. Future Generation Computer Systems, 17(3):227–245, November 1, 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/26/abstract.
html.

Hey:2002:USC

[HT02] Tony Hey and Anne E. Trefethen. The UK e-Science Core Programme and the Grid. Future Generation Computer Systems, 18(8): 1017–1031, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hoarau:2007:FFV

[HTV07] William Hoarau, Sébastien Tixeuil, and Fabien Vauchelles. FAIL–FCI: Versatile fault injection. Future Generation Computer Systems, 23(7): 913–919, August 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huang:2005:GPI

[Hua05] Yan Huang. GSiB: PSE infrastructure for dynamic serviceoriented Grid applications. Future Generation Computer Systems, 21(6):868–877, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hullar:1989:GCC

[Hul89] Chancellor T. Hullar. Grand challenges to computational sciences: fundamental problems in science progress and future: Molokai, Hawaii, 3 January 1989. Future Generation Computer Systems, 5(2– 3):169–170, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hummel:1992:WHS

[Hum92] Susan Flynn Hummel. Workstation hardware and software support for parallel applications. Future Generation Computer Systems, 8(1– 3):49–65, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hertzberger:1984:PFG

[HV84] L. O. Hertzberger and R. P. Van de Riet. Progress in the fifth generation inference architectures. Future Generation Computer Systems, 1(2): 93–102, November 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hofman:1992:DHS

[HV92] Rutger Hofman and Willem G. Vree. Distributed hierarchical scheduling with explicit grain size control. Future Generation Computer Systems, 8(1– 3):111–119, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huper:2003:NAI

[HV03] K. Hüper and P. Van Dooren.

New algorithms for the iterative refinement of estimates of invariant subspaces. Future Generation Computer Systems, 19(7):1231–1242, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hoekstra:2004:UFE

[HvHAS04] A. G. Hoekstra, Jos van’t Hoff, A. M. Artoli, and P. M. A. Sloot. Unsteady flow in a 2D elastic tube with the LBGK method. Future Generation Computer Systems, 20(6):917–924, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hillmann:1995:SMF

[HW95] M. Hillmann and J. Weiher. Simulation of metal forming processes with respect to MPP-systems. Future Generation Computer Systems, 11(4– 5):431–437, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hamilton-Wright:2007:CGL

[HWS07] Andrew Hamilton-Wright and Deborah Stacey. ‘cheap grid’: Leveraging system failure using stochastic computation. Future Generation Computer Systems, 23(4):525–535, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hsu:2004:GOS

[HWW04] Chien-Lung Hsu, Tzong-Sun Wu, and Tzong-Chen Wu. Group-oriented signature scheme with distinguished signing authorities. Future Generation Computer Systems, 20(5):865–873, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huang:2008:OBP

[HWZL08] Kun Huang, Li’e Wang, Dafang Zhang, and Yongwei Liu. Optimizing the Bit-Torrent performance using an adaptive peer selection strategy. Future Generation Computer Systems, 24(7):621–630, July 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hu:1990:ADC

[HXL90] Yueming Hu, Zhiliang Xie, and Xinda Lu. Approaches to decentralized control of job scheduling for homogeneous and heterogeneous parallel computer systems. Future Generation Computer Systems, 6(1):91–96, June 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Holmgren:2003:SNG

[HY03] Sverker Holmgren and Anders Ynnerman. The Swedish National Graduate School in Scientific Computing (NGSSC). Future Generation Computer Systems, 19(8):1275– 1283, November 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Han:2009:NGR

[HY09] Youngjoo Han and Chan-Hyun Youn. A new grid resource management mechanism with resource-aware policy administrator for SLAconstrained applications. Future Generation Computer Systems, 25(7):768–778, July 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Huang:2004:STI

[HYC04] Bo Huang, Shanzhen Yi, and Weng Tat Chan. Spatiotemporal information integration in XML. Future Generation Computer Systems, 20(7): 1157–1170, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hongjian:2004:FRA

[HYS04] You Hongjian, Shao Yun, and Li Shukai. Fast rectifying airborne infrared scanning image based on GPS and INS. Future Generation Computer Systems, 20(7): 1209–1214, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Han:2008:RAS

[HZC^{+}08] Songqiao Han, Shensheng Zhang, Jian Cao, Ye Wen, and Yong Zhang. A resource aware software partitioning algorithm based on mobility constraints in pervasive grid environments. Future Generation Computer Systems, 24 (6):512–529, June 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Issman:1998:NOP

[ID98] E. Issman and G. Degrez. Non-overlapping preconditioners for a parallel implicit Navier–Stokes solver. Future Generation Computer Systems, 13(4–5):303–313, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/19/23/abstract.
                                        html.

Indrusiak:2001:BWI

[IdLR01] Leandro Soares Indrusiak and Ricardo Augusto da Luz Reis. Best of Websim99: 3D integrated circuit layout visualization using VRML. Future Generation Computer Systems, 17(5):503– 511, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/26/abstract.
html.

Iglesias:2004:FNB

[IEG04] A. Iglesias, G. Echevarría, and A. Gálvez. Functional networks for B-spline surface reconstruction. Future Generation Computer Systems, 20(8): 1337–1353, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Iglesias:2004:CGW

[Igl04a] A. Iglesias. Computer graphics for water modeling and rendering: a survey. Future Generation Computer Systems, 20(8):1355–1374, November 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Iglesias:2004:I

[Igl04b] Andrés Iglesias. Introduction. Future Generation Computer Systems, 20(8):1235– 1239, November 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Iglesias:2007:SSC

[Igl07] Andrés Iglesias. Special section: Computer algebra systems and their applications, CASA’2003–06: Selected papers. Future Generation Computer Systems, 23(5):714–715, June 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ierotheou:2003:UIP

[IJLC03] C. S. Ierotheou, S. P. Johnson, P. F. Leggett, and M. Cross. Using an interactive parallelisation toolkit to parallelise an ocean modelling code. Future Generation Computer Systems, 19(5):789–801, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Iosup:2008:GWA

[ILJ^{+}08] Alexandru Iosup, Hui Li, Mathieu Jan, Shanny Anoep, Catalin Dumitrescu, Lex Wolters, and Dick H. J. Epema. The Grid Workloads Archive. Future Generation Computer Systems, 24 (7):672–686, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Inui:1989:DMB

[IMKB89] Masahiro Inui, Nobuji Miyasaka, Kazuhiko Kawamura, and John R. Bourne. Development of a model-based intelligent training system. Future Generation Computer Systems, 5 (1):59–69, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Iannello:1990:PSD

[IMSV90] G. Iannello, A. Mazzeo, C. Savy, and G. Ventre. Parallel software development in the DISC programming environment. Future Generation Computer Systems, 5(4): 365–372, January 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Inamuro:2004:NSB

[IOO04] T. Inamuro, T. Ogata, and F. Ogino. Numerical simulation of bubble flows by the lattice Boltzmann method. Future Generation Computer Systems, 20(6):959–964, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Islas:2003:MSM

[IS03] A. L. Islas and C. M. Schober.

Multi-symplectic methods for generalized Schrödinger equations. Future Generation Computer Systems, 19(3): 403–413, April 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Islas:2006:CPM

[IS06] A. L. Islas and C. M. Schober. Conservation properties of multisymplectic integrators. Future Generation Computer Systems, 22(4): 412–422, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Irony:2004:PFR

[IST04] Dror Irony, Gil Shklarski, and Sivan Toledo. Parallel and fully recursive multifrontal sparse Cholesky. Future Generation Computer Systems, 20 (3):425–440, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Iamnitchi:2005:PCI

[IT05] Adriana Iamnitchi and Domenico Talia. P2P computing and interaction with grids. Future Generation Computer Systems, 21(3):331–332, March 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Javadi:2007:ACN

[JAA07] Bahman Javadi, Mohammad K. Akbari, and Jemal H. Abawajy. Analytical communication networks model for enterprise Grid computing. Future Generation Computer Systems, 23(6):737–747, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Javadi:2009:MCC

[JAA09] Bahman Javadi, Mohammad K. Akbari, and Jemal H. Abawajy. Multi-cluster computing interconnection network performance modeling and analysis. Future Generation Computer Systems, 25 (7):737–746, July 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jezequel:1994:PMP

[JBA94] J.-M. Jézéquel, F. Bergheul, and F. André. Programming massively parallel architectures with sequential object oriented languages. Future Generation Computer Systems, 10(1):59–70, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jin:2008:SES

[JC08] Hai Jin and Hanhua Chen.

SemreX: Efficient search in a semantic overlay for literature retrieval. Future Generation Computer Systems, 24 (6):475–488, June 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Joung:2009:OOB

[JC09] Yuh-Jzer Joung and Feng-Yuan Chuang. OntoZilla: An ontology-based, semistructured, and evolutionary peer-to-peer network for information systems and services.

Future Generation Computer Systems, 25(1):53–63, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jedrzejowicz:2001:EBS

[JCSS01] Piotr Jȩdrzejowicz, Ireneusz Czarnowski, Aleksander Skakowski, and Henryk Szreder. Evolutionbased scheduling of multiple variant and multiple processor programs. Future Generation Computer Systems, 17 (4):405–414, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/31/abstract.
html.

Jacquet:1994:S

[JD94] Jean-Marie Jacquet and Koenraad De Bosschere. On the semantics of μLog. Future Generation Computer Systems, 10 (1):93–135, April 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jugravu:2005:JPM

[JF05] Alexandru Jugravu and Thomas Fahringer. JavaSymphony, a programming model for the Grid. Future Generation Computer Systems, 21 (1):239–246, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Janzadeh:2009:SCB

[JFDF09] Hamed Janzadeh, Kaveh Fayazbakhsh, Mehdi Dehghan, and Mehran S. Fallah. A secure credit-based cooperation stimulating mechanism for MANETs using hash chains. Future Generation Computer Systems, 25 (8):926–934, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jo:2006:IHV

[JHL^{+}06] Jinyong Jo, Wontaek Hong, SeungJoo Lee, Dongkyun Kim, JongWon Kim, and OkHwan Byeon. Interactive 3D HD video transport for e-science collaboration over UCLP-enabled GLO-RIAD lightpath. Future Generation Computer Systems, 22 (8):884–891, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Janicki:1992:IPC

[JK92] Ryszard Janicki and M. Koutny.

Invariants and paradigms of concurrency theory. Future Generation Computer Systems, 8(4):423–435, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Janssen:1995:PVP

[JL95] Jack Janssen and Hai Xiang Lin. A parallel version of the phase space evolution model. Future Generation Computer Systems, 11 (2):115–122, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jenq:1998:FBA

[JL98] John Jingfu Jenq and Wingning Li. Feedforward backpropagation artificial neural networks on reconfigurable meshes. Future Generation Computer Systems, 14(5–6): 313–319, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jia:2003:PMT

[JL03] Zhidong Jia and Ben Leimkuhler. A parallel multiple timescale reversible integrator for dynamics simulation. Future Generation Computer Systems, 19(3):415–424, April 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jia:2008:SSS

[JL08] Xiaohua Jia and Francis C. M. Lau. Special section: Scalable information systems. Future Generation Computer Systems, 24(8):833, October 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jung:2000:SSP

[JLMR00] Bernd Jung, Hans-Peter Lenhof, Peter Müller, and Christine Rüb. Simulating synthetic polymer chains in parallel. Future Generation Computer Systems, 16(5): 513–522, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/34/abstract.
html.

Jacob:2003:PBC

[JLU03] Eduardo Jacob, Fidel Liberal, and Juanjo Unzilla. PKIXbased certification infrastructure implementation adapted to non-personal end entities. Future Generation Computer Systems, 19(2):263– 275, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jovanovic:2001:HAD

[JM01] Zoran Jovanovic and Slavko Maric. A heuristic algorithm for dynamic task scheduling in highly parallel computing systems. Future Generation Computer Systems, 17 (6):721–732, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/31/abstract.
html.

Jimenez-Morales:2002:ICB

[JM02] F. Jiménez-Morales. Intermittent collective behavior in totalistic cellular automata with high connectivity. Future Generation Computer Systems, 18(7):921–929, August 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jones:2006:UCC

[JNPY06] Mark Jones, Zahi Nakad, Paul Plassmann, and Yanhua Yi. The use of configurable computing for computational kernels in scientific simulations. Future Generation Computer Systems, 22(1–2):67–79, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Japhet:2001:OOM

[JNR01] Caroline Japhet, Frederic Nataf, and François Rogier. The optimized order 2 method — application to convection-diffusion problems. Future Generation Computer Systems, 18(1):17– 30, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/27/29/abstract.
                                        html.

Johnson:1989:EPC

[Joh89] Gary M. Johnson. Exploiting parallelism in computational science. Future Generation Computer Systems, 5(2– 3):319–337, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Johannsen:1992:TMF

[Joh92] Wolfgang Johannsen. Transaction models for federative distributed database systems. Future Generation Computer Systems, 7(2–3): 329–337, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Johnston:2002:CDG

[Joh02] William E. Johnston. Computational and data Grids in large-scale science and engineering. Future Generation Computer Systems, 18 (8):1085–1100, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jonkers:2000:UVE

[Jon00] Hugo Jonkers. On the use of VRML in educational software — Experiences from the project: JIMM Problem Solver. Future Generation Computer Systems, 17(1):49– 56, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/24/30/abstract.
html.

Joita:2007:CMD

[JRF^{+}07] L. Joita, Omer F. Rana, Felix Freitag, Isaac Chao, Pablo Chacin, Leandro Navarro, and Oscar Ardaiz. A catallactic market for data mining services. Future Generation Computer Systems, 23 (1):146–153, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jansen:1989:PBB

[JS89] J. M. Jansen and F. W. Sijstermans. Parallel branchand-bound algorithms. Future Generation Computer Systems, 4(4):271–279, March 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jarvis:2006:PPU

[JSK^{+}06] Stephen A. Jarvis, Daniel P. Spooner, Helene N. Lim Choi Keung, Junwei Cao, Subhash Saini, and Graham R. Nudd. Performance prediction and its use in parallel and distributed computing systems. Future Generation Computer Systems, 22(7):745–754, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jovanov:1999:EAT

[JSS^{+}99] Emil Jovanov, Dušan Starčević, Aleksandar Samardžić, Andy Marsh, and Željko Obrenović. EEG analysis in a telemedical virtual world. Future Generation Computer Systems, 15(2):255–263, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X98000685.

Kurokawa:1988:DVT

[KA88] Takakazu Kurokawa and Hideo Aiso. 3-D VLSI technology in Japan and an example: a syndrome decoder for double error correction. Future Generation Computer Systems, 4 (2):145–155, September 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Khanli:2008:AGR

[KA08] L. Mohammad Khanli and M. Analoui. An approach to grid resource selection and fault management based on ECA rules. Future Generation Computer Systems, 24 (4):296–316, April 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kalantari:2009:PSS

[KA09] Mohammad Kalantari and Mohammad Kazem Akbari. A parallel solution for scheduling of real time applications on Grid environments. Future Generation Computer Systems, 25(7):704–716, July 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kaandorp:1998:GEE

[Kaa98]  J. A. Kaandorp.  Guest

editorial: EURO-VR Miniconference ’97: Virtual reality in industry and research. Future Generation Computer Systems, 14(3–4): 143–145, ???? 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kaandorp:1999:GET

[Kaa99] J. A. Kaandorp. Guest Editorship: Thematic issue on virtual reality in industry and research. Future Generation Computer Systems, 15 (4):467–468, July 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kacsuk:2000:SMD

[Kac00]  P. Kacsuk.  Systematic

macrostep debugging of message passing parallel programs. Future Generation Computer Systems, 16(6): 609–624, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/29/29/abstract.
                                        html.

Kagayama:1989:FES

[Kag89] Shigeru Kagayama. The fundamentals of expert systems on tort in Japan. Future Generation Computer Systems, 5(1):143–149, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kalamboukis:1994:CPM

[Kal94] T. Z. Kalamboukis. Communication performance of a message passing MIMD computer. Future Generation Computer Systems, 10(4): 403–409, November 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kamae:1985:VTU

[Kam85] T. Kamae. Visual terminals and user interfaces. Future Generation Computer Systems, 1(5):257–278, September 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kari:2001:DCV

[Kar01] Lila Kari. DNA computing in vitro and in vivo. Future Generation Computer Systems, 17 (7):823–834, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/34/26/abstract.
html.

Kashiwagi:1985:JSS

[Kas85]  Hiroshi Kashiwagi.  The

Japanese super-speed computer project. Future Generation Computer Systems, 1(3):

153–160, February 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kashiwagi:1986:O

[Kas86]  H. Kashiwagi.  Obituary.

Future Generation Computer Systems, 2(1):1, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Katsinis:2004:MSM

[Kat04] Constantine Katsinis. Merging, sorting and matrix operations on the SOME-Bus multiprocessor architecture. Future Generation Computer Systems, 20(4):643–661, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kawabe:1992:SQP

[Kaw92] Ushio Kawabe. The status quo and prospect of Josephson junction device technology. Future Generation Computer Systems, 7(2–3): 275–282, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kumar:1992:GPC

[KB92] Manoj Kumar and Yurij Baransky. The GF11 parallel computer: Programming and performance. Future Generation Computer Systems, 7(2– 3):169–179, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kotsis:2000:GBC

[KB00] Gabriele Kotsis and Markus Braun. Graph based characterization of distributed applications. Future Generation Computer Systems, 16 (6):597–607, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/28/abstract.
html.

Kosar:2009:NPD

[KB09a] Tevfik Kosar and Mehmet Balman. A new paradigm: Data-aware scheduling in grid computing. Future Generation Computer Systems, 25(4): 406–413, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Krishnan:2009:SSA

[KB09b] Sriram Krishnan and Karan Bhatia. SOAs for scientific applications: Experiences and challenges. Future Generation Computer Systems, 25(4): 466–473, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Krefting:2009:MTU

[KBB^{+}09] Dagmar Krefting, Julian Bart, Kamen Beronov, Olga Dzhimova, Jürgen Falkner, Michael Hartung, Andreas Hoheisel, Tobias A. Knoch, Thomas Lingner, Yassene Mohammed, Kathrin Peter, Erhard Rahm, Ulrich Sax, Dietmar Sommerfeld, Thomas Steinke, Thomas Tolxdorff, Michal Vossberg, Fred Viezens, and Anette Weisbecker. MediGRID: Towards a user friendly secured Grid infrastructure. Future Generation Computer Systems, 25(3):326–336, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kielmann:2002:PEH

[KBM^{+}02] Thilo Kielmann, Henri E. Bal, Jason Maassen, Rob van Nieuwpoort, Lionel Eyraud, Rutger Hofman, and Kees Verstoep. Programming environments for high-performance Grid computing: the Albatross Project. Future Generation Computer Systems, 18 (8):1113–1125, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Karkowski:1998:OLT

[KC98] Ireneusz Karkowski and Henk Corporaal. Overcoming the limitations of the traditional loop parallelization. Future Generation Computer Systems, 13(4–5):407–416, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/32/abstract.
html.

Kim:2004:NVC

[KCK04] Deok-Soo Kim, Youngsong Cho, and Hyun Kim. Normal vector compression of 3D mesh model based on clustering and relative indexing. Future Generation Computer Systems, 20(8):1241– 1250, November 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kier:1999:CAM

[KCT99] Lemont B. Kier, Chao-Kun Cheng, and Bernard Testa. Cellular automata models of biochemical phenomena. Future Generation Computer Systems, 16(2–3): 273–289, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/34/abstract.
html.

Kotapati:2000:BMW

[KD00] Kamala Kotapati and Sivarama P.

Dandamudi. Buffer management in wormhole-routed torus multicomputer networks. Future Generation Computer Systems, 16(5): 483–491, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

                                        elsevier.com/gej-ng/10/
                                        19/19/41/28/31/abstract.
                                        html.

Karnik:2004:CSD

[KDE04] Madhuri Karnik, Bhaskar Dasgupta, and Vinayak Eswaran. A comparative study of Dirichlet and Neumann conditions for path planning through harmonic functions. Future Generation Computer Systems, 20(3):441–452, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kacsuk:1999:GGE

[KDFL99] Péter Kacsuk, Gábor Dózsa, Tibor Fadgyas, and Róbert Lovas. The GRED graphical editor for the GRADE parallel program development environment. Future Generation Computer Systems, 15(3): 443–452, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/19/29/abstract.
                                        html.

Kuwahara:1985:MIP

[KE85] Michiyoshi Kuwahara and Shigeru Eiho. Medical image processing: an overview of and case study in the diagnosis of cardiac diseases. Future Generation Computer Systems, 1 (5):353–375, September 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Keane:1993:PFS

[Kea93] J. A. Keane. Parallelising a financial system. Future Generation Computer Systems, 9 (1):41–51, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Keahey:1999:PPL

[Kea99] Katarzyna Keahey. PAR-DIS: Programmer-level abstractions for metacomputing. Future Generation Computer Systems, 15(5–6):637– 647, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/25/abstract.
html.

Kueter:2000:BIC

[KF00] Derek Kueter and Robbert Fisher. Business insights in e-commerce and trusted services. Future Generation Computer Systems, 16(4): 373–378, February 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/32/abstract.
html.

Kondo:2007:CRA

[KFC^{+}07] Derrick Kondo, Gilles Fedak, Franck Cappello, Andrew A. Chien, and Henri Casanova. Characterizing resource availability in enterprise desktop grids. Future Generation Computer Systems, 23(7): 888–903, August 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Keshi:1989:KBF

[KFF89] Ikuo Keshi, Naoyuki Fukuda, and Yoshiji Fujimoto. A knowledge-based framework in an intelligent assistant system for making documents. Future Generation Computer Systems, 5(1):51–58, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Keahey:2002:CGA

[KFP^{+}02] K. Keahey, T. Fredian, Q. Peng, D. P. Schissel, M. Thompson, I. Foster, M. Greenwald, and D. Mc-Cune. Computational Grids in action: the National Fusion Collaboratory. Future Generation Computer Systems, 18 (8):1005–1015, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kaandorp:2001:EPB

[KG01] Jaap A. Kaandorp and Shay Gueron. Editorial: Particle based modelling methods applied in biology. Future Generation Computer Systems, 17 (7):v–vi, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
19/19/45/34/25/abstract.
html.

Kuo:1985:USN

[KGLA85] Franklin F. Kuo and Jose J.

Garcia-Luna-Aceves. USER-NET: a supercomputer network architecture. Future Generation Computer Systems, 1(3):161–168, February 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kumar:1995:CIE

[KGW95] Padam Kumar, J. P. Gupta, and S. C. Winter. CTD-Net III — an eager reduction model with laziness features. Future Generation Computer Systems, 11(3):273–282, June 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Keane:1995:CDM

[KGX95] J. A. Keane, A. J. Grant, and M. Q. Xu. Comparing distributed memory and virtual shared memory parallel programming models. Future Generation Computer Systems, 11(2):233–243, March 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kozai:1989:IIS

[KH89] Toyoki Kozai and Takehiko Hoshi. Intelligent information systems for production management in agriculture and horticulture. Future Generation Computer Systems, 5(1): 131–136, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kononenko:1997:ASM

[KH97] Igor Kononenko and Se June Hong. Attribute selection for modelling. Future Generation Computer Systems, 13(2–3): 181–195, November 14, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/22/abstract.
html.

Kurozumi:1989:TRP

[KI89] Takashi Kurozumi and Takao Ichiko. Target and results from phase one and two of the Fifth Generation Computer Systems study. Future Generation Computer Systems, 4 (4):307–325, March 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2007:LWK

[Kim07a] Kwang-Hoon Kim. A layered workflow knowledge Grid/ P2P architecture and its models for future generation workflow systems. Future Generation Computer Systems, 23 (3):304–316, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2007:SSI

[Kim07b] Sangkyun Kim. Special section: Information engineering and enterprise architecture in distributed computing environments. Future Generation Computer Systems, 23 (2):262, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kukla:1997:ISD

[KK97] Robert Kukla and Jon Kerridge. Intelligent storage devices for scalable information, management systems. Future Generation Computer Systems, 12(5): 335–344, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/18/abstract.
html.

Kacsuk:2000:GED

[KK00] P. Kacsuk and G. Kotsis.

Guest editorial: Distributed and parallel systems. Future Generation Computer Systems, 16(6):v–vii, April 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/25/abstract.
html.

Kandhai:2001:IMR

[KKHS01] D. Kandhai, A. Koponen, A. Hoekstra, and P. M. A.

Sloot. Iterative momentum relaxation for fast lattice-Boltzmann simulations. Future Generation Computer Systems, 18(1):89–96, September 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/35/abstract.
html.

Kim:2007:HPA

[KKK07] Minhyung Kim, Sangkyun Kim, and Arcy J. Kong. High performance AAA architecture for massive IPv4 networks. Future Generation Computer Systems, 23(2): 275–279, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Koo:2006:NSS

[KKL06] Simon G. M. Koo, Karthik Kannan, and C. S. George Lee. On neighbor-selection strategy in hybrid peer-to-peer networks. Future Generation Computer Systems, 22(7): 732–741, August 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2009:SDT

[KKL09a] Seoksoo Kim, Soongohn Kim, and Geuk Lee. Structure design and test of enterprise security management system with advanced internal security. Future Generation Computer Systems, 25 (3):358–363, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2009:SVN

[KKL09b] Soongohn Kim, Seoksoo Kim, and Geuk Lee. Secure verifiable non-interactive oblivious transfer protocol using RSA and bit commitment on distributed environment. Future Generation Computer Systems, 25(3):352–357, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Krawczyk:2000:STC

[KKP00] Henryk Krawczyk, Bartosz Krysztop, and Jerzy Proficz. Suitability of the time controlled environment for race detection in distributed applications. Future Generation Computer Systems, 16 (6):625–635, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/30/abstract.
html.

Kim:2005:CRA

[KKP^{+}05] Hyun Taek Kim, Bo Yeon Kim, Eun Hye Park, Jong Woo Kim, Eui Whan Hwang, Seung Kee Han, and Sunyoung Cho. Computerized recognition of Alzheimer disease-EEG using genetic algorithms and neural network. Future Generation Computer Systems, 21(7):1124–1130, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kacsuk:2008:SGI

[KKS08] Peter Kacsuk, Tamas Kiss, and Gergely Sipos. Solving the grid interoperability problem by P-GRADE portal at workflow level. Future Generation Computer Systems, 24 (7):744–751, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kofidis:1999:WBM

[KKV^{+}99] Eleftherios Kofidis, Nicholas Kolokotronis, Aliki Vassilarakou, Sergios Theodoridis, and Dionisis Cavouras. Wavelet-based medical image compression. Future Generation Computer Systems, 15 (2):223–243, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167739X98000661.

Kim:2004:IRI

[KKYK04] Won Kim, Il-Ju Ko, Jin-Sung Yoon, and Gye-Young Kim. Inference of recommendation information on the Internet using improved FAM. Future Generation Computer Systems, 20(2):265–273, February 16, 2004. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Kutrib:2002:MPF

[KL02] Martin Kutrib and Jan-Thomas Löwe. Massively parallel fault tolerant computations on syntactical patterns. Future Generation Computer Systems, 18(7):905–919, August 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2005:RTI

[KLC05] Jin Ok Kim, Bum Ro Lee, and Chin Hyun Chung. Realtime interactive motion transitions by a uniform posture map. Future Generation Computer Systems, 21(7): 1106–1116, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2004:LPR

[KLH^{+}04] Jin Ok Kim, Woongjae Lee, Jun Hwang, Kyong Seok Baik, and Chin Hyun Chung. Lip print recognition for security systems by multi-resolution architecture. Future Generation Computer Systems, 20(2): 295–301, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Khan:2003:OQS

[KLM^{+}03] Shahadat Khan, Kin F. Li, Eric G. Manning, Robert Watson, and G. C. Shoja. Optimal Quality of Service routing and admission control using the Utility Model. Future Generation Computer Systems, 19(7):1063–1073, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kupczyk:2005:ADE

[KLM^{+}05] Mirosław Kupczyk, Rafał Lichwała, Norbert Meyer, Bartosz Palak, Marcin Płóciennik, and Paweł Wolniewicz. “applications on demand” as the exploitation of the Migrating Desktop. Future Generation Computer Systems, 21(1):37– 44, January 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Klopotek:2005:VLB

[Kło05] Mieczysław A. Kłopotek. Very large Bayesian multinets for text classification. Future Generation Computer Systems, 21(7):1068–1082, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kunszt:2005:FBR

[KLSS05] Peter Kunszt, Erwin Laure, Heinz Stockinger, and Kurt Stockinger. File-based replica management. Future Generation Computer Systems, 21 (1):115–123, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kraeva:2001:ATP

[KM01] M. A. Kraeva and V. E.

Malyshkin. Assembly technology for parallel realization of numerical models on MIMDmulticomputers. Future Generation Computer Systems, 17 (6):755–765, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/34/abstract.
html.

Kost:2003:ACL

[KMCH03] Corrie Kost, Steven Mc-Donald, Bryan Caron, and Wade Hong. ATLAS Canada lightpath data transfer trial. Future Generation Computer Systems, 19(6):1051– 1062, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Korkhov:2009:DWB

[KMK09] Vladimir V. Korkhov, Jakub T.

Moscicki, and Valeria V. Krzhizhanovskaya. Dynamic workload balancing of parallel applications with userlevel scheduling on the Grid. Future Generation Computer Systems, 25(1):28–34, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kuksheva:2005:SSS

[KMN^{+}05] E. A. Kuksheva, V. E. Malyshkin, S. A. Nikitin, A. V. Snytnikov, V. N. Snytnikov, and V. A. Vshivkov. Supercomputer simulation of self-gravitating media. Future Generation Computer Systems, 21(5):749–757, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Knupfer:2006:CMD

[KN06] Andreas Knüpfer and Wolfgang E. Nagel. Compressible memory data structures for event-based trace analysis. Future Generation Computer Systems, 22(3):359– 368, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Knight:1989:NSD

[Kni89] Doyle D. Knight. Numerical simulation of 3-D shock waveturbulent boundary layer interactions. Future Generation Computer Systems, 5(2– 3):275–281, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kim:2008:TOM

[KNK^{+}08] Jik-Soo Kim, Beomseok Nam, Peter Keleher, Michael Marsh, Bobby Bhattacharjee, and Alan Sussman. Trade-offs in matching jobs and balancing load for distributed Desktop Grids. Future Generation Computer Systems, 24 (5):415–424, May 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kobayashi:1992:POS

[Kob92] Yoshikazu Kobayashi. A perspective of OSI standardization: Object-oriented architecture for distributed processing. Future Generation Computer Systems, 7(2–3): 311–316, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kollman:1989:GCC

[Kol89]  Peter Kollman.  Grand

challenges in computational science: Simulations on biological macromolecules. Future Generation Computer Systems, 5(2–3):207– 211, September 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Komatsu:1989:UES

[Kom89a] Osamu Komatsu. The underwriting expert system in Nippon life. Future Generation Computer Systems, 5(1): 119–122, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Komiya:1989:APC

[Kom89b] Seiichi Komiya. Automatic programming by composing program components and its realization method. Future Generation Computer Systems, 5(1):151–161, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Koski:1995:STL

[Kos95] Kimmo Koski. A step towards large scale parallelism: Building a parallel computing environment from heterogeneous resources. Future Generation Computer Systems, 11(4–5): 491–498, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kosch:2000:MOO

[Kos00]  Harald Kosch.  Managing

the operator ordering problem in parallel databases. Future Generation Computer Systems, 16(6):665– 676, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/29/33/abstract.
                                        html.

Kowalski:1984:SEA

[Kow84] Robert Kowalski. Software engineering and artificial intelligence in new generation computing. Future Generation Computer Systems, 1 (1):39–49, July 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kowalski:1985:SEA

[Kow85] R. Kowalski. Software engineering and artificial intelligence in new generation computing. Future Generation Computer Systems, 1(1): 39–49, ???? 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kacsuk:2000:LEM

[KP00] Peter Kacsuk and Norbert Podhorszki. Logicflow execution model for parallel databases. Future Generation Computer Systems, 16 (6):677–692, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/34/abstract.
html.

Karonis:2003:HRR

[KPB^{+}03] Nicholas T. Karonis, Michael E.

Papka, Justin Binns, John Bresnahan, Joseph A. Insley, David Jones, and Joseph M. Link. High-resolution remote rendering of large datasets in a collaborative environment. Future Generation Computer Systems, 19(6):909–917, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

King:2001:AIS

[KRLR01] Roger L. King, Samuel H. Russ, Aric B. Lambert, and Donna S. Reese. An artificial immune system model for intelligent agents. Future Generation Computer Systems, 17 (4):335–343, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/29/25/abstract.
                                        html.

Kobayashi:2002:CVD

[KS02] Kei Kobayashi and Kokichi Sugihara. Crystal Voronoi diagram and its applications. Future Generation Computer Systems, 18(5): 681–692, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/36/35/abstract.
                                        html.

Khonsari:2003:AMA

[KSAOK03] A. Khonsari, H. Sarbazi-Azad, and M. Ould-Khaoua. An analytical model of adaptive wormhole routing with time-out. Future Generation Computer Systems, 19 (1):1–12, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kiasari:2008:AMP

[KSAOK08] A. E. Kiasari, H. Sarbazi-Azad, and M. Ould-Khaoua. An accurate mathematical performance model of adaptive routing in the star graph. Future Generation Computer Systems, 24(6):461–474, June 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kryza:2007:EGO

[KSM^{+}07a] Bartosz Kryza, Renata Słota, Marta Majewska, Jan Pieczykolan, and Jacek Kitowski. Erratum to “Grid organizational memory — provision of a highlevel Grid abstraction layer supported by ontology alignment” [Future Gener. Comput. Syst. 23 (2007) 348–358]. Future Generation Computer Systems, 23(6):773, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). See [KSM^{+}07b].

Kryza:2007:GOM

[KSM^{+}07b] Bartosz Kryza, Renata Słota, Marta Majewska, Jan Pieczykolan, and Jacek Kitowski. Grid organizational memory — provision of a high-level Grid abstraction layer supported by ontology alignment. Future Generation Computer Systems, 23(3):348–358, March 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). See erratum [KSM^{+}07a].

Kim:1992:OOD

[KST92] Won Kim, Mark Scheevel, and Chris Tomlinson. Objectoriented databases for new applications. Future Generation Computer Systems, 7(2– 3):317–327, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kodama:1992:PHP

[KSY92] Yuetsu Kodama, Shuichi Sakai, and Yoshinori Yamaguchi. A prototype of a highly parallel dataflow machine EM-4 and its preliminary evaluation. Future Generation Computer Systems, 7(2–3):199–209, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kacsuk:2008:SSG

[KT08] Peter Kacsuk and Gabor Terstyanszky. Special section: Grid-enabling legacy applications and supporting end users. Future Generation Computer Systems, 24 (7):709–710, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kyriazis:2008:IWM

[KTM^{+}08] Dimosthenis Kyriazis, Konstantinos Tserpes, Andreas Menychtas, Antonis Litke, and Theodora Varvarigou. An innovative workflow mapping mechanism for Grids in the frame of Quality of Service. Future Generation Computer Systems, 24(6):498–511, June 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kranzlmuller:2003:EDL

[KTV03] Dieter Kranzlmüller, Nam Thoai, and Jens Volkert. Error detection in large-scale parallel programs with long runtimes. Future Generation Computer Systems, 19 (5):689–700, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kaneda:2003:VPG

[KTY03] Kenji Kaneda, Kenjiro Taura, and Akinori Yonezawa. Virtual private grid: a command shell for utilizing hundreds of machines efficiently. Future Generation Computer Systems, 19(4):563–573, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kutil:2001:PAW

[KU01] R. Kutil and A. Uhl. Parallel adaptive wavelet analysis. Future Generation Computer Systems, 18(1):97– 106, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/36/abstract.
html.

Kuntz:1994:PEC

[Kun94] Jean-Marc Kuntz. Performance evaluation of cache architectures in tightly coupled multiprocessor systems. Future Generation Computer Systems, 10(1):15–27, April 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kuo:1986:RVI

[Kuo86] Franklin F. Kuo. A return visit to ICOT. Future Generation Computer Systems, 2 (1):61–63, March 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kranzlmuller:2003:TPD

[KV03] Dieter Kranzlmüller and Jens Volkert. Tools for program development and analysis. Future Generation Computer Systems, 19(5):597–598, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kokkinos:2009:FPH

[KV09] Panagiotis Kokkinos and Emmanouel A. Varvarigos. A framework for providing hard delay guarantees and user fairness in Grid computing. Future Generation Computer Systems, 25(6):674–686, June 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kato:1985:OND

[KY85] Masao Kato and Yutaka Yoshida. An overview of NTTs digital transmission networks — existing and planned. Future Generation Computer Systems, 1(5):309– 323, September 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kwon:2004:GDR

[KY04] Jin B. Kwon and Heon Y.

Yeom. Generalized data retrieval for pyramid-based periodic broadcasting of videos. Future Generation Computer Systems, 20(1):157–170, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kokol:1999:NIE

[KZBK99] Peter Kokol, Damjan Zazula, Viljem Brumec, and Ljudmila Kolenc. Nursing informatics education for the next millennium. Future Generation Computer Systems, 15(2): 211–216, March 11, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/17/23/abstract.
html.

Kang:2004:PIP

[KZC04] Ning Kang, Jun Zhang, and Eric S. Carlson. Performance of ILU preconditioning techniques in simulating anisotropic diffusion in the human brain. Future Generation Computer Systems, 20(4): 687–698, May 3, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kale:2006:SAM

[KZLK06] Laxmikant V. Kalé, Gengbin Zheng, Chee Wai Lee, and Sameer Kumar. Scaling applications to massively parallel machines using projections performance analysis tool. Future Generation Computer Systems, 22(3): 347–358, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2007:FTP

[LAM07] Lu Liu, Nick Antonopoulos, and Stephen Mackin. Fault-tolerant peer-to-peer search on small-world networks. Future Generation Computer Systems, 23(8): 921–931, November 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lang:2000:CSW

[Lan00] Ulrich Lang. CORBA security on the Web — an overview. Future Generation Computer Systems, 16(4): 417–421, February 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/36/abstract.
html.

Laue:1992:AAB

[Lau92] Herwig A. Laue. ATOS — an AI-based space mission operations system. Future Generation Computer Systems, 7 (4):439–451, May 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Laure:2001:OJF

[Lau01]  Erwin Laure.  OpusJava:

A Java framework for distributed high performance computing. Future Generation Computer Systems, 18(2): 235–251, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/31/abstract.
html.

Lan:2003:ECA

[LB03] Zhiling Lan and Valerie E.

Taylorand Greg Bryan. Exploring cosmology applications on distributed environments. Future Generation Computer Systems, 19(6): 839–847, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2009:MBS

[LB09] Hui Li and Rajkumar Buyya.

Model-based simulation and performance evaluation of Grid scheduling strategies. Future Generation Computer Systems, 25(4):460–465, April 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lorenz:2009:JMS

[LBB^{+}09] D. Lorenz, S. Borovac, P. Buchholz, H. Eichenhardt, T. Harenberg, P. Mättig, M. Mechtel, R. Müller-Pfefferkorn, R. Neumann, K. Reeves, Ch. Uebing, W. Walkowiak, Th. William, and R. Wismüller. Job monitoring and steering in D-Grid’s High Energy Physics Community Grid. Future Generation Computer Systems, 25(3):308–314, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2002:GEB

[LBR02] Craig A. Lee, Rajkumar Buyya, and Paul Roe. Guest editorial: The best papers from CCGrid 2001. Future Generation Computer Systems, 18(4):v–vii, March 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/33/27/abstract.
html.

Li:2008:SKM

[LBYL08] Jason H. Li, Bobby Bhattacharjee, Miao Yu, and Renato Levy. A scalable key management and clustering scheme for wireless ad hoc and sensor networks. Future Generation Computer Systems, 24 (8):860–869, October 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lin:2001:GEC

[LC01] Hai Xiang Lin and Tony Chan.

Guest editorial: Combining the power of high speed computer systems and efficient algorithms. Future Generation Computer Systems, 18 (1):1, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/27/abstract.
html.

Lu:2003:DIF

[LC03] Eric Jui-Lin Lu and Rai-Fu Chen. Design and implementation of a fine-grained menu control processor for Web-based information systems. Future Generation Computer Systems, 19(7): 1105–1119, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Latt:2004:VCM

[LC04] Jonas Lätt and Bastien Chopard. VLADYMIR — a C++ matrix library for dataparallel applications. Future Generation Computer Systems, 20(6):1023–1039, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2005:SAA

[LC05] Jiageng Li and David Cordes. A scalable authorization approach for the Globus grid system. Future Generation Computer Systems, 21(2):291–301, February 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2004:DEB

[LCP04] Chang-Doo Lee, Bong-Jun Choi, and Kyoo-Seok Park. Design and evaluation of a block encryption algorithm using dynamic-key mechanism. Future Generation Computer Systems, 20(2): 327–338, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lorenzetto:2004:ALT

[LD04] Gian Paolo Lorenzetto and Amitava Datta. An almost linear-time algorithm for trapezoidation of GIS polygons. Future Generation Computer Systems, 20(7):1145– 1155, October 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lindemann:2006:UCM

[LDS06] J. Lindemann, O. Dahlblom, and G. Sandberg. Using CORBA middleware in finite element software. Future Generation Computer Systems, 22 (1–2):158–193, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lischka:1995:PCQ

[LDSH95] Hans Lischka, Holger Dachsel, Ron Shepard, and Robert J. Harrison. Parallel computing in quantum chemistry — Message passing and beyond for a general ab initio program system. Future Generation Computer Systems, 11(4– 5):445–450, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2004:VSQ

[Lee04]  Yong Woo Lee.  A vir-

tual server queueing network method for component based performance modelling of metacomputing. Future Generation Computer Systems, 20(1):145–155, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lenz:2001:MPS

[Len01] Jürgen Lenz. Many-particle simulation of ameboid motility. Future Generation Computer Systems, 17(7):

863–872, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/34/29/abstract.
html.

Leopold:1998:ASD

[Leo98] Claudia Leopold. Arranging statements and data of program instances for locality. Future Generation Computer Systems, 14(5–6):293– 311, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Leopold:2001:SSS

[Leo01] Claudia Leopold. Structuring statement sequences in instance-based locality optimization. Future Generation Computer Systems, 17 (4):425–440, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/33/abstract.
html.

Lanteri:1995:UCC

[LF95a] Stéphane Lanteri and Charbel Farhat. Unstructured CFD computations on the KSR-1: Preliminary results. Future Generation Computer Systems, 11(1):27–33, February 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Loriot:1995:FFC

[LF95b] Mark Loriot and Loula Fezoui. FEM/FVM calculations of compressible flows on a Meiko system. Future Generation Computer Systems, 11 (1):7–18, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2001:BMR

[LF01] Kangsun Lee and Paul A.

Fishwick. Building a model for real-time simulation. Future Generation Computer Systems, 17(5):585–600, March 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.elsevier.com/gej-ng/
10/19/19/45/30/32/abstract.
html.

Lendasse:2005:VQW

[LFWV05] A. Lendasse, D. François, V. Wertz, and M. Verleysen. Vector quantization: a weighted version for timeseries forecasting. Future Generation Computer Systems, 21 (7):1056–1067, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lefevre:2008:IIA

[LG08] Laurent Lefèvre and Jean-Patrick Gelas. IAN^{2}: Industrial Autonomic Network Node architecture for supporting personalized network services in the industrial context.

Future Generation Computer Systems, 24(1):58–65, January 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Larkin:1997:LSD

[LGH97] Steve Larkin, Andrew J. Grant, and W. T. Hewitt. Libraries to support distribution and processing of visualization data sets. Future Generation Computer Systems, 12(5): 431–440, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/25/abstract.
html.

Luchnikov:2002:VDA

[LGMV02] V. A. Luchnikov, M. L. Gavrilova, N. N. Medvedev, and V. P. Voloshin. The Voronoi–Delaunay approach for the free volume analysis of a packing of balls in a cylindrical container. Future Generation Computer Systems, 18 (5):673–679, April ??, 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/34/abstract.
html.

Lloyd:2007:IBC

[LGS^{+}07] Sharon Lloyd, Dave Gavaghan, Andrew Simpson, Matthew Mascord, Clint Seneurine, Geoff Williams, Joe Pitt-Francis, David Boyd, Damian Mac Randal, Lakshmi Sastry, Srikanth Nagella, Katie Weeks, Ronald Fowler, Daniel Hanlon, James Handley, and Gianni de Fabritiis. Integrative biology — the challenges of developing a collaborative research environment for heart and cancer modelling. Future Generation Computer Systems, 23 (3):457–465, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2007:MPD

[LGW07] Hui Li, David Groep, and Lex Wolters. Mining performance data for metascheduling decision support in the Grid. Future Generation Computer Systems, 23(1):92–99, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2007:DBP

[LH07] Jianxun Liu and Jinmin Hu.

Dynamic batch processing in workflows: Model and implementation. Future Generation Computer Systems, 23(3): 338–347, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Libeskind-Hadas:1995:OBF

[LHB95] Ran Libeskind-Hadas and Eli Brandt. Origin-based faulttolerant routing in the mesh. Future Generation Computer Systems, 11(6):603–615, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lin:2003:NKA

[LHC03] Iuon-Chang Lin, Min-Shiang Hwang, and Chin-Chen Chang. A new key assignment scheme for enforcing complicated access control policies in hierarchy. Future Generation Computer Systems, 19(4):457–462, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lin:2003:NRU

[LHL03] Iuon-Chang Lin, Min-Shiang Hwang, and Li-Hua Li. A new remote user authentication scheme for multi-server architecture. Future Generation Computer Systems, 19 (1):13–22, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Leal:2009:DMS

[LHL09] Katia Leal, Eduardo Huedo, and Ignacio M. Llorente. A decentralized model for scheduling independent tasks in Federated Grids. Future Generation Computer Systems, 25 (8):840–852, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:1990:TTA

[Li90] Wei Li. A type-theoretic approach for program development. Future Generation Computer Systems, 6 (1):49–63, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liddell:1999:HPC

[Lid99] Heather M. Liddell. High-Performance Computing and Networking Europe 1998. Future Generation Computer Systems, 15(3):307– 308, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/31/abstract.
html.

Lindamood:1984:SJF

[Lin84] George E. Lindamood. The structure of the Japanese fifth generation computer project — then and now. Future Generation Computer Systems, 1 (1):51–55, July 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Little:2003:CSM

[Lit03] Leigh J. Little. The computational science major at SUNY Brockport. Future Generation Computer Systems, 19(8): 1285–1292, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2007:WFJ

[LJ07] Gun Ho Lee and Junsu Jung. Web framework with Java and XML in multi-tiers for productivity. Future Generation Computer Systems, 23(2): 263–268, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2005:RBV

[LJPS05] Hua Liu, Lian Jiang, Manish Parashar, and Deborah Silver. Rule-based visualization in the Discover computational steering collaboratory. Future Generation Computer Systems, 21(1):53–59, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2008:OPM

[LJW08] Qin Liu, Xiaohua Jia, and Chanle Wu. Optimal precomputation for mapping service level agreements in Grid computing. Future Generation Computer Systems, 24(8): 849–859, October 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2004:DIM

[LJY04] Bong-Hwan Lee, Il-Hong Jung, and Chan-Hyun Youn. Design and implementation of MPλS network simulator. Future Generation Computer Systems, 20(2):339–347, February 16, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2008:GQO

[LK08] Shuo Liu and Hassan A.

Karimi. Grid query optimizer to improve query processing in Grids. Future Generation Computer Systems, 24 (5):342–353, May 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Litke:2008:MSL

[LKA^{+}08] Antonios Litke, Kleopatra Konstanteli, Vassiliki Andronikou, Sotirios Chatzis, and Theodora Varvarigou. Managing service level agreement contracts in OGSAbased Grids. Future Generation Computer Systems, 24 (4):245–258, April 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2007:ESA

[LKG07] Seokcheon Lee, Soundar Kumara, and Natarajan Gautam. Efficient scheduling algorithm for component-based networks. Future Generation Computer Systems, 23 (4):558–568, May 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liakopoulos:2008:GEH

[LKG08] P. I. K. Liakopoulos, I. C. Kampolis, and K. C. Giannakoglou. Grid enabled, hierarchical distributed metamodel-assisted evolutionary algorithms for aerodynamic shape optimization. Future Generation Computer Systems, 24(7):701–708, July 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lockemann:1991:FDT

[LKM91] Peter C. Lockemann, Alfons Kemper, and Guido Moerkotte. Future database technology: driving forces and directions. Future Generation Computer Systems, 7(1): 41–54, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2003:CCA

[LL03] Chunlin Li and Layuan Li. Combine concept of agent and service to build distributed object-oriented system. Future Generation Computer Systems, 19(2):161– 171, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lagana:2004:F

[LL04a] Antonio Laganà and Noelia Faginas Lago. Foreword. Future Generation Computer Systems, 20(5):701–702, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2004:DIS

[LL04b] Wongoo Lee and Jaekwang Lee. Design and implementation of secure e-mail system using elliptic curve cryptosystem. Future Generation Computer Systems, 20(2): 315–326, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2004:CPR

[LL04c] Chunlin Li and Layuan Li.

Competitive proportional resource allocation policy for computational grid. Future Generation Computer Systems, 20(6):1041–1054, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2003:GTF

[LLH^{+}03] David Lee, Abel W. Lin, Tom Hutton, Toyokazu Akiyama, Shimojo Shinji, Fang-Pang Lin, Steven Peltier, and Mark H. Ellisman. Global telescience featuring IPv6 at iGrid2002. Future Generation Computer Systems, 19(6): 1031–1039, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2009:LBC

[LLKF09] Byoung-Hoon Lee, Sung-Hwa Lim, Jai-Hoon Kim, and Geoffrey C. Fox. Lease-based consistency schemes in the Web environment. Future Generation Computer Systems, 25 (1):8–19, January 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Linden:1994:SAP

[LLRS94] J. Linden, G. Lonsdale, H. Ritzdorf, and A. Schüller. Scalability aspects of parallel multigrid. Future Generation Computer Systems, 10 (4):429–439, November 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Luo:2007:DDMb

[LLSH07] Ping Luo, Kevin Lü, Zhongzhi Shi, and Qing He. Distributed data mining in grid computing environments. Future Generation Computer Systems, 23(1): 84–91, January 1, 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ludwig:2002:TEC

[LLSR02] T. Ludwig, M. Lindermeier, A. Stamatakis, and G. Rackl. Tool environments in CORBA-based medical high-performance computing. Future Generation Computer Systems, 18(6):841–847, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2004:ENN

[LLWN04] Zhengjun Liu, Aixia Liu, Changyao Wang, and Zheng Niu. Evolving neural network using real coded genetic algorithm (GA) for multispectral image classification. Future Generation Computer Systems, 20(7):1119– 1129, October 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2007:SHP

[LLZ07] Hongjun Liu, Ping Luo, and Zhifeng Zeng. A structured hierarchical P2P model based on a rigorous binary tree code algorithm. Future Generation Computer Systems, 23(2): 201–208, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Livezey:1990:ACS

[LM90a] Brian K. Livezey and Richard R.

Muntz. ASPEN: a concurrent stream processing environment. Future Generation Computer Systems, 6(3):197– 207, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lue:1990:GSB

[LM90b] K. Lue and K. Miyai. Graphics supercomputer benchmark: Basic performance of two graphics supercomputers: Stellar GS1000 and Ardent Titan-2. Future Generation Computer Systems, 5(4):387– 394, January 1, 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liao:2007:TBG

[LM07] Lijun Liao and Mark Manulis. Tree-based group key agreement framework for mobile ad-hoc networks. Future Generation Computer Systems, 23 (6):787–803, July 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Levin:1989:DME

[LMBCC89] Simon A. Levin, Kirk Moloney, Linda Buttel, and Carlos Castillo-Chavez. Dynamical models of ecosystems and epidemics. Future Generation Computer Systems, 5(2– 3):265–274, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lynden:2009:DIO

[LMH^{+}09] Steven Lynden, Arijit Mukherjee, Alastair C. Hume, Alvaro A. A. Fernandes, Norman W. Paton, Rizos Sakellariou, and Paul Watson. The design and implementation of OGSA–DQP: a servicebased distributed query processor. Future Generation Computer Systems, 25(3): 224–236, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Louter-Nool:1994:PMC

[LN94] Margreet Louter-Nool. A parallel multigrid code with a fast vectorized ILU-relaxation. Future Generation Computer Systems, 10(2–3):309–313, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2004:LUA

[LNJ04] T. H. Li, J. R. Ni, and W. X. Ju. Land-use adjustment with a modified soil loss evaluation method supported by GIS. Future Generation Computer Systems, 20(7): 1185–1195, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lim:2007:EHP

[LOJ^{+}07] Dudy Lim, Yew-Soon Ong, Yaochu Jin, Bernhard Sendhoff, and Bu-Sung Lee. Efficient Hierarchical Parallel Genetic Algorithms using Grid computing. Future Generation Computer Systems, 23 (4):658–670, May 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Loucif:2009:PAD

[LOK09] S. Loucif and M. Ould-Khaoua. Performance analysis of deterministically-routed bi-directional torus with nonuniform traffic distribution. Future Generation Computer Systems, 25(5):489–498, May 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lopriore:1993:DCP

[Lop93] Lanfranco Lopriore. A data cache for Prolog architectures. Future Generation Computer Systems, 9(3):219– 234, September 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

LopezdeMantaras:1996:RML

[Lop96] Ramon Lopez de Mantaras. Reflection and meta-level artificial intelligence architectures. Future Generation Computer Systems, 12(2–3): 119–121, September 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lopez:2003:SDS

[Lop03] Luciano Lopez. Structural dynamical systems in linear algebra and control: computational aspects. Future Generation Computer Systems, 19 (7):1123–1124, October 2003.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lowe:2001:HPC

[Low01]  Christopher P. Lowe.  A

hybrid particle/continuum model for micro-organism motility. Future Generation Computer Systems, 17 (7):853–862, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/34/28/abstract.
html.

Low:2005:DAP

[Low05] Colin Low. Decentralised application placement. Future Generation Computer Systems, 21(2):281–290, February 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Launay:2001:EPP

[LP01] Pascale Launay and Jean-Louis Pazat. Easing parallel programming for clusters with Java. Future Generation Computer Systems, 18(2): 253–263, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/32/abstract.
html.

Lagana:2004:PSQ

[LPB04] A. Laganà, L. Pacifici, and D. Bellucci. Parallelization strategies for quantum reactive scattering codes. Future Generation Computer Systems, 20(5):829–840, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lang:1995:PCS

[LPC^{+}95] Ulrich Lang, J. P. Peltier, Paul Christ, Stefan Rill, Dirk Rantzau, Harald Nebel, Andreas Wierse, Ruth Lang, Sylvain Causse, Frédéric Juaneda, Michel Grave, and Peter Haas. Perspectives of collaborative supercomputing and networking in European aerospace research and industry. Future Generation Computer Systems, 11(4–5): 419–430, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lloret:2008:SCB

[LPE08] Jaime Lloret, Carlos Palau, and Manuel Esteve. Structuring connections between content delivery servers groups. Future Generation Computer Systems, 24(3):191– 201, March 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lourens:1994:LSN

[LPK94] T. Lourens, N. Petkov, and P. Kruizinga. Large scale natural vision simulations. Future Generation Computer Systems, 10(2–3):351–358, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lecomber:2001:EPA

[LR01] David Lecomber and Mike Rudgyard. Efficient parallel algorithms for numerical simulation. Future Generation Computer Systems, 17 (8):961–967, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/32/abstract.
html.

Le:2006:DMC

[LR06] Thuy T. Le and Jalel Rejeb. A detailed MPI communication model for distributed systems. Future Generation Computer Systems, 22(3): 269–278, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Leigh:2006:GLV

[LRJ^{+}06] Jason Leigh, Luc Renambot, Andrew Johnson, Byungil Jeong, Ratko Jagodic, Nicholas Schwarz, Dmitry Svistula, Rajvikram Singh, Julieta Aguilera, Xi Wang, Venkatram Vishwanath, Brenda Lopez, Dan Sandin, Tom Peterka, Javier Girado, Robert Kooima, Jinghua Ge, Lance Long, Alan Verlo, Thomas A. DeFanti, Rollin Feld, Jacob Balser, Steve Morris, Trevor Henthorn, Greg Dawe, Peter Otto, and Larry Smarr. The Global Lambda Visualization Facility: An international ultra-high-definition wide-area visualization collaboratory. Future Generation Computer Systems, 22(8): 964–971, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Luque:1994:SPP

[LRMC94] E. Luque, A. Ripoll, T. Margalef, and A. Cortés. Scheduling of parallel programs including dynamic loops. Future Generation Computer Systems, 10(2–3):301–304, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2001:WMB

[LRW01] Maozhen Li, Omer F. Rana, and David W. Walker. Wrapping MPI-based legacy codes as Java/CORBA components. Future Generation Computer Systems, 18(2): 213–223, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/31/29/abstract.
                                        html.

Loh:2001:GBF

[LS01] Peter K. K. Loh and Venson Shaw. A genetic-based fault-tolerant routing strategy for multiprocessor networks. Future Generation Computer Systems, 17(4): 415–423, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/32/abstract.
html.

Li:2005:NWN

[LS05] Shuyou Li and Binheng Song.

Normalized workflow net (NWF-net): Its definition and properties. Future Generation Computer Systems, 21(7): 1004–1014, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Laccetti:2007:FMG

[LS07a] G. Laccetti and G. Schmid.

A framework model for Grid security. Future Generation Computer Systems, 23(5): 702–713, June 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lamehamedi:2007:DDM

[LS07b] Houda Lamehamedi and Boleslaw K. Szymanski. Decentralized data management framework for Data Grids. Future Generation Computer Systems, 23(1): 109–115, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Luckow:2008:MFT

[LS08] André Luckow and Bettina Schnor. Migol: a faulttolerant service framework for MPI applications in the Grid. Future Generation Computer Systems, 24(2): 142–152, February 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Luque:1994:PET

[LSF^{+}94] E. Luque, M. A. Senar, D. Franco, P. Hernández, E. Heymann, and J. C. Moure. Programming environment for a transputer based computer. Future Generation Computer Systems, 10(2– 3):295–299, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Leangsuksun:2005:AHA

[LSLS05] Chokchai Box Leangsuksun, Lixin Shen, Tong Liu, and Stephen L. Scott. Achieving high availability and performance computing with an HA-OSCAR cluster. Future Generation Computer Systems, 21(4):597–606, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Luque:1994:SPS

[LSS94] Emilio Luque, Remo Suppi, and Joan Sorribes. Simulation of parallel systems: PSEE (Parallel System Evaluation Environment). Future Generation Computer Systems, 10(2–3):291–294, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Litke:2007:ETR

[LSTV07] Antonios Litke, Dimitrios Skoutas, Konstantinos Tserpes, and Theodora Varvarigou. Efficient task replication and management for adaptive fault tolerance in Mobile Grid environments. Future Generation Computer Systems, 23(2): 163–178, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Laskowski:2007:BCS

[LTOT07] Eryk Laskowski, Marek Tudruj, Richard Olejnik, and Bernard Toursel. Byte-code scheduling of Java programs with branches for Desktop Grid. Future Generation Computer Systems, 23(8): 977–982, November 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lukkien:1989:SED

[Luk89] Johan J. Lukkien. Some experience in distributed programming using shortest path spanning trees. Future Generation Computer Systems, 4 (4):281–284, March 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Luksch:2000:PDI

[Luk00] Peter Luksch. Parallel and distributed implementation of large industrial applications. Future Generation Computer Systems, 16(6): 649–663, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/29/32/abstract.
                                        html.

Lei:2008:LRS

[LVH08] Ming Lei, Susan V. Vrbsky, and Xiaoyan Hong. An online replication strategy to increase availability in Data Grids. Future Generation Computer Systems, 24(2):85– 98, February 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2004:SSO

[LvSW^{+}04] M. Li, P. van Santen, D. W. Walker, O. F. Rana, and M. A. Baker. SGrid: a serviceoriented model for the Semantic Grid. Future Generation Computer Systems, 20(1):7– 18, January 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lin:2008:GPB

[LW08] Maria Lin and David W.

Walker. GECEM: a portalbased Grid application for computational electromagnetics. Future Generation Computer Systems, 24(1):66– 72, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2007:BSS

[LWHC07] Sung Lee, Taowei David Wang, Nada Hashmi, and Michael P. Cummings. Bio-STEER: a Semantic Web workflow tool for Grid computing in the life sciences. Future Generation Computer Systems, 23(3):497– 509, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Luo:2007:DDMa

[LWHS07] Jiewen Luo, Maoguang Wang, Jun Hu, and Zhongzhi Shi. Distributed data mining on Agent Grid: Issues, platform and development toolkit. Future Generation Computer Systems, 23(1):61–68, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liang:2007:GES

[LWSC07] Tyng-Yeu Liang, Chun-Yi Wu, Ce-Kuen Shieh, and Jyh-Biau Chang. A grid-enabled software distributed shared memory system on a wide area network. Future Generation Computer Systems, 23 (4):547–557, May 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lee:2009:SSS

[LXLS09] Wang-Chien Lee, Jianliang Xu, Jianzhong Li, and Fabrizio Silvestri. Special section: Scalable information systems. Future Generation Computer Systems, 25(1):51–52, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liangliang:1990:CRC

[LY90a] Li Liangliang and Ci Yungui. Clause representations in a compiler-based Prolog database. Future Generation Computer Systems, 6 (1):17–23, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liangliang:1990:IIP

[LY90b] Li Liangliang and Ci Yungui. An integrated implementation of Prolog database operations. Future Generation Computer Systems, 6 (1):11–16, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2009:HLB

[LYMZ09] Yajun Li, Yuhang Yang, Maode Ma, and Liang Zhou. A hybrid load balancing strategy of sequential tasks for grid computing environments. Future Generation Computer Systems, 25(8): 819–828, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Li:2006:PPG

[LYQ06] Maozhen Li, Bin Yu, and Man Qi. PGGA: a predictable and grouped genetic algorithm for job scheduling. Future Generation Computer Systems, 22 (5):588–599, April 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:2005:VKC

[LYT^{+}05] Yunfeng Liu, Dongqing Yang, Shiwei Tang, Tengjiao Wang, and Jun Gao. Validating key constraints over XML document using XPath and structure checking. Future Generation Computer Systems, 21 (4):583–595, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Maeda:1989:KBS

[Mae89] Kazuo Maeda. A knowledgebased system for the wastewater treatment plant. Future Generation Computer Systems, 5(1):29–32, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Maiocchi:1991:SE

[Mai91] Marco Maiocchi. Software engineering. Future Generation Computer Systems, 7(1): 23–29, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Male:1994:PCC

[Mal94] Jean-Michel Malé. Parallel calculations on the CM-2: BBS test programs and spectral methods. Future Generation Computer Systems, 10(4): 381–389, November 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Malyshkin:2001:EPC

[Mal01] Victor Malyshkin. Editorial: Parallel computing technologies. Future Generation Computer Systems, 17 (6):667–668, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/33/25/abstract.
                                        html.

Malyshkin:2002:PCT

[Mal02]  Victor Malyshkin.  Paral-

lel computing technologies. Future Generation Computer Systems, 18(6):v–vi, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Malyshkin:2005:PCT

[Mal05]  Victor Malyshkin.  Paral-

lel computing technologies. Future Generation Computer Systems, 21(5):625–626, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mambretti:2009:OEA

[Mam09] Joe Mambretti. OptIPuter:

Enabling advanced applications with novel optical control planes and backplanes. Future Generation Computer Systems, 25(2): 137–141, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Marchand:1986:CES

[Mar86] H. Marchand. On commercial expert systems projects. Future Generation Computer Systems, 2(4):213–216, December 1986. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Martinez:1990:SMA

[Mar90] Tony R. Martinez. Smart memory architecture and methods. Future Generation Computer Systems, 6(2):145– 162, November 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Marsh:1998:CWB

[Mar98a] Andy Marsh. A 21st Century Web-based telemedical information society. Future Generation Computer Systems, 14 (1–2):1–2, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/17/abstract.
html.

Marsh:1998:CCT

[Mar98b]  Andy Marsh.  The cre-

ation of a 21st Century telemedical information society. Future Generation Computer Systems, 14(1–2): 3–22, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/18/abstract.
html.

Marsh:1999:CVM

[Mar99a] Andy Marsh. The creation of a virtual medical worlds community. Future Generation Computer Systems, 15(2): 131–132, March 11, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/17/17/abstract.
html.

Marsh:1999:EPT

[Mar99b] Andy Marsh. The establishment of a pilot telemedical information society. Future Generation Computer Systems, 15(2):133–156, March 11, 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/17/18/abstract.
html.

Marchese:2002:DDA

[Mar02] Fabio M. Marchese. A directional diffusion algorithm on cellular automata for robot path-planning. Future Generation Computer Systems, 18(7):983–994, August 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Matsumoto:1988:NNM

[Mat88] Gen Matsumoto. Neurocomputing — neurons as microcomputers. Future Generation Computer Systems, 4(1): 39–51, August 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Matsumoto:1989:ESS

[Mat89]  Shunji Matsumoto.  ES/

SDEM software development engineering methodology for expert systems. Future Generation Computer Systems, 5 (1):33–39, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Marinescu:2001:BMD

[MB01] Dan C. Marinescu and Ladislau Bölöni. Biological metaphors in the design of complex software systems. Future Generation Computer Systems, 17(4): 345–360, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/29/26/abstract.
                                        html.

Merlin:1999:MDP

[MBFC99] John Merlin, Scott Baden, Stephen Fink, and Barbara Chapman. Multiple data parallelism with HPF and KeLP. Future Generation Computer Systems, 15(3): 393–405, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/19/24/abstract.
                                        html.

McPhillips:2009:SWD

[MBZL09] Timothy McPhillips, Shawn Bowers, Daniel Zinn, and Bertram Ludäscher. Scientific workflow design for mere mortals. Future Generation Computer Systems, 25(5):541–551, May 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Maniezzo:2000:AHF

[MC00] Vittorio Maniezzo and Antonella Carbonaro. An ANTS heuristic for the frequency assignment problem. Future Generation Computer Systems, 16(8): 927–935, June 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/31/30/abstract.
html.

Min:2004:AIT

[MC04] Byoung Joon Min and Joong Sup Choi. An approach to intrusion tolerance for missioncritical services using adaptability and diverse replication. Future Generation Computer Systems, 20(2):303–313, February 16, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Martin:2002:FAC

[MCA02] Marc Martin, Bastien Chopard, and Paul Albuquerque. Formation of an ant cemetery: swarm intelligence or statistical accident? Future Generation Computer Systems, 18(7): 951–959, August 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

McColl:1996:SPP

[McC96] W. F. McColl. Scalability, portability and predictability: The BSP approach to parallel programming. Future Generation Computer Systems, 12 (4):265–272, December 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Maad:2007:TCG

[MCQ^{+}07] Soha Maad, Brian Coghlan, Geoff Quigley, John Ryan, Eamonn Kenny, and David O’Callaghan. Towards a complete grid filesystem functionality. Future Generation Computer Systems, 23 (1):123–131, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Malony:2000:CEU

[MCSS00] Allen D. Malony, Janice E. Cuny, Jenifer L. Skidmore, and Matthew J. Sottile. Computational experiments using distributed tools in a Webbased electronic notebook environment. Future Generation Computer Systems, 16(5): 453–464, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/28/28/abstract.
                                        html.

Mastroianni:2009:SSP

[MCT^{+}09] Carlo Mastroianni, Pasquale Cozza, Domenico Talia, Ian Kelley, and Ian Taylor. A scalable super-peer approach for public scientific computation. Future Generation Computer Systems, 25(3):

213–223, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Magee:1992:CAP

[MD92a] Jeff Magee and Naranker Dulay. A configuration approach to parallel programming. Future Generation Computer Systems, 8(4):337– 347, September 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Masini:1992:ISC

[MD92b] Andrea Masini and Marco Danelutto. Implementation of a synchronous communication in a loosely coupled system: a correctness proof. Future Generation Computer Systems, 8 (1–3):137–147, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

McKoy:1989:SCS

[MDD89] Vincent McKoy, Richard L.

Dubs, and S. N. Dixit. Some computational studies in molecular physics. Future Generation Computer Systems, 5(2–3):213– 223, September 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Messina:2002:FSI

[Mes02] Paul Messina. Foreword to special issue of FGCS on Grid computing. Future Generation Computer Systems, 18(8): vii–ix, October 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Meuwly:2005:RCB

[Meu05] Markus Meuwly. Reactions in complex biologically relevant systems: challenges for computational approaches. Future Generation Computer Systems, 21(8):1285–1297, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mitchell:1993:MMD

[MF93] P. J. Mitchell and D. Fincham. Multicomputer molecular dynamics. Future Generation Computer Systems, 9(1):5–10, May 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mirto:2008:BGA

[MFE^{+}08] Maria Mirto, Sandro Fiore, Italo Epicoco, Massimo Cafaro, Silvia Mocavero, Euro Blasi, and Giovanni Aloisio. A Bioinfomatics Grid Alignment Toolkit. Future Generation Computer Systems, 24 (7):752–762, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Martin-Flatin:2005:HSN

[MFP05] Jean-Philippe Martin-Flatin and Pascale Vicat-Blanc Primet. High-speed networks and services for data-intensive Grids: The DataTAG Project. Future Generation Computer Systems, 21(4):439–442, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Meirosu:2005:NGE

[MGH^{+}05] Catalin Meirosu, Piotr Golonka, Andreas Hirstius, Stefan Stancu, Bob Dobinson, Erik Radius, Antony Antony, Freek Dijkstra, Johan Blom, and Cees de Laat. Native 10 Gigabit Ethernet experiments over long distances. Future Generation Computer Systems, 21 (4):457–468, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Morin:2004:TES

[MGLV04] Christine Morin, Pascal Gallard, Renaud Lottiaux, and Geoffroy Vallée. Towards an efficient single system image cluster operating system. Future Generation Computer Systems, 20(4):505–521, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mambretti:2006:ACA

[MGYC06] Joe Mambretti, Rachel Gold, Fei Yeh, and Jim Chen. AM-ROEBA: Computational astrophysics modeling enabled by dynamic lambda switching. Future Generation Computer Systems, 22(8):949–954, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Malony:2001:TAA

[MH01] Allen D. Malony and B. Robert Helm. A theory and architecture for automating performance diagnosis. Future Generation Computer Systems, 18 (1):189–200, September 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/27/44/abstract.
                                        html.

Molto:2008:SOW

[MHA08] G. Moltó, V. Hernández, and J. M. Alonso. A serviceoriented WSRF-based architecture for metascheduling on computational Grids. Future Generation Computer Systems, 24(4):317–328, April 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Molto:2009:ARW

[MHA09] G. Moltó, V. Hernández, and J. M. Alonso. Automatic replication of WSRF-based Grid services via operation providers. Future Generation Computer Systems, 25 (8):876–883, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Matsuoka:2001:TPE

[MI01] Satoshi Matsuoka and Shigeo Itou. Towards performance evaluation of highperformance computing on multiple Java platforms. Future Generation Computer Systems, 18(2):281–291, October 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/34/abstract.
html.

Michaud:1997:CT

[Mic97] Pierre Michaud. Clustering techniques. Future Generation Computer Systems, 13(2–3): 135–147, November 14, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/19/abstract.
html.

Middelham:2001:TSP

[Mid01]  Frans Middelham.  Traf-

fic simulation: Predictability: some thoughts on modeling. Future Generation Computer Systems, 17(5): 627–636, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/35/abstract.
html.

Misra:1992:LCP

[Mis92]  Jayadev Misra.  Loosely-

coupled processes. Future Generation Computer Systems, 8(4):269–286, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mizoguchi:1989:BOC

[Miz89a] Riichiro Mizoguchi. A brief overview of the current status of expert systems in Japan. Future Generation Computer Systems, 5(1):7–10, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mizoguchi:1989:P

[Miz89b] Riichiro Mizoguchi. Preface. Future Generation Computer Systems, 5(1):3–6, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Marovic:1998:VFM

[MJ98] Branko Marovic and Zoran Jovanovic. Visualization of 3D fields and medical data and using VRML. Future Generation Computer Systems, 14(1– 2):33–49, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/29/17/20/abstract.
                                        html.

Milde:2000:EUV

[MJ00] Jan-Torsten Milde and Bernhard Jung. Educational use of VRML and Java in agent-based AI and computer graphics. Future Generation Computer Systems, 17(1):79– 87, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/24/34/abstract.
html.

Marovic:2006:WBG

[MJ06] Branko Marović and Zoran Jovanović. Web-based gridenabled interaction with 3D medical data. Future Generation Computer Systems, 22 (4):385–392, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Murai:1988:CSJ

[MK88] Jun Murai and Akira Kato.

Current status of JUNET. Future Generation Computer Systems, 4(3):205–215, October 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Murthy:1995:PPP

[MK95] V. K. Murthy and E. V.

Krishnamurthy. Probabilistic parallel programming based on multiset transformation. Future Generation Computer Systems, 11(3):283–293, June 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mun:2004:LMS

[MK04] Youngsong Mun and Youngyuk Kim. A location management scheme to provide IP mobility over wireless ATM. Future Generation Computer Systems, 20(2):205–219, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mathis:2006:PMN

[MKH06] Mark M. Mathis, Darren J. Kerbyson, and Adolfy Hoisie. A performance model of nondeterministic particle transport on large-scale systems. Future Generation Computer Systems, 22(3): 324–335, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Munthe-Kaas:2003:EPL

[MKK03] H. Munthe-Kaas and S. Krogstad. On enumeration problems in Lie–Butcher theory. Future Generation Computer Systems, 19(7):1197–1205, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Menychtas:2009:RTR

[MKT09] Andreas Menychtas, Dimosthenis Kyriazis, and Konstantinos Tserpes. Real-time reconfiguration for guaranteeing QoS provisioning levels in Grid environments. Future Generation Computer Systems, 25(7):779–784, July 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

McGough:2008:SBA

[MLD08] A. Stephen McGough, William Lee, and Shikta Das. A standards based approach to enabling legacy applications on the Grid. Future Generation Computer Systems, 24 (7):731–743, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Modersitzki:2001:ERB

[MLSO01] J. Modersitzki, G. Lustig, O. Schmitt, and W. Obelöer. Elastic registration of brain images on large PC-clusters. Future Generation Computer Systems, 18(1):115– 125, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/38/abstract.
html.

Munz:2000:PAP

[MLZ^{+}00] F. Munz, T. Ludwig, S. Ziegler, P. Bartenstein, M. Schwaiger, and A. Bode. Performance assessment of parallel spectral analysis: towards a practical performance model for parallel medical applications. Future Generation Computer Systems, 16(5): 553–562, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/28/38/abstract.
                                        html.

monga-Made:1995:IPB

[mM95] Magolu monga Made. Implementation of parallel block preconditionings on a transputerbased multiprocessor. Future Generation Computer Systems, 11(2):167–173, March 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Michailidis:2003:PEL

[MM03] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Performance evaluation of load balancing strategies for approximate string matching application on an MPI cluster of heterogeneous workstations. Future Generation Computer Systems, 19 (7):1075–1104, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Meo:2008:QCM

[MM08] Michela Meo and Fabio Milan. QoS content management for P2P file-sharing applications. Future Generation Computer Systems, 24(3): 213–221, March 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Martin:2005:DTT

[MMFM^{+}05] O. Martin, J. P. Martin-Flatin, E. Martelli, P. Moroni, H. Newman, S. Ravot, and D. Nae. The DataTAG transatlantic testbed. Future Generation Computer Systems, 21(4):443–456, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mostefaoui:2002:IOA

[MMR02] Achour Mostefaoui, Eric Mourgaya, and Michel Raynal. An introduction to oracles for asynchronous distributed systems. Future Generation Computer Systems, 18 (6):757–767, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mancini:2008:GAM

[MMVV08] Emilio P. Mancini, Sonya Marcarelli, Igor Vasilyev, and Umberto Villano. A gridaware MIP solver: Implementation and case studies. Future Generation Computer Systems, 24(2):133–141, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Min:2006:SSS

[MOK06] Geyong Min and Mohamed Ould-Khaoua. Special section: Systems performance analysis and evaluation. Future Generation Computer Systems, 22(7):742–744, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Moore:1999:ET

[Moo99] Mary Moore. The evolution of telemedicine. Future Generation Computer Systems, 15(2):245–254, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X98000673.

Morale:2001:MSA

[Mor01] D. Morale. Modeling and simulating animal grouping

— Individual-based models. Future Generation Computer Systems, 17(7): 883–891, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/34/31/41/show/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/45/34/31/abstract.
html.

Mauri:2002:PAP

[MP02] Giancarlo Mauri and Giulio Pavesi. A parallel algorithm for pattern discovery in biological sequences. Future Generation Computer Systems, 18 (6):849–854, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mailund:2007:EGM

[MPB^{+}07] Thomas Mailund, Christian N. S. Pedersen, Jonas Bardino, Brian Vinter, and Henrik H. Karlsen. Experiences with GeneRecon on MiG. Future Generation Computer Systems, 23(4):580–586, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Meijer:1996:PMC

[MPG96] P. M. Meijer, S. Poedts, and J. P. Goedbloed. Parallel magnetohydrodynamics on the Cray T3D. Future Generation Computer Systems, 12 (4):307–323, December 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mainguet:2000:FRB

[MPH00] Jean-François Mainguet, Marc Pégulu, and John B. Harris. Fingerprint recognition based on silicon chips. Future Generation Computer Systems, 16 (4):403–415, February 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/35/abstract.
html.

Moralis:2009:KSA

[MPPM09] Athanasios Moralis, Vassiliki Pouli, Symeon Papavassiliou, and Vasilis Maglaris. A Kerberos security architecture for Web services based instrumentation grids. Future Generation Computer Systems, 25 (7):804–818, July 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

McLachlan:2003:LGF

[MPQ03] R. I. McLachlan, M. Perlmutter, and G. R. W. Quispel. Lie group foliations: dynamical systems and integrators. Future Generation Computer Systems, 19 (7):1207–1219, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Monrose:2000:KDB

[MR00] Fabian Monrose and Aviel D. Rubin. Keystroke dynamics as a biometric for authentication. Future Generation Computer Systems, 16(4): 351–359, February 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/30/abstract.
html.

Moore:2003:MSI

[MR03a] Brian E. Moore and Sebastian Reich. Multi-symplectic integration methods for Hamiltonian PDEs. Future Generation Computer Systems, 19 (3):395–402, April 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). See erratum [MR04b].

Mori:2003:CSH

[MR03b] Paolo Mori and Laura Ricci.

Computational science in high schools: defining curricula and environments. Future Generation Computer Systems, 19(8): 1349–1361, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Melnik:2004:CMM

[MR04a] R. V. N. Melnik and A. H.

Roberts. Computational models for multi-scale coupled dynamic problems. Future Generation Computer Systems, 20 (3):453–464, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Moore:2004:EMS

[MR04b] Brian E. Moore and Sebastian Reich. Erratum to: “Multisymplectic integration methods for Hamiltonian PDEs” [Future Gen. Comput. Sys. 19 (2003) 395–402]. Future Generation Computer Systems, 20 (4):699–700, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). See [MR03a].

Mazzeo:1992:UCL

[MRV92] A. Mazzeo, S. Russo, and G. Ventre. Using CSP languages to program parallel workstation systems. Future Generation Computer Systems, 8(1–3):149–163, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mostefaoui:2001:LIC

[MRV01] Achour Mostéfaoui, Michel Raynal, and Paulo Veríssimo. The logically instantaneous communication mode: a communication abstraction. Future Generation Computer Systems, 17(6):669– 678, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/26/abstract.
html.

Migliardi:2001:PIL

[MS01] Mauro Migliardi and Vaidy Sunderam. Plug-ins, layered services and behavioral objects — Application programming styles in the Harness metacomputing system. Future Generation Computer Systems, 17(6): 795–811, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/38/abstract.
html.

Martin:2003:IIG

[MS03] José-Luis Fernández-Villacañas Martín and Mark Shackleton. Investigation of the importance of the genotype– phenotype mapping in information retrieval. Future Generation Computer Systems, 19 (1):55–68, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Matijasevic:2003:DEM

[MSK03] Maja Matijasevic and Lea Skorin-Kapov. Design and evaluation of a multi-user virtual audio chat. Future Generation Computer Systems, 19 (2):229–239, February 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Maran:2007:MCI

[MSKT07] Uko Maran, Sulev Sild, Iiris Kahn, and Kalev Takkis. Mining of the chemical information in GRID environment. Future Generation Computer Systems, 23(1):76–83, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Marino:1993:NNA

[MSLP93] G. Marino, G. Succi, G. Levo, and R. Pavesio. NAUTA: a network administration utility for transputer architectures. Future Generation Computer Systems, 9(1):63–72, May 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Marsh:1998:VMV

[MSR98] A. Marsh, F. Simistira, and R. Robb. VR in medicine: Virtual colonoscopy. Future Generation Computer Systems, 14(3–4):253–264, July 31, 1998. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/18/29/abstract.
html.

Mishra:2002:ORS

[MSS02] S. Mishra, K. Sikdar, and M. Satpathy. Optimizing register spills for eager functional languages. Future Generation Computer Systems, 18(5): 699–708, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/37/abstract.
html.

Miller:2000:JWB

[MSX00] John A. Miller, Andrew F. Seila, and Xuewei Xiang. The JSIM Web-based simulation environment. Future Generation Computer Systems, 17(2): 119–133, October 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/27/27/abstract.
html.

Michopoulos:2005:DDE

[MTH^{+}05] J. Michopoulos, P. Tsompanopoulou, E. Houstis, C. Farhat, M. Lesoinne, J. Rice, and A. Joshi. On a data-driven environment for multiphysics applications. Future Generation Computer Systems, 21(6):953–968, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Matsuda:2000:NIA

[MTKS00] Motohiko Matsuda, Yoshio Tanaka, Kazuto Kubota, and Mitsuhisa Sato. Network interface active messages for low overhead communication on SMP PC clusters. Future Generation Computer Systems, 16 (5):493–502, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/32/abstract.
html.

Machida:2008:IDS

[MTNM08] Yuya Machida, Shin’ichiro Takizawa, Hidemoto Nakada, and Satoshi Matsuoka. Intelligent data staging with overlapped execution of Grid applications. Future Generation Computer Systems, 24 (5):425–433, May 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mastroianni:2005:SPM

[MTV05] Carlo Mastroianni, Domenico Talia, and Oreste Verta. A super-peer model for resource discovery services in largescale Grids. Future Generation Computer Systems, 21 (8):1235–1248, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mullender:1992:DMS

[Mul92]  Sape J. Mullender.  Dis-

tributed multimedia systems. Future Generation Computer Systems, 8(1–3):243–247, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mun:2004:MSS

[Mun04]  Youngsong Mun.  Model-

ing and simulation in supercomputing and telecommunications. Future Generation Computer Systems, 20(2): 179–180, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Muraki:1986:VTP

[Mur86] Kazunori Muraki. VENUS: Two-phase machine translation system. Future Generation Computer Systems, 2(2): 117–119, June 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Muraszkiewicz:1988:CAA

[Mur88] Mieczyslaw R. Muraszkiewicz.

Cellular array architecture for relational database implementation. Future Generation Computer Systems, 4(1): 31–38, August 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Murphy:1995:PHR

[Mur95] J. A. Murphy. A perspective of HPCN requirements in the European aerospace industry. Future Generation Computer Systems, 11(4–5): 409–418, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Moreno-Vozmediano:2009:HMR

[MV09] Rafael Moreno-Vozmediano.

A hybrid mechanism for resource/service discovery in adhoc Grids. Future Generation Computer Systems, 25 (7):717–727, July 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

McHugh:1989:CMM

[MVAS89] P. E. McHugh, A. G. Varias, R. J. Asaro, and C. F. Shih. Computational modeling of microstructures. Future Generation Computer Systems, 5(2–3):295–318, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

MagolumongaMade:2001:GDD

[MvdV01] M. Magolu monga Made and H. A. van der Vorst. A generalized domain decomposition paradigm for parallel incomplete LU factorization preconditionings. Future Generation Computer Systems, 17 (8):925–932, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/28/abstract.
html.

Mulder:1998:CSC

[MvLvW98] Jurriaan D. Mulder, Robert van Liere, and Jarke J. van Wijk. Computational steering in the CAVE. Future Generation Computer Systems, 14(3–4):199–207, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/29/18/23/28/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/29/18/23/abstract.
html.

Mancini:2008:SBO

[MVRM08] E. P. Mancini, U. Villano, M. Rak, and F. Moscato. Simulation-based optimization of multiple-task GRID applications. Future Generation Computer Systems, 24 (6):594–604, June 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Monmarche:2000:HPA

[MVS00] N. Monmarché, G. Venturini, and M. Slimane. On how Pachycondyla apicalis ants suggest a new search algorithm. Future Generation Computer Systems, 16(8): 937–946, June 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/31/31/abstract.
html.

Meliones:1999:MED

[MVT^{+}99] A. Meliones, T. Varvarigou, P. Tsagronis, A. Emmen, and I. Barosan. A metacomputing environment for demanding applications: design, implementation, experiments and business benefit. Future Generation Computer Systems, 15(5–6):787– 806, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/37/abstract.
html.

Mulder:1999:SCS

[MvWvL99] Jurriaan D. Mulder, Jarke J.

van Wijk, and Robert van Liere. A survey of computational steering environments. Future Generation Computer Systems, 15(1): 119–129, February 12, 1999.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/19/19/30/18/26/29/
                                        index.htt;   http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/18/26/abstract.
                                        html.

Mambretti:2003:PTE

[MWC^{+}03] Joe Mambretti, Jeremy Weinberger, Jim Chen, Elizabeth Bacon, Fei Yeh, David Lillethun, Bob Grossman, Yunhong Gu, and Marco Mazzucco. The Photonic TeraStream: enabling next generation applications through intelligent optical networking at iGRID2002. Future Generation Computer Systems, 19(6): 897–908, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mateos:2008:JAE

[MZC08] Cristian Mateos, Alejandro Zunino, and Marcelo Campo. JGRIM: An approach for easy gridification of applications. Future Generation Computer Systems, 24(2):99–118, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nagao:1986:CRI

[Nag86a] M. Nagao. Cooperative R&D of information technologies between the government and private sector in Japan. Future Generation Computer Systems, 2(1):39–44, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nagao:1986:CSF

[Nag86b] Makota Nagao. Current status and future trends in machine translation. Future Generation Computer Systems, 2 (2):77–82, June 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Narjes:1986:PEC

[Nar86]  Karl-Heinz Narjes.  Per-

spectives for European cooperation. Future Generation Computer Systems, 2(1): 13–17, March 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Naidoo:2004:AKL

[NB04] R. Naidoo and S. Baboolal.

Application of the Kurganov– Levy semi-discrete numerical scheme to hyperbolic problems with nonlinear source terms. Future Generation Computer Systems, 20(3): 465–473, April 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Neary:1999:JPC

[NCCS99] Michael O. Neary, Bernd O.

Christiansen, Peter Cappello, and Klaus E. Schauser.

Javelin: Parallel computing on the Internet. Future Generation Computer Systems, 15(5–6):659–674, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/27/abstract.
html.

Nguyen:2004:HPF

[NCS04] Hung Phi Nguyen, Bastien Chopard, and Serge Stoll. Hydrodynamic properties of fractal aggregates in 2D using Lattice Boltzmann simulation. Future Generation Computer Systems, 20(6):981–991, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nedoma:2006:SCP

[Ned06] Jiří Nedoma. On a solvability of contact problems with visco-plastic friction in the thermo-visco-plastic Bingham rheology. Future Generation Computer Systems, 22(4): 484–499, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nieuwstadt:1994:DLE

[NEJP94] F. T. M. Nieuwstadt, J. G. M. Eggels, R. J. A. Janssen, and M. B. J. M. Pourquié. Direct and large-eddy simulations of turbulence in fluids. Future Generation Computer Systems, 10(2–3):189–205, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nemeth:2000:AMD

[Ném00] Zsolt Németh. Abstract machine design on a multithreaded architecture. Future Generation Computer Systems, 16(6):705–716, April 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.elsevier.com/gej-ng/
10/19/19/41/29/36/abstract.
html.

Naiksatam:2007:ERE

[NF07] Sumit Naiksatam and Silvia Figueira. Elastic reservations for efficient bandwidth utilization in LambdaGrids. Future Generation Computer Systems, 23(1):1–22, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Natrajan:2002:LSA

[NHG02] Anand Natrajan, Marty A.

Humphrey, and Andrew S. Grimshaw. The Legion support for advanced parameterspace studies on a grid. Future Generation Computer Systems, 18(8):1033–1052, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). See erratum [NHG03].

Natrajan:2003:ELS

[NHG03] Anand Natrajan, Marty A.

Humphrey, and Andrew S.

Grimshaw. Erratum to: “The Legion support for advanced parameter-space studies on a grid” [Future Generation Computer Systems 18 (2002) 1033–1052]. Future Generation Computer Systems, 19 (7):1121–1122, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). See [NHG02].

Neubauer:2006:WBG

[NHG06] Falk Neubauer, Andreas Hoheisel, and Joachim Geiler. Workflow-based Grid applications. Future Generation Computer Systems, 22(1–2): 6–15, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Norden:2006:OVM

[NHT06] Markus Nordén, Sverker Holmgren, and Michael Thuné. OpenMP versus MPI for PDE solvers based on regular sparse numerical operators. Future Generation Computer Systems, 22(1–2):194–203, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nishio:1993:EFD

[Nis93] Shojiro Nishio. An evaluation of the FGCS data & knowledge base system — expectations and achievements. Future Generation Computer Systems, 9(2):153–158, July 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nitta:1986:PMT

[Nit86] Yoshihiko Nitta. Problems of machine translation systems: Effect of cultural differences on sentence structure. Future Generation Computer Systems, 2(2):101–115, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Niwa:1989:UAI

[Niw89] Katsuhiko Niwa. Use of artificial intelligence confirming compliance to building codes during process of making architectural drawings. Future Generation Computer Systems, 5(1):129–130, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ng:2006:BGE

[NJW^{+}06] Muan Hong Ng, Steven Johnston, Bing Wu, Stuart E. Murdock, Kaihsu Tai, Hans Fangohr, Simon J. Cox, Jonathan W. Essex, Mark S. P. Sansom, and Paul Jeffreys. BioSimGrid: Gridenabled biomolecular simulation data storage and analysis. Future Generation Computer Systems, 22(6):657–664, May 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Neruda:2005:LMR

[NK05] Roman Neruda and Petra Kudová. Learning methods for radial basis function networks. Future Generation Computer Systems, 21(7): 1131–1142, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nguyen:2007:PGP

[NK07] Tuan-Anh Nguyen and Pierre Kuonen. Programming the Grid with POP-C++. Future Generation Computer Systems, 23(1):23–30, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Na:2009:MDC

[NKX09] Yun Ji Na, Il Seok Ko, and Shiying Xu. A multilayered digital content distribution using a group-key based on Web. Future Generation Computer Systems, 25(3): 371–377, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Naka:2000:WWB

[NMA00] Toshiya Naka, Yoshiyuki Mochizuki, and Shigeo Asahara. WonderSpace: Web based humanoid animation. Future Generation Computer Systems, 17(1):57– 64, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/24/31/abstract.
html.

North:2005:OBA

[NMC05] M. North, C. Macal, and P. Campbell. Oh behave! Agent-based behavioral representations in problem solving environments. Future Generation Computer Systems, 21(7): 1192–1198, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nan:2006:TPD

[NMZC06] Kai Nan, Yongzheng Ma, Hongmei Zhang, and Gang Chen. Transfer, processing and distribution of cosmic ray data from Tibet. Future Generation Computer Systems, 22 (8):852–855, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nossal:1998:EAM

[Nos98] Roman Nossal. An evolutionary approach to multiprocessor scheduling of dependent tasks. Future Generation Computer Systems, 14(5–6): 383–392, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nikolopoulos:2003:ASU

[NP03] Dimitrios S. Nikolopoulos and Constantine D. Polychronopoulos. Adaptive scheduling under memory constraints on non-dedicated computational farms. Future Generation Computer Systems, 19(4):505–519, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nam:2006:DEE

[NP06] Young Jin Nam and Chanik Park. Design and evaluation of an efficient proportionalshare disk scheduling algorithm. Future Generation Computer Systems, 22(5): 601–610, April 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

NSF:1987:FRP

[NSF87]  NSF.  FY 1986 research

projects sponsored by the National Science Foundation. Future Generation Computer Systems, 3(2):117–135, May 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nakashima:1988:TCI

[NSHP88] Hideyuki Nakashima, Hiroyuki Suzuki, Per-Kristian Halvorsen, and P. Stanley Peters. Towards a computational interpretation of situation theory. Future Generation Computer Systems, 3(??):??, ???? 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nakamura:1984:ALU

[NSI84] Kiyohiko Nakamura, Andrew P. Sage, and Sosuke Iwai. Associative learning using similarity knowledge bases for relational database search. Future Generation Computer Systems, 1(2):123– 133, November 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nolte:2002:ECN

[NSI02] Jörg Nolte, Mitsuhisa Sato, and Yutaka Ishikawa. Exploiting cluster networks for distributed object groups and collective operations. Future Generation Computer Systems, 18(4):461– 476, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/33/30/abstract.
html.

Natarajan:2007:GBA

[NSP07] Ramesh Natarajan, Radu Sion, and Thomas Phan. A Grid-based approach for enterprise-scale data mining. Future Generation Computer Systems, 23(1):48–54, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nakada:1999:DIN

[NSS99] Hidemoto Nakada, Mitsuhisa Sato, and Satoshi Sekiguchi.

Design and implementations of Ninf: towards a global computing infrastructure. Future Generation Computer Systems, 15(5–6):649–658, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/26/abstract.
html.

Nagao:1986:STA

[NTN86] Makoto Nagao, Junichi Tsujii, and Junichi Nakamura. Science and technology agency’s Mu machine translation project. Future Generation Computer Systems, 2(2):125–139, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Naji:2004:CAE

[NWE04] Hamid Reza Naji, B. Earl Wells, and Letha Etzkorn. Creating an adaptive embedded system by applying multiagent techniques to reconfigurable hardware. Future Generation Computer Systems, 20 (6):1055–1081, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nijim:2007:SQS

[NZQ07] Mais Nijim, Ziliang Zong, and Xiao Qin. StReD: a quality of security framework for storage resources in Data Grids. Future Generation Computer Systems, 23(6):816–824, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Okun:2004:AWD

[OB04] Michael Okun and Amnon Barak. Atomic Writes for data integrity and consistency in shared storage devices for clusters. Future Generation Computer Systems, 20(4): 539–547, May 3, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

ODonnell:1988:VIA

[OBK88] John T. O’Donnell, Timothy Bridges, and Sidney W. Kitchel. A VLSI implementation of an architecture for applicative programming. Future Generation Computer Systems, 4(3):245–254, October 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Oliveira:2007:TAC

[OCdAM07] Marcelo Costa Oliveira, Walfredo Cirne, and Paulo M. de Azevedo Marques. Towards applying content-based image retrieval in the clinical routine. Future Generation Computer Systems, 23(3): 466–474, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Oh:2007:OWS

[OF07] Sangyoon Oh and Geoffrey C. Fox. Optimizing Web service messaging performance in mobile computing. Future Generation Computer Systems, 23 (4):623–632, May 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ochsenbein:1999:VSU

[OFO^{+}99] F. Ochsenbein, P. Fernique, P. Ortiz, D. Egret, and F. Genova. The VizieR system, a unified interface to astronomical catalogs. Future Generation Computer Systems, 16(1):39– 48, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/32/30/abstract.
                                        html.

Olejnik:2009:WOD

[OFT09] Richard Olejnik, Teodor-Florin Fortiş, and Bernard Toursel. Webservices oriented data mining in knowledge architecture. Future Generation Computer Systems, 25 (4):436–443, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ohyama:1989:KBD

[Ohy89] Minoru Ohyama. A knowledgebased directory assistance system. Future Generation Computer Systems, 5(1):109– 117, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Oldfield:2002:APF

[OK02] Ron Oldfield and David Kotz.

Armada: a parallel I/O framework for computational grids. Future Generation Computer Systems, 18(4): 501–523, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/33/33/abstract.
html.

Okuno:1992:EPA

[Oku92] Hiroshi G. Okuno. Experience of parallel AI programming with parallel Lisp. Future Generation Computer Systems, 7(2–3):211–219, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Olejnik:2007:SSG

[Ole07] Richard Olejnik. Special section: Grid-like distributed computing in amorphous networks. Future Generation Computer Systems, 23(8): 956, November 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Omondi:1991:FLP

[Omo91] Amos R. Omondi. On function languages and parallel computers. Future Generation Computer Systems, 6(4):355– 372, September 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ohno:1985:E

[ON85] Yutaka Ohno and Jurg Nievergelt. Editorial. Future Generation Computer Systems, 1 (5):255, September 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Oki:1989:SOS

[ONHT89] Michio Oki, Toshiro Nishimori, Minoru Hiyoshi, and Yoshiyuki Takaoka. Substation operation support system with event-driven processing. Future Generation Computer Systems, 5(1):41–49, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Omondi:1995:MPE

[OP95] Amos R. Omondi and David A. Plaisted. A model for the parallel execution of subsetequational languages. Future Generation Computer Systems, 11(3):295–320, June 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Onbasioglu:1997:CWB

[OP97] E. Onbasioglu and Y. Paker. A comparative workloadbased methodology for performance evaluation of parallel computers. Future Generation Computer Systems, 12(6): 521–545, June 15, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/18/19/abstract.
html.

Oppliger:2000:PPA

[Opp00] Rolf Oppliger. Privacy protection and anonymity services for the World Wide Web (WWW). Future Generation Computer Systems, 16(4): 379–391, February 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/33/abstract.
html.

Ohlendorf:1992:NNE

[OS92] G. Ohlendorf and N. Schielow.

NEPTUNE: a new expert planning tool for users in a network environment. Future Generation Computer Systems, 7(4):391–401, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Overeinder:2001:EWA

[OS01] B. J. Overeinder and H. J.

Sips. Editorial: Widearea distributed applications in high performance computing. Future Generation Computer Systems, 17(6):

767–768, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/33/35/abstract.
                                        html.

Ouillon:2006:LRS

[OS06] G. Ouillon and D. Sornette. Long-range static directional stress transfer in a cracked, nonlinear elastic crust. Future Generation Computer Systems, 22(4):500–520, March 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Oyang:1993:MHM

[OSCY93] Yen-Jen Oyang, David Jinsung Sheu, Chih-Yuan Cheng, and Cheng-Zen Yang. The M^{2} hierarchical multiprocessor. Future Generation Computer Systems, 9(3):235– 240, September 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Overeinder:1996:DLB

[OSHH96] B. J. Overeinder, P. M. A. Sloot, R. N. Heederik, and L. O. Hertzberger. A dynamic load balancing system for parallel cluster computing. Future Generation Computer Systems, 12(1):101–115, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ostholm:1992:EMD

[Öst92]  Stig Östholm.  Electronic

mail in a distributed heterogeneous system. Future Generation Computer Systems, 8 (1–3):261–262, July 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ochi:1998:PEA

[OVDV98] Luiz S. Ochi, Dalessandro S.

Vianna, Lúcia M. A. Drummond, and AndréO. Victor. A parallel evolutionary algorithm for the vehicle routing problem with heterogeneous fleet. Future Generation Computer Systems, 14(5–6):285– 292, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Olmedo:2009:NMS

[OVK^{+}09] Vicente Olmedo, Víctor A.

Villagrá, Kleopatra Konstanteli, Juan E. Burgos, and Julio Berrocal. Network mobility support for Web Service-based Grids through the Session Initiation Protocol. Future Generation Computer Systems, 25 (7):758–767, July 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Papadopoulos:2001:CDR

[PA01a] George A. Papadopoulos and Farhad Arbab. Configuration and dynamic reconfiguration of components using the coordination paradigm. Future Generation Computer Systems, 17(8):1023– 1038, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/38/abstract.
html.

Priol:2001:CSA

[PA01b] Thierry Priol and Guillaume Alléon. A client/server approach for HPC applications within a networking environment. Future Generation Computer Systems, 17 (6):813–822, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/39/abstract.
html.

Padua:1992:PSE

[Pad92] David A. Padua. Problemsolving environments for parallel computers. Future Generation Computer Systems, 7(2–3):221–229, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pape:2003:YFC

[PADD03] Dave Pape, Josephine Anstey, Margaret Dolinsky, and Edward J. Dambik. Ygdrasil — a framework for composing shared virtual worlds. Future Generation Computer Systems, 19(6):1041–1049, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pagendarm:1999:VWE

[Pag99] H.-G. Pagendarm. Visualization within environments supporting human communication. Future Generation Computer Systems, 15(1): 109–117, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/18/25/29/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/18/25/abstract.
html.

Palsson:2001:TDM

[Pal01]  E. Palsson.  A three-

dimensional model of cell movement in multicellular systems. Future Generation Computer Systems, 17 (7):835–852, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/34/27/43/show/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/45/34/27/abstract.
html.

Palmieri:2006:GBS

[Pal06] Francesco Palmieri. GMPLSbased service differentiation for scalable QoS support in all-optical Grid applications. Future Generation Computer Systems, 22(6):688–698, May 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Palmieri:2009:NAS

[Pal09] Francesco Palmieri. Networkaware scheduling for realtime execution support in data-intensive optical Grids. Future Generation Computer Systems, 25(7):794–803, July 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pan:1995:OSL

[Pan95a] Yi Pan. Order statistics on a linear array with a reconfigurable bus. Future Generation Computer Systems, 11 (3):321–327, June 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Panda:1995:FBS

[Pan95b] Dhabaleswar K. Panda. Fast barrier synchronization in wormhole k-ary n-cube networks with multidestination worms. Future Generation Computer Systems, 11(6): 585–602, October 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Papadopoulos:2005:MIA

[Pap05] George A. Papadopoulos. Modelling and implementing asynchronous timed multimedia frameworks using coordination principles. Future Generation Computer Systems, 21 (5):687–698, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Partridge:1987:SLF

[Par87] Derek Partridge. The scope and limitations of first generation expert systems. Future Generation Computer Systems, 3(1):1–10, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Parrow:1990:EPP

[Par90] Joachim Parrow. The expressive power of parallelism. Future Generation Computer Systems, 6(3):271– 285, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Paris:1991:P

[Par91] E. Paris. Preface. Future Generation Computer Systems, 7 (1):1, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Parsons:1994:HAF

[Par94] D. J. Parsons. HPCN applications in finance. Future Generation Computer Systems, 10(2–3):169–172, June 1994.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Park:2004:PEL

[Par04] Gyung-Leen Park. Performance evaluation of a list scheduling algorithm in distributed memory multiprocessor systems. Future Generation Computer Systems, 20(2): 249–256, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Parker:2006:CBA

[Par06]  Steven G. Parker.  A

component-based architecture for parallel multi-physics PDE simulation. Future Generation Computer Systems, 22(1–2): 204–216, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pham:2005:VFS

[PB05] Binh Pham and Ross Brown. Visualisation of fuzzy systems: requirements, techniques and framework. Future Generation Computer Systems, 21(7): 1199–1212, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pullen:2005:UWS

[PBB^{+}05] J. Mark Pullen, Ryan Brunton, Don Brutzman, David Drake, Michael Hieb, Katherine L. Morse, and Andreas Tolk. Using Web services to integrate heterogeneous simulations in a grid environment. Future Generation Computer Systems, 21(1):97– 106, January 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pickles:2001:MAI

[PBC^{+}01] S. M. Pickles, J. M. Brooke, F. C. Costen, Edgar Gabriel, Matthias Müller, Michael Resch, and S. M. Ord. Metacomputing across intercontinental networks. Future Generation Computer Systems, 17 (8):911–918, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/26/abstract.
html.

Plaat:2001:SPA

[PBHK01] Aske Plaat, Henri E. Bal, Rutger F. H. Hofman, and Thilo Kielmann. Sensitivity of parallel applications to large differences in bandwidth and latency in two-layer interconnects. Future Generation Computer Systems, 17 (6):769–782, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/36/abstract.
html.

Pfenning:1995:VSM

[PBM95] Jörg-Thomas Pfenning, Achim Bachem, and Ronald Minnich. Virtual shared memory programming on workstation clusters. Future Generation Computer Systems, 11(4–5): 387–399, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Penenko:2002:MST

[PBT02] Vladimir Penenko, Alexander Baklanov, and Elena Tsvetova. Methods of sensitivity theory and inverse modeling for estimation of source parameters. Future Generation Computer Systems, 18(5): 661–671, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/36/33/abstract.
                                        html.

Popescu:1999:MDF

[PCB99] Mihai Popescu, Paul Cristea, and Anastasios Bezerianos. Multiresolutional distributed filtering: a novel technique that reduces the amount of data required in high resolution electrocardiography. Future Generation Computer Systems, 15(2):195– 209, March 11, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

19/19/30/17/22/abstract.
html.

Plank:1999:DFT

[PCBD99] James S. Plank, Henri Casanova, Micah Beck, and Jack J. Dongarra. Deploying faulttolerance and task migration with NetSolve. Future Generation Computer Systems, 15(5–6):745–755, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/34/abstract.
html;   http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/netsolve-ft-tm.pdf.

Perez:2006:MFM

[PCG^{+}06] María S. Pérez, Jesús Carretero, Félix García, José M. Peña, and Víctor Robles. MAPFS: a flexible multiagent parallel file system for clusters. Future Generation Computer Systems, 22(5):620–632, April 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Plante:1999:NAD

[PCM99] Raymond L. Plante, Richard M.

Crutcher, and Robert E. Mc-Grath. The NCSA astronomy digital image library: from data archiving to data publishing. Future Generation Computer Systems, 16(1):49– 61, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/31/abstract.
html.

Post:1999:GGF

[PdLS^{+}99] Frits H. Post, Wim C. de Leeuw, I. Ari Sadarjoen, Freek Reinders, and Theo van Walsum. Global, geometric, and feature-based techniques for vector field visualization. Future Generation Computer Systems, 15 (1):87–98, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/18/23/29/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/18/23/abstract.
html.

Primet:2005:EED

[PEG05] Pascale Vicat-Blanc Primet, Franois Echantillac, and Mathieu Goutelle. Experiments with equivalent differentiated services in a grid context. Future Generation Computer Systems, 21(4):515–524, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Perry:1986:NSS

[Per86] Dennis G. Perry. Network support of supercomputers. Future Generation Computer Systems, 2(1):65–67, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Peterson:1989:CCA

[Pet89] Victor L. Peterson. Computational challenges in aerospace. Future Generation Computer Systems, 5(2–3):243– 258, September 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Petkov:1995:BMC

[Pet95] Nikolay Petkov. Biologically motivated computationally intensive approaches to image pattern recognition. Future Generation Computer Systems, 11(4–5):451–465, August 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Perumalla:2001:BWI

[PF01] Kalyan S. Perumalla and Richard M. Fujimoto. Best of Websim99: Interactive parallel simulations with the Jane framework. Future Generation Computer Systems, 17(5): 525–537, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/28/abstract.
html.

Pfister:1999:ART

[Pfi99] Hanspeter Pfister. Architectures for real-time volume rendering. Future Generation Computer Systems, 15 (1):1–9, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/18/17/29/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/18/17/abstract.
html.

Perez:2004:LCS

[PFMC04] M. Perez, M. Fernandez, P. Morillo, and I. Coma. Locally constrained synthetic LoDs generation for natural terrain meshes. Future Generation Computer Systems, 20 (8):1375–1387, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Piro:2009:UHA

[PGPW09] Rosario M. Piro, Andrea Guarise, Giuseppe Patania, and Albert Werbrouck. Using historical accounting information to predict the resource usage of Grid jobs. Future Generation Computer Systems, 25 (5):499–510, May 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pairot:2005:TNL

[PGSM05] Carles Pairot, Pedro García, Antonio F. Gómez Skarmeta, and Rubén Mondéjar. Towards new load-balancing schemes for structured peerto-peer grids. Future Generation Computer Systems, 21 (1):125–133, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Potma:1994:BPL

[PH94] Kitty Potma and Walter Hoffmann. Boosting the performance of the linear algebra part in an ODE solver for shared memory systems. Future Generation Computer Systems, 10(2–3):315–319, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Peddemors:1999:HPD

[PH99] A. J. H. Peddemors and L. O.

Hertzberger. A high performance distributed database system for enhanced Internet services. Future Generation Computer Systems, 15(3): 407–415, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/25/abstract.
html.

Perez:2007:SSD

[PH07] María S. Pérez and Pilar Herrero. Special section: Data analysis, access and management on Grids. Future Generation Computer Systems, 23 (1):107–108, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pan:1998:ESQ

[PHL98] Yi Pan, Mounir Hamdi, and Keqin Li. Efficient and scalable quicksort on a linear array with a reconfigurable pipelined bus system. Future Generation Computer Systems, 13(6):501– 513, May 20, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/20/22/abstract.
html.

Prins:1999:VES

[PHM^{+}99] Jan F. Prins, Jan Hermans, Geoffrey Mann, Lars S. Nyland, and Martin Simons. A virtual environment for steered molecular dynamics. Future Generation Computer Systems, 15(4):485–495, July 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/20/19/29/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/20/19/abstract.
html.

Peper:2002:STC

[PIKM02] F. Peper, T. Isokawa, N. Kouda, and N. Matsui. Self-Timed Cellular Automata and their computational ability. Future Generation Computer Systems, 18(7):893–904, August 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pinkston:1987:VIP

[Pin87] John T. Pinkston. From visionary ideas to products. Future Generation Computer Systems, 3(4):233– 243, December 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pitrat:1996:IRS

[Pit96] Jacques Pitrat. Implementation of a reflective system. Future Generation Computer Systems, 12(2–3):235– 242, September 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pavlopoulos:1999:IPM

[PK99] S. Pavlopoulos and D. Koutsouris. An image processing and management system for radiology with telemedicine services. Future Generation Computer Systems, 15 (2):293–299, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X98000740.

Polak:2008:IVV

[PK08] Martin Polak and Dieter Kranzlmüller. Interactive videostreaming visualization on Grids. Future Generation Computer Systems, 24(1):

39–45, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Park:2004:IPL

[PKC04] Seungmin Park, Daeyoung Kim, and Gihwan Cho. Improving prediction level of prefetching for locationaware mobile information service. Future Generation Computer Systems, 20(2):197–203, February 16, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Parashar:2005:AGE

[PKC^{+}05] Manish Parashar, Hector Klie, Umit Catalyurek, Tahsin Kurc, Wolfgang Bangerth, Vincent Matossian, Joel Saltz, and Mary F. Wheeler. Application of Grid-enabled technologies for solving optimization problems in data-driven reservoir studies. Future Generation Computer Systems, 21 (1):19–26, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Puzone:2002:IFM

[PKSC02] R. Puzone, B. Kohler, P. Seiden, and F. Celada. IMMSIM, a flexible model for in machina experiments on immune system responses. Future Generation Computer Systems, 18(7): 961–972, August 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pruyne:1996:ICP

[PL96] Jim Pruyne and Miron Livny.

Interfacing Condor and PVM to harness the cycles of workstation clusters. Future Generation Computer Systems, 12 (1):67–85, May 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Polemi:2000:IRS

[PM00] Despina Polemi and Andy Marsh. Internet race is on but security holds the key. Future Generation Computer Systems, 16(4):v, February 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/27/25/abstract.
html.

Pal:2004:AES

[PM04] Mahesh Pal and P. M.

Mather. Assessment of the effectiveness of support vector machines for hyperspectral data. Future Generation Computer Systems, 20(7): 1215–1225, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Prodan:2009:PBR

[PN09] Radu Prodan and Vlad Nae. Prediction-based realtime resource provisioning for massively multiplayer online games. Future Generation Computer Systems, 25 (7):785–793, July 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pancake:1999:SPS

[PNH99] Cherri M. Pancake, Mark Newsome, and F. Joe Hanus. ‘split personalities’ for scientific databases: targeting database middleware and interfaces to specific audiences. Future Generation Computer Systems, 16(1): 135–152, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/32/38/abstract.
html.

Page:2000:IAW

[PO00] Ernest H. Page and Jeffrey M. Opper. Investigating the application of Web-based simulation principles within the architecture for a next-generation computer generated forces model. Future Generation Computer Systems, 17(2): 159–169, October 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/27/30/abstract.
html.

Pohl:1987:SRN

[Poh87] Ira Pohl. Should robots have nuclear arms? AI technology and SDI software. Future Generation Computer Systems, 3 (2):111–115, May 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Polemi:1998:TTP

[Pol98] Despina Polemi. Trusted third party services for health care in Europe. Future Generation Computer Systems, 14(1–2): 51–59, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/21/abstract.
html.

Polemi:1999:TBS

[Pol99] Despina Polemi. TTPs and biometrics for securing the payment of telemedical services. Future Generation Computer Systems, 15(2): 265–276, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X98000697.

Porcher:1995:BPC

[Por95]  Thierry Porcher.  Bench-

marking the POMPC compiler on the Connection Machine CM-2. Future Generation Computer Systems, 11(1):

19–26, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Politi:2006:NMC

[PP06] T. Politi and A. Pugliese. Numerical methods for computing SVD in the D-orthogonal group. Future Generation Computer Systems, 22(4): 423–429, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Parsa:2007:CDM

[PP07] Saeed Parsa and Fereshteh-Azadi Parand. Cooperative decision making in a knowledge grid environment. Future Generation Computer Systems, 23(8):932–938, November 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Prentza:1999:IHC

[PPAK99] A. Prentza, S. Palamas, A. Anagnostaki, and D. Koutsouris. Intranet health clinic

— a Web-based interactive communication environment for the continuation in health care. Future Generation Computer Systems, 15(2): 277–285, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X98000727.

Plantikow:2009:GDM

[PPH^{+}09] Stefan Plantikow, Kathrin Peter, Mikael Högqvist, Christian Grimme, and Alexander Papaspyrou. Generalizing the data management of three community grids. Future Generation Computer Systems, 25 (3):281–289, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pouzet:1995:PAD

[PPJ95] P. Pouzet, J. Paris, and V. Jorrand. Parallel application design: The simulation approach with HASTE. Future Generation Computer Systems, 11(4–5):363– 373, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Power:2006:SWS

[PPSS06] D. J. Power, E. A. Politou, M. A. Slaymaker, and A. C. Simpson. Securing Web services for deployment in health grids. Future Generation Computer Systems, 22(5):547–570, April 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X05001202.

Peakall:1995:SIB

[PR95] A. Peakall and G. Robertson.

Strategies for the implementation of the BBS communication tests on the supernode and their implications for efficiency. Future Generation Computer Systems, 11(1): 35–47, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pringle:1995:ETM

[Pri95]  Gavin J. Pringle.  Em-

bedding a ‘treecode’ on a MIMD parallel computer using a domain decomposition paradigm. Future Generation Computer Systems, 11(2): 183–192, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Prodan:2007:SCS

[Pro07] Radu Prodan. Specificationcorrect and scalable coordination of Grid applications. Future Generation Computer Systems, 23(4):587–605, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Przybylska:2003:ILF

[Prz03] Maria Przybylska. Isospectrallike flows and eigenvalue problem. Future Generation Computer Systems, 19(7):1165– 1175, October 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Perez:2009:AAM

[PSA^{+}09] María S. Pérez, Alberto Sánchez, Jemal H. Abawajy, Víctor Robles, and José M. Peña. An agent architecture for managing data resources in a Grid environment. Future Generation Computer Systems, 25(7):747–755, July 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Peterka:2006:PVA

[PSG^{+}06] Tom Peterka, Daniel J.

Sandin, Jinghua Ge, Javier Girado, Robert Kooima, Jason Leigh, Andrew Johnson, Marcus Thiebaux, and Thomas A. DeFanti. Personal Varrier: Autostereoscopic virtual reality display for distributed scientific visualization. Future Generation Computer Systems, 22(8): 976–983, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Park:2004:HSC

[PSL^{+}04] Sang-Joon Park, Ji-Young Song, Jongchan Lee, Kwan-Joong Kim, and Byung-Gi Kim. A handover scheme in clustered cellular networks. Future Generation Computer Systems, 20(2): 221–227, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ploski:2009:GBD

[PSP^{+}09] Jan Ploski, Guido Scherp, Thomas I. Petroliagis, Otto Büchner, and Wilhelm Hasselbring. Grid-based deployment and performance measurement of the Weather Research & Forecasting model. Future Generation Computer Systems, 25(3):346–350, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Perez:2007:DID

[PSR^{+}07] María S. Pérez, Alberto Sánchez, Víctor Robles, Pilar Herrero, and José M. Peña. Design and implementation of a data mining grid-aware architecture. Future Generation Computer Systems, 23(1):42– 47, January 1, 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Praehofer:2001:BWC

[PSS01] Herbert Praehofer, Johannes Sametinger, and Alois Stritzinger. Best of Websim99: Concepts and architecture of a simulation framework based on the JavaBeans component model. Future Generation Computer Systems, 17(5): 539–559, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/29/abstract.
html.

Pascoe:2002:MEM

[PSVL02] J. S. Pascoe, V. S. Sunderam, U. Varshney, and R. J. Loader. Middleware enhancements for metropolitan area wireless Internet access. Future Generation Computer Systems, 18(5):721– 735, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/39/abstract.
html.

Pham:2005:GIP

[PT05] Congduc Pham and Bernard Tourancheau. Grid infrastructures: practice and perspectives. Future Generation Computer Systems, 21(2): 247–248, February 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pudner:1987:DPA

[Pud87] A. Pudner. DLM — a powerful ai computer for embedded expert systems. Future Generation Computer Systems, 3 (4):299–306, December 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pudov:2001:LCN

[Pud01]  Sergey Pudov.  Learn-

ing of cellular neural networks. Future Generation Computer Systems, 17(6):

689–697, April 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/33/28/abstract.
html.

Popov:2005:EIM

[PVBH05] K. Popov, V. Vlassov, P. Brand, and S. Haridi. An efficient incremental marshaling framework for distributed systems. Future Generation Computer Systems, 21(5):717–724, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pavani:2009:CSL

[PW09] Gustavo Sousa Pavani and Helio Waldman. Co-scheduling in Lambda Grid systems by means of Ant Colony Optimization. Future Generation Computer Systems, 25(3): 257–265, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Park:2003:ERS

[PWY03] Taesoon Park, Namyoon Woo, and Heon Y. Yeom. An efficient recovery scheme for fault-tolerant mobile computing systems. Future Generation Computer Systems, 19 (1):37–53, January 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Perez:2007:SSS

[PX07] María S. Pérez and Bin Xiao.

Special section: Security on grids and distributed systems. Future Generation Computer Systems, 23(6):774–775, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Park:2000:ELR

[PY00] Taesoon Park and Heon Y.

Yeom. An efficient logging and recovery scheme for lazy release consistent distributed shared memory systems. Future Generation Computer Systems, 17(3): 265–278, November 1, 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/28/abstract.
html.

Qin:2007:DAL

[Qin07] Xiao Qin. Design and analysis of a load balancing strategy in Data Grids. Future Generation Computer Systems, 23 (1):132–137, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Quiroz:2008:FDC

[QP08] Andres Quiroz and Manish Parashar. A framework for distributed content-based Web services notification in Grid systems. Future Generation Computer Systems, 24 (5):452–459, May 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Qu:2004:POL

[Qu04] Yuzhong Qu. A predicateordered logic for knowledge representation on the web. Future Generation Computer Systems, 20(1):19–26, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ramme:1995:BVM

[Ram95] Friedhelm Ramme. Building a virtual machine-room — a focal point in metacomputing. Future Generation Computer Systems, 11(4–5): 477–489, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Raynal:2005:WFC

[Ray05] Michel Raynal. Wait-free computing: an introductory lecture. Future Generation Computer Systems, 21(5):655–663, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reynolds:1988:BPL

[RBC^{+}88] T. J. Reynolds, A. J. Beaumont, A. S. K. Cheng, S. A. Delgado-Rannauro, and L. A. Spacek. BRAVE — a parallel logic language for artificial intelligence. Future Generation Computer Systems, 4 (1):69–75, August 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reale:1993:PDH

[RBS93] F. Reale, M. Barbera, and S. Sciortino. A parallel 2d hydrodynamic FORTRAN code for astrophysical applications on a Meiko computing surface. Future Generation Computer Systems, 9 (1):19–24, May 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ronsse:2003:DSM

[RCD03] Michiel Ronsse, Mark Christiaens, and Koen De Bosschere. Debugging shared memory parallel programs using record/replay. Future Generation Computer Systems, 19(5):679–687, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reed:2006:RCL

[RdLM06] Daniel A. Reed, Charng da Lu, and Celso L. Mendes. Reliability challenges in large systems. Future Generation Computer Systems, 22(3): 293–302, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rackl:2000:ASU

[RdSH^{+}00] Günther Rackl, Filippo de Stefani, François Héran, Antonello Pasquarelli, and Thomas Ludwig. Airport simulation using CORBA and DIS. Future Generation Computer Systems, 16(5): 465–472, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/28/29/abstract.
                                        html.

Refenes:1987:EII

[Ref87] Apostolos N. Refenes. Nexpression implementations for integrated symbolic and numeric processing. Future Generation Computer Systems, 3(3):161–187, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rehn:2006:DMC

[Reh06] Christian Rehn. Dynamic mapping of cooperating tasks to nodes in a distributed system. Future Generation Computer Systems, 22(1–2): 35–45, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ruiz:2004:LOE

[REM04] José-Vicente Pitarch Ruiz, Stefano Evangelisti, and Daniel Maynau. Local orbitals for excited states. Future Generation Computer Systems, 20 (5):821–828, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reussner:2003:ACP

[Reu03a] Ralf H. Reussner. Automatic component protocol adaptation with the CoConut/J tool suite. Future Generation Computer Systems, 19 (5):627–639, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reussner:2003:USD

[Reu03b] Ralf H. Reussner. Using SKaMPI for developing highperformance MPI programs with performance portability. Future Generation Computer Systems, 19(5):749–759, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rough:2004:DEC

[RG04] J. T. Rough and A. M. Goscinski. The development of an efficient checkpointing facility exploiting operating systems services of the GENE-SIS cluster operating system. Future Generation Computer Systems, 20(4):523–538, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rosswog:2001:TSC

[RGH^{+}01] S. Rosswog, C. Gawron, S. Hasselberg, R. Böning, and P. Wagner. Traffic simulation: Computational aspects in traffic simulation problems. Future Generation Computer Systems, 17(5): 659–665, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/30/38/abstract.
                                        html.

Ranjan:2008:CCI

[RHB08] Rajiv Ranjan, Aaron Harwood, and Rajkumar Buyya. A case for cooperative and incentive-based federation of distributed clusters. Future Generation Computer Systems, 24(4):280–295, April 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rhoades:1989:PCG

[Rho89] Clifford E. Rhoades, Jr. Proceedings of the conference on grand challenges to computational science: January 3–6, 1989 Molokai. Future Generation Computer Systems, 5(2– 3):167–168, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rivas:2000:PAR

[RICW00] R. Rivas, M. B. Ibáñez, Y. Cardinale, and P. Windyga.

A parallel algorithm for 3D reconstruction of angiographic images. Future Generation Computer Systems, 16(5): 533–539, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/36/abstract.
html.

Renambot:2009:EHR

[RJH^{+}09] Luc Renambot, Byungil Jeong, Hyejung Hur, Andrew Johnson, and Jason Leigh. Enabling high resolution collaborative visualization in display rich virtual organizations. Future Generation Computer Systems, 25(2): 161–168, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rizk:2008:PGO

[RKSU08] Phil Rizk, Cameron Kiddle, Rob Simmonds, and Brian Unger. Performance of a GridFTP overlay network. Future Generation Computer Systems, 24(5):442–451, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rantzau:1998:SVE

[RL98] Dirk Rantzau and Ulrich Lang. A scalable virtual environment for large scale scientific data analysis. Future Generation Computer Systems, 14(3–4):215–222, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/29/18/25/28/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/29/18/25/abstract.
html.

Rudolph:1997:TSI

[RM97] George L. Rudolph and Tony R. Martinez. A transformation strategy for implementing distributed, multilayer feedforward neural networks: Backpropagation transformation. Future Generation Computer Systems, 12 (6):547–564, June 15, 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/18/20/abstract.
html.

Rasmussen:1998:VTB

[RMM^{+}98] Mary Rasmussen, Theodore P. Mason, Alan Millman, Ray Evenhouse, and Daniel Sandin. The virtual temporal bone, a tele-immersive educational environment. Future Generation Computer Systems, 14(1–2): 125–130, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/27/abstract.
html.

Rickert:2001:TSD

[RN01] Marcus Rickert and Kai Nagel. Traffic simulation: Dynamic traffic assignment on parallel computers in TRAN-SIMS. Future Generation Computer Systems, 17(5): 637–648, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/36/abstract.
html.

Raney:2004:IRP

[RN04] Bryan Raney and Kai Nagel.

Iterative route planning for large-scale modular transportation simulations. Future Generation Computer Systems, 20(7):1101–1118, October 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Roshanaei:2009:VNS

[RNJK09] V. Roshanaei, B. Naderi, F. Jolai, and M. Khalili. A variable neighborhood search for job shop scheduling with set-up times to minimize makespan. Future Generation Computer Systems, 25 (6):654–661, June 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Robenack:2005:ADN

[Röb05] Klaus Röbenack. Automatic differentiation and nonlinear controller design by exact linearization. Future Generation Computer Systems, 21 (8):1372–1379, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Roskies:1989:SBS

[Ros89] Ralph Roskies. Supercomputing and biomedical science. Future Generation Computer Systems, 5(2–3):197– 205, September 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rossi:1994:SIH

[Ros94] Pietro Rossi. Science and industry in HPCN. Future Generation Computer Systems, 10(2–3):183–188, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rourk:2000:VBC

[Rou00] Will Rourk. Virtual biochemistry — a case study. Future Generation Computer Systems, 17(1):7–14, September 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/19/19/45/24/26/41/
                                        index.htt;   http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/24/26/abstract.
                                        html.

Renaut:2003:E

[RR03] Rosemary A. Renaut and Ulrich Ruede. Editorial. Future Generation Computer Systems, 19(8):1265, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Resch:1999:MET

[RRS99] Michael M. Resch, Dirk Rantzau, and Robert Stoy. Metacomputing experience in a transatlantic wide area application test-bed. Future Generation Computer Systems, 15(5–6):807–816, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/38/abstract.
html.

Riley:1994:SPS

[RS94] C. P. Riley and J. Simkin.

The serial performance standards for the BBS test program. Future Generation Computer Systems, 10(4): 419–428, November 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rogers:1998:UBC

[RS98] R. O. Rogers and D. B.

Skillicorn. Using the BSP cost model to optimise parallel neural network training. Future Generation Computer Systems, 14(5–6):409– 424, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rossi:1999:ISC

[RS99] Louis F. Rossi and George Sohos. Interactive simulation of contaminant evolution through porous media. Future Generation Computer Systems, 15(4):477–484, July 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/30/20/18/29/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/30/20/18/abstract.
html.

Ribler:2001:APD

[RSR01] Randy L. Ribler, Huseyin Simitci, and Daniel A. Reed. The Autopilot performancedirected adaptive control system. Future Generation Computer Systems, 18(1):175– 187, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/43/abstract.
html.

Ravikanth:1988:RAO

[RSRV88] K. Ravikanth, P. S. Sastry, K. R. Ramakrishnan, and Y. V. Venkatesh. A reduction architecture for the optimal scheduling of binary trees. Future Generation Computer Systems, 4(3):225–233, October 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reinefeld:2002:GTC

[RSSD02] Alexander Reinefeld, Hinnerk Stüben, Florian Schintke, and George Din. GuiGen: a toolset for creating customized interfaces for grid user communities. Future Generation Computer Systems, 18(8):1075–1084, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ravikanth:1990:SSP

[RSV90] K. Ravikanth, P. S. Sastry, and Y. V. Venkatesh. Simulation studies on the performance of an organizational model for graph reduction. Future Generation Computer Systems, 6(2):163– 180, November 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reuter:2006:LKE

[RT06] Jürgen Reuter and Walter F.

Tichy. Logging kernel events on clusters. Future Generation Computer Systems, 22(3): 313–323, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rumpf:1999:RNM

[Rum99]  M. Rumpf.  Recent nu-

merical methods — a challenge for efficient visualization. Future Generation Computer Systems, 15(1):43–58, February 12, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/18/20/abstract.
html.

Ruschitzka:1990:P

[Rus90a] Manfred Ruschitzka. Preface. Future Generation Computer Systems, 5(4):351–352, January 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ruschitzka:1990:TLG

[Rus90b] Manfred Ruschitzka. Twolevel grammars for data conversions. Future Generation Computer Systems, 5(4):373– 380, January 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Roest:1995:UGS

[RV95] Mark R. T. Roest and Edwin A. H. Vollebregt. Using the GCel for simulation of flow in the continental shelf region. Future Generation Computer Systems, 11(2): 135–144, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Renambot:2003:GER

[RvdSB^{+}03] Luc Renambot, Tom van der Schaaf, Henri E. Bal, Desmond Germans, and Hans J. W. Spoelder. Griz: experience with remote visualization over an optical grid. Future Generation Computer Systems, 19(6):871–882, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rokicki:2001:PPO

[RŻDM01] J. Rokicki, J. Żółtak, D. Drikakis, and J. Majewski. Parallel performance of overlapping mesh technique for compressible flows. Future Generation Computer Systems, 18(1): 3–15, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/28/abstract.
html.

Srikant:1997:MGA

[SA97] Ramakrishnan Srikant and Rakesh Agrawal. Mining generalized association rules. Future Generation Computer Systems, 13(2–3):161–180, November 14, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/21/abstract.
html.

Salter:2007:OTT

[SA07] James Salter and Nick Antonopoulos. An optimized two-tier P2P architecture for contextualized keyword searches. Future Generation Computer Systems, 23(2):241– 251, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sarbazi-Azad:2003:AAC

[SAKOK03] H. Sarbazi-Azad, A. Khonsari, and M. Ould-Khaoua. Analysis of k-ary n-cubes with dimension-ordered routing. Future Generation Computer Systems, 19(4):493–502, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Seidel:2002:GGA

[SAMN02] Ed Seidel, Gabrielle Allen, André Merzky, and Jarek Nabrzyski. GridLab — a grid application toolkit and testbed. Future Generation Computer Systems, 18 (8):1143–1153, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sapaty:1988:WNI

[Sap88] Peter S. Sapaty. WAVE-1: a new ideology of parallel and distributed processing on graphs and networks. Future Generation Computer Systems, 4(1):1–14, August 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sarmenta:2002:STM

[Sar02] Luis F. G. Sarmenta. Sabotagetolerance mechanisms for volunteer computing systems. Future Generation Computer Systems, 18(4): 561–572, March 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/33/37/abstract.
html.

Sasaki:1985:TVJ

[Sas85] Hajime Sasaki. Trends of VLSI in Japan. Future Generation Computer Systems, 1 (6):397–401, December 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schiemann:1997:NAL

[SB97] B. Schiemann and L. Borrmann. A new approach for load balancing in highperformance decision support systems. Future Generation Computer Systems, 12(5): 345–355, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/19/abstract.
html.

Szalay:1999:AAF

[SB99] A. S. Szalay and R. J. Brunner. Astronomical archives of the future: a Virtual Observatory. Future Generation Computer Systems, 16(1):63– 72, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/32/32/abstract.
                                        html.

Sudholt:2005:AGC

[SBA^{+}05] Wibke Sudholt, Kim K. Baldridge, David Abramson, Colin Enticott, Slavisa Garic, Chris Kondric, and Duy Nguyen. Application of grid computing to parameter sweeps and optimizations in molecular modeling. Future Generation Computer Systems, 21(1):27–35, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Smarr:2009:SSO

[SBdL09] Larry Smarr, Maxine Brown, and Cees de Laat. Special section: OptIPlanet — the OptIPuter global collaboratory. Future Generation Computer Systems, 25(2): 109–113, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Scholl:2009:SCD

[SBG^{+}09] Tobias Scholl, Bernhard Bauer, Benjamin Gufler, Richard Kuntschke, Angelika Reiser, and Alfons Kemper. Scalable communitydriven data sharing in escience grids. Future Generation Computer Systems, 25 (3):290–300, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schulz:2008:CMB

[SBHD08] Sven Schulz, Wolfgang Blochinger, Markus Held, and Clemens Dangelmayr. COHESION — a microkernel based Desktop Grid platform for irregular task-parallel applications. Future Generation Computer Systems, 24(5):354–370, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Steiger:2005:UAD

[SBLT05] Rainer Steiger, Christian H.

Bischof, Bruno Lang, and Walter Thiel. Using automatic differentiation to compute derivatives for a quantum-chemical computer program. Future Generation Computer Systems, 21 (8):1324–1332, October 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Seinstra:1998:PSI

[SBS98] Frank J. Seinstra, Henri E.

Bal, and Hans J. W. Spoelder. Parallel simulation of ion recombination in nonpolar liquids. Future Generation Computer Systems, 13(4–5):261– 268, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/19/19/abstract.
                                        html.

Scarpa:2006:HID

[SBSdL06] M. Scarpa, R. G. Belleman, P. M. A. Sloot, and C. T. A. M. de Laat. Highly interactive distributed visualization. Future Generation Computer Systems, 22(8):896–900, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shakhov:2004:DMD

[SCB04] Vladimir V. Shakhov, Hyunseung Choo, and Young-Cheol Bang. Discord model for detecting unexpected demands in mobile networks. Future Generation Computer Systems, 20(2):181–188, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schoemaker:1985:AIS

[Sch85] S. Schoemaker. Artificial intelligence in simulation. Future Generation Computer Systems, 1(4):245–247, June 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schlumberger:1994:SDI

[Sch94] Maurice Schlumberger. Software developments for industrial research. Future Generation Computer Systems, 10(2– 3):263–267, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schoffel:1998:DRS

[Sch98] Frank Schöffel. Dynamic radiosity shadows for interactive virtual environments. Future Generation Computer Systems, 14(3–4):223–229, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/29/18/26/28/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/29/18/26/abstract.
html.

Schroeder:2000:TVA

[Sch00] Michael Schroeder. Towards a visualization of arguing agents. Future Generation Computer Systems, 17(1):15– 26, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/24/27/abstract.
html.

Schonfisch:2001:SIB

[Sch01]  Birgitt Schönfisch.  Sim-

ple individual based models of movement, alignment and schooling behaviour. Future Generation Computer Systems, 17 (7):873–882, May 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/34/30/abstract.
html.

Schneider:2003:SBH

[Sch03] Johannes Schneider. Searching for Backbones — a highperformance parallel algorithm for solving combinatorial optimization problems. Future Generation Computer Systems, 19(1):121–131, January 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shimojo:2000:SMD

[SCK^{+}00] Fuyuki Shimojo, Timothy J. Campbell, Rajiv K. Kalia, Aiichiro Nakano, Priya Vashishta, Shuji Ogata, and Kenji Tsuruta. A scalable moleculardynamics algorithm suite for materials simulations: designspace diagram on 1024 Cray T3E processors. Future Generation Computer Systems, 17(3):279–291, November 1, 2000. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/29/abstract.
html.

Soudan:2009:FSE

[SCP09] Sebastien Soudan, Bin Bin Chen, and Pascale Vicat-Blanc Primet. Flow scheduling and endpoint rate control in GridNetworks. Future Generation Computer Systems, 25 (8):904–911, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sheu:2001:MSC

[SCY01] Ruey-Kai Sheu, Yue-Shan Chang, and Shyan-Ming Yuan. Managing and sharing collaborative files through WWW. Future Generation Computer Systems, 17(8): 1039–1049, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/39/abstract.
html.

Shevchenko:2002:TCM

[SD02] R. Shevchenko and A. Doroshenko.

A time cost model for distributed objects parallel computation. Future Generation Computer Systems, 18 (6):807–812, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Soleimany:2003:PDA

[SD03] Cyrus Soleimany and Sivarama P.

Dandamudi. Performance of a distributed architecture for query processing on workstation clusters. Future Generation Computer Systems, 19 (4):463–478, May 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shi:2006:SWA

[SD06] Zhiao Shi and Jack J. Dongarra. Scheduling workflow applications on processors with different capabilities. Future Generation Computer Systems, 22(6):665–675, May 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stankovski:2007:SSD

[SD07] Vlado Stankovski and Werner Dubitzky. Special section: Data mining in grid computing environments. Future Generation Computer Systems, 23(1):31–33, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Smarr:2006:SSI

[SDBdL06] Larry Smarr, Thomas A. De-Fanti, Maxine D. Brown, and Cees de Laat. Special section: iGrid 2005: The Global Lambda Integrated Facility. Future Generation Computer Systems, 22(8):849–851, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stevens:2009:MCJ

[SDD^{+}09] T. Stevens, M. De Leenheer, C. Develder, B. Dhoedt, K. Christodoulopoulos, P. Kokkinos, and E. Varvarigos. Multicost job routing and scheduling in Grid networks. Future Generation Computer Systems, 25(8):912–925, September 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schoneveld:1999:PCF

[SdR99] A. Schoneveld and J. F.

de Ronde. P-CAM: a framework for parallel complex systems simulations. Future Generation Computer Systems, 16(2–3):217–234, December 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/30/abstract.
html.

Surmas:2004:LBS

[SdSP04] Rodrigo Surmas, Luís O. E.

dos Santos, and Paulo C. Philippi. Lattice Boltzmann simulation of the flow interference in bluff body wakes. Future Generation Computer Systems, 20(6):951–958, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sander:1999:HPC

[SEH99] Volker Sander, Dietmar Erwin, and Valentina Huber. High-performance computer management based on Java. Future Generation Computer Systems, 15(3):425– 432, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/27/abstract.
html.

Sergot:1995:CFT

[Ser95] Marek Sergot. Contributions of FGCS technology to applications in legal reasoning. Future Generation Computer Systems, 11(3):329–343, June 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Seredynski:1998:STP

[Ser98]  F. Seredyński.  Schedul-

ing tasks of a parallel program in two-processor systems with use of cellular automata. Future Generation Computer Systems, 14(5–6): 351–364, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Smith:2006:BSU

[SF06] Scott F. Smith and James F. Frenzel. Bioinformatic searches using a single-chip sharedmemory multiprocessor. Future Generation Computer Systems, 22(1–2):80–87, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sloot:1995:GEM

[SG95] Peter M. A. Sloot and Andrew Grant. Guest editorial: Massive parallel computing. Future Generation Computer Systems, 11(2): 113–114, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schenk:2004:SUS

[SG04] Olaf Schenk and Klaus Gärtner. Solving unsymmetric sparse systems of linear equations with PARDISO. Future Generation Computer Systems, 20(3):475–487, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Selikhov:2005:CMB

[SG05] A. Selikhov and C. Germain.

A Channel Memory based fault tolerance for MPI applications. Future Generation Computer Systems, 21 (5):709–715, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sierra:1996:DDL

[SGdMM96] Carles Sierra, Lluis Godo, Ramon López de Màntaras, and Mara Manzano. Descriptive dynamic logic and its application to reflective architectures. Future Generation Computer Systems, 12(2– 3):157–171, September 1996.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schenk:2001:PHP

[SGFS01] Olaf Schenk, Klaus Gärtner, Wolfgang Fichtner, and Andreas Stricker. PARDISO: a high-performance serial and parallel sparse linear solver in semiconductor device simulation. Future Generation Computer Systems, 18(1):69– 78, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/33/abstract.
html.

Sodan:2008:TSA

[SGH^{+}08] Angela C. Sodan, Garima Gupta, Lin Han, Lun Liu, and Benjamin Lafreniere. Time and space adaptation for computational grids with the ATOP-grid middleware. Future Generation Computer Systems, 24(6):561–581, June 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sarti:1999:PBM

[SGL99] Alessandro Sarti, Roberto Gori, and Claudio Lamberti. A physically based model to simulate maxillofacial surgery from 3D CT images. Future Generation Computer Systems, 15(2): 217–221, March 11, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/17/24/abstract.
html.

Smarr:2009:BOC

[SGP^{+}09] Larry Smarr, Paul Gilna, Phil Papadopoulos, Thomas A. DeFanti, Greg Hidley, John Wooley, E. Virginia Armbrust, Forest Rohwer, and Eric Frost. Building an OptIPlanet collaboratory to support microbial metagenomics. Future Generation Computer Systems, 25(2): 124–131, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shen:2007:ABP

[SGY^{+}07] Jun Shen, Georg Grossmann, Yun Yang, Markus Stumptner, Michael Schrefl, and Thomas Reiter. Analysis of business process integration in Web service context. Future Generation Computer Systems, 23(3):283–294, March 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shi:1990:ATL

[SH90] Zhongzhi Shi and Jianchao Han. Attribute theory in learning systems. Future Generation Computer Systems, 6 (1):65–69, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sarmenta:1999:BBS

[SH99] Luis F. G. Sarmenta and Satoshi Hirano. Bayanihan: building and studying Web-based volunteer computing systems using Java. Future Generation Computer Systems, 15(5–6):675–686, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/21/28/abstract.
                                        html.

Stutzle:2000:MMA

[SH00] Thomas Stützle and Holger H. Hoos. MAX–MIN Ant System. Future Generation Computer Systems, 16 (8):889–914, June 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/31/28/abstract.
                                        html.

Smarr:1989:SLB

[SHB89] Larry Smarr, David Hobill, and David Bernstein. Shedding light on black holes. Future Generation Computer Systems, 5(2–3): 225–242, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shen:2000:DES

[She00] Chien-Chung Shen. Discreteevent simulation on the Internet and the Web. Future Generation Computer Systems, 17 (2):187–196, October 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/27/32/abstract.
html.

Sheka:2004:FP

[She04] Elena F. Sheka. Fullerenes as polyradicals. Future Generation Computer Systems, 20(5): 749–762, June 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shirai:1985:RV

[Shi85]  Y. Shirai.  Robot vi-

sion. Future Generation Computer Systems, 1(5):325– 352, September 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shibusawa:1992:CPI

[Shi92]  Susumu Shibusawa.  The

Cubemat: a parallel interconnection model consisting of hypercube and global bus connections. Future Generation Computer Systems, 7(2– 3):151–160, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shim:2004:PES

[Shi04] Young-Chul Shim. Performance evaluation of scheduling schemes for NOW with heterogeneous computing power. Future Generation Computer Systems, 20(2):229–236, February 16, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stack:2006:MAF

[SHJ06] Brian M. Stack, Gene Hsiao, and Stephen F. Jenks. A middleware architecture to facilitate distributed programming: DAROC: Data-Activated Replicated Object Communications. Future Generation Computer Systems, 22 (1–2):88–101, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sierra:2004:LCC

[SHJR04] José M. Sierra, Julio C. Hernández, Narayana Jayaram, and Arturo Ribagorda. Low computational cost integrity for block ciphers. Future Generation Computer Systems, 20(5):857–863, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sun:2008:RDT

[SHLB08] Hailong Sun, Jinpeng Huai, Yunhao Liu, and Rajkumar Buyya. RCT: a distributed tree for supporting efficient range and multi-attribute queries in Grid computing. Future Generation Computer Systems, 24(7):631–643, July 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Simmons:1986:TMT

[Sim86] R. F. Simmons. Technologies for machine translation. Future Generation Computer Systems, 2(2):83–94, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sinclair:1984:SCS

[Sin84] Sir Clive Sinclair. Sir Clive Sinclair on the third industrial revolution. Future Generation Computer Systems, 1 (2):119–122, November 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sinkwitz:1992:PMS

[Sin92] R. C. Sinkwitz. Particle motion simulation — a parallel distributed workstation application for real-time. Future Generation Computer Systems, 8(1–3):43–47, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sinnott:2007:AIM

[Sin07] Richard O. Sinnott. From access and integration to mining of secure genomic data sets across the Grid. Future Generation Computer Systems, 23 (3):447–456, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Strizh:2007:SBG

[SJTG07] Irina Strizh, Alexei Joutchkov, Nikolay Tverdokhlebov, and Sergey Golitsyn. Systems biology and Grid technologies: Challenges for understanding complex cell signaling networks. Future Generation Computer Systems, 23 (3):428–434, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schneier:1997:RAS

[SK97] Bruce Schneier and John Kelsey. Remote auditing of software outputs using a trusted coprocessor. Future Generation Computer Systems, 13(1):9–18, June 20, 1997. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
counterpane.com/remote_
auditing.html;   http://
www.elsevier.com/gej-ng/
10/19/19/28/17/17/abstract.
html.

Sarfraz:2004:AAA

[SK04] M. Sarfraz and M. A. Khan. An automatic algorithm for approximating boundary of bitmap characters. Future Generation Computer Systems, 20(8):1327–1336, November 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Subramani:2005:GSD

[SK05] K. Subramani and L. Kovalchick. A greedy strategy for detecting negative cost cycles in networks. Future Generation Computer Systems, 21 (4):607–623, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Spooner:2006:PFI

[SK06] Daniel P. Spooner and Darren J. Kerbyson. Performance feature identification by comparative trace analysis. Future Generation Computer Systems, 22(3):369–380, February 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shirai:2009:RTS

[SKF^{+}09] Daisuke Shirai, Tetsuo Kawano, Tatsuya Fujii, Kunitake Kaneko, Naohisa Ohta, Sadayasu Ono, Sachine Arai, and Terukazu Ogoshi. Real time switching and streaming transmission of uncompressed 4K motion pictures. Future Generation Computer Systems, 25(2): 192–197, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Seredynski:2001:DMS

[SKJ01] F. Seredyński, J. Koronacki, and C. Z. Janikow. Distributed multiprocessor scheduling with decomposed optimization criterion. Future Generation Computer Systems, 17(4):387–396, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/29/abstract.
html.

Sirakoulis:2002:CAM

[SKT02] G. Ch. Sirakoulis, I. Karafyllidis, and A. Thanailakis. A cellular automaton methodology for the simulation of integrated circuit fabrication processes. Future Generation Computer Systems, 18(5): 639–657, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/31/abstract.
html.

Sakurai:2008:PML

[SKT^{+}08] Tetsuya Sakurai, Yoshihisa Kodaki, Hiroto Tadano, Daisuke Takahashi, Mitsuhisa Sato, and Umpei Nagashima. A parallel method for large sparse generalized eigenvalue problems using a GridRPC system. Future Generation Computer Systems, 24(6):613–619, June 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sheu:1987:EPI

[SL87] Phillip C. Sheu and W. S.

Lee. Efficient processing of integrity constraints in deductive databases. Future Generation Computer Systems, 3 (3):201–215, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sloot:1997:HGA

[SL97] Peter Sloot and Heather Liddell. HPCN96: General applications, computational and computer science. Future Generation Computer Systems, 12(5):331– 333, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/17/abstract.
html.

Singh:2003:THR

[SLDK03] Rajvikram Singh, Jason Leigh, Thomas A. DeFanti, and Fotis Karayannis. TeraVision: a high resolution graphics streaming device for amplified collaboration environments. Future Generation Computer Systems, 19(6): 957–971, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sobieski:2006:DPL

[SLJ^{+}06] Jerry Sobieski, Tom Lehman, Bijan Jabbari, Chester Ruszczyk, Rick Summerhill, and Alan Whitney. Dynamic provisioning of LightPath services for radio astronomy applications. Future Generation Computer Systems, 22(8):984–992, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sloot:1996:RMD

[Slo96] Peter Sloot. Resource management in distributed systems. Future Generation Computer Systems, 12(1):1–2, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sloot:2005:P

[Slo05a] P. M. A. Sloot. Preface. Future Generation Computer Systems, 21(1):1–2, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stamatakis:2005:DDP

[SLO^{+}05b] A. Stamatakis, M. Lindermeier, M. Ott, T. Ludwig, and H. Meier. DRAxML@home: a distributed program for computation of large phylogenetic trees. Future Generation Computer Systems, 21 (5):725–730, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sloot:2006:RFGa

[Slo06a] Peter Sloot. To the readers of Future Generation Computer Systems. Future Generation Computer Systems, 22 (1–2):1, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sloot:2006:RFGb

[Slo06b] Peter Sloot. To the readers of Future Generation Computer Systems. Future Generation Computer Systems, 22 (3):253, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sommerfeld:2009:AMA

[SLS^{+}09] Dietmar Sommerfeld, Thomas Lingner, Mario Stanke, Burkhard Morgenstern, and Harald Richter. AUGUSTUS at MediGRID: Adaption of a bioinformatics application to Grid computing for efficient genome analysis. Future Generation Computer Systems, 25 (3):337–345, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sun:2001:DPS

[SLW01] Yudong Sun, Zhengyu Liang, and Cho-Li Wang. Distributed particle simulation method on adaptive collaborative system. Future Generation Computer Systems, 18(1):

79–87, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/34/abstract.
html.

Steinfeld:1995:FAP

[SLZ95] P. Steinfeld, L. Lequette, and E. Znaty. First attempt to parallelise a CFD application software package: The CAL-IFE code. Future Generation Computer Systems, 11(1): 71–86, February 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Skobeltsyn:2009:QDI

[SLŽ^{+}09] Gleb Skobeltsyn, Toan Luu, Ivana Podnar Žarko, Martin Rajman, and Karl Aberer. Query-driven indexing for scalable peer-to-peer text retrieval. Future Generation Computer Systems, 25(1):89– 99, January 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sunderam:1996:PDS

[SM96] V. S. Sunderam and Steven A. Moyer. Parallel I/O for distributed systems: Issues and implementation. Future Generation Computer Systems, 12 (1):25–38, May 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sanchez:2001:BWA

[SM01a] Miguel Sánchez and Pietro Manzoni. Best of Websim99: ANEJOS: a Java based simulator for ad hoc networks. Future Generation Computer Systems, 17(5): 573–583, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/31/abstract.
html.

Shudo:2001:AME

[SM01b] Kazuyuki Shudo and Yoichi Muraoka. Asynchronous migration of execution context in Java Virtual Machines. Future Generation Computer Systems, 18(2): 225–233, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/30/abstract.
html.

Scariot:2003:NMB

[SM03] Jean-François Scariot and Bernard Martinet. NetSEC: metrology-based application for network security. Future Generation Computer Systems, 19(2):303–311, February 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sharifian:2008:CBL

[SMA08] Saeed Sharifian, Seyed A. Motamedi, and Mohammad K. Akbari. A content-based load balancing algorithm with admission control for cluster Web servers. Future Generation Computer Systems, 24(8): 775–787, October 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schelkens:1999:WBC

[SMC99] Peter Schelkens, Adrian Munteanu, and Jan Cornelis. Waveletbased compression of medical images: Protocols to improve resolution and quality scalability and region-of-interest coding. Future Generation Computer Systems, 15 (2):171–184, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/17/20/abstract.
html.

Smidt:1986:UIC

[Smi86] Palle F. Smidt. U.S. industrial cooperation in R&D. Future Generation Computer Systems, 2(1):45–50, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sena:2001:IPG

[SMI01] Giuseppe A. Sena, Dalila Megherbi, and Germinal Isern. Implementation of a parallel Genetic Algorithm on a cluster of workstations: Traveling Salesman Problem, a case study. Future Generation Computer Systems, 17 (4):477–488, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/37/abstract.
html.

Schreiber:2005:PSE

[SMK05] Andreas Schreiber, Thijs Metsch, and Hans-Peter Kersken. A problem solving environment for multidisciplinary coupled simulations in computational grids. Future Generation Computer Systems, 21(6):942–952, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Solvberg:1992:KBI

[SNA92] Ingeborg Sølvberg, Inge Nordbø, and Agnar Aamodt. Knowledgebased information retrieval. Future Generation Computer Systems, 7(4):379–390, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schreckenberg:2001:TSS

[SNW01] M. Schreckenberg, L. Neubert, and J. Wahle. Traffic simulation: Simulation of traffic in large road networks. Future Generation Computer Systems, 17(5): 649–657, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/30/37/abstract.
                                        html.

Stohr:1998:FFS

[SO98] E. A. Stöhr and M. F. P. O’Boyle. First Fast Sink: a compiler algorithm for barrier placement optimisation. Future Generation Computer Systems, 13(4–5):397– 406, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/19/31/abstract.
                                        html.

Sas:2005:VET

[SOR05] Corina Sas, Gregory O’Hare, and Ronan Reilly. Virtual environment trajectory analysis: a basis for navigational assistance and scene adaptivity. Future Generation Computer Systems, 21(7): 1157–1166, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fogelman-Soulie:1991:NNC

[Sou91] Françoise Fogelman Soulié. Neural networks and computing. Future Generation Computer Systems, 7(1):69– 77, October 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Smith:1993:EOK

[SP93] H. N. Smith and K. J. Poulter.

The elements of an open KBS infrastructure. Future Generation Computer Systems, 9(4): 349–369, December 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sulistio:2007:IDL

[SPBT07] Anthony Sulistio, Gokul Poduval, Rajkumar Buyya, and Chen-Khong Tham. On incorporating differentiated levels of network service into Grid-Sim. Future Generation Computer Systems, 23(4):606–615, May 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Smirnov:2004:KLI

[SPCL04] Alexander Smirnov, Mikhail Pashkin, Nikolai Chilov, and Tatiana Levashova. Knowledge logistics in information grid environment. Future Generation Computer Systems, 20(1):61–79, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Spinnato:2009:SSN

[SPEW09] Piero Spinnato, Pascale Vicat-Blanc Primet, Chris Edwards, and Michael Welzl. Special section on networks for grid applications. Future Generation Computer Systems, 25 (8):893–894, September 2009.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sanchez:2007:MDE

[SPK^{+}07] Alberto Sánchez, María S. Pérez, Konstantinos Karasavvas, Pilar Herrero, and Antonio Pérez. MAPFS–DAI, an extension of OGSA–DAI based on a parallel file system. Future Generation Computer Systems, 23(1):138– 145, January 1, 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Small:1986:PAI

[SPM86] M. Small, A. Pinkerton, and I. Meyer. Practical applications involving uncertainty. Future Generation Computer Systems, 2(3):153– 159, September 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Seguel:2003:DPC

[SR03] Jaime Seguel and Domingo Rodríguez. The doctoral program in Computing and Information Sciences and Engineering of the University of Puerto Rico. Future Generation Computer Systems, 19(8): 1293–1298, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Strumpen:1997:PHP

[SRCR97] V. Strumpen, B. Ramkumar, T. L. Casavant, and S. M. Reddy. Perspectives on high performance network computing. Future Generation Computer Systems, 12(5): 451–459, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/27/abstract.
html.

Stiller:2003:CAH

[SRG^{+}03] Burkhard Stiller, Peter Reichl, Jan Gerke, Hasan, and Placi Flury. Charging and accounting in high-speed networks. Future Generation Computer Systems, 19(1):101–109, January 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schwister:1990:SEM

[SS90] Bernd Schwister and Karl Solchenbach. SUPRENUM — a European made supercomputer. Future Generation Computer Systems, 5(4):381– 385, January 1, 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sofroniou:2003:IFR

[SS03] Mark Sofroniou and Giulia Spaletta. Increment formulations for rounding error reduction in the numerical solution of structured differential systems. Future Generation Computer Systems, 19 (3):375–383, April 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shen:2004:IRT

[SS04] Haifeng Shen and Chengzheng Sun. Improving real-time collaboration with highlighting. Future Generation Computer Systems, 20(4):605–625, May 3, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stuer:2005:TOC

[SSB05] Gunther Stuer, Vaidy Sunderam, and Jan Broeckhove. Towards OGSA compatibility in the H2O metacomputing framework. Future Generation Computer Systems, 21 (1):221–226, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sosonkina:2004:UPA

[SSC04] M. Sosonkina, Y. Saad, and X. Cai. Using the parallel algebraic recursive multilevel solver in modern physical applications. Future Generation Computer Systems, 20(3): 489–500, April 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shao:2009:EDC

[SSC09] Qihong Shao, Peng Sun, and Yi Chen. Efficiently discovering critical workflows in scientific explorations. Future Generation Computer Systems, 25 (5):577–585, May 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Smith:2009:SDG

[SSF^{+}09] M. Smith, M. Schmidt, N. Fallenbeck, T. Dörnemann, C. Schridde, and B. Freisleben. Secure on-demand Grid computing. Future Generation Computer Systems, 25(3): 315–325, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stankovski:2008:GED

[SSK^{+}08] Vlado Stankovski, Martin Swain, Valentin Kravtsov, Thomas Niessen, Dennis Wegener, Jörg Kindermann, and Werner Dubitzky. Gridenabling data mining applications with DataMiningGrid: An architectural perspective. Future Generation Computer Systems, 24(4):259–279, April 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schafers:1995:TGP

[SSKF95] Lorenz Schäfers, Christian Scheidler, and Ottmar Krämer-Fuhrmann. Trapper: a graphical programming environment for parallel systems. Future Generation Computer Systems, 11(4–5):351–361, August 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sterling:1995:IEC

[SSMG95] Thomas Lawrence Sterling, Daniel F. Savarese, Phillip R. Merkey, and Jeffrey P. Gardner. An initial evaluation of the Convex SPP-1000 for earth and space science applications. Future Generation Computer Systems, 11(6): 567–583, October 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schmidt:2002:HAB

[SSS02] Bertil Schmidt, Heiko Schröder, and Manfred Schimmler. A hybrid architecture for bioinformatics. Future Generation Computer Systems, 18 (6):855–862, May 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shimizu:2006:IRT

[SST^{+}06] Takashi Shimizu, Daisuke Shirai, Hirokazu Takahashi, Takahiro Murooka, Kazuaki Obana, Yoshihide Tonomura, Takeru Inoue, Takahiro Yamaguchi, Tetsuro Fujii, Naohisa Ohta, Sadayasu Ono, Tomonori Aoyama, Laurin Herr, Natalie van Osdol, Xi Wang, Maxine D. Brown, Thomas A. DeFanti, Rollin Feld, Jacob Balser, Steve Morris, et al. International realtime streaming of 4K digital cinema. Future Generation Computer Systems, 22(8): 929–939, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Spezzano:1998:DPM

[ST98] Giandomenico Spezzano and Domenico Talia. Designing parallel models of soil contamination by the CARPET language. Future Generation Computer Systems, 13(4–5): 291–302, March 11, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/19/22/abstract.
html.

Spezzano:1999:PCA

[ST99] Giandomenico Spezzano and Domenico Talia. Programming cellular automata algorithms on parallel computers. Future Generation Computer Systems, 16(2–3):203– 216, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/29/abstract.
html.

Steels:1985:SGE

[Ste85] Luc Steels. Second generation expert systems. Future Generation Computer Systems, 1 (4):213–221, June 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Steiner:1992:EMD

[Ste92] Peter Steiner. Extending multiprogramming to a DMPP. Future Generation Computer Systems, 8(1–3):93–109, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stevens:1994:HPC

[Ste94] Rick Stevens. High-performance computing and communications. Future Generation Computer Systems, 10(2–3): 159–167, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sato:1998:NPL

[STH^{+}98] Mitsuhisa Sato, Hiroshi Tezuka, Atsushi Hori, Yutaka Ishikawa, Satoshi Sekiguchi, Hidemoto Nakada, Satoshi Matsuoka, and Umpei Nagashima. Ninf and PMCommunication libraries for global computing and high-performance cluster computing. Future Generation Computer Systems, 13(4– 5):349–359, March 11, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/28/19/27/abstract.
                                        html.

Stickel:1993:ATP

[Sti93] Mark E. Stickel. Automated theorem-proving research in the Fifth Generation Computer Systems Project: Model generation theorem provers. Future Generation Computer Systems, 9(2):143–152, July 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Scheetz:2005:GTC

[STP^{+}05] Todd E. Scheetz, Nishank Trivedi, Kevin T. Pedretti, Terry A. Braun, and Thomas L. Casavant. Gene transcript clustering: a comparison of parallel approaches. Future Generation Computer Systems, 21(5):731–735, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schulz:2003:SIM

[STTK03] Martin Schulz, Jie Tao, Carsten Trinitis, and Wolfgang Karl. SMiLE: an integrated, multi-paradigm software infrastructure for SCIbased clusters. Future Generation Computer Systems, 19 (4):521–532, May 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stamatakos:1998:SSM

[SUD^{+}98] Georgios S. Stamatakos, Nikolaos K. Uzunoglu, Konstantinos Delibasis, Mersini Makropoulou, Nikolaos Mouravliansky, and Andy Marsh. A simplified simulation model and virtual reality visualization of tumour growth in vitro. Future Generation Computer Systems, 14(1–2): 79–89, June 15, 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/29/17/24/abstract.
html.

Sunderam:1992:HNB

[Sun92] V. S. Sunderam. Heterogeneous network-based concurrent computing environments. Future Generation Computer Systems, 8(1–3):191–203, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Suzuki:1989:DDM

[Suz89] Yasuo Suzuki. DYMOS (Dynamic MOnitoring System): Monitoring and real time fault diagnosis system for a ship’s propelling plant. Future Generation Computer Systems, 5 (1):71–76, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Spinnato:2001:PBC

[SvAS01] P. F. Spinnato, G. D. van Albada, and P. M. A. Sloot. Performance of N-body codes on hybrid machines. Future Generation Computer Systems, 17 (8):951–959, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/31/abstract.

html. Accepted.

Stuer:2007:CMA

[SVB07] Gunther Stuer, Kurt Vanmechelen, and Jan Broeckhove. A commodity market algorithm for pricing substitutable Grid resources. Future Generation Computer Systems, 23(5):688–701, June 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Scherson:2007:FGC

[SVC^{+}07] Isaac D. Scherson, Daniel Valencia, Enrique Cauich, John Duselis, and Richert Wang. Federated grid clusters using service address routed optical networks. Future Generation Computer Systems, 23(8): 957–967, November 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Silver:1999:VES

[SW99] Deborah Silver and Xin Wang.

Visualizing evolving scalar phenomena. Future Generation Computer Systems, 15 (1):99–108, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-

                                        ng/10/19/19/30/18/24/29/
                                        index.htt;   http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/18/24/abstract.
                                        html.

Stratmann:2002:LED

[SW02] Michael Stratmann and Thomas Worsch. Leader election in d-dimensional CA in time diam log(diam). Future Generation Computer Systems, 18(7):939–950, August 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schamberger:2005:PFE

[SW05] Stefan Schamberger and Jens-Michael Wierum. Partitioning finite element meshes using space-filling curves. Future Generation Computer Systems, 21(5):759–766, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stolk:2006:BMG

[SW06] Bram Stolk and Paul Wielinga. Building a 100 Mpixel graphics device for the OptIPuter. Future Generation Computer Systems, 22(8):972–975, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Saulsbury:1995:ASC

[SWCL95] Ashley Saulsbury, Tim Wilkinson, John Carter, and Anders Landin. An argument for simple COMA. Future Generation Computer Systems, 11(6):553–566, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.cs.utah.edu/~retrac/
papers/hpca95.ps.Z.

Schmidt:2003:GSV

[SWCP03] Thomas C. Schmidt, Matthias Wählisch, Hans L. Cycon, and Mark Palkow. Global serverless videoconferencing over IP. Future Generation Computer Systems, 19(2): 219–227, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shen:2004:ERD

[SY04] Jun Shen and Yun Yang. Extending RDF in distributed knowledge-intensive applications. Future Generation Computer Systems, 20(1):27– 46, January 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Shih:2009:OBC

[SYT09] Wen-Chung Shih, Chao-Tung Yang, and Shian-Shyong Tseng. Ontology-based content organization and retrieval for SCORM-compliant teaching materials in Data Grids. Future Generation Computer Systems, 25(6):687–694, June 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Salleh:1998:MSU

[SZ98] Shaharuddin Salleh and Albert Y. Zomaya. Multiprocessor scheduling using meanfield annealing. Future Generation Computer Systems, 14(5–6):393–408, December 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sanna:2005:DJB

[SZC05] A. Sanna, C. Zunino, and L. Ciminiera. A distributed JXTA-based architecture for searching and retrieving solar data. Future Generation Computer Systems, 21(3): 349–359, March 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sun:2004:ASC

[SZGbC04] Chengzhung Sun, Wanlei Zhou, Andrzej Goscinski, and Xue bin Chi. Advanced services for Clusters and Internet computing. Future Generation Computer Systems, 20(4): 501–503, May 3, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sarjoughian:2000:CDN

[SZP00] Hessam S. Sarjoughian, Bernard P. Zeigler, and Sunwoo Park. Collaborative distributed network system: a lightweight middleware supporting collaborative DEVS modeling. Future Generation Computer Systems, 17(2):89– 105, October 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/27/25/abstract.
html.

Szuba:1998:MQR

[Szu98] Tadeusz Szuba. A molecular quasi-random model of computations applied to evaluate collective intelligence. Future Generation Computer Systems, 14(5–6): 321–339, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Szuba:2001:FDP

[Szu01] Tadeusz Szuba. A formal definition of the phenomenon of collective intelligence and its IQ measure. Future Generation Computer Systems, 17 (4):489–500, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/38/abstract.
html.

Tyugu:1996:DRT

[TA96] Enn Tyugu and Mattin Addibpour. Declarative reflection tools for agent shells. Future Generation Computer Systems, 12(2–3):203– 215, September 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tabata:2006:FEA

[Tab06]  Masahisa Tabata.  Fi-

nite element approximation to infinite Prandtl number Boussinesq equations with temperature-dependent coefficients — thermal convection problems in a spherical shell. Future Generation Computer Systems, 22(4): 521–531, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Takahashi:1989:ATE

[Tak89a] Riichi Takahashi. Automobile troubleshooting expert system “ATREX”. Future Generation Computer Systems, 5(1): 97–101, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Takamura:1989:ESC

[Tak89b]  Jun Takamura.  An ex-

pert system for computer operation and user assistance. Future Generation Computer Systems, 5(1):87–95, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Takatsuka:2005:COS

[Tak05]  Masahiro Takatsuka.  A

component-oriented software authoring system for exploratory visualization. Future Generation Computer Systems, 21(7):1213–1222, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tan:2002:GEC

[Tan02a] C. J. Kenneth Tan. Guest editorial: Computational science. Future Generation Computer Systems, 18(5): 659, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/32/abstract.
html.

Tan:2002:PPP

[Tan02b] Chih Jeng Kenneth Tan.

The PLFG parallel pseudorandom number generator. Future Generation Computer Systems, 18(5):693– 698, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/36/abstract.
html.

Trivedi:2002:PCN

[TBD^{+}02] Nishank Trivedi, Jared Bischof, Steve Davis, Kevin Pedretti, Todd E. Scheetz, Terry A. Braun, Chad A. Roberts, Natalie L. Robinson, Val C. Sheffield, M. Bento Soares, and Thomas L. Casavant. Parallel creation of nonredundant gene indices from partial mRNA transcripts.

Future Generation Computer Systems, 18(6):863–870, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Turek:2006:HON

[TBK06] S. Turek, Chr. Becker, and S. Kilian. Hardware-oriented numerics and concepts for PDE software. Future Generation Computer Systems, 22 (1–2):217–238, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Truong:2009:DDP

[TBNF09] Hong-Linh Truong, Peter Brunner, Vlad Nae, and Thomas Fahringer. DIPAS: a distributed performance analysis service for Grid servicebased workflows. Future Generation Computer Systems, 25 (4):385–398, April 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tsang:1992:AAA

[TC92] Jean Patrick Tsang and Gilles Cardonne. An AI approach to automate the preliminary design phase of electronic equipment for satellites. Future Generation Computer Systems, 7(4):353–364, May 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Thaker:2006:STS

[TC06] Darshan D. Thaker and Vipin Chaudhary. Simulation tools to study a distributed shared memory for clusters of symmetric multiprocessors. Future Generation Computer Systems, 22(1–2):57–66, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Temam:1995:SAD

[TD95] O. Temam and N. Drach.

Software assistance for data caches. Future Generation Computer Systems, 11(6): 519–536, October 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Truong:2007:PMO

[TDF07] Hong-Linh Truong, Schahram Dustdar, and Thomas Fahringer. Performance metrics and ontologies for Grid workflows. Future Generation Computer Systems, 23(6):760–772, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Travostino:2006:SLM

[TDG^{+}06] Franco Travostino, Paul Daspit, Leon Gommans, Chetan Jog, Cees de Laat, Joe Mambretti, Inder Monga, Bas van Oudenaarde, Satish Raghunath, and Phil Yonghui Wang. Seamless live migration of virtual machines over the MAN/WAN. Future Generation Computer Systems, 22(8): 901–907, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tran:2005:SMG

[TDL05] Minh Tran, Amitava Datta, and Nick Lowe. A simple model generation system for computer graphics. Future Generation Computer Systems, 21(7):1223–1234, July 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Thysebaert:2008:SDR

[TDV^{+}08] Pieter Thysebaert, Marc De Leenheer, Bruno Volckaert, Filip De Turck, Bart Dhoedt, and Piet Demeester. Scalable dimensioning of resilient Lambda Grids. Future Generation Computer Systems, 24 (6):549–560, June 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tebra:1986:PPF

[Teb86] Hans Tebra. Prolog for programmers: by F. Kluzniak and S. Szpakowicz (Academic Press, 400 pages, $47.50, 1985). Future Generation Computer Systems, 2(4):261– 262, December 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tse:2004:TMC

[TG04] Rebecca O. C. Tse and Christopher Gold. TIN meets CAD — extending the TIN concept in GIS. Future Generation Computer Systems, 20(7):1171–1184, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Taniar:2007:CCI

[TG07] David Taniar and Sushant Goel. Concurrency control issues in Grid databases. Future Generation Computer Systems, 23(1):154–162, January 1, 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Talbi:1998:FTP

[THKG98] E.-G. Talbi, Z. Hafidi, D. Kebbal, and J.-M. Geib. A fault-tolerant parallel heuristic for assignment problems. Future Generation Computer Systems, 14(5–6):425– 438, December 1998. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Takefusa:2006:GLC

[THN^{+}06] Atsuko Takefusa, Michiaki Hayashi, Naohide Nagatsu, Hidemoto Nakada, Tomohiro Kudoh, Takahiro Miyamoto, Tomohiro Otani, Hideaki Tanaka, Masatoshi Suzuki, Yasunori Sameshima, Wataru Imajuku, Masahiko Jinno, Yoshihiro Takigawa, Shuichi Okamoto, Yoshio Tanaka, and Satoshi Sekiguchi. G-lambda: Coordination of a Grid scheduler and lambda path service over GMPLS. Future Generation Computer Systems, 22(8): 868–875, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Thomas:2006:ANP

[Tho06] Nigel Thomas. Approximation in non-product form finite capacity queue systems. Future Generation Computer Systems, 22(7):820–827, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tick:1993:APP

[Tic93] Evan Tick. Appraisal of parallel processing research at ICOT. Future Generation Computer Systems, 9(2): 127–136, July 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tiskin:2007:CEP

[Tis07] Alexander Tiskin. Communicationefficient parallel generic pairwise elimination. Future Generation Computer Systems, 23 (2):179–188, February 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tierney:2000:DID

[TJLT00] Brian Tierney, William Johnston, Jason Lee, and Mary Thompson. A data intensive distributed computing architecture for “Grid” applications. Future Generation Computer Systems, 16(5): 473–481, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/30/abstract.
html.

Tarczynski:2008:AGC

[TKT^{+}08] Andrzej Tarczynski, Tamas Kiss, Gabor Tersztyanszki, Thierry Delaitre, Dongdong Qu, and Stephen Winter. Application of Grid computing for designing a class of optimal periodic nonuniform sampling sequences. Future Generation Computer Systems, 24 (7):763–773, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tang:2006:IDR

[TLTY06] Ming Tang, Bu-Sung Lee, Xueyan Tang, and Chai-Kiat Yeo. The impact of data replication on job scheduling performance in the Data Grid. Future Generation Computer Systems, 22(3): 254–268, February 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tang:2005:DRA

[TLYT05] Ming Tang, Bu-Sung Lee, Chai-Kiat Yeo, and Xueyan Tang. Dynamic replication algorithms for the multi-tier Data Grid. Future Generation Computer Systems, 21 (5):775–790, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Treebushny:2005:CRR

[TM05] Dimitri Treebushny and Henrik Madsen. On the construction of a reduced rank square-root Kalman filter for efficient uncertainty propagation. Future Generation Computer Systems, 21(7): 1047–1055, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tantar:2007:PHG

[TMT^{+}07] A.-A. Tantar, N. Melab, E.-G. Talbi, B. Parent, and D. Horvath. A parallel hybrid genetic algorithm for protein structure prediction on the computational grid. Future Generation Computer Systems, 23 (3):398–409, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tezuka:2005:MCG

[TMTY05] Shu Tezuka, Hiroki Murata, Shuji Tanaka, and Shoji Yumae. Monte Carlo grid for financial risk management. Future Generation Computer Systems, 21(5):811–821, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Taffoni:2007:EGT

[TMV^{+}07] Giuliano Taffoni, Davide Maino, Claudio Vuerli, Giuliano Castelli, Riccardo Smareglia, Andrea Zacchei, and Fabio Pasian. Enabling Grid technologies for Planck space mission. Future Generation Computer Systems, 23(2):189– 200, February 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Toffoli:1999:PMM

[Tof99]  Tommaso Toffoli.  Pro-

grammable matter methods. Future Generation Computer Systems, 16(2–3):187– 201, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/28/abstract.
html.

Torrens:2004:ESD

[Tor04] Francisco Torrens. Effect of size and deformation on polarizabilities of carbon nanotubes from atomic increments. Future Generation Computer Systems, 20(5):763–772, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Treleaven:1985:FGV

[TR85] Philip C. Treleaven and Apostolos N. Refenes. Fifth generation and VLSI architectures. Future Generation Computer Systems, 1(6):387– 396, December 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Trendafilov:2003:PP

[Tre03] Nickolay T. Trendafilov. On the ℓ_{1} Procrustes problem. Future Generation Computer Systems, 19(7):1177– 1186, October 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Talbi:2001:PAC

[TRFR01] E.-G. Talbi, O. Roux, C. Fonlupt, and D. Robillard. Parallel Ant Colonies for the quadratic assignment problem. Future Generation Computer Systems, 17(4):441– 449, January 1, 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/29/34/abstract.
                                        html.

Talia:1999:CAP

[TS99] Domenico Talia and Peter M. A. Sloot. Cellular automata: Promise and prospects in computational science. Future Generation Computer Systems, 16(2–3): v–vii, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/25/abstract.
html.

Trnkoczy:2008:IPF

[TS08] Jernej Trnkoczy and Vlado Stankovski. Improving the performance of Federated Digital Library services. Future Generation Computer Systems, 24(8):824–832, October 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tao:2003:AAR

[TSK03] Jie Tao, Martin Schulz, and Wolfgang Karl. ARS: an adaptive runtime system for locality optimization. Future Generation Computer Systems, 19 (5):761–776, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tomassini:1999:GHQ

[TSZP99] Marco Tomassini, Moshe Sipper, Mosé Zolla, and Mathieu Perrenoud. Generating high-quality random numbers in parallel by cellular automata. Future Generation Computer Systems, 16(2–3): 291–305, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        19/19/41/25/35/abstract.
                                        html.

Trunfio:2007:PPR

[TTP^{+}07] P. Trunfio, D. Talia, H. Papadakis, P. Fragopoulou, M. Mordacchini, M. Pennanen, K. Popov, V. Vlassov, and S. Haridi. Peer-to-Peer resource discovery in Grids: Models and systems. Future Generation Computer Systems, 23(7):864–878, August 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

tenTeije:1996:URT

[tTvH96] Annette ten Teije and Frank van Harmelen. Using reflection techniques for flexible problem solving (with examples from diagnosis). Future Generation Computer Systems, 12(2–3):217– 234, September 1996. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tulub:2004:CTS

[Tul04] Alexander A. Tulub. Coherent triplet and singlet states in tubulin dynamics. Future Generation Computer Systems, 20(5):773–780, June 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tikar:2008:ERR

[TV08] Sandip Tikar and Sathish Vadhiyar. Efficient reuse of replicated parallel data segments in computational grids. Future Generation Computer Systems, 24(7):644–657, July 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Taesombut:2006:CDV

[TWC^{+}06] Nut Taesombut, Xinran (Ryan)

Wu, Andrew A. Chien, Atul Nayak, Bridget Smith, Debi Kilb, Thomas Im, Dane Samilo, Graham Kent, and John Orcutt. Collaborative data visualization for Earth Sciences with the OptIPuter. Future Generation Computer Systems, 22(8):955–963, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tomiyama:1985:KEC

[TY85] Tetsuo Tomiyama and Hiroyuki Yoshikawa. Knowledge engineering and CAD. Future Generation Computer Systems, 1(4):237–243, June 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tzeng:2004:NTM

[TYH04] Shiang-Feng Tzeng, Cheng-Ying Yang, and Min-Shiang Hwang. A nonrepudiable threshold multi-proxy multisignature scheme with shared verification. Future Generation Computer Systems, 20(5): 887–893, June 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ufuktepe:2007:AM

[UB07] Ünal Ufuktepe and Goksen Bacak. Applying Mathematica. Future Generation Computer Systems, 23(5):716–720, June 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Uchida:1986:FMT

[Uch86] Hiroshi Uchida. Fujitsu machine translation system: AT-LAS. Future Generation Computer Systems, 2(2):95–100, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Uchida:1987:PIM

[Uch87]  Shunichi Uchida.  Paral-

lel inference machines at ICOT. Future Generation Computer Systems, 3(4):245– 252, December 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Uehara:1989:ILH

[Ueh89]  Kuniaki Uehara.  An in-

telligent on-line help system: ASSIST. Future Generation Computer Systems, 5(1): 11–20, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Umeo:2002:DTT

[UM02] Hiroshi Umeo and Giancarlo Mauri. A duality theorem for two connectivity-preserving parallel shrinking transformations. Future Generation Computer Systems, 18(7): 931–937, August 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Uhrmacher:2000:MSM

[UTT00] Adelinde M. Uhrmacher, Petra Tyschler, and Dirk Tyschler. Modeling and simulation of mobile agents. Future Generation Computer Systems, 17(2):107–118, October 2000. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/27/26/abstract.
html.

Umeo:1992:PGB

[UWV92] Hiroshi Umeo, Thomas Worsch, and Roland Vollmar. On the power of global-bus in mesh-connected architectures. Future Generation Computer Systems, 7(2–3): 161–168, April 1992. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

VanDeRiet:1987:PES

[Van87a] Reind P. Van De Riet. Problems with expert systems? Future Generation Computer Systems, 3(1):11–16, February 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

VanDeRiet:1987:SAK

[Van87b] Reind P. Van De Riet. The state of the art of knowledge engineering at five Japanese Research Institutes — a travel report. Future Generation Computer Systems, 3(1):37– 49, February 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

VanDongen:1992:MUR

[Van92] Vincent Van Dongen. Mapping uniform recurrences onto small size arrays. Future Generation Computer Systems, 8 (4):349–361, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Varvitsiotis:2000:SIL

[Var00]  A. P. Varvitsiotis.  Scal-

ing issues in large PKI communities. Future Generation Computer Systems, 16(4): 361–372, February 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/27/31/abstract.
                                        html.

Varga:2003:NRA

[Var03] A. Varga. A numerically reliable approach to robust pole assignment for descriptor systems. Future Generation Computer Systems, 19 (7):1221–1230, October 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Verhoeven:1995:POA

[VAS95] M. G. A. Verhoeven, E. H. L.

Aarts, and P. C. J. Swinkels. A parallel 2-opt algorithm for the Traveling Salesman Problem. Future Generation Computer Systems, 11(2): 175–182, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vaughan:1993:LTB

[Vau93] John G. Vaughan. A logicaltime-based approach to decentralised resource allocation in distributed systems. Future Generation Computer Systems, 9(3):241–257, September 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vishwanath:2009:ATC

[VBLS09] Venkatram Vishwanath, Robert Burns, Jason Leigh, and Michael Seablom. Accelerating tropical cyclone analysis using LambdaRAM, a distributed data cache over wide-area ultra-fast networks. Future Generation Computer Systems, 25(2):184– 191, February 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vicat-Blanc-Primet:2003:GHP

[VBP03] Pascale Vicat-Blanc-Primet. Grid high performance networking in the DataGRID project. Future Generation Computer Systems, 19(2): 199–208, February 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vanderHam:2006:URD

[vdHDT^{+}06] Jeroen J. van der Ham, Freek Dijkstra, Franco Travostino, Hubertus M. A. Andree, and Cees T. A. M. de Laat. Using RDF to describe networks. Future Generation Computer Systems, 22(8):862–867, October 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vanderster:2009:RAC

[VDPHS09] Daniel C. Vanderster, Nikitas J. Dimopoulos, Rafael Parra-Hernandez, and Randall J. Sobie. Resource allocation on computational grids using a utility model and the knapsack problem. Future Generation Computer Systems, 25(1):35–50, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1985:W

[vdR85] R. P. van de Riet. Welcome. Future Generation Computer Systems, 1(3):151, February 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1986:EDS

[vdR86a] R. P. van de Riet. Expert database systems. Future Generation Computer Systems, 2(3):191–199, September 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1986:PPT

[vdR86b] R. P. van de Riet. Parallel processing, the technology of fifth generation computers: R. K. Miller Published by: SEAI, Technical Publications, Madison, Georgia, USA. Number of pages: 123. Future Generation Computer Systems, 2 (3):203–204, September 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:AAIb

[vdR87a] R. P. van de Riet. Advances in Artificial Intelligence: by J. Hallam and C. Mellish, Eds. (Wiley, New York, 1987) 290 pp., price: £25. Future Generation Computer Systems, 3 (3):223, September 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:AAIa

[vdR87b] R. P. van de Riet. Advances in artificial intelligence CIIAM86: (Kogan Page, London, 1987) 328 pp., price: £40. Future Generation Computer Systems, 3(3): 222, September 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:AII

[vdR87c] R. P. van de Riet. Artificial intelligence II. Methodology, systems, applications: by Ph. Jorrand and V. Sgurev, Eds. (North-Holland, Amsterdam, 1987) 403 pp. Future Generation Computer Systems, 3 (3):219–220, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:ESS

[vdR87d] R. P. van de Riet. Expert systems: Strategic implications and applications: by A. C. Beerel (Ellis Horwood and Wiley, Chichester, 1987) 173 pp., price: £22.50. Future Generation Computer Systems, 3 (3):220–221, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:OBT

[vdR87e] R. P. van de Riet. From optical bistability towards optical computing, the EJOB project: by P. Mandel, S. D. Smith and B. S. Wherrett, Eds. (North-Holland, Amsterdam, 1987) 362 pp. Future Generation Computer Systems, 3 (3):225, September 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:FSD

[vdR87f] R. P. van de Riet. Fuzzy sets:

Decision making, and expert systems: by H. J. Zimmermann (Kluwer Academic Publishers, Dordrecht, 1987) 335 pp., price: 57.50. Future Generation Computer Systems, 3 (3):221–222, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:KSP

[vdR87g] R. P. van de Riet. Knowledge systems and Prolog: by A. Walker, M. McCord, J. F. Sowa and W. G. Wilson (Addison-Wesley, Reading, MA, 1987) 475 pp. Future Generation Computer Systems, 3(3):217–218, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:LPF

[vdR87h] R. P. van de Riet. Logic programming functions, relations and equations: by D. de Groot and G. Lindstrom (Prentice-Hall, Englewood Cliffs, NJ, 1986) 533 pp. Future Generation Computer Systems, 3 (3):218–219, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:PPF

[vdR87i] R. P. van de Riet. Prolog for programmers: by F. Kluzniak and S. Szpakowicz [and J. St Bień] (Academic Press, New York, 1985) 299 pp., price: £14.95. Future Generation Computer Systems, 3 (3):221, September 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:PPT

[vdR87j] R. P. van de Riet. Prolog programming for tomorrow: by J. Doores, A. R. Reiblein and S. Vadera (Sigma Press, 1987) 149 pp., price: £10.95. Future Generation Computer Systems, 3(3):223– 224, September 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:SVD

[vdR87k] R. P. van de Riet. Statistics with vague data: by R. Kruse and K. D. Meyer (Reidel, Dordrecht/Boston, 1987) 279 pp. Future Generation Computer Systems, 3(3):224, September 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1987:IRA

[vdR87l] Reind P. van de Riet. An impression of the research activities of MCC in the area of data- and knowledge base systems — a travel report. Future Generation Computer Systems, 3(2):137–145, May 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1993:GEF

[vdR93a] R. P. van de Riet. Guest editorial: Fifth generation computer systems: Success or failure? Future Generation Computer Systems, 9 (2):79–81, July 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeRiet:1993:OAF

[vdR93b] R. P. van de Riet. An overview and appraisal of the Fifth Generation Computer System project. Future Generation Computer Systems, 9 (2):83–103, July 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vanderSman:2004:DUT

[vdS04] R. G. M. van der Sman. Diffusion on unstructured triangular grids using Lattice Boltzmann. Future Generation Computer Systems, 20(6): 965–971, August 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vandeVorst:1989:SLS

[vdV89a] Johannes G. G. van de Vorst. Solving the least squares problem using a parallel linear algebra library. Future Generation Computer Systems, 4(4): 293–297, March 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vanderVorst:1989:PAP

[vdV89b] Henk A. van der Vorst. Practical aspects of parallel scientific computing. Future Generation Computer Systems, 4 (4):285–291, March 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vengerov:2008:GBR

[Ven08] David Vengerov. A gradientbased reinforcement learning approach to dynamic pricing in partially-observable environments. Future Generation Computer Systems, 24 (7):687–693, July 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vengerov:2009:RLF

[Ven09] David Vengerov. A reinforcement learning framework for utility-based scheduling in resource-constrained systems. Future Generation Computer Systems, 25(7):728–736, July 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vuik:2001:CGA

[VF01] C. Vuik and J. Frank.

Coarse grid acceleration of a parallel block preconditioner. Future Generation Computer Systems, 17(8): 933–940, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/29/abstract.
html.

Vuik:2001:PBP

[VFS01] C. Vuik, J. Frank, and A. Segal. A parallel blockpreconditioned GCR method for incompressible flow problems. Future Generation Computer Systems, 18(1):31– 40, September 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/27/30/abstract.
html.

Vega-Gorgojo:2006:SAD

[VGBLGS^{+}06] Guillermo Vega-Gorgojo, Miguel L.

Bote-Lorenzo, Eduardo Gómez-Sánchez, Yannis A. Dimitriadis, and Juan I. Asensio-Pérez. A semantic approach to discovering learning services in Grid-based collaborative systems. Future Generation Computer Systems, 22 (6):709–719, May 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Venkataraman:2003:KFS

[VLC03] Shalini Venkataraman, Jason Leigh, and Tom Coffin. Kites flying in and out of space — distributed physically based art on the grid. Future Generation Computer Systems, 19(6):973–982, August 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Varvarigou:2009:SSR

[VLK09] Theodora Varvarigou, Antonios Litke, and Dimosthenis Kyriazis. Special section: Real-time attributes in Grids. Future Generation Computer Systems, 25(7):756–757, July 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vo:1993:UCB

[VM93] Dinh Phuoc Vo and David Macchion. A use of case-based reasoning technique in building expert systems. Future Generation Computer Systems, 9(4):311–319, December 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vanMourik:1994:VPM

[vM94] P. A. van Mourik. Vectorization and parallelization of a multi-block Navier–Stokes flow solver on different computer architectures. Future Generation Computer Systems, 10(2–3):327–329, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

VanLiere:1997:CS

[VMvW97] Robert Van Liere, Jurriaan D.

Mulder, and Jarke J. van Wijk. Computational steering. Future Generation Computer Systems, 12(5):441– 450, April 1, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/17/26/abstract.
html.

Visser:2001:ETS

[VN01] Arnoud Visser and Kai Nagel.

Editorial: Traffic simulation. Future Generation Computer Systems, 17(5): 625–626, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/34/abstract.
html.

vanOpheusden:1995:IFC

[vOB95] J. H. J. van Opheusden and M. T. A. Bos. Induced flocculation of casein micelles: a Brownian dynamics simulation on the Parsytec GCel MPP. Future Generation Computer Systems, 11(2): 123–133, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vanOudenaarde:2005:DPM

[vOHD^{+}05] S. van Oudenaarde, Z. Hendrikse, F. Dijkstra, L. Gommans, C. de Laat, and R. J. Meijer. Dynamic paths in multi-domain optical networks for grids. Future Generation Computer Systems, 21 (4):539–548, April 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Violard:1994:PSU

[VP94] Eric Violard and Guy-René Perrin. PEI: a simple unifying model to design parallel programs. Future Generation Computer Systems, 10(2– 3):269–272, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Veit:2000:FDP

[VR00] Holger Veit and Gernot Richter. The FTA design paradigm for distributed systems. Future Generation Computer Systems, 16(6): 727–740, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/29/38/abstract.
                                        html.

Varadarajan:2005:NRS

[VR05] Srinidhi Varadarajan and Naren Ramakrishnan. Novel runtime systems support for adaptive compositional modeling in PSEs. Future Generation Computer Systems, 21 (6):878–895, June 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vree:1988:ECG

[Vre88] W. G. Vree. Experiments with coarse-grain parallel graph reduction. Future Generation Computer Systems, 4(??):299– 306, ???? 1988. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vree:1989:ECG

[Vre89] W. G. Vree. Experiments with coarse-grain parallel graph reduction. Future Generation Computer Systems, 4(4): 299–306, March 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vecchia:1988:RSN

[VS88] G. Della Vecchia and C. Sanges.

A recursively scalable network VLSI implementation. Future Generation Computer Systems, 4(3):235–243, October 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vecchia:1990:OBT

[VS90] G. Della Vecchia and C. Sanges.

An optimized broadcasting technique for WK-recursive topologies. Future Generation Computer Systems, 5(4): 353–357, January 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vazquez:2004:BRR

[VS04] Pere-Pau Vázquez and Mateu Sbert. Bandwidth reduction for remote navigation systems through view prediction and progressive transmission. Future Generation Computer Systems, 20(8):1251– 1262, November 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vazhkudai:2002:PDI

[VSM02] Sudharshan Vazhkudai, Jeelani Syed, and Tobin Maginnis. PODOS — the design and implementation of a performance oriented Linux cluster. Future Generation Computer Systems, 18(3): 335–352, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/29/abstract.
html.

Vuurpijl:1995:PPL

[VSV95] Louis Vuurpijl, Theo Schouten, and Jan Vytopil. Performance prediction of large MIMD systems for parallel neural network simulations. Future Generation Computer Systems, 11 (2):221–232, March 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Voogd:1994:CS

[VSvD94] J. M. Voogd, P. M. A. Sloot, and R. van Dantzig. Crystallization on a sphere. Future Generation Computer Systems, 10(2–3):359–361, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Voogd:1995:CVP

[VSvD95] J. M. Voogd, P. M. A.

Sloot, and R. van Dantzig. Comparison of vector and parallel implementations of the simulated annealing algorithm. Future Generation Computer Systems, 11(4–5): 467–475, August 1995. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
fwi.uva.nl/fwi/research/
vg4/pwrs/papers/archive/
Voogd95_1.ps.gz.

vanVliet:1985:SS

[vV85] J. C. van Vliet. STARS and stripes. Future Generation Computer Systems, 1(6):411– 416, December 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Vasell:1992:FPD

[VV92] Jesper Vasell and Jonas Vasell. The function processor: a data-driven processor array for irregular computations. Future Generation Computer Systems, 8 (4):321–335, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.ce.chalmers.
se/Documents/fgcs.ps.Z.

VandenBerghe:2003:DPB

[VVC^{+}03] S. Van den Berghe, P. Van Heuven, J. Coppens, F. De Turck, and P. Demeester. Distributed policy-based management of measurement-based traffic engineering: design and implementation. Future Generation Computer Systems, 19 (2):291–302, February 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vanVeen:1998:NTV

[vVDBB98] Hendrik A. H. C. van Veen, Hartwig K. Distler, Stephan J. Braun, and Heinrich H. Bülthoff. Navigating through a virtual city: Using virtual reality technology to study human action and perception. Future Generation Computer Systems, 14 (3–4):231–242, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/29/18/27/28/
index.htt;   http://www.
elsevier.com/gej-ng/10/

19/19/29/18/27/abstract.
html.

Verstappen:1994:DNS

[VWCV94] R. Verstappen, J. G. Wissink, W. Cazemier, and A. E. P. Veldman. Direct numerical simulations of turbulent flow in a driven cavity. Future Generation Computer Systems, 10(2–3):345–350, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Volckaert:2008:GCA

[VWD^{+}08] Bruno Volckaert, Tim Wauters, Marc De Leenheer, Pieter Thysebaert, Filip De Turck, Bart Dhoedt, and Piet Demeester. Gridification of collaborative audiovisual organizations through the MediaGRID framework. Future Generation Computer Systems, 24(5):371–389, May 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wilson:1989:ISC

[WAD^{+}89] William D. Wilson, Robert J.

Asaro, Robert W. Dutton, Juan M. Sanchez, David J. Srolovitz, Richard H. Boyd III, William A. Goddard, John R. Smith, and Wilhelm G. Wolfer. The impact of supercomputing capabilities on U.S. materials science and technology. Future Generation Computer Systems, 5(2– 3):283–293, September 1989.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wieners:2006:DPO

[WAE06] Christian Wieners, Martin Ammann, and Wolfgang Ehlers. Distributed Point Objects: a new concept for parallel finite elements applied to a geomechanical problem. Future Generation Computer Systems, 22(4): 532–545, March 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wahlster:1984:CAS

[Wah84] W. Wahlster. Cooperative access systems. Future Generation Computer Systems, 1(2): 103–111, November 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Walker:1986:KRT

[Wal86] Donald E. Walker. Knowledge resource tools for information access. Future Generation Computer Systems, 2 (3):161–171, September 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wallace:1994:HEP

[Wal94] D. J. Wallace. HPCN in Europe: a personal perspective. Future Generation Computer Systems, 10(2–3):

153–158, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wybranietz:1990:LPE

[WB90] Dieter Wybranietz and Peter Buhler. The LADY programming environment for distributed operating systems. Future Generation Computer Systems, 6(3):209– 223, December 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wismuller:2008:HLA

[WBF08] Roland Wismüller, Marian Bubak, and Włodzimierz Funika. High-level applicationspecific performance analysis using the G-PM tool. Future Generation Computer Systems, 24(2):121–132, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Williams:1999:FQR

[WBMP99] Roy Williams, Julian Bunn, Reagan Moore, and James C. T. Pool. Findings, questions and recommendations from the ISDA workshop. Future Generation Computer Systems, 16(1):1– 8, November 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        19/19/41/32/26/abstract.
                                        html.

Wohrer:2005:NMA

[WBT05] Alexander Wöhrer, Peter Brezany, and A. Min Tjoa. Novel mediator architectures for Grid information systems. Future Generation Computer Systems, 21(1): 107–114, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wozniak:2008:MBB

[WBT^{+}08] J. M. Wozniak, P. Brenner, D. Thain, A. Striegel, and J. A. Izaguirre. Making the best of a bad situation: Prioritized storage management in GEMS. Future Generation Computer Systems, 24(1): 10–16, January 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Walshaw:2001:MMP

[WC01] C. Walshaw and M. Cross. Multilevel mesh partitioning for heterogeneous communication networks. Future Generation Computer Systems, 17 (5):601–623, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/30/33/abstract.
                                        html.

Wang:2006:CTG

[WC06a] Libin Wang and Kefei Chen.

Comments on a theorem on Grid access control. Future Generation Computer Systems, 22(4):381–384, March 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2006:CAC

[WC06b] Libin Wang and Kefei Chen.

Comments on an access control model in Semantic Grid. Future Generation Computer Systems, 22(1–2):3–5, January 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wu:2009:WBR

[WCC^{+}09] BangYu Wu, Chi-Hung Chi, Zhe Chen, Ming Gu, and JiaGuang Sun. Workflow-based resource allocation to optimize overall performance of composite services. Future Generation Computer Systems, 25 (3):199–212, March 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wismuller:2000:EMG

[WDD00] Roland Wismüller, Gábor Dózsa, and Dániel Drótos. Enhanced monitoring in the GRADE programming environment by using OMIS. Future Generation Computer Systems, 16(6):637– 648, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/31/abstract.
html.

Wesche:1999:TDV

[Wes99]  Gerold Wesche.  Three-

dimensional visualization of fluid dynamics on the Responsive Workbench. Future Generation Computer Systems, 15 (4):469–475, July 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/20/17/abstract.
html.

Wei:2007:TED

[WFC07] Baohua Wei, Gilles Fedak, and Franck Cappello. Towards efficient data distribution on computational desktop grids with BitTorrent. Future Generation Computer Systems, 23 (8):983–989, November 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wiendahl:1991:TCI

[WG91] H.-P. Wiendahl and R. Garlichs. Trends in computerintegrated manufacturing. Future Generation Computer Systems, 7(1):97–107, October 1991. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Walton:2000:AMM

[WG00] Chris Walton and Dilsun Kirliand Stephen Gilmore. An abstract machine model of dynamic module replacement. Future Generation Computer Systems, 16(7): 793–808, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/30/29/abstract.
html.

Waite:1992:PAC

[WGL92] Martin E. Waite, Bret Giddings, and Simon H. Lavington. Parallel associative combinator evaluation II. Future Generation Computer Systems, 8(4):303–319, September 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Walker:2005:CPS

[WH05] David W. Walker and Elias Houstis. Complex problemsolving environments for Grid computing. Future Generation Computer Systems, 21(6): 841–842, June 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wieczorek:2009:TGM

[WHP09] Marek Wieczorek, Andreas Hoheisel, and Radu Prodan. Towards a general model of the multi-criteria workflow scheduling on the grid. Future Generation Computer Systems, 25(3):237–256, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wiederhold:1985:KB

[Wie85]  Gio Wiederhold.  Knowl-

edge bases. Future Generation Computer Systems, 1(4): 223–235, June 1985. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wiechert:2003:RMC

[Wie03] W. Wiechert. The role of modeling in computational science education. Future Generation Computer Systems, 19(8): 1363–1374, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wiig:1984:MTA

[Wii84] K. Wiig. Market trends in artificial intelligence in the United States and Japan. Future Generation Computer Systems, 1(2):113– 118, November 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wilmot:1986:MP

[Wil86] R. W. Wilmot. The market perspective. Future Generation Computer Systems, 2 (1):27–31, March 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wilson:1989:GCC

[Wil89] Kenneth G. Wilson. Grand challenges to computational science. Future Generation Computer Systems, 5(2–3): 171–189, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Williams:2000:GEH

[Wil00] Roy Williams. Guest editorial: High Performance Computing and Networking Europe 1999. Future Generation Computer Systems, 16 (5):v–vi, March 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/28/25/abstract.
html.

Witten:1994:RHP

[Wit94] Matthew Witten. The role of high performance computing in medicine and public health. Future Generation Computer Systems, 10(2–3): 223–232, June 1994. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wilde:2003:IVC

[WKF03] Torsten Wilde, James A.

Kohl, and Raymond E. Flanery, Jr. Immersive and 3D viewers for CUMULVS: VTK/ CAVE^{T M} and AVS/Express. Future Generation Computer Systems, 19(5):701–719, July 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wernhart:2000:RCD

[WKT00] Heidemarie Wernhart, Eva Kühn, and Georg Trausmuth. The replicator coordination design pattern. Future Generation Computer Systems, 16 (6):693–703, April 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/29/35/abstract.
html.

Wang:2003:GLP

[WKZ^{+}03] Kai Wang, Sang-Bae Kim, Jun Zhang, Kengo Nakajima, and Hiroshi Okuda. Global and localized parallel preconditioning techniques for large scale solid Earth simulations. Future Generation Computer Systems, 19(4):443–456, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Weng:2005:HSB

[WL05] Chuliang Weng and Xinda Lu. Heuristic scheduling for bagof-tasks applications in combination with QoS in the computational grid. Future Generation Computer Systems, 21 (2):271–280, February 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wagner:2000:AAN

[WLB00] Israel A. Wagner, Michael Lindenbaum, and Alfred M. Bruckstein. ANTS: Agents on Networks, Trees, and Subgraphs. Future Generation Computer Systems, 16(8): 915–926, June 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/31/29/abstract.
html.

Wang:2009:APS

[WLF^{+}09] Liqiang Wang, Shiyong Lu, Xubo Fei, Artem Chebotko, H. Victoria Bryant, and Jeffrey L. Ram. Atomicity and provenance support for pipelined scientific workflows. Future Generation Computer Systems, 25(5):568–576, May 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wurst:2007:DFE

[WM07] Michael Wurst and Katharina Morik. Distributed feature extraction in a P2P setting — a case study. Future Generation Computer Systems, 23(1): 69–75, January 1, 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wasiewicz:2001:DCI

[WMN^{+}01] Piotr Wa̧siewicz, Artur Malinowski, Robert Nowak, Jan J. Mulawka, Piotr Borsuk, Piotr Wȩgleński, and Andrzej Płucienniczak. DNA computing: Implementation of data flow logical operations. Future Generation Computer Systems, 17(4): 361–378, January 1, 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/29/27/abstract.
html.

Worsch:1999:SCA

[Wor99] Thomas Worsch. Simulation of cellular automata. Future Generation Computer Systems, 16(2–3):157– 170, December 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/41/25/26/abstract.
html.

Wille:1994:TFP

[WRBG94] L. T. Wille, J. L. Rogers, C. P. Burmester, and R. Gronsky. Towards first-principles theories of materials and biological systems — The need for massive parallelism. Future Generation Computer Systems, 10(2–3):331–338, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wendler:2005:EOC

[WS05] Jan Wendler and Florian Schintke. Executing and observing CFD applications on the Grid. Future Generation Computer Systems, 21(1):11– 18, January 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wolski:1999:NWS

[WSH99] Rich Wolski, Neil T. Spring, and Jim Hayes. The network weather service: a distributed resource performance forecasting service for metacomputing. Future Generation Computer Systems, 15(5– 6):757–768, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/35/abstract.
html.

Wegener:2009:GRB

[WSS^{+}09] Dennis Wegener, Thierry Sengstag, Stelios Sfakianakis, Stefan Rüping, and Anthony Assi. GridR: An R-based tool for scientific data analysis in grid environments. Future Generation Computer Systems, 25(4):481–488, April 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wherrett:1987:OCD

[WSTW87] Brian S. Wherrett, S. Desmond Smith, Frank A. P. Tooley, and Andrew C. Walker. Optical components for digital optical circuits. Future Generation Computer Systems, 3(4): 253–259, December 1987. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2002:DPC

[WTC^{+}02] Cho-Li Wang, Anthony T. C. Tam, Benny W. L. Cheung, Wenzhang Zhu, and David C. M. Lee. Directed Point: a communication subsystem for commodity supercomputing with Gigabit Ethernet. Future Generation Computer Systems, 18(3): 401–420, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/33/abstract.
html.

Wang:2007:DPP

[WTK07] Hong Wang, Hiroyuki Takizawa, and Hiroaki Kobayashi. A dependable Peer-to-Peer computing platform. Future Generation Computer Systems, 23(8):939–955, November 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wachter:2005:LSN

[WVC05] Andreas Wächter, Chandu Visweswariah, and Andrew R. Conn. Large-scale nonlinear optimization in circuit tuning. Future Generation Computer Systems, 21(8):1251– 1262, October 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:1997:USC

[WWC^{+}97] Fu-Ching Wang, Chun-Hung Wen, Chih-Yuan Cheng, Meng-Huang Lee, Tzu-How Lin, Szu-Chi Wang, and Yen-Jen Oyang. Upgrading the service capacity of video-on-demand servers with memory buffer. Future Generation Computer Systems, 12(6): 565–577, June 15, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/27/18/21/abstract.
html.

Ward:1998:DAS

[WWSM98] J. W. Ward, D. P. M. Wills, K. P. Sherman, and A. M. M. A. Mohsen. The development of an arthroscopic surgical simulator with haptic feedback. Future Generation Computer Systems, 14 (3–4):243–251, July 31, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/29/18/28/28/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/29/18/28/abstract.
html.

Wang:2002:CFS

[WX02] Jianyong Wang and Zhiwei Xu. Cluster file systems: a case study. Future Generation Computer Systems, 18(3): 373–387, January 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/32/31/abstract.
                                        html.

Waheed:1999:PNB

[WYJ99] Abdul Waheed, Jerry Yan, and Haoqiang Jin. Parallelization of NAS benchmarks for shared memory multiprocessors. Future Generation Computer Systems, 15(3): 353–363, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/19/21/abstract.
                                        html.

Wu:1990:COT

[WYN^{+}90] Zhaohui Wu, Tao Yang, Feng Ni, Zhijun He, and Rui Zhao Yu. A C-oriented tool for building second generation expert systems. Future Generation Computer Systems, 6 (1):77–83, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wang:2008:ACM

[WZC08] Hua Wang, Yanchun Zhang, and Jinli Cao. Access control management for ubiquitous computing. Future Generation Computer Systems, 24(8): 870–878, October 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xuejun:1990:PSS

[XHY^{+}90] Yang Xuejun, Chen Haibo, Ci Yungui, Chen Fujie, and Chen Lijie. Processor selfscheduling for parallel loops in preemptive environments. Future Generation Computer Systems, 6(1):97–103, June 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xiang:2006:FTR

[Xia06] Dong Xiang. Fault-tolerant routing in hypercubes using partial path set-up. Future Generation Computer Systems, 22(7):812–819, August 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xue:2004:G

[XSJ04] Yong Xue, Xiangyu Sheng, and Narayana Jayaram. Geocomputation. Future Generation Computer Systems, 20(7): 1097–1099, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xue:2004:RTD

[XSM04] Yong Xue, Min Sun, and Ainai Ma. On the reconstruction of three-dimensional complex geological objects using Delaunay triangulation. Future Generation Computer Systems, 20(7):1227– 1234, October 1, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xie:2005:RTR

[XYZ05] Kai Xie, Jie Yang, and Y. M. Zhu. Real-time rendering of 3D medical data sets. Future Generation Computer Systems, 21(4):573–581, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yang:2002:DDS

[YA02] Shuang H. Yang and James L. Alty. Development of a distributed simulator for control experiments through the Internet. Future Generation Computer Systems, 18(5): 595–611, April ??, 2002. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/36/28/abstract.
html.

Yanami:2007:MPS

[YA07] Hitoshi Yanami and Hirokazu Anai. The Maple package SyNRAC. Future Generation Computer Systems, 23(5): 721–726, June 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yamaguchi:1989:TAE

[Yam89] Takahira Yamaguchi. A technical analysis expert system in the stock market. Future Generation Computer Systems, 5 (1):21–27, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yamanouchi:1992:PEP

[Yam92] Nagatsugu Yamanouchi. Performance effects of program structures on a snoop-cached multiprocessor system. Future Generation Computer Systems, 7(2–3):247–258, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yatagai:1988:OCJ

[Yat88] Toyohiko Yatagai. Optical computing in Japan. Future Generation Computer Systems, 4(3):177–187, October 1988. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yang:2007:DRA

[YBQ07] Juan Yang, Yun Bai, and Yuhui Qiu. A decentralized resource allocation policy in minigrid. Future Generation Computer Systems, 23(3): 359–366, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yaun:2003:OPS

[YCAS03] Garrett Yaun, Christopher D. Carothers, Sibel Adali, and David Spooner. Optimistic parallel simulation of a largescale view storage system. Future Generation Computer Systems, 19(4):479–492, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yu:2005:PGB

[YCX05] Wei Yu, Sriram Chellappan, and Dong Xuan. P2P/Gridbased overlay architecture to support VoIP services in largescale IP networks. Future Generation Computer Systems, 21(1):209–219, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

YarKhan:2005:BSA

[YD05] Asim YarKhan and Jack J. Dongarra. Biological sequence alignment on the computational grid using the GrADS framework. Future Generation Computer Systems, 21(6): 980–986, June 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yero:2005:JIJ

[YdOLS^{+}05] Eduardo Javier Huerta Yero, Fabiano de Oliveira Lucchese, Francisco Sérgio Sambatti, Miriam von Zuben, and Marco Aurélio Amaral Henriques.

JoiN: The implementation of a Java-based massively parallel grid. Future Generation Computer Systems, 21 (5):791–810, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yang:2005:CCG

[YHJC05] X. Yang, M. Hayes, K. Jenkins, and S. Cant. The Cambridge CFD grid for largescale distributed CFD applications. Future Generation Computer Systems, 21(1):45– 51, January 1, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yang:2003:IDC

[YJA03] Ge Yang, Ruoming Jin, and Gagan Agrawal. Implementing data cube construction using a cluster middleware: algorithms, implementation experience, and performance evaluation. Future Generation Computer Systems, 19 (4):533–550, May 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yan:2007:ASL

[YKL^{+}07] Jun Yan, Ryszard Kowalczyk, Jian Lin, Mohan B. Chhetri, Suk Keong Goh, and Jianying Zhang. Autonomous service level agreement negotiation for service composition provision. Future Generation Computer Systems, 23 (6):748–759, July 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yu:2006:MDL

[YLC^{+}06] Oliver Yu, Anfei Li, Yuan Cao, Leping Yin, Ming Liao, and Huan Xu. Multi-domain Lambda Grid data portal for collaborative Grid applications. Future Generation Computer Systems, 22(8): 993–1003, October 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yoshida:2000:WEK

[YMM00] Ryo Yoshida, Takaaki Murao, and Tatsuo Miyazawa. 3D Web environment for knowledge management. Future Generation Computer Systems, 17(1):73–78, September 2000. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/24/33/abstract.
                                        html.

Yoshimura:1989:RBA

[Yos89] Takeshi Yoshimura. A rulebased and algorithmic approach for logic synthesis. Future Generation Computer Systems, 5(1):123–127, August 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Youn:2005:BPS

[YPF05] Choonhan Youn, Marlon E.

Pierce, and Geoffrey C. Fox. Building problem-solving environments with Application Web Service toolkits. Future Generation Computer Systems, 21(6):856–867, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yamaguchi:1984:ELB

[YTHY84] Yoshinori Yamaguchi, Kenji Toda, Jayanta Herath, and Toshitsugu Yuba. EM-3: a Lisp-based data-driven machine. Future Generation Computer Systems, ??(??): 524–532, November 1984. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yui:1989:AES

[YWA^{+}89] Katsuhiko Yui, Satoshi Watanabe, Shigeru Amano, Tsuyoshi Takarabe, and Takashi Nakamori. Application of an expert system to blast furnace operation. Future Generation Computer Systems, 5(1):137– 142, August 1989. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yang:2009:RAC

[YYW^{+}09] Chao-Tung Yang, I-Hsien Yang, Shih-Yu Wang, Ching-Hsien Hsu, and Kuan-Ching Li. A Recursively-Adjusting Co-allocation scheme with a Cyber-Transformer in Data Grids. Future Generation Computer Systems, 25(7): 695–703, July 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zadeh:1987:MMP

[Zad87] Lotfi A. Zadeh. Mind over machine: The power of human intuition and expertise in the era of the computer: Hubert L. Dreyfus and Stuart E. Dreyfus. Published by: New York, The Free Press, 1986. Future Generation Computer Systems, 3(2):149–151, May 1987. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zudilova:2005:SSI

[ZAP05] Elena Zudilova, Tony Adriaansen, and Binh Pham. Special section: “Interaction and visualisation techniques for problem solving environments”. Future Generation Computer Systems, 21(7): 1143–1144, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhao:2009:SSW

[ZBB09] Zhiming Zhao, Adam Belloum, and Marian Bubak. Special section on workflow systems and applications in e-science. Future Generation Computer Systems, 25 (5):525–527, May 2009. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zeng:2004:CDD

[ZCT^{+}04] Yi Zeng, Wentong Cai, Stephen J. Turner, Suiping Zhou, and Bu-Sung Lee. Characterization and delivery of directly coupled causal messages in distributed systems. Future Generation Computer Systems, 20(1): 171–178, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhu:2004:PDM

[ZCW^{+}04] Huabing Zhu, Tony Kai-Yun Chan, Lizhe Wang, Wentong Cai, and Simon See. A prototype of distributed molecular visualization on computational grids. Future Generation Computer Systems, 20(5): 727–737, June 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zou:2007:DLK

[ZDR07] Xukai Zou, Yuan-Shun Dai, and Xiang Ran. Dual-Level Key Management for secure Grid communication in dynamic and hierarchical groups. Future Generation Computer Systems, 23(6):776–786, July 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zemanek:1986:BES

[Zem86] Heinz Zemanek. Building expert systems: edited by F. Hayes-Roth, D. A. Waterman, D. B. Lenat (Addison Wesley, 1983, xvi + 443 pp.). Future Generation Computer Systems, 2(4):262– 263, December 1986. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zomaya:1998:BIS

[ZEO98] Albert Y. Zomaya, Fikret Ercal, and Stephan Olariu. Bioinspired solutions to parallel processing problems. Future Generation Computer Systems, 14(5–6):271–273, December 1998. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zomaya:2001:GEP

[ZEO01] Albert Y. Zomaya, Fikret Ercal, and Stephan Olariu. Guest editorial: Parallel computing problems and natureinspired solutions. Future Generation Computer Systems, 17(4):v–vii, January 1, 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/29/24/abstract.
                                        html.

Ziavras:2000:NGP

[ZGCM00] Sotirios G. Ziavras, Haim Grebel, Anthony T. Chronopoulos, and Florent Marcelli. A new-generation parallel computer and its performance evaluation. Future Generation Computer Systems, 17(3): 315–333, November 1, 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/31/abstract.
html.

Zhang:1993:EPL

[Zha93] Kang Zhang. Exploiting ORparallelism in logic programs: a review. Future Generation Computer Systems, 9(3):259– 280, September 1993. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2004:SRG

[Zhu04] Hai Zhuge. Semantics, resource and Grid. Future Generation Computer Systems, 20 (1):1–5, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2007:SSS

[Zhu07] Hai Zhuge. Special section:

Semantic Grid and Knowledge Grid. Future Generation Computer Systems, 23(3): 281–282, March 2007. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zinzani:2000:VCP

[Zin00]  Carlo Zinzani.  Virtual

computer project: “Delivering content in a VRML world”. Future Generation Computer Systems, 17 (1):27–37, September 2000. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/19/19/45/24/28/41/
index.htt;   http://www.
elsevier.com/gej-ng/10/
19/19/45/24/28/abstract.
html.

Zhou:2004:TCO

[ZJWZ04] Feng Zhou, Chao Jin, Yinghui Wu, and Weimin Zheng. TODS: cluster object storage platform designed for scalable services. Future Generation Computer Systems, 20(4): 549–563, May 3, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2004:SPB

[ZL04a] Hai Zhuge and Yanyan Li. Semantic profile-based document logistics for cooperative research. Future Generation Computer Systems, 20(1):47– 60, January 15, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2004:FCA

[ZL04b] Hai Zhuge and Jie Liu. A fuzzy collaborative assessment approach for Knowledge Grid. Future Generation Computer Systems, 20(1): 101–111, January 15, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:2003:TDV

[ZLD^{+}03] Chong Zhang, Jason Leigh, Thomas A. DeFanti, Marco Mazzucco, and Robert Grossman. TeraScope: distributed visual data mining of terascale data sets over photonic networks. Future Generation Computer Systems, 19(6): 935–943, August 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zait:1997:CSC

[ZM97] Mohamed Zaït and Hammou Messatfa. A comparative study of clustering methods. Future Generation Computer Systems, 13(2–3):149– 159, November 14, 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/28/18/20/abstract.
html.

Zingirian:1999:ESS

[ZMN99] N. Zingirian, M. Maresca, and S. Nalin. Efficiency of standard software architectures for Java-based access to remote databases. Future Generation Computer Systems, 15(3): 417–424, April 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/26/abstract.
html.

Zhang:2006:ELB

[ZMS^{+}06] Feng Zhang, Andryas Mawardi, Eugene Santos, Jr., Ranga Pitchumani, and Luke E. K. Achenie. Examination of loadbalancing methods to improve efficiency of a composite materials manufacturing process simulation under uncertainty using distributed computing. Future Generation Computer Systems, 22(5):571–587, April 2006. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Znaty:1994:BEW

[Zna94] E. Znaty. BECAUSE European workshop Sophia-Antipolis (France) 14–16 October 1992. General presentation of the BECAUSE Benchmark Set: BBS. Future Generation Computer Systems, 10 (4):365–379, November 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:2009:EPP

[ZQZZ09] Rong Zhang, Weining Qian, Aoying Zhou, and Minqi Zhou. An efficient peer-to-peer indexing tree structure for multidimensional data. Future Generation Computer Systems, 25(1):77–88, January 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhiyi:1990:CAE

[ZS90] Hwang Zhiyi and Hu Shouren.

A compiling approach for exploiting AND-parallelism in logic programs. Future Generation Computer Systems, 6 (1):35–41, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zanikolas:2005:TGM

[ZS05a] Serafeim Zanikolas and Rizos Sakellariou. A taxonomy of grid monitoring systems. Future Generation Computer Systems, 21 (1):163–188, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zudilova:2005:BCI

[ZS05b] E. V. Zudilova and P. M. A.

Sloot. Bringing combined interaction to a problem solving environment for vascular reconstruction. Future Generation Computer Systems, 21 (7):1167–1176, July 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:2008:PTR

[ZSI08] Yuanyuan Zhang, Wei Sun, and Yasushi Inoguchi. Predict task running time in grid environments based on CPU load predictions. Future Generation Computer Systems, 24 (6):489–497, June 2008. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:1990:DDM

[ZT90] K. Zhang and R. Thomas. DIALOG: a dataflow model for parallel execution of logic programs. Future Generation Computer Systems, 6(??): 373–388, ???? 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:1991:DDM

[ZT91] Kang Zhang and Ray Thomas. DIALOG — a dataflow model for parallel execution of logic programs. Future Generation Computer Systems, 6(4):373– 388, September 1991. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhou:2004:AUO

[ZWJ04] Wanlei Zhou, Li Wang, and Weijia Jia. An analysis of update ordering in distributed replication systems. Future Generation Computer Systems, 20(4):565–590, May 3, 2004. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic).

Zhiying:1990:DER

[ZY90] Wang Zhiying and Ci Yungui.

Design and evaluation of a relational knowledge base prototype machine. Future Generation Computer Systems, 6 (1):71–75, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:2004:ACS

[ZY04] Jian J. Zhang and Lihua You.

Analytical C ^{2} smooth blending surfaces. Future Generation Computer Systems, 20(8): 1317–1326, November 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhuge:2005:ERS

[ZYXL05] Hai Zhuge, Erlin Yao, Yunpeng Xing, and Jie Liu. Extended resource space model. Future Generation Computer Systems, 21(1):189–198, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhang:1990:HSH

[ZZ90] Bo Zhang and Ling Zhang. Hierarchy and statistical heuristic search. Future Generation Computer Systems, 6 (1):43–47, June 1990. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zhou:2009:MGS

[ZZ09] Jing Zhou and Guosun Zeng. A mechanism for grid service composition behavior specification and verification. Future Generation Computer Systems, 25(3):378–383, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Zheng:2004:LDT

[ZZN04] Hong Zheng, Jingxin Zhang, and Saeid Nahavandi. Learning to detect texture objects by artificial immune approaches. Future Generation Computer Systems, 20(7): 1197–1208, October 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).